A Complete Bibliography of Publications in Technology and Culture: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 September 2023 Version 1.00

Title word cross-reference

e [But92, Dre95, Lav93, Smi92a, vR90].

100 [Lai96a]. 14 [Dou91a]. 14th [Mue91b]. 14th-Century [Mue91b]. 15-18 [Ano99m]. 150th [Lup91]. 1571 [Bac90]. 1600 [Por98]. 1620 [HV98]. 16th [Lon91b]. 16th-Century [Lon91b]. 1715 [Lyn98]. 1718 [Mac96]. 1750 [War98a]. 1770s [She94b]. 1793 [Lan94b]. 1795 [Kip97]. 17th [Arr94]. 1800 [Bed91, Gei90, Haf90, HD90, Rob90b]. 1800s [Sta97b]. 1802 [Sta97a]. 1809 [Tan93]. 1815 [Nye93b, Sha95c]. 1820 [Eat90, Mau93]. 1825 [Rob93]. 1828 [Jan91]. 1830 [War92b]. 1832 [Gli94]. 1840 [Kah96, Lin94b]. 1848 [Tod95]. 1850 [Cam95, Ler98, Mic94, O’H96, Pas96, Pet92b]. 1851 [Lan95a]. 1853 [Bro95c]. 1856 [Paq97]. 1857 [Hea92]. 1860 [Lea98, Rob92a, Win90, Ros98]. 1860er [All96a]. 1864 [Smi91a]. 1865 [Joh91, Sum93]. 1870 [Ell92b, Fit93b, Hår98a, Rab96]. 1871 [Hea91b]. 1873 [Kli91]. 1874 [Smi97a]. 1875 [Mou99]. 1876 [Har93b]. 1878 [Fri96]. 1880 [Hal94, Wor93a].

1880s [Tra91]. 1885 [Dud92]. 1889 [Hon97]. 1890 [Kea98, Myl97]. 1890s [Jac96]. 1895 [Pay95]. 1897 [Sch95]. 18th [Gui93a, Mac91, Gou98, HP91, Kry90]. 18th- [Mac91]. 18th-Century [Gou98, HP91, Kry90]. 1900 [Zac99, Hom94, Ros94b, See98a, Wen92]. 1900-talets [Har99a]. 1902 [Rei96, Rob94b]. 1903 [Cro99]. 1906 [Ser92]. 1910 [Gre91, Min95b]. 1913 [Dre97, Gro90a, Rob90a]. 1914 [All96c, Bra90, Brü94, Goo90, Gra98b, Gut95b, Jos95a, Lay91, Owe91a, Tho91a]. 1915 [Bla90b]. 1917 [Ber98, Dou95, Rho93]. 1918 [Ben92b, Bry95, McN93, Omi92, Pis96, Sch94c]. 1919 [Gus91, Til90]. 1920 [Ber97a, Egg94, Gal92a, Gal92b, Gra95, Hac92a, Leo96, Lew97, Mue91a, Ros93, Roz98, Gro98]. 1920s [Rot99a, Sti91, Lai93]. 1921 [Hol94, War91]. 1922 [New91]. 1923 [Big91]. 1924 [Mea98]. 1925 [Sim94]. 1926 [Wei94]. 1929 [Bla96b, Vol98b]. 1930 [Bra96, Bro97b, Cas98b, Cla92, Kli90, Lun90a, Mer97, Scr90, She93, Sta94, Lan99]. 1930er [All96a]. 1930s [Sti91]. 1932 [Dal90]. 1933 [Dan90, Sta90b]. 1934 [Coo98c, Kra95, McC93]. 1935 [Sch94d, Aik98]. 1936 [Oli98]. 1939 [Chr95b, Edg98, Eva96, Gar90, Hil97, Ken92, Kni95, Liu97, Rei93a, Smi92b, Rou96, Whi98]. 1940 [Tri93, Ada96, Bro93b, Con96b, Coo90, Dav96, Eis93, Lin91, Sha91, Spe95, Vol94a, vL90]. 1940s [Jac96]. 1941 [Jos96, Liv92, Min95c, Nuw90, Sch91]. 1944 [Mor93]. 1945 [All98b, Bad91, Buc97, Con96a, Fin93, Hoy92, Kay92, Mor91d, Nel92, Øve97, See95b]. 1946 [Maz90a, Kal98]. 1950 [Coo93c, Hor94, Tho98, SD98]. 1950s [Bod99]. 1951 [Kra90]. 1954 [Con94b]. 1955 [Jer94b, Tob92]. 1956 [Kid94]. 1958 [Eva93a]. 1960 [Scr91b, Sei96b, Was92]. 1960s [Bau99]. 1961 [Tra95]. 1964 [Pri98]. 1969 [Sha95a]. 1970 [Dan95, Hec92, Jer93, Les93, Mor91a, Leu98]. 1971 [Hac94a]. 1973 [Dol97]. 1974 [Sei96a]. 1975 [Gol93, Mei90, Nas90]. 1980 [Bed90a, Bir90]. 1980s [Rol93a]. 1982 [Bud96, Mer95]. 1983 [Sei93]. 1984 [Cra97, Mah90]. 1985 [Qui98, Rey90b, Sha90a, Fit98]. 1986 [Hof91, LaL90, Mis90, Cer98]. 1987 [Rey91, Yea90]. 1988 [Hau90, Pye90]. 1989 [Ano90h, CS90, Lar91]. 1990 [Ano91c, LM91, Cer97, Gui95b, Ver97, Zel97]. 1990s [She94b]. 1991 [Ano92i, Fit97]. 1992 [Ano93l, Ham98a]. 1993 [Ano94i, Cut97, Mea98, Bau99]. 1994 [Ano95i, Wor95]. 1995 [Ano96b, Gis98, Lan98b, Hur98]. 1996 [Ano97i, Hor97a]. 1997 [Ano98e]. 1998 [Ano99m]. 19th [Bro92b, Bur92, Coo95, Doo91, Sil93, Smi90a, Eve92, Ler97, Mac91, Mor99, Smi91b, Ste89, Uss91]. 19th-Century [Bur92, Eve92, Mac91, Smi91b, Ste89, Uss91].

2 [Har96b, Neu96b]. 200 [See95a]. 20th [Fel97, Nor90, Ste93]. 20th-Century [Ste93]. 21st [Kel93d]. 21st-Century [Kel93d]. 29 [Bil97]. 2d [Les94].

320 [Law92]. 34 [Car98].

’40 [Les94]. 41 [Pos98a]. 49 [Jar90].

50th [Kay91]. 586 [Hal98].

60 [Uss90]. 68 [Sal92a].

757 [Mac93a].

A.D [Gei90]. Abbot [Ano97a]. Abbott [Ano90a, Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Rhe94]. Abbri [Tra97b]. Abel [Rei93b]. Abigail [Ber97a]. Aboard [Min95a]. Abraham [Lin98]. Abramson [Oga90, Cla97b]. Abroad [Bod99]. Absolutism [Lyn98]. Absolutist [Kuc94]. Abstract [Cap90b]. Academic [Abb96, Gei94, Mor98, Lie90]. Academische [vL90]. Academy [Dan95, Lar91, Ser96, Cla96]. Accidents [Bra91b, Moh95]. Accounts [Har99b]. Accuracy [DeV92]. Acheson [Lun90b]. Achievement [Bla97b]. Achievements [Nye90]. Ackerman [Ada93]. Acknowledging [Bla97c]. Acoustics [AC95]. Acquisition [Mil91a]. Act [Ser92]. Actes [Lan98b]. Action [Fri99, Cha93]. Activities [Wor98]. Adams [Smi92b, Cut93, Dav92b, Lon95b]. Adaptation [Bro92b, GK93]. Adaption [Hea98]. Adas [Hea90, Hea98]. Addington [DeV91]. Adele [Kay93]. Adjusting [Maz94]. Adjusts [Bat90]. Adler [Car98]. Administration [Fie94a, Hei95a, Jan91]. Administrators [Gus91]. Admiral [Min97, Hol95]. Adoption [Hea98, Lau98a]. Adrian [Tan93]. Advanced [Vol91b]. advantage [Hei99]. Advent [Bul91]. Adventure [Mei90]. Adventuring [Rei93a]. Advertising [Bla99b, BS93, Lai93, Lai96c, Mat92a]. Advisers [Her93]. Advising [Smi93a]. Aer [Gra97b]. Aerial [Edm96]. Aero [Nah97]. Aerodynamic [Vin97]. Aerodynamics [Lau99]. Aeronautical [Flo92, Nah97]. Aeronautics [Bra95]. Aeroplane [Woo92]. Aeroplanes [Gol91b]. Aerospace [Meu98, Sch92a]. Aesthetics [Lev97b]. af [Myl97]. Afar [Elv99]. Affair [Hal98]. Africa [Kri97, GK93, Kri99]. African [Har97, Car96, Gre95b, Kri98]. After [Cha99a, All98b, Hac94b, Les94, Spe95, Fri98, Sum97]. AG [Mer98]. Again [Ror91]. Against [Rot99c, Ald95, Her94, Wal94]. Agatha [Ben92a]. Age [Big96, Cal92b, Chr95a, Gei90, Gom96, Har95a, Kei94, Kip97, Mic94, Mil94, Min95c, Nei98, Rab96, Rob94a, Sta90c, Sti91, Ten92, Ter94, Vin93, Web90a, Wil90, Wil93a, Wos91, Yat92, vK92a, Alv99, Gib99, See99b, Tei99, Bar92a, Buc94, Bul91, Cor92, Dav93, Her94, Pin90, Wos92, Ale99a, Chr95b]. Agencies [Rob92b]. Agency [Con97]. Agenda [Kat97b]. Agent [Neu92]. Agents [KP96]. Ages [Rol95a, Bin93]. Aggression [Rol90]. Agricola [Rob95]. Agricultural [Fin95, Sha95c, Zel97, Ham98b]. Agriculture [Eva92, Fit91, Fit95, Fit98, Hur99, Pol91, Sha93, Rug96]. Agro [yH98, Tri90a]. Agro-Industrial [Tri90a]. Agro-Industries [yH98]. Ahl [Lon97]. Ahlström [Ols97]. Ahmet [Ter94]. Ainley [Och92a]. Air [Bat90, Bil97, Cla96, CK95, Coo94, Day99, Dou91a, Gol91b, Hol94, Ken92, Lew90, Mer95, Mor91d, Mor93, New94, Omi92, Slo98, Smu93, Wil90, Ken92].

Aircraft [Fro92b, Hig92, Tri93, Zei95]. Aircrew [Bil97]. Airfields [Sch91]. Airlines [Van97]. Airplane [Con94a, LV95, Sch94a, Vin94]. Airpower [Kue99]. Airshipmen [Tri93]. Aitken [Dou94]. Akademgorodok [Bal99]. akademiker [Har99a]. Akos [Tod96]. aktuelle [Lun90a]. al [Rot92]. al. [Ahv90, And92, Bin93, Bla90b, But92, Coo97, Fit95, For97, Gor92a, Hac91b, Har93b, Hei95a, Hei91, Hig96, Isr92a, Isr94a, Jac96, Jac97, Kid91, Kli91, Kri98, Ler98, Lig97a, McG97b, MM92, Mil97, Mit96, Neu92, Nye97, Oga90, Per90, Rol93b, Roz90, See95b, See98b, Sel97, Sha93, Sta97b, Str90a, Til94, Tri90a, Web90c, Wos92]. Alabama [Lew97, Rey91, Sum96]. Alain [Smi92a, Smi95b]. Alamos [Hal94, See95b]. Alan [App96, Bed90c, Dou91b, Goo90, Hig97a, Kel93c, Log95, Rei96, Sol98]. Alaska [Ste94b, Coa99, Les90a]. Alaskan [She93]. Alastair [Nob97]. Albert [Cla97b, Cut94b, Jac91, Mor97]. Albrecht [Bra90]. Albro [Hil92b]. Album [Rob90a]. alchemie [Hom94]. Alchemist [Eam97]. Alcoa [Mis90]. Aldcroft [Ang90]. Aldous [O’H98]. Alexander [Koh96, Leo96, Mag94]. Alexanderson [Ros94a]. Alfred [Dou91a, Mil95, Scr91a, Kle91]. Algae [Bel97]. Alice [Cap90a]. Alkali [Don94]. Alla [For97]. Allan [Ber93, Rhe94]. alle [Tod99]. Allen [Whi92b, Eis93, Rue90, Wri92]. Alliance [Lau95]. Allied [Bil97, Kra90]. Allocating [Ait94]. Alloys [Wes90]. Alois [Lue90]. Alone [Con92a]. Alternating [Hec92]. Alternative [Big95a, Dou95]. Alumbrado [PO99]. Alvah [Fer90]. älvar [Kai98]. Always [Gol95, Col93]. am [Pet98, Pur93a]. Amann [Eva92]. Ambiguous [Cor97]. Ambitions [Laz99, Vol93]. Ambitious [Tra98a]. Amdam [Qui98]. America [Big95a, Bin91, Bro96, Cam92, Dou91b, Ell92b, Fit93b, Fit98, Gal92a, Gos90, Han93, Hea96, Hir92, Joh91, Lan94c, Lay97, Led93, Les93, Lin91, Mea98, Por93, Sch95, Sha95b, Smi91a, Whi92a, You93, Ano99n, Bro95b, Con95, Cru99, Dic92, Dre99, Gam92, GK93, Hil99, Hor97b, Jag99, Kar99, Mac91, Mei99a, Nic99b, Rei99a, See99b, SS99, Smi91b, Ste93, Ste89, Tap99, Uss91, Bol90, Bor97, Cor95, Daw91, Ebn92, Hug96b, Jac95b, Kat94a, Lai91, Lai92, Lan95b, Lay96, Lie90, Mac93b, Mis93, Nye92, Pye92, Sch93a, Sel90, Skr93, Sla95, Uss97, Wri94, D’A97, Big91, Con96a, Kin99, Lie93, Scr98a, Tho97a, Vol94a]. American [Ale99b, Bas99, Ber94a, Ber97a, Big97, Bla97a, Bog98, Bra96, Bro90, Bry96, Bud96, Cal92a, Cam93, Chr91, Chr95b, Coa99, Cod92, Col93, Cut99, Dav92b, Dou95, Eam97, Fie98, Fit95, Fle98, Fos98, Gar91, Gar92, Gei94, Gol92, Gol93, Gor95a, Gor95b, Gra95, Hac90a, Hac90b, Ham98a, Ham92a, Hil92b, Hil95, Hod93, Hof91, Hol93, Hor92, Hug95, Hur99, Hyd90, Ing94, Jac91, Jud90, Kea98, Kel95, Kie90b, Kle95, Kle99, Koh96, Kra95, Lai96a, Lai97, Lay92a, Lea92, Leo96, Lub91a, Mal91, Mal98, Mar90, Mat90, Mat92a, McC93, Mei90, Mei93, Mel96, Meu98, Miy94, Moe94, Nas91, Neu92, O’M93, Pas96, Pet97, Rab98, Ras92a, Rei93b, Ros93, Ros94a, Rot94a, Rum92, Sch92a, Sch99b, See90, Sha95b, Sha96b, Sim94, Sin99c, Sla92, Sla97, SD98]. American [Smi99, Spe95, Sta94, Sta92a, Sta97b, Str90b, Swe93, Ten92, Til90, Tri93, Vol93, Vol98b, Wes95, Whi93, Whi98, Wil92a, Wor91, Zac98, Zem93, Ale99b, Bod99, Bro99b, CC95, Col90, Dou99, Fel99, Fit91, Gam95, Gre95b, Hen99, Her99b, HS96, Hui99, Kli97b, Lea99, Lev97b, Mat99c, Och92b, Owe91b, Per91, Rot94b, Scr94, See93, Sil99, Smi90c, Smu94, Tho91b, Wil99a, Wil99b, Blo90, Coz93b, Dan90, Ler98, Mei95, Mor91a, Nuw90, Rhe98, Sal92b, See98a, SS97, Ube99]. Americanism [Scr91b]. Americans [Eat90, Ing91]. amerikanischen [All96a]. Amerikas [Wen92]. Amish [Fis98]. Amiya [Pfa98]. Ammann [DB94]. Amnesia [Ald97]. among [Bro92b, Was93]. Amsterdam [Win98]. Amusement [Dav92b]. Amusements [Bro97b, Ada95a]. Analysen [Lun90a]. Analysis [Bro92a, CK95, Gal92b, LMO97b, Mau90, Vol93]. Analytical [Flo92]. Anatoly [Kry92b]. Anatomy [Hal98, Owe91b]. Ancestor [Tra90]. Ancient [Ber97b, Chi97, Fer92, Hei95a, Mul94, Sim95a, The90, Vas98, You95]. Anden [Sun97]. Anderson [Fie94b, Goo95, Wor94]. andinos [Gua98]. Andre [Isr91, Tho97a, Gei90, Kah96, Lev91b]. André-Georges [Lev91b]. Andreas [Wen92]. Andrei [Jos99]. Andrew [Con97, Cut94a, How93, Mau91, Nob97, Smi95a, Wil93a, Moy96]. Andrews [Gro90b]. Angeles [Rot99a, Gra91]. Anglo [BA90, Wal95]. Anglo-Saxon [BA90, Wal95]. Angus [Lev97a]. Aniline [Tra94]. Ann [Bra96, Hof92]. Anna [Kni95, O’D94b]. Anne [Hat90]. années [Smi91d]. Anneli [Gra97b]. Anniversary [Fel99, Kay91, Lup91]. Annotated [Adr94, Coo91d, FG98, Fin95, Hig96, Kri97, Low91, Moh94, Sta92a, Svi93]. Anomaly [Con94a, Vin94]. ans [Smi95b]. Antebellum [Kel95, Rol93b, For93, Lub91b]. Anthology [Ker99]. Anthony [Bed91, Lan91a, Roc94, Tur95]. Anthracite [Sal92a]. Anthropologist [Ing99]. Anthropology [Ell94a, NP94]. antike [Lon93b]. Antiquity [Kel93a, Tod95]. Antisubmarine [Hig97b]. Antisystems [Ham92b]. Antoine [Lan90a, Lan95a, Lan98b, JJ98]. Anton [Bro99b]. Antonio [Coo91a, Lon95b]. Anxieties [Cam95]. Anxious [Sha95c]. Any [Hor99]. ANZAAS [Pye90]. Apartments [Con92a]. Apollo [Joh97a, Swe92, Vin93, Log91]. Appalachian [Qui93]. Apparatus [ML98]. Appetites [Skr93]. Applications [Cer97, Bat98, Kal95]. apports [Hal93a]. Appreciation [Hal93b]. Apprenticeship [Ror91]. Approach [Hoo99, Jac97, McE95]. Approaches [Joh97b, Kal95]. Appropriate [Pur93b]. Appropriating [Kie90a]. April [Ano90k, Ano91n, Ano91o, Ano92n, Ano93o, Ano94n, Ano95n, Ano95o, Ano96k, Ano97n, Ano98g, Ano99r]. Aqueducts [Lan94a]. Arabic [Bre99]. Arbeit [All96a]. Arbeits [All96a]. Arbeits- [All96a]. Arbeitssicherheit [Lun90a]. Archaeological [Lan95b]. Archaeology [Buc98a, Gib91, Gor92a, Kri99, Rab98, Rol93b, Wil99a, Sta90a]. Archaic [Hei95a]. Archeology [Jac93a, See95a, Jac93a]. Archeomaterials [Way90]. Archimedes [BS93, Sim91]. Architect [Hol95]. Architectes [Lan90a]. Architects [Ols98, MM92]. Architectural [AC95, Har96b, Lon95b, Rue90, San94]. Architecture [Buc94, Cla96, Kuc94, Lee96, McG99, Ozd98, Pet96, San92, See99b, Ada93, Gui91, Lon95b, Mat90, San97, Fie94b]. Arctic [Coa99]. Arden [Gui93b]. Area [Cla92, Hal94]. Arid [Hol97b]. Arie [Jac97]. Arizona [Jac96, Rey98]. Arkwrights [Jer90a]. Armi [Hyd90]. Armour [Mor98]. Armoured [Edg98]. Arms [Bel98b, Rab96, Rol90, Wes95, Moy94]. Armstrong [Bas92b]. Army [Hac92a, Mer95, Mer96, Reu99a, Sch91, Sch94c, Sla98, Sha90b]. Arne [Gie96]. Arno [You95]. Arnold [Bal97, Bas92a, Fit97, Les94]. Arnoldi [Kri98]. Arrhenius [Kra98]. Arrow [See98b]. Arsenal [Ham98a, Kuc92, Rab96]. Art [Ada91, Ada93, Cou94, Cro94, Gor92a, Gui91, Gui93a, Hyd90, Isr92b, Kat94a, Lan94b, Lon97, Mil94, Per91, Ree98, Reu99a, San94, Wes93a, Wos91, Gol91a, Kas99, Lev91a]. Artefacts [Tra97a]. Artful [Pla96]. Arthur [Cap90a, Kid91, Mei90, Mit96]. Artifact [Pin92]. Artifacts [JM97]. Artifice [Bes95]. Artificial [Ben94, Kid90, Wal99b]. artillerie [Lan98b]. Artillery [Bas92b, DeV90]. Artisan [Gut93]. Artisans [Ler98, Whi92c, Bro92b]. Artist [Far99, Mat92b, Smi99]. Artistry [Wal99b]. Artists [Wos91]. Arts [Bra96, Gom97, Kel93a, Mil94, Øve97, Pan98c, Rhe94, Sum97, Ube94, Rhe94, Sin96]. Arturo [Dol97]. Asa [Web90b]. Asia [Jer97]. Asian [Lin94a, Rho93, Sel97, Wig96]. ASME [Kay91]. Aspects [Gib91, Bra91a]. Asphalt [Hug95]. Aspirations [Cam95]. Aspray [Kid92, Mei98a]. Assembly [See95b]. Assessment [Car97, Jac97, Skr91, Str90a, Sti98]. Association [Gra98b]. Astrid [Lon93b]. Astro [Dek98]. Astro-Technik [Dek98]. Astrolabe [Bed90c]. Astrolabes [Bed90c]. Astronomer [Gin90, Smi91c]. Astronomical [Aga98, Rot92, War98b]. Astronomie [Dek98]. astronomischen [Dek98]. Astronomy [Huf92, Sul98b]. Atanasoff [Cap90a, Cap90a]. Atlantic [Bjö97, Roz98, Hig98]. Atlantis [Bal99]. Atmospheric [New97]. Atom [Hac94a]. Atomic [Hac91a, Maz90b, Maz94, Rot99c, Sei96a, Sil99, Smi93a, Wal99a, Bad91, Wal92]. Atoms [Hac91a]. Attachments [Leo96]. Attack [Mor91d]. Attention [Lai92]. Attitudes [Nye95]. Audiences [Bro98a]. Audio [Con92c]. Aufbruch [Neu94]. Auftrag [Fin98]. August [Ano93l, Ano97i, CS90, Hor97a, Ber95b]. Augusta [Hay90]. Auschwitz [All96b, All96d]. Australasia [Pye90]. Australasian [Ben91b]. Australia [Jar90, Pye92, Sal95, CF92, Ras99, Dun93]. Author [Ano91b, Ano91j, Ano92b, Ano93b, Ano94b, Ano95b, Ano97b, Ano99a, Ano99b, LCH^{+}96a]. Author-Title [Ano91j]. Authority [Kra90, Mit91, Smi92b]. Authorship [Lon91a]. Auto [Bor99, Cor95, Pet90, Lai96b]. Autobiography [Dan93, Gol95, Lan90b]. Autogiros [Tri90b]. Autographs [Lon93a]. Automatic [Ald93, Ben91c]. Automating [CK95]. Automation [Gre95a]. Automobile [Big91, Dan90, Hug95, Hug96b, Kir95, KP96, Lai96b, Lai96c, Vol98a, Whi92a, Hyd90]. Automobiles [Sta90b]. Automobility [Vol96]. Automotive [Vol95]. autrefois [Lau98a]. aux [Coh90, Lav93, Smi92a]. av [Kai98]. Ava [Cam93]. Avant [Bra97]. Avant-Garde [Bra97]. Avante [Tho94]. Avante-Garde [Tho94]. Aveni [Lan91a]. Aviation [Edm96, Eze90, Hol94, Jos99, Neu93, Pis96, Rei93a, She93, Hol95, Lew93, Sti91]. Avionics [Lee97]. Avital [MM93]. Award [Ano90c, Ano91e, Ano92d, Ano93d, Ano94d, Ano95d, Ano96d, Ano97d]. Awards [Ano98a, Ano99c, Ano99d]. Awareness [Ste89]. Axe [Jag99].

B [Bau93, Bor97, Cal92a, Ebn92, Edw95, Gor92a, Hil95, Hol90, Jar90, Kel93d, Lai96a, Lan95b, Lee97, Ler98, Lie93, Lin94a, Mil91b, Mul94, Oga90, Scr96, The90, Vol91a, Wig96, Bil97]. B&O [Bro95b]. B-1 [Hol90]. B-29 [Bil97]. B. [Cap93, Cap95]. Bâtir [Kah96]. Baader [Gam99]. Baatz [Dan95]. Baber [Ada98]. Bache [Koh96]. Bachrach [Hal98]. Back [Hil92a, Jud99, Mil91a, Ror91, Sin95, Lee96]. Back-to-the-Land [Jud99]. Backwoods [Jud90]. Badaracco [Tas97]. Bagchi [Pfa98]. Bailey [Ben91a, Daw96]. Bairoch [Con90]. Bakelites [Mei96]. Baker [Koc98]. Bakewell [Coo91a]. Bakker [Rob94a]. Balance [Wos92]. Baldwin [Mal98]. Bale [Cou94]. Bali [Pfa92]. Ball [Arm96]. Ballard [Hei91]. Ballistic [Buc98b, Hev95]. Ballistics [Ste94a]. Balloon [Zie94]. Ballooning [Daw91]. Balloons [Zie92]. Balogh [Gol93]. Baltimore [Bro95c, Uss93, Ano99m]. Bamford [Ben90]. Banda [Ell92a]. bande [Lav93]. Bank [CK98, Mor91b]. Banta [Big96]. Barbara [Ber94a, Pet97, Pla96, Pri98, Ros92]. Barcelona [Gli94]. Barger [Lie93]. Barnett [Vas98]. Baron [Rei96, Cam93]. Barons [BS94, Sch92a]. Baroque [Bjö97]. Barrel [Ell93]. Barrie [Buc98a]. Barrier [Vin97]. Barry [Cha93, Die97]. Barth [Vie90]. Barton [Sei96a]. Basalla [May91]. Based [Kön96]. Bases [Mer95]. BASF [Tra98a]. Basic [Cha99b]. Basin [Qui93]. Bat [Joe90]. Batchelor [O’D97]. Bath [Ber95b, dlP99]. Baton [Pos95, Sta96b]. Batschelet [Sta92a]. Batstone [Roz90]. Battery [Kir95]. Battery-Only [Kir95]. Battle [Hig98, Mor93, Reu93b, Sch94d, Sum91, Wal95]. Battlefield [Mor91d]. Battleship [McB92, McB97]. Battling [Hof91]. Baucom [Sei93]. Bauer [Shi97]. Baumwollspinnerei [Lin93a]. Bavarian [Tho90]. Bay [LaL90, Mil91a]. Bayer [Sim90]. Bayern [Gam99]. Bayerns [Gam99]. Be [Bix98]. Beacon [Gam92]. Beautiful [Con91, Mor90, Rob93]. Bedini [Dan92, Lan93, Ste96]. Beedle [Mor91b]. Beef [App96]. Been [Mol91]. Beer [Gui93a]. Before [Har95a, Sta98, Swe92, Blu99, Kön96, Low97, Mit95, Kid94, Mil93, Phi93, Tan93]. Beginnings [Ser96, Wes90]. Begriffe [Tod96]. Behavior [Cap93, DB94]. Behavioral [Cow93, Edw95]. behind [Hew94, Hyd90, Mau91]. Beispiele [Lun90a]. Beitrag [Wen92]. Belhoste [But92, Lan98b]. Belief [Cal92b]. Beliefs [Con99]. Bell [Mue91a, Smi90b, Cla93, Gre95a, Gre95b]. Belt [Rug96]. Beltran [Smi92a, Smi95b]. Benad [All96c, Tod96]. Benad-Wagenhoff [All96c, Tod96]. Bending [Bro92b]. Bengt [O’D94a]. Benhamou [Lan94b]. Benjamin [Har96b, Maz94, Kry90, Ste94a, Kry92a]. Bennett [Nas90, War92b]. Bensaude [Tra97b]. Bensaude-Vincent [Tra97b]. Beretta [Rob95]. Berk [Dou95]. Berkeley [Kri92]. Berlanstein [Tal94]. Berliet [Lau98b]. Bernadette [Tra97b]. Bernard [Die97, Dre95, Hal98, Kry92a, Ros94b, Smi91c, Svi93]. Bernice [Old98]. Bernstein [Rei91]. beroep [Wac94]. Berthold [Hal96]. Bertrand [Wal92].

Bertrum [Goo90]. Berucksichtigung [Dek98]. Berzelius [Frä91]. Besserat [Phi93]. Best [Min99, Led91]. Beton [Pet98]. Better [Ber90]. Bettyann [Her98]. Between [Ale99b, Pal93, Caw95, Hil95, Isr92a, Jar95, Jer94a, Lau95, Les90b, Mit92, Moh90, Pye92, Wos92, Bix99]. Beyond [Fit91, Gom96, Pur99, Wil90, Sel97, Tra97a, Day99, Lev91c, Mou99, Rot93, Sta94]. Beziehungen [All96a]. BFGoodrich [Daw99]. bianco [Sel92]. Bias [Lin98]. Bibliographic [Cer97, Wes99]. Bibliography [Ano99e, Ano99f, Bal97, Ber95a, Coo91d, Goo90, LAC^{+}90, LAC^{+}91, LAC^{+}92, LAC^{+}93, LAC^{+}94, LCH^{+}95, LCH^{+}96b, LCH^{+}97, Low99a, Low99b, Sta92a, FG98, Fin95, Hig96, Kri97, Low91, Moh94, Svi93]. Bibliotheca [Adr94]. Bicentennial [Kry90]. Bicycles [Mei96]. Biddick [Fle90]. Biddle [Sch92a]. Big [Joh97b, Lee96, Ols98, Rot94a, WH96, Tal94, Wes93b]. Bijker [Gie94, Mei96]. Bild [dV95]. Bildung [Bra90]. Bilge [Ron99]. Bill [Coo92b, McN93]. Billington [Isr92b]. Bindlestiffs [Ras92a]. Bindocci [Moh94]. Biographical [Por91, Pur97]. Biographie [Neu94, vL99]. Biography [Cor95, Hew94, Min97, Lie98, Mei98b]. Biological [yH98, Lan98a, Ras99]. Biology [yH98, Bud94, Lin97a]. Biomedical [Cow93]. Biotechnology [Ham98b, Tri90a, GS98, Kay94, Shi97, Gus98]. Birmingham [Gor96, Lew97]. Birth [Ale99a, Neu98]. bis [All96a, All96c, Gam99, Har99d, Lun90a]. Bisbee [Qui94]. Bishir [MM92]. Bishop [Gol91c]. Bitted [Jag99]. Bjerknes [Kie90a]. Bjork [Cap95]. Björn [Mah97]. bk [Lin94a, Wig96]. Black [Tra94, Yat96, Har95b]. Blacksmith [Coo98a, Min95a]. Blake [Mil94]. Blanchard [Mac93b, Coo91c]. Blauer [Web96]. Bleaching [Adr94]. Bless [Oli98]. Blessings [Sla95]. bleu [Smi95b]. Blewett [Gre91, Per93]. Bloch [End97]. Blondheim [Sch95]. Blood [Eva93b, Tap99]. Bloody [Bin91]. Bloomery [GK93]. Bloomington [Hou90]. Blue [Hol90, Par97]. Blueprint [Kra93]. Bluff [Rol93b]. Blurred [GS99]. Board [Car99b]. Boat [Hig97b]. Boats [Les90a, Moy96]. Body [Mei99b, Old97]. Boeing [Mac93a]. Boiler [Daw96]. Bolivia [Gua98]. Bomb [Bad91, Sil99, Smi93a, Wal99a, Hew94, Sut90]. Bomber [Zie94, Hol90]. Bombing [Con96a, Kue99]. Bombs [Rot99c]. Bond [Ste89, Scr96]. Bone [Her98]. Book [Chr95a, Hea99b, Lyn98, MM93]. Bookkeeping [Hei95a]. Books [Har99b, Lon96a, Sta92a, Tas97]. Boott [Jer94b, Lea99]. Border [Gre99]. Borders [Mul95]. Boris [Thr91, Kry92b]. Born [Bra97, Bor97]. Borst [You95]. Boston [Gam92]. Bothwell [Maz90b]. Bottle [Sch94c]. Boudriot [Gui91]. Bougnoux [Lav93]. Boulez [Bra97]. Boundaries [Bar92a, Mou99, Pan98b, Sum97]. Bourgeois [Gut93]. Bourgeoisie [Car98]. Bowker [Con96b]. Bowler [Bri92, Sim95b]. Bowman [Sin93]. Bowron [Dan93]. Box [Yat96]. Boyne [Hyd90]. Boys [Gol91c, Old97]. Braak [Rot97]. Bradford [Dre97]. Bradley [Wes95]. Brahe [War98b]. Brainchildren [Ben99]. Branch [Con95, Sla98, Str91]. Brashear [Lan90b]. Brass [Kid91]. Braun [Neu94, Dav92a]. Braund [Les90a]. Braunschweig [Bra90, Gis98]. Brazil [Pri98, Was92]. Breaking [Bro92b, Ube94]. Breaks [Vin97]. Breakthrough [Cha99a]. Breechloaders [Bas92b]. Breeding [Pal93, Sha91]. Bretagne [Kah96]. Breton [Coh90]. Brewing [Coo95]. Brian [Fer92, Gol93, Hil92a, Hug96b, Mei93, Sol98]. Brick [Ham92b]. Brickmaking [Sta90a]. Bricks [Sta90a]. Bridge [DB94, Dre99, Gro90a]. Bridges [Wal99b, Smi90a]. Bridget [Ble91, Nye95]. Brief [Dou91b]. Briggs [Web90b]. Brinckerhoff [Rom97a]. Bringing [Rot95]. Britain [Eva96, Gen92, Lay91, Mau93, Moh90, Pye92, Atk99, Hil99, Mac91, Pur93a, Zei95, Ang90, Bro92a, Coo92b]. British [Blo90, Bud96, Cas98a, Dou91a, Edg98, Gol91b, Har95a, Har96c, Hig92, HC98, Jer94a, Kli90, Leu98, Liu97, Owe91a, Pay95, Sam99, BP91, Cru99, Div94, Don94, Kub90, Kue99, Nah97, New97, Pal93, Rog95, TJ96, Pol91, Voe90, Was92]. Brittain [Ros94a]. Britton [Cou94]. Broadaxe [And99b]. Broadcast [Slo95]. Broadcasting [Kra95, Mau91, Sch94d, Slo98, Smu90]. Broader [Tho91b]. Broadsides [Sta92a]. Broken [Ald99]. Bromberg [Les93]. Broncano [MA97]. Bronze [Gui93a]. Brooke [Dup92]. Brooks [Tri90b]. Brose [Tod95]. Brothers [Cor91, Gor95a]. Brown [Por93, Joh91, Pan95, Sch91]. Brownell [Har96b]. Brownfields [Gor97b]. Bruce [Bro93a, Coz93b, Gru90, Hac90b, Her93, Kat94b, Rol93b, SS97, Wes93b]. Brucken [Pet98, Pet98]. Bruegmann [Cla96]. Brüggemeier [Web96]. Bruland [Hår93a, Voe90]. Brunanburh [Wal95]. Brunelleschi [Ada91]. Bruno [But92, Lan98b, Rei91, War97]. Bryan [Mag94]. Bryant [Coo93b]. Brynes [Nee96]. BTechnologies [Fer99]. Bucciarelli [Mag95]. Buchanan [Lay91, Lev97a, May94]. Buchheim [Blu92]. Bucholz [Gui93b]. Buchwald [Kli97a]. Bud [Maz93, Kay94]. Budapest [Hor97a]. Buffalo [Scr96]. Bugos [Hou97]. Builder [Far99, Jac91]. Builders [MM92, Tri91]. Building [Ber94a, Bro95c, Coa99, Gie94, Ham95, Hur98, Lan90b, Mat92b, Mor91b, Reu99b, Sha90b, Tri93, Wer97, Whi98, MM92, Zak92, Big97, Bil97, Edw96a, Mer95, Nye98a, Rue90, Smi95a, Sul98a]. Buildings [Bla97a, San94]. Builds [Sut90]. Built [Gol91b, JM97]. Buisseret [Cod92]. Bulbs [Mei96]. Bulkeley [Mac92]. Bull [Ras92a]. Bunch [Mag94]. Bundesrepublik [Rüd92]. Bungalows [Whi93]. Bunha [Has93]. Bureau [Ald95]. Bureaucracy [Sha96a, Kie93]. Burgers [Bel97]. Burghal [BA90]. Burghard [All98b]. Burghardt [Mer97]. Burke [Mul90b, Rei95b]. Burks [Cap90a, Cap90a]. Burnett [War91]. Burning [Joh99b, Sim91, Dun93]. Burr [Bel98b]. Burroughs [Kid94]. Burton [Bel98b]. Bus [Min95c]. Buschman [Kal95]. Bush [Bas99, Dun93, McM94a, Rei95b, Lai92]. Business [Asp93, Ben90, Bla90a, Coo98d, GS98, Gor95a, Gor95b, Gut90, Han99, Hig91, Isr91, Ler97, Mue91a, Nor90, Rey94, Sha91, Tal94, Web90a, Whi98, Mau91, Lam93, Lip98]. Businessmen [Ser96, Tri93]. Butterfly [Joe90]. Buttresses [Cut93]. Buy [Lai93]. Buzet [Eva92]. Bynum [Mar96]. Byzantium [Rol92].

C [Ben92a, Bry95, Con96b, Coo93b, Cor90, Gra91, Gui95b, Hin94, Hyd90, Jan91, Joh99a, Kei94, Kel93b, Kie93, Kuc92, Mac93b, Mar91, Mul95, Neu98, New91, Nye98a, O’D96a, Oli90, Ors97, Pfa97, Rhe94, Rug96, Sei96a, Sel90, Sta97a, Tob92, War97, Was92, Wil92b, Wor93b, Zel97, vL90]. C&A [Hou90]. ca [Lon93a]. Cabinet [War99]. Cable [Kra94b, Smi90a]. Cahaba [Sum96]. Cahalin [Ano90f, Ano91h, Ano92g, Ano93g, Ano94g, Ano95g, Ano96f, Ano97g]. cakmak [Ter94]. calcul [Coh90]. calculateurs [Coh90]. Calculating [Fel97]. Calculation [Nor90, Vin97]. Calendars [Lan91a]. California [Ano90h, Cor93, Low93a, Min95c, Pos98b, Bau93, Ors97, Rob92b, Lan98a]. Californians [Min95b, She94b]. Calling [Vol94a]. Caltech [Bud94]. Came [Mar90, Wac99]. Camera [Sam99]. Cameron [Bar90]. Campaign [Ald95]. Campanella [Hal93b]. Campbell [Cla90]. Camps [Jac96]. Can [Hor99]. Canada [Hig90, Maz90b, Pry93, Par97, Str91, Eze90]. Canadensis [Pry93]. Canadian [Arm96, Bal97, Goo90, McN93, Och92a]. Canal [Elv99, Low97, Rey90b, Rob90a, Rob92a, Sha90a]. Canals [Sha96a, Rob92a]. Cancer [App96]. Candida [Sel92]. Cannery [NP94]. Canney [Dud92]. Canning [O’B91]. Cannon [Gui93a, Mau90]. Can’t [Ben96]. Capabilities [Sch99b]. Capacity [Rus99]. Capital [Day92, Hil99, Sal92a, Scr98a, Kuc94]. Capitalism [Cam92, Fro92b, Sto99, Gib99, Gui95a, Kel93d, Scr91a, Har98b]. Capsule [Arr94]. Captain [Lan90b]. Capturing [Buc92]. Car [Hyd90, Kle95, Lai96c, Pos98b, Wri92]. Card [Nor90]. Cardinal [Smi93a]. Cardiology [SS97]. Care [Gam97, Gor94, MKR93, SS99, Koc98]. Career [Hol97b, Rus99, Smi94, Ste98]. Careers [Fit98, Vol91b]. Carel [Gui93a]. Carey [Joh92]. Carl [Bec91, Caw95, Cut98, Dre97, Ile96, Jar90, Lue90, Iri97]. Carla [End97]. Carlebach [Wri94]. Carlisle [Ham98a]. Carlo [Dav93]. Carlos [Qui94]. Carlson [Ros94b]. Carnegie [Ber97a]. Caro [Tra98a]. Carol [Hac92a]. Carolina [MM92, Whi92b, Whi92b]. Carolinum [Gis98]. Carolyn [Mac93b, Por94, Ste91a]. Carr [Sha93]. Carriers [Ald99]. Carroll [Pac95]. Cars [All90, Whi90, Kir95]. Carstensen [Fit95]. carta [Sel92]. Cartography [Bed90b, Kon97, Kuc94, Lin94a, The90, Wig96, Kie90b]. Cary [Hut92]. Case [Ben91b, Cut94a, Fer95, For93, Gor97b, IR91, Jam90, Kal98, Law92, Mil91a, Och92b, Par92, Sch93b, Sim90, Tho90, Zie94, vk92b]. Casey [Pin91]. caso [Sel92]. Casper [Kip97]. Cassedy [Led93]. Castaneda [Con94b]. Castilla [Gli91]. Casting [Gui93a]. Cat [All99a]. Catalá [D’A97]. Catalog [Adr94]. Catalogue [O’M93, Tur95, War99]. Catapults [All99a]. Catastrophic [HS96]. Cathedrals [Mue99, Sun99]. Catherine [Log91, MM92]. Cathy [Bla90b]. Cats [Rei99a]. Cattle [Whi90]. Cautionary [Mis93]. Cavities [AC95]. Celebration [Led91, Nic99b]. Center [Wes93a, Bil98, Joh99a]. Centers [Rot92]. Central [Fis92, Gre95a, Jac96, Alv99]. centrales [Lan91b]. Centralizing [Dic92]. Centre [Hei99]. Centuries [Aga98, Bro90, Low93b, Man99, O’H96, Haf97]. Century [Bac98, Ben90, Ble91, Bro92a, Bro92b, Bur92, Cla96, Coo91a, Coo95, Day99, Doo91, FO92, Gro98, Gut95a, Haf92, Haf99, Kay93, Kel93d, Kin99, Kyn99, Lay93, Lie90, Lie93, Lin97a, Mac93b, Mor91a, Mor91b, ML98, Nor90, Per93, Pet92a, Rhe94, Rum97, Ryc92, San97, Sch99b, Sel90, Sil93, Smi90a, Ste96, Tal94, Tho93, Ube99, Vol96, Wes95, Wor95, Arr94, Dre99, Eve92, Gou98, HP91, Jos95b, Kry90, Ler97, Lon91b, Mac91, Mor99, Mue91b, Nic99b, Reu99b, Smi91b, Ste93, Ste89, Tap99, Uss91, Bla90a, Bro96, Buc92, Coo98b, Coo91d, DeV91, Fel97, Gam97, Gui93a, Gut90, Gut92, Gut93, Her98, Hil95, Hof92, Kat92, Kat97a, Kel93a, LeC94, Lüt98, Mar96, Opi90, Rot92, Ube94, Voe90, Vol91a, War92a, vK92a, Bas99, Don91]. Century-of-Progress [Rum97]. Ceramic [Pet97, Rey98]. Ceramics [Bla92b]. Ceruzzi [Sma95]. Cesare [Mac96]. CF&I [LeC94]. Chadwick [Ham92b]. Chaffee [Wes93a]. Chagyongnu [Hon98]. Chain [Gol93]. Chaisson [Spi95]. Challenge [Bul94, Cor97, Fri98, Hil97]. Challenger [Coo91d, Nea98]. Chance [Smi91c]. Chandler [Scr91a]. Change [Arn97, Bau99, Big91, Bla92a, CK98, Can95, Chu99, Fit90, Hac94c, Haf97, Har98b, HS96, Kel99, KP96, Kra94a, Par92, Per90, Por94, Rab98, Ryd98, Scr98b, Tod95, Tra95, Ahv90, Str90b, Cha91, Edw95, Edw97, Gie94, Hur90, Mei96, Wil95a]. Changed [Ano99n]. Changes [Ger96]. Changing [Div95, Don94, Gra95, Hac94a, Nas90, Rot95, Ser94, Wos92, Lau95, Min96, Old98]. Channel [Rog95, Tra91, Sta96a]. Channell [Low91, May92]. Chaplin [Sha95c]. Chapman [Fro92b]. Chappell [Fie94b]. Chaptal [Gou98, JJ98]. Chaptalization [Gou98]. Chapter [Jac93a]. Character [Cow96]. Charcoal [Rol94, See95a]. Charge [Gre94, Hug96b]. Charles [Bat98, Har96b, Hea96, Log91, Mil93, Scr96, Ste98, Tho93, BP91, For93, Rei95a, Kat97a, Sta92b]. Charlotte [Rue90]. Charlottesville [Ano96b]. Charts [Cod92, Dic92]. Chasing [Mel96]. Chassagne [Lin94b]. Chauffeur [Bor99]. Chaussées [Lan95a]. Checklist [Kuc94, Lon93a]. Chemical [Col90, Div94, Gar99, Mor99, Mul96b, Set99, tZS90, Tra90, Tra95, Von99, Ful95]. Chemicals [Cal98]. chemicus [Wac94]. Chemie [Hal96, Hom94]. Chemiker [Wac94]. chemins [Frä91]. chemische [Hom94]. Chemistry [Don91, Gou98, Mau90, Mau93, Mor99, Mul96b, Rob95, Ste98, tZS90, Tra98b, Tra97b]. Chemists [Don94]. Chenango [Dup92]. Chernobyl [For97]. Chesapeake [Sha90a, Wor93b]. Chet [Hac98]. Chiba [Mat99a]. Chicago [Cla92, Con92b, Mea98, Pet92a, Ste90]. Chicagoans [Smi97a]. Children [Gor98]. China [Bla95, yH98, Jam90, Mul96b, tZS90, Jam90, Per99a, Sel99, Mul94, Sut90]. Chinese [Ber97b]. Chiropractic [Mar93b, Big95a]. Chlorella [Bel97]. Chlorinate [Gal92b]. Choice [Par97, Per90, Sch94a, Leo92, Log92]. Choices [Smi93a]. Chris [Cha98]. Christ [Poo99]. Christel [Lin93a]. Christian [yH98, Pri95, Wal94]. Christine [Qui95, Rob90b]. Christoph [Lon95b]. Christopher [Ben91a, Bro92a, Con94b, Cor95, Jer97, Les90b]. Chronique [Smi91d]. Chronology [Cap90b, Mag94]. Chronometers [Tur95]. Church [Poo99]. Ciencia [Gua98, D’A97]. Cierva [Tri90b]. Ciesla [All98b]. Cigar [Til90]. Cigarettes [Mat92a]. Cincinnati [Gal92b]. Cinema [Kop91, Pot99]. Cipolla [Dav93]. Circle [Lon95b]. Circuit [Lie91]. Circumstance [Gor92b]. Cities [Cod91, Ebn92, Kue99, Ros93, Vie90, Con90, O’D96a]. Citizens [Her94]. Citoyen [War92b]. City [Bal99, Bol90, Bro97b, Cla92, Con91, Hal93b, Har99c, Hof92, Hur98, KRT90, Kuc94, Lie92, Rei96, Ros93, Scr91b, Smi97a, Wil92b, Olw96, Qui95, Hug95, Kuc92]. Cityscapes [Spe95]. Civil [Atk99, For93, Hau90, Sal92a]. Civilian [Bad94]. Civilisation [Bin93, yH98, Mul96b, tZS90]. Civilization [Ada98, Bas92a, Kon97, Coo97]. Civilizing [Ada96]. Claire [Tas97]. Clangor [Min95a]. Clark [Cap90a, New91, Per90, Maz90a, Ryc92]. Clarke [Kay93, Kid91, Fit96]. Clary [Han93]. Class [Fro92b, Gre91, Gut95a, Gut95b, Kal98, Len91, Lew97, Pri98, Sta94, Whi92c, Scr91b]. Claude [All96b, Bin93, Vol94a]. Claudius [Smi91a]. Clay [Gol92, Hug95]. Cleanliness [Mel96]. Clepsydra [Hon98]. Clercq [War98a]. Clerical [Pri95]. Cleveland [Ano91c, McS98, McS98]. Clinton [Mol95a]. Clock [FF96, Hen99, Lan93]. Clocks [Hon98, Lan91a]. Clockwork [Bra98]. Closed [Uss97]. Closing [Jam90]. Cloud [Sil99]. Co [Kat94b, Qui98]. Co-Evolution [Kat94b]. Coachmen [Gut95a]. Coad [Pet92b]. Coal [Ald95, Bin91, Cla95, Gra98b, Har95a, Jer94a, San90, Har95a]. Coast [Bur92, Fit90, Koh96]. Coastal [Han93]. Coates [Ste94b]. Cockburn [Rot95, Old92]. Cognition [San92, Coo98a]. Cohen [Die93, Ell93, Har96b, Kry92a, Lon96b]. Cold [And94, Die97, Gei94, Olw98, Uss97, See99a]. Cole [Lon97, Sha93]. Colin [Rei95b, Tra98b]. Collapse [Cla90]. Collection [Adr94, Kid91, Ste94c]. Collections [Rob96b, War97, Eze90]. Colleen [Ber95c, Gal92b]. College [SS97]. Colleges [See93]. Collegium [Gis98]. Collins [Swe93]. Collision [Ald93]. Colonel [Bel98b]. Colonial [Ada98, Ken92, Lay97, Pas96]. Colonialism [Ada97, Set99, Cha96, Wil93b]. Colonisation [Hea92]. Colonization [Hur99, Pfa90, Rho93, Sel90]. Colorado [Smi92b, Och92b]. Colors [Chi97]. Colossus [Wor91]. Colten [Low97, Miy94]. Columbia [Sca99]. Columbian [Pan95]. Columbus [Kei94]. Colvin [Uss90]. Combating [Ald93]. Comes [Har94]. Comienzos [PO99]. Coming [Buc98b, Cor92, Ing94, Kle95]. Command [Mor93]. Commemoration [Por94]. Comment [Bar91b, BS94, Con94a, KM92, Lay92b, LV95, MB92, Scr91c, Scr00]. Commerce [Ald93]. Commercial [Bro97b, Gol93, Slo98]. Commercialization [Kra95]. Commission [Maz94, Sei96a, Ald93, Slo95, Hac91a]. Committee [Hor97a, LM91, Nah97, Owe91b]. Commodities [Rey98]. Common [Hug99, Rey98]. Commonwealth [Bla90a, Oga90, Pye90]. Communication [Kyn99, Lub91a, Moh90, SD98, Sta99, Tym99, Lav93, Lav93, Joh92, Nei98]. Communicationen [Bro99b]. Communications [LHTA99, Ran97, Yat92]. Communities [Rot99b, Wil92b]. Community [Lew97, Tan93, Ing94]. Commuter [Van97]. comp [Goo90]. Companies [Ald99]. Company [Big97, Blo91, Buc94, Cal98, Gor95a, Hir96, Kal98, Nas90, New91, Qui98, Was92, Bla95, Chu99, Cor90, Mil93, Tal94, Cor90]. Comparative [Cha98, Eve92]. Comparison [Bra92, Wal94]. Competition [Mis90, Sin99c, Mer98]. Complex [Gei94, Rot97, Rau98]. Compton [Swe92]. Computation [San92]. Computer [Cap90a, Cer97, Cer98, Cor97, Fri98, Jam90, Joe90, Kid94, Sei96b, Wil90, Abb93, Cap90a, You95, Edw96b, Sei98]. Computerization [Sin99b]. Computerizing [Kai96]. Computers [Abb93, All90, Lig99, Lue90, Uss97, Coo98d, Pfa91]. Computing [Coh90, Sma95, Yat92, Abb96, Kid92]. Computus [You95]. Concepteurs [But92]. Conception [Bry95, Joh99b]. Conceptions [Kip97]. Concepts [Cap90b]. Conceptual [CK95, Lon91a]. Concerning [Ree98]. Concevoir [Fer99]. Concise [Bar90]. Concrete [Wal99b, Isr92b]. Condit [Dre97, Iri97]. Condition [Rey94]. Conditioning [Cap93, Coo94]. Conditions [Mic94]. Confederate [Bel98b, Sum93, Col99]. Conference [Fit97, HC98]. Confidence [App96]. Conflict [Tal94, Rol98a]. Confrontation [Gol91a]. Confronting [dlP99]. Confronts [Sil99]. Congress [Sti98, War97]. Congressional [Lew96]. Congreve [vK92a]. connaissances [Fer99]. Connecticut [Vic92]. Connections [Maz93]. Conquest [Bul91, Kid90, Sel90]. Conrey [New91]. Consciousness [Gut95a]. Consensus [All96d]. Consequences [Sin99b]. Conservation [Hug99, Ste94b]. Conspicuous [Sta90b]. Conspiracy [Mil94, Tra98b]. Constant [Scr00]. Constitution [Dou95]. Construct [Moh97b]. constructeurs [But92]. Constructing [Dre99, Fit97, Smi99, Ros93]. Construction [Abb96, Bar92b, Buc99, Coo91c, Jac96, Kie90a, KP96, Mat92b, Pet96, San97, Wer97, Kah96]. Constructions [Cut98, Tod97]. Constructive [Jac97]. construire [Fer99]. Consultants [Don94]. Consulting [Wor95]. Consumer [App96, Lai93, Scr94]. Consuming [Hor99, Wil99b]. Consumption [Bla97c, Hes99]. Containers [Rey98]. Containing [Hac94a]. Contemporary [Mer98, Moh97a, SS99, Sin99a, Thr91]. Contesting [Pan95]. Context [Arn97, Gam92, Gra95, Har98b, Lon91b, Mar93b, Mar91, Pfa91, Tra97b, Zie94, Wan96, KM92]. Contexts [Pan98b]. Continental [Mer98]. Continuities [Car96]. Contract [Kie92]. Contracts [Cap90b]. Contradictions [Cla90, Gut95b]. Contrasts [Qui98]. Contributions [Bed90b]. Control [Ald93, Bad94, Ben91c, Ben92c, Cla90, Hon98, Ken92, Lan98a, Law92, Neu98, Nye98a, Reu91, Sch94d, Ste93, Lub91a]. Controlling [Elv99]. Controversy [Eva93b, Lea92, Lin98, Sei96a, Ste94b, Tra98b]. Conveniences [Kea98]. Converter [Gor92c]. Cooper [Cha93, Mac93b, Til90]. Cooperation [Hei91]. Cooperative [Lie90]. Coopersmith [Wei94]. Coordination [Tho90, Lip97]. Coos [LaL90]. Coosa [Sum96]. Copies [Lon93a]. Copp [Cut94a]. Copper [Dub93, Har91, Mou99, New97, Qui93]. Cord [Tri93]. Corn [Fit93a, Rug96, Sha91]. Corncribs [Eva92]. Corporate [Bir90, Bla90a, Gal92a, Mer90, Pul92]. Corporation [Mea98, Nas90, Scr98a, Vol91c, Fos98]. Corporations [Les90b]. Corps [Bel98a, Han93, McN93, Mer95, Mer96, Reu99a, Sla98]. Correspondence [Hug96a]. Corruption [Rei96]. Cortada [Cer97, Kid94, Sei96b]. Cost [Eva93a]. Costs [Lea92]. coton [Lin94b]. Cotterell [Fer92]. Cotton [Bla90b, Cou94, Gli94, Jer93, Jer94b, Lea92, Lea99, Win90]. Cottonopolis [Web90a]. Could [Chu99]. Council [Jar90, Rol93b]. Count [Goo95]. Counter [Mil94]. Counter-Arts [Mil94]. Counting [Phi93]. Countries [Ile96]. Country [Bor97]. County [O’H96]. Courage [Bil97]. Course [Jer94b, Kel99]. Cox [Log91]. Coy [Ror91]. Cozzens [Maz93]. Craddock [Gor97a]. Cradle [Lay97, Dub93]. Craft [Bix99, Len91, Pal93, Til97]. Crafting [Ler98]. Craftsman [Old97]. Craftsmanship [McG99]. Craftsmen [Lon95a]. Craig [Hir96, Lin91, Low97, Miy94]. Crary [Kat92]. Cravens [Sta97b]. Crawford [Big97, Kra98, Vol91b]. Created [Hac94b, Kid94]. Creating [See99a, Slo95, Cra97, Ste90]. Creation [Bau93, Bro98a, Kli97a, Kuc94]. Creativity [Hea91a, Hyd90, Pre97]. Credit [Lai93]. Creek [Hir96]. Creese [Seg94]. Creighton [Roz98]. Creveld [Hac92b]. Crisis [Mac92, Mit95, Elv99, Ell94b, Rob96a]. Critical [Ber95a, Bla97c, Fit97, Pin92, Skr91, Kal95, See95b]. Critique [Mac92, Slo98]. Cromley [Con92a]. Cronon [Con92b]. Crosbie [Por91]. Cross [Nel90]. Crossing [Min95b]. Crossover [Tho91b]. Crossroads [Isr92a, Has93]. Crouch [Gol91c]. Crow [Str90a]. Crowd [Lai92]. Crown [Pet97]. Crozet [Smi91a]. Crusading [Ste98]. Crystal [Gut93]. Cuba [PO99]. Cuban [Alv99]. Cudahy [Hil92a]. Cueto [Gua98]. Cuisine [Bel97]. Cult [Sta92b]. Cultivated [Moh90]. Cultivation [Car96]. Cultural [Bla97b, Ell94b, Har98b, JJ98, Rhe98, Sch93b, Skr93, Sum97, Skr93]. Culture [Ano93k, Ber97a, Blu99, Bod99, Bra97, Cre99, Dan95, Dre99, Fis92, Fit93b, Fit97, Fro93, GK93, Gui95b, Has93, Hof91, How98, Hur90, Hur99, Joh92, Koh96, Kon97, Kuc94, Mau93, Moh90, Nea98, Neu90, Pos98b, Pri95, Reu93b, Ros93, Rot99a, San97, Sil99, Sim94, Spe95, Sta97b, Til90, Vol91a, Wil99a, Wil93a, Ker99, Fer92, Hod93, Kri98, Lai96a, Lon96a, Was91, Whi92a, Wil95c, Ano90e, Ano91g, Ano92f, Ano93f, Ano94f, Ano95f, Ano97f, Bog98, Bro98a, Pan98a, Rab98, Wal99a]. Cultures [Bla99a, Gol90, Koc93, Lev91c, Lan91a]. Cumbria [Har91]. Cummins [Bry95, Chu99]. Cuneiform [Phi93]. Cunningham [How93]. Cup [Cas99]. Curators [Fit93b]. Current [Ell94b, Gie96, LAC^{+}90, LAC^{+}91, LAC^{+}92, LAC^{+}93, LAC^{+}94, LCH^{+}95, LCH^{+}96b, LCH^{+}97, Low99a, Low99b]. Currents [Hec92, Wos91]. Curtis [Zie92, Mah90]. Custom [Coo94, Tan93]. Customs [Pos98b, Sch99b]. Cutcliffe [Mar91, Til94]. CyberCities [Nei98]. Cybernation [Fea99]. Cybernetics [Kap92]. Cyberpunk [Sla97]. Cyberspace [Bow98]. Cyborg [Cap91]. Cynthia [Moh94, Old92, Rot95, Thr91]. Cyril [Hug93a].

D [Bar92c, Bas96, Ber93, Bro97a, Bro95c, Fit95, GS98, Gam97, Gol92, Hat90, Hof92, Hol97b, Joh91, Kid96, Kuc94, Mic94, Moh90, Moh95, Mul96b, Ran97, Ras92a, Rhe94, Scr91a, Sha90a, Sha95a, Swe93, Vin93, War91, Wor94]. D.C [Pos98a, Rei96]. Daguerreotype [Lie93, Lie93]. Dahl [Bjö97]. Dahlgren [Min97]. Daily [Jac96]. Dallas [Koh96]. Dalton [Cra97]. d’aluminium [Hal93a]. Dam [Jac96, Nye98a, Smi95a]. Damming [Smi92b]. Dams [Jac95a, Jac96]. Danbom [Bor97]. Danby [Rol94]. Danger [Jac96]. Dangerous [Mit95]. Daniel [Big94, Bla97c, Bro90, Cap95, Fin93, Jan91, Lav93, Lip97, Nuw90, O’H96, Har91, Uss93]. Daniels [yH98, Thr91]. dans [Kah96]. dansk [Hår98a]. danske [Myl97]. d’anthropologie [Fer99]. Daoguang [Elv99]. d’application [Lan98b]. Dark [Gre94]. Darkness [Sun97]. Daryl [Til97]. Dasgupta [Pre97]. Data [CK98, Coo94, Mah97]. Databases [Bat98]. David [Ada95a, All90, Bir90, Bor97, Cal92b, Cam95, Cod92, Coo97, Coo98c, Daw91, Fle92, Gor96, Hac94b, Hea96, Hir92, How98, Isr92b, Kel93b, Ken92, Kid90, Lai96b, Lea92, Lin94a, Low91, Low93b, May92, Mei95, Mis90, Pri95, Rei96, Rou96, Sum91, Swe92, The90, Tra97a, Tri90a, Uss93, Whi93, Wig96, Wil95c, Wor93b, Fie98]. Davies [Van97, Wos92]. Davis [Coo91d, Gib93, Kuc92, Sta90b]. Davison [Sal95]. Dawn [Bro98b, Con90, Per91]. Dawson [Hol93]. Day [Pur97, May94]. Days [NP94, Smu90]. DDT [Rus99]. Deal [Mat99c]. Dear [Jes97]. Death [Dub93, Fos98, Kry90]. Debate [Gol93, Ryc93]. Debeir [Bin93]. Debora [Lev91a]. Deborah [Haf96, Sha91, Wor95]. Decade [Wes91]. Decades [Gru90]. December [Cro99, Isr94a]. Decision [Gal92b, Log92, Mel99, Nea98, Slo96]. Decision-Making [Slo96]. Decision-Support [Mel99]. Decline [Ber93, Cas98a, Gre95a, Jer94b, Jer94a, Leu98, Mar90, Sch94a, See90]. Deep [Zac98]. Defense [Ber90, Fri91b, Owe91a, Rol92, Han93]. defining [Leo92]. Definition [Rob95]. Deforestation [Eve92]. d’el [Was92, PO99]. Delaware [Sha90a]. della [Sel92]. Demise [Cla90, Whi93]. Democracy [Moh90, Sch94d, Sin96, Mit97]. Democratic [Cla90, Oga90]. Denis [Jam90]. Denise [Phi93]. Denmark [Hey98]. Dennis [All97, Smi95a]. Denver [Vie90]. Department [Col99]. Depiction [Sil93]. Depot [Dic92]. Depression [Lin97b, Smi95a]. d’équipement [Smi91d]. Derek [Ang90, Ben92b, Gin90]. Descartes [Cow93]. Description [Bra91a]. Descriptive [War99]. Design [Ben91a, Big97, Blu99, Coo94, Div94, Dre99, Gor92b, Hyd90, Kat94a, Kat97b, Kub90, Kuc94, Law92, McG99, Mei90, San92, Sch99a, Ube99, Wer97, Whi92a, Fri91a, Lai96b, O’D97, Pul92, Sha94, Fer95]. Designer [Far99]. Designers [Kat97a]. Designing [Ben99, Mag95, Øve97]. Designs [Hec94]. Desire [Joh99b]. Deskilled [Fit93a]. Despite [Och92a]. Dessalles [Paq97]. Destombes [Bed90b]. Destroyer [Sum93]. Destruction [Fer90, Rot99c]. Determination [Lan93]. Determinism [McB92, Scr95, Wen98]. Dethloff [Oli90]. d’ethnologie [Lev91b]. Detlef [Jam90]. Detroit [Sta90b, Olw96]. d’étude [Lan98b]. deutsche [Mul90a]. Deutschen [Fin98, All96a, All96c, Lun90a]. deutscher [Brü94]. Deutsches [And92]. Deutschland [Die93, Pet98, Rüd92]. Developing [Bud96, Rot99b, Sul98a]. Development [Car98, Car99a, Con90, Dot99, Har98b, Hev95, Jam90, Jer97, Kel95, Kie90b, Koc93, Kra94b, Law92, Moo98, Nah97, O’B91, Ols97, Rei99a, Rei93b, Rob94a, Ron99, Sch93b, Sch91, Str90a, Tri90b, Vol98a, Yat92, Zac98, Zie94, Car99b, Car99b, Tri90a]. développement [Hal93a]. Deverell [Min95b]. Deviance [Nea98]. Device [Wor90]. Devil [Cro99]. DeVorkin [Daw91, Hac94b]. Devoted [Qui95]. DeVries [Bac98, Par94]. Dew [Scr96]. Dewey [Gol91a, Pol91]. Dexter [Ano90b, Ano91d, Ano92c, Ano93c, Ano94c, Ano95c, Ano96c, Ano97c]. d’histoire [Lev91b, Smi95b]. Diagnostic [Koc93]. Dialects [Bel91, MB92].

Diamond [Fel99]. Diane [All98a, Fis98, Lon97]. Dianne [Gib91, O’B91]. Diary [Paq97]. Dibner [Ano90c, Ano91e, Ano92d, Ano93d, Ano94d, Ano95d, Ano96d, Ano97d]. Dick [Zie92]. Dictionary [Mal91, Pur97]. Diderot [Pan98c]. Dienel [Brü94]. Diesel [HK99, Sch99b, Bry95]. Dieter [Neu96b, tZS90]. Difference [Hew94]. difficile [Lau98a]. Diffusion [Hea98, Joh91, Mac91, TJ96, Uss91, Mul90a]. Digital [Sma95]. Dilemma [Wen98]. Dilić [Rot95]. Dilsaver [Miy94]. Dilts [Bro95c]. Dintenfass [Jer94a]. Diplomatic [Hal98]. Diplomats [Nic99a]. Direction [Hig98]. Directions [LMO97a]. Dirt [Hoy92, Mel96]. Disappearing [Vol91a]. Disarmament [McB97]. Disaster [Har99c]. Disciplinary [Sla92]. Discipline [Bac98, Cow96, Fit97, Wal91, Rei93b, Jes97]. Disciplines [Hig97a]. Discontents [Sin97]. Discontinuity [Kat94b]. Discourse [Ger96, Owe91b, Uss97]. Discovering [Smu94, Rei91]. Discovery [Bil98, Tra97a, vHT96, Cut94a, Smi98]. Discrimination [Jac96]. Discursive [Mit91]. Disease [Aik98, Dav93, Tap99]. Disenchanted [Opi90]. Display [Gol90]. Distance [Ada93]. Distillers [Cal98]. Distinctions [GS99]. Distinctive [Gli94]. Distribution [Rey98]. District [Gor96, Har94]. Districts [You93]. Disunity [Pan98b]. Diversification [Cal98]. Diversity [Scr94, Hut92]. Divide [Kra90]. Divided [Rob94b, Leo92]. Divine [Bul94]. Division [Hor97b, Liu97, Kay91]. Dix [San90]. Dneprostroi [Hat90]. Do [Ben96, San90]. Dockyards [Pet92b]. Does [Rol97, Wen98]. Dogs [Rei99a]. Dohme [Car99a]. Dohrn [LF93]. Dohrn-van [LF93]. Doing [Pas96, vHT96]. Domain [Old97]. Domestic [Bla97c, FF96, Hoy92, Moh97b, Rue90, Sha95b, Wan96]. Domesticity [Hoy92]. Dominance [Old92, Hea90]. Domingue [Wil93b]. Don [Gol91a, Mit92, Ras90]. Donald [DeV92, Dud92, Edw97, Fri90, Fro92a, Hug90, Mul94, Nye98a, O’M93, Reu95, Rob94b, Rot92, Rot94a, Sei93, Sim95a, Sta90b, Wer97, Zem93]. Donaldson [Die97]. Done [Ols98]. Dooley [Smi91a]. Doomsday [McN90]. Dora [Ale99a]. Dorf [Lin93a]. Dorn [Tod95, Hug93b]. Dorothy [Chi97]. Double [Jag99]. Double-Bitted [Jag99]. Douglas [Bad94, Ber94b, Cra97, Fin95, Tho94, Smu90]. Down [Hug95]. Downs [Liu97]. Draft [Kub90]. Draftsmanship [McG99]. Drag [Gui95b]. Draining [Cla95]. Drawing [Bel91, MB92, Tap99, Has93, Bed90a]. Drawings [Har96b, Lon93a, Lon95b]. Dream [Ber94a, Cro94, Kel99]. Dreams [Sch93a, Sta96a, Isr94b]. dredging [New91]. Drehautomaten [All96c]. Drengson [Hig97a]. Dressmaking [Gam95, Lew99]. Dritten [Neu96a]. Drive [Wen98, Con95]. Driven [Chu99]. Driving [Vol93]. Drug [Lin98]. Drugs [Ser92]. Dryden [Bil98]. Dublin [Coo96, Kon98]. Due [Man99]. Duffy [Ham92a]. Dufour [Smi90a]. Duitsland [Wac94]. Dunar [Smi95a]. Duncan [Tra97a, Wos92, Wal91]. d’une [Hal93a]. Dunlap [Gos90]. Dunlavy [Ber95c]. Dupin [BP91]. d’Upsal [Frä91]. Dupuy [Bol90]. Durables [Lai93]. durante [Gli91]. during [Hau90, Jan91, See95b, vHT96].

Düsenflugzeuge [Neu96a]. dustri [Hår98a]. Dutch [HC98, Smi98]. Duties [Ler97]. Dwarfs [Cor97]. Dyeing [Adr94]. Dyestuffs [Roc94]. Dyke [Hac98]. Dynamic [NP94]. Dynamics [Bla97b, McB97, Scr91a, Sto97].

E. [Daw96]. Eagles [Sch91]. Eakin [Was92]. Eames [Kat97a]. Eamon [Lon96a]. Ear [Con92c]. Early [Bac98, Bal98, Ber95c, Bla97b, Buc94, Con92a, Eno98, Gam97, Haf90, Hoc98, Joh91, Kra94b, Kuc94, Lai96c, Lea92, Lig97b, Lon95a, Lon96a, Mac92, Mah97, McG95, McG99, Mil97, Nor90, Phi99, Rab98, Rey99, Sch92a, Sha90b, Sha93, Smi90a, Smu90, Sta97b, Sto97, Tar95, Tho93, Wes99, Bor99, Bro99b, Bed91, Gor97a, Pas96, Sta92a]. Earth [Bar92b, Gol91a, Cal92b]. East [Wig96, And99b, Bla92b, Kri99, Hei95a, Sel97]. Eastern [Mea98]. Eastman [Lie98]. Easton [Mat90]. Eastwood [Nye98a]. Eaton [Cod91]. eau [Rei99b]. Eaves [Mil94]. Ebner [Ste90]. Eckert [Rüd92]. Eclipse [Pla96]. École [Lan95a]. écoles [Lan91b]. Ecological [Bia97, Cal92b, Jud90, McE95, Reu95, Ste91a]. Ecology [Bes95, Fit98, Gre99]. Economic [Bac99, Bar90, Bla97b, Cam92, Con90, Har98b, Har96c, Hea91a, Hei95a, Kel95, Nye93b]. Economics [Alv99, Gol99, Gol97, Mea98, Yat96]. Economy [Fle90, Haf90, Lew99, Neu92, Ryc92, Tho91b, Tra95, Bar92c, Lam93, Nye93a, Tas97]. Ecophilosophy [Hig97a]. Ecopolitics [Gra97a]. Ecuador [Gua98]. ed [Abb96, Alb97, All98a, All98b, Ang90, Asp93, Bad91, Bas97, Bec91, Ben92a, Ben91a, Ben91b, Big94, Bla97b, Bla97c, Blu92, Bol90, Bra96, Bro97a, Bro90, Bry92, Buc94, Cam93, Cap91, Cha90, Cla96, Cod92, Col93, Coo97, Coz93a, Cur91, Cut94b, Cut98, Dav93, Die93, Ebn92, End97, Fit95, Fit97, Fri96, Ful95, Gar91, Gar90, Gib93, Gie96, Gie94, Gis98, Gol90, Gom97, Goo95, Gor91, Gor92a, Gua98, Gui93a, Gus91, Hår93a, Hac91b, Haf97, Hal93a, Har91, Har93b, Har96c, Har96b, Hea91b, Hea98, Hig96, Hil91a, Hod93, Hof92, Hou90, How98, How93, Hug96a, Hun96, Ile96, Isr92a, Isr94a, Jac97, Jan91, Kal95, Kay93, Kay91, Kel93b, Kel96, Kid96, Kli91, Kni95, Kra93, Kri97, Kri98, Kry92c, Lan98b, Lar91, Lar98, Ler98, Les94, Lev91c]. ed [Lev97a, Lie93, Lin94a, Lin93a, Lip97, Lon95b, Lon95a, Lon97, Low96, Low92, Lue90, Lun92, Lup91, Mal91, Mar93a, Mar91, Maz93, McG97b, McM94a, McM94b, Mil97, Mil95, Min96, Mit96, MA97, Miy94, Moh97a, Mor91b, Mor97, Mul96a, Nel90, Nob97, Nye90, Nye97, O’B91, O’D94a, Och92a, Øve97, Paq97, Pas96, Pet92a, Pet90, Pet97, Pfa91, Pfa98, Pin91, Pin92, Por93, Por94, Pry93, Pur97, Pye90, Pye92, Qui94, Qui95, Ras92b, Rhe94, Rob94a, Rol95a, Ror91, Ros92, Rot93, Rot95, Roz98, San94, Sch93a, Scr98b, See98b, Sel97, Sha95a, Shi97, Sim90, Sim95b, Sin93, Sin97, Skr91, Sol98, Sta97a, Sta97b, Ste95, Sti91, Swe93, Ter94, The90, Tho94, Thr91, Til94, Til97, Tod96, Tra97a, Tra97b, Tri94, Vas98, Wan96, War92a, War93, Way90]. ed [Web90c, Wer92, Wes90, Wes93b, Whi92a, Wig96, Wil95a, Wil95c, vL90]. Edelstein [Adr94, Tra96]. Eden [Hur98]. Edge [Fri99]. Edgerton [Woo92]. Edison [Cha90, Har93b, Kli91, Mil93, Mil97, Nye92, IR91, Isr91]. Edition [Cha90, Lea92]. Editor [Sta98, Ste91b]. Editorial [Fri95]. Editorship [Ano93k]. Edson [Hyd90]. Education [Abb96, Div94, Hac93, Lan94b, Rey94, See95c, Has93, Pla96, Kni95, See98b]. Edward [Ber90, Cla90, Edm96, Fri92b, Gol92, Hoc98, Bla99b, Mol95a, Scr00, Tra98b]. Edwards [Les90b, Uss97]. Edwin [Kay91, Mil93, Smi91a, Ham92b]. Effect [Hew94, Kra98]. Effects [Kli97a, Kli97b]. Efficiency [Min99]. efter [Hår98a]. Eggert [Ing94]. Eiffel [Wil91]. Eighteenth [Ben90, Ble91, DeV91, Gut90, Haf92, Lay93, Low93b, Rhe94, War92a, Don91]. Eighteenth-Century [Ben90, Ble91, Haf92, Rhe94, Don91]. Eileen [Thr91]. Einstein [Har95b, Hol91]. Eisen [All96a, Pet98]. Eisen- [All96a]. Eisenbahn [Gam99]. Eisenhower [Bul94, Hac91a]. Eisler [Lie91]. Elborg [Paq97]. Eldoret [Har97]. Eleazar [Ano90i, Ano91l, Ano92j, Ano93i, Ano94j, Ano95j, Ano96h, Ano97j]. Electric [Bri90, Chr91, Cla92, Cor97, HS96, Hir96, Hug96b, Kir95, Kli97a, McB92, MM93, Ros94b]. Electrical [Ano90d, Ano91f, Ano92e, Ano93e, Ano94e, Ano95e, Ano96e, Ano97e, Arm96, Kli90, Mat99c, Myl91, Ros94a, Svi93, Kön96]. Électricité [Smi91d]. Electricity [Coo92b, Fin99, Neu92]. Electrico [PO99]. Electrification [Eva93a, Wei94]. Electrifying [Hir92, Nye93a]. Electron [Kun95, Ras99]. Electronic [Cap90a, Nei98, Oga90, Phi91, Sin96, Wos91]. Electronics [Jam90]. électroniques [Coh90]. Elektrizitätsversorgung [Buc97]. Elektrizitätswirtschaft [Buc97]. Elements [Ell94a, Sei96a]. Elevator [Mil91a]. Eli [Vic92]. Elihu [Ros94b]. Elisabeth [Kra98]. Elites [Sta90b]. Elizabeth [Con92a, Rot93]. Ellen [Fin95]. Ellis [Bri93]. Elspeth [Kel93a]. Elting [MMH96]. Elwood [Smi94]. Embankment [Bow99b]. Embree [Pin91]. Emergence [Bac99, Cha99a, Kon98, Mea98, Poo99, Vas98, Win90]. Emerson [Edw96a]. Emigrants [Jer90b]. Emilio [Gol95]. Emperor [Elv99]. Empire [Ada98, Jac91, Kei94, Kub90, Por91, Ran97, Sam99, Sam99, Smu93]. Empires [Lig97b, Lan91a]. Empirical [CK95]. Employers [Div94, Gar90]. Empty [Lay97]. Enabling [Kle99]. Encyclopaedia [Low92]. Encyclopedia [Bre99]. Encyclopédie [Pan98c]. End [Cha95, Con92c, Olw98, Szo96]. Ende [All96c]. Endless [Bas99]. Energies [Wil99b]. Energy [Bin93, Cla90, Cla92, Gie96, Hac91a, Har94, Kar99, Maz90b, Maz94, Ryc92, Sei96a, Str91, Wal94, FG98, Maz90a, Ors97, Por91, Wil95b]. Engagement [Pin92]. Engaging [Dur98]. Engendered [Cam93]. Engine [Bry95, Chu99, Law92, Luk99, Nah97, Tar95, Bow92]. Engineer [Ada95b, Bas99, Chr99, Chr93, For93, McC93, McW95, Pur93a, Smi91a, Cut93, Pfa97, Hol97a]. Engineered [Big95b, Pfa92]. Engineering [Ald97, Atk99, Bel91, Ben92b, Bud96, Cal92a, Cap93, Car97, Caw95, Chr95b, Coo94, Div94, DB94, Dre99, Gar91, GA98, Hac93, HK99, Har96b, Hau90, Jac93b, Joh97b, Kle99, Kön96, Lay91, Lev97a, Lew96, Low91, Mol91, Pur93a, Pur99, Rey91, Rey94, Rot99a, See93, See95c, SD98, Smi90a, Ste92, Sun99, Tra91, Vin97, Wor91, MB92, Pfa97, Fer95, Lon97, Ros94a, Bry92, Cha98, Hon98, Hou97, Kra94c, Mai98]. Engineers [Atk99, Flo92, Ham92b, Han93, HM96, Lay91, McC95, Mer95, Och92b, Sch99a, Ser96, Slo95, Tho91a, Vin93, Wor95, Reu99a, Mag95, Mer96, Vol91b, Lev97a, Mei98a]. Engines [Hoo99, Win98, Str90b, Hol93]. England [Bla97b, Bra92, Cam95, Coo96, Gre91, Lay93, Lea98, Was93, Cla95, Hug99, Phi99, Ste89, Ble91, Mue90, Per93, Smi93b, Ste91a, Tra98b, Whi92c, Woo92]. English [Gar90, Rob90b, Tan93]. Enigma [Eno98, Hon97, Hun95, Min99]. Enlightenment [Hau90, Mau93, Nye93b, Pla96, Rob95]. Enola [Pre98, May98]. Enseignement [Lan98b, Lan91b]. Enterprise [Asp93, Hea98, New91, Pye90, Sha96a, Was92, Ben96, Bry92, Don91, Lip97, Meu98, Pay95]. Enterprises [Con94b]. Enters [Wil90]. Entertainment [Bro97b, Kra94a, Pla96, Gom97]. Enthusiasm [Mor91a, Sch93a, Vol93]. Enthusiasts [Isr94b]. Entomology [Fit98]. Entrepreneur [Dia99, Car99b, Mai96]. Entrepreneurial [DB94, Les90b]. Entrepreneurs [Mic94, Mor98, Sha96b]. Entrepreneurship [Coo91a, Jer94a]. Entreprise [Hea99a]. Entropy [Cut93]. Entstehung [Buc97]. Entwicklung [All96c, Dek98, Die93, Lin93b]. Environment [Bow99b, Coo92b, Dav93, Gei90, GK93, Hac94a, Ham95, Ham98b, JM97, Low96, Miy94, Nas90, Qui93, Qui95, Rei95a, ST98, Cla91, Qui95]. Environmental [Aik98, Ell94b, Reu95, Sim95a, Sim95b, Bra91b, Qui95]. Environmentalism [Hur98, Rus99]. Environments [McE95]. Envisioning [Fri92b]. EPA [Low97]. Epic [Hil95, Gor96]. Epidemic [Far96]. Epidemics [Dav93]. Equipment [McS98]. Era [Kle95, Pas96, Rob90a, Rob92a, Bor99, Jac93b, Jos99, Hir96, Tho93, Wos94, Buc98b, Rab98]. Eras [Mah90]. Erasing [Sum97]. Erfindergeist [Wen92]. Eric [Bri90, Hig91, Isr90, Roz90, Spi95, Tod95]. Eriksson [Bjö97]. Ermenc [Bad91]. Ernest [Kir95]. Ernst [Neu94, Wac94]. Erratum [Ano99g]. Error [Fer95]. Ersten [All96c]. Espace [Lav93]. Essay [Cha93, Kat97b, Sla91, Woo92]. Essays [Ada93, Ben99, Cam92, Dav93, Edw97, Fel99, Gom97, Goo95, Har98b, Joh92, Kay91, Kid96, Lev91c, Mar91, Och92a, Pet90, Por93, Rot93, Sim95b, Wil95c, Cod91, Ebn92, Lay93]. Essence [Doo91]. Essex [O’H96]. Estate [Ber93, Poo99, Fle90]. Estrange [Goo95]. Età [Sel92]. Etableringen [Myl97]. Etel [Wan96]. Ethan [Phi91]. Ether [Mac93a]. Ethical [Hil91a]. Ethics [Bec91, Chr99, Kue99]. Ethiopia [Zel97]. Ethnic [Jud90]. Ethnography [Bix99]. Ethnological [Sun97]. Ethnologies [All98a]. Etienne [War92b]. Etnologiska [Sun97]. études [Frä91]. Eugene [Kra94c, Lew93, Nye90]. Euler [Kip97, Ste94a]. Europe [Bed91, Bol90, Haf90, Haf97, Kni95, The90, Tra97a, Dor91, HV99, Mue99, Rot95, Fle92, Lon95a, Roc94, Cam92, Dol97]. European [Hea98, Kuc94, Tra97b, Voe90, All99b, For93, Til97]. Evan [Ben91b]. Evans [Jac93a]. Evelyn [Lin97a]. Events [Mag94]. Everett [Hac91b]. Everyday [Dav92a, Fri90]. Evolution [Bry96, FF96, Hyd90, Kat94b, Kid90, Lee97, May91, Sch94b, Sha95b]. Evolutionary [Con92c, GS98]. Exact [Ada96]. examined [Ros92].

Excavating [Wor90]. Excavator [Stu94]. Excursions [Skr93]. Executed [Ada95b]. Executives [Mei98a]. Exemple [Lan91b]. Exeter [Cla97a]. Exhibit [Por97, Bra98]. Exhibiting [Bod99, Fin99, Gro93]. Exhibition [Por97]. Exhibitions [Ols97]. Existential [Pfa97]. Exotic [Cha96]. Expanding [Ber94a]. Expansion [Ada96, Hoc98, Kai98, Sha96b]. Experience [Bil97, Hea92, Jes97, Min95a, Nye95, Ran97, Tho93, Col93]. Experiment [Mil95]. Experimental [Coo93a]. experimentele [Dav96]. Experiments [Mit91, Sok93]. Expert [Gol93]. Expertise [Gus91, Sti98]. Experts [Rus99]. Exploitation [Mor91c]. Exploration [Dre99, Kei94, Pat95, Swe92, Gru90]. Explorations [Kyn99]. Exploratorium [LaF91]. Exploring [Hig97a, Har92, Yat96]. Explosions [Ald95]. Exposition [Pan95]. Expositions [Rum97]. Extraordinary [Gol91b]. Exxon [Nas90]. Eye [Eis93, Wos94, Kra94c, Rot92]. Ezell [Mol95a]. Ezio [Fri91a]. Ezrahi [Nye97].

F [Cal92a, Cap90b, Cap93, Cap95, Chr91, Fer90, Fri91b, Hol95, Hol91, Hou97, Jer94b, Kli90, Lan91a, Les90b, Low91, Mar96, Mau91, May92, Mil91b, Pri95, Pur90, Smi90a, Smi91a, Vol93, Wor93a]. F-4 [Hou97]. Fabric [Lüt98]. Fabrics [Per99a]. Fabrik [Lin93a]. Fachgebietes [Buc97]. Fact [Kra93]. Factories [Ing94, Til90]. Factory [All96d, Big95b, Bra91a, Cam95, Fea99, Gro98, Lea98, Mue91b, See99b, Vic92, Web90a, Coo92a]. Facts [Pfa90]. Fad [Neu90]. Fahey [FO92]. Fail [Gor98, Mor98]. Failed [CS90]. Failure [Car97, Kai96, Kun95]. Failures [Coo91d]. Fair [Ste93, Rou96]. Fairbanks [Ebn92]. Fairs [Rum97]. Fall [Ada95b, Ada95a, Bel97, O’D96a, Pur93b, Smi90b, Sti98]. Fallout [Hac90a]. Falls [Ber98, Sha94]. Familiar [O’M90]. Family [Ben90, Bla90b, Gut95b, Paq97, Kle91]. Fan [Bro98a]. Fantasies [Bla99b]. Fantastic [Tra98a]. Farben [Kra90]. Farley [Rab96]. Farm [Fit96, Kli97b]. Farmer [Fit93a]. Farmers [Fit93a, O’H96]. Farming [Han99, Tri90a]. Farther [Rei93a]. Fascism [Car98]. Faster [Rei93a]. Fate [Ald97]. Father [Cap90a]. faubourg [Bel98a]. Faust [Bas96]. Feaver [Bad94]. February [Kli91]. Fed [Ras90]. Federal [App96, Maz90a, Ser92, Rus99, Ser94, Slo95]. Feeding [Gor94, Rol95a]. Feeling [Ste93]. Feenberg [Nob97]. Feldenkirchen [Mai96]. Fellows [Lev97a]. Fellowship [Lew96]. Female [Cow96, Lew99]. Feminism [Rue90]. Feminist [Her99a, McG97a]. Ferdinando [Tra97b]. Ferguson [Kas99, Kra94c]. Ferkiss [Ell94b]. Fermilab [WH96]. Fernando [MA97]. Ferranti [Arm96, Kli90]. Ferranti-Packard [Arm96]. Ferré [Bec91, Sel91]. Ferries [Kin99]. Ferris [Wor94]. Ferryboats [Hil92a]. Festival [Lon95b]. Festschrift [Tod96]. Fetal [Ree98]. Fever [Bri93]. Fiasco [May98]. Fiction [Kra93, Sla92]. Fictive [Kel95]. Field [Ada97, Cha99a, Fit95, Lon95a, Rey91]. Fighting [Win98]. Figured [Liv92]. Fin [Lev91a]. Fin-de [Lev91a]. Finally [Mar90]. Finance [Owe91a]. Finch [Cor95]. Finding [Hig98]. Findlay [Spe95]. Fine [Con92c]. Fine-Tuning [Con92c]. Finishing [Adr94]. Finland [Nye93a, Myl91]. Finlay [Sta90b, Fit90]. Finn [Svi93]. Finne [Hal90]. Finnish [Ahv90]. Fire [Dun93, Eam97, Ing99, Rol92, Sim94, Smi97a, Win98]. Fires [Win98]. Firm [Moe94]. First [Bro95c, Cap90a, Coo98b, DB94, Gam92, Gib93, Gol91b, Gru90, Haf96, Kyn99, Omi92, Pol91, Rot92, Ste92, Vol91c, Whi98, vK92a, Blu99]. Fischer [Vol94a]. Fishbein [Lee97]. Fisher [Old97]. Fishermen [O’H96]. Fitton [Jer90a]. Fitzgerald [Sha91]. FitzSimmons [Rob92b]. Fix [Gro93]. Flame [Sim94]. Flames [Hoo99]. Flamming [Cra97]. Fleet [Hev95]. Fleming [Low96, Ell92b, Eva93a]. Flexibility [Zei95]. Flexible [Bla95]. Flichy [Lav93]. Fliers [Neu93]. Flight [Bil98, Mei99b, Sch92a, Wil90, Tri90b, Lee97]. Flights [Bil98]. Flood [Ano99n]. Florentine [Kel99, Mue91b]. Florida [Hur98]. Floris [Lon96b]. Florman [Pfa97, Pfa97]. Flour [Mue91b, TJ96]. Flow [Eve92, Sch95]. Floyd [Lan90b]. Flugabwehrraketen [Neu96a]. Flying [And99b, Cor91, Lau99, Cut93]. FM [Slo96]. Fokker [Mei98b]. Folly [Bla96a]. fondation [Frä91]. Food [App96, Hut92, Ser92, Sel90]. forandringer [Hår98a]. Forbidden [For97]. Force [Cla96, Day99, Ken92, Mic94, Hil92b]. Forced [Lau95]. Forces [Edg98]. Forcing [Fea99]. Ford [Bra98, Isr94b, Lan94c, O’D97, Zem92, Big96]. Förderung [Lon93b]. Fordism [Vol98a]. Fordyce [dlP99]. Forest [And99b]. Forester [Pfa91]. Forestry [yH98]. Forests [Ing91]. Forge [Scr96]. Forging [Jos95a]. Forgotten [Cap90a]. Formation [Fis92]. Formative [Ste99]. Forms [Cla91]. Forschung [Die93]. forskning [Car99b]. Forskningscenter [Hei99]. forsta [Har99a]. Forster [Paq97, Paq97]. Fort [Fer90, Yea90]. Fortification [Kuc94]. Fortifications [Lon95b, DeV93]. Fortress [Cor93, Han93]. Fortschritt [Tod99]. Fortune [Jer90a, Kel99]. Fossil [Maz90a]. Foster [Jac91, Mor91b]. Foundation [Bud94, Mil95, Mor98, Ste92, Kle99]. Foundations [Kal95, Kay92]. Founding [WH96]. Four [Gui91, O’D96b]. Fournier [Lüt98]. Fourteenth [Bac98, Pos98a]. Fourth [Kat94b]. Fowler [Hut92]. Fox [Kni95, Lin97a, Ald96, Scr98b]. Foy [Bra96]. Fradkin [Hac90a]. Fragile [Gro90a]. Frame [Ube94]. Frampton [San97]. fran [Car99b]. francais [Hal93a]. Francaviglia [You93]. France [But92, Eat90, Eis93, Eva96, Gei90, Gut92, Gut95b, Hea99a, Hec92, HV98, Kah96, Lan91b, Lan94b, Lev91a, Lin94b, Ros98, Tal94, War92b, vR90, Ald97, Ben94, Bes95, Dre99, Gou98, Hec94, HP91, Mac91, Mit95, Nic99a, War92b, Ald96, Ben90, Bra92, Lay93, Rei91, Smi91d, Smi95b, Gut93]. Frances [Scr90]. Francesco [Lon93a]. Francis [Ben92c, Hac98, Hol97b, Kid96, Wal91]. Francisco [Vie90]. Frängsmyr [War92a, Lar91, Lun92]. Frank [Cur91, Hug96a, Kid91, Kry92c, Lon95a, Mer95, Ste95, Tri94, Wer92, vK92a]. Frankel [Lig97a]. Frankford [Rab96]. Frankland [Tra98b]. Franklin [Car98, Hod93, Kry92a, Lai97, Hac90b, Kry90]. Frankreich [Die93, Pet98]. Franz [Web96, Bro99b]. Franz-Josef [Web96]. Fraser [O’M90]. Fraud [Hin94]. Fray [Min95a]. Frazier [Kle91]. Fred [Jam90, Sha95b, Uss90]. Frederic [Don91]. Frederick [Bec91, Fit95, Kra93, Min99, Neu94, Sel91, Jac93b]. Frederik [Fel97]. Free [Ber97a]. Freedom [Hen99, Mat99c]. Freeman [Ang90]. Freese [Wor95]. Freight [Kle95]. French [Ada96, Cha96, Eat90, Eva92, Fro92b, Gut90, Haf92, Liu97, Lyn98, Sha96a, Smi91a, Vol91b, Fie94a, Fri92a, HK99, JJ98, Lev96, Cla91]. Frequency [Hig98]. Friday [Ozd98]. Friedel [Fit97, Pat95]. Friedman [Kie90a, Wer97]. Friedrichshafen [De 99]. Friedricks [Bau93]. Fries [Swe93, Vin93]. Friess [dV95]. Frigates [Dud92]. Fritz [vL99]. Fritzsche [Neu93]. Frommel [Lon95b]. Front [And94, Sha95b]. Frontier [Bas99, Jud90, Qui94, Ste94b]. Frontiers [Fit97, Owe91b, Rei91]. Frost [Hec92]. Frozen [Coo95]. Fruhe [Pet98]. Fruhen [Har99d]. Fruhzeit [Gam99]. Fuel [Maz90a, Str90a, Str91]. Fujimura [Kay93]. Functions [Mue91b]. Furman [Wes91]. Furnace [Rol93b]. Furnaces [Gor96]. Fürst [Rot95]. Fusion [Bow99c]. Future [Fea99, Jud99, Sta90c, Szo96, Yat92, Cut94b, Lup91, McM94b, End97, Sla95]. Futuristic [Neu90]. Fye [SS97]. fysiker [Har99a].

G [Bur92, Cla90, Cou94, Dav96, Daw96, Dou94, Eat90, Ell93, Flo92, Goo95, Hac91a, Hol97b, Hou90, Hug96a, Ing94, Jud90, Kie92, Kra90, LaL90, Les90b, Lip97, Maz90a, Mul96a, Mul90b, Och92a, Pet92b, Por93, Pos98a, Pye92, Rhe94, Rho93, Ryc92, Sha96a, Sha93, Tra95, Tur95, Van97, Vol94b, Was91]. Gabriel [Bol90]. Gaimster [Tra97a]. Galambos [Bla90a, Smi90b]. Galen [Sim91]. Galilei [Lan93]. Galileo [Lan93, Mac96, Har92]. Galison [Wes93b]. Gallery [Jer90b, Bra95]. Galloping [Whi93]. Galveston [Les99]. Game [Mit95, O’B91]. Gamel [Bra91a]. Gangs [Lun90b]. Ganley [Ran97]. Gannon [Hac98]. Gap [Jam90]. Garber [Rot93]. García [Gli91]. Garde [Bra97, Tho94]. Garden [Gol91a]. Gardens [Laz99]. Garner [Buc94]. Garnert [Sun97]. Garrison [BA90]. Gartman [Lai96b]. Gary [Nel90, Scr91b, Tod97]. Gas [Con94b, Hir96, Jac95b, Mea98, Tal94, Wor93a, Myl97, Mea98]. Gaskins [Bra91b]. gassystem [Myl97]. Gatekeeper [Von99]. Gateway [Cod91]. Gauge [Sal92b]. Gay [Moh94, Pre98, May98]. Gayle [Mit91, Pin92]. Gaz [Smi95b]. Gear [Con94a, Vin94, LV95]. Geburtstag [Tod96]. Geddes [Hig91, Hug96a]. Gehennical [Eam97]. Geiger [Sha96a]. Gelernter [Rou96]. Gemery [Low96]. Gender [Fis92, Gam95, Gre91, Gre99, Hes99, Hor97b, LMO97b, Liu97, Moh97a, Moh97b, Per99a, Rot95, Roz98, Sta94, Ste91a, Par97, Old98, Rot95]. Gender-Technology [Moh97a]. Gendered [Her99a]. General [Cor97, Ros94b, Sum93]. Generation [Hat90]. Genesis [Dek98, Mor91a]. Genetic [Bud96, Hut92]. Genetics [Pal93]. Geneva [Pur90]. Genie [Sun97, Lan98b]. Genius [Hew94]. Geoffrey [Con96b, HD90]. Geographical [Rug96]. Geographies [Miy94]. Geography [Buc92, Gui95a, Hug93b, Bry96, Ing91, Bac90]. Geomagnetic [Man99]. geomagnetici [Man99]. Geophysics [Con96b]. George [Cur91, Eam97, Har91, May91, Mis90, Oli98, Por93, Sal92b, Sch94c, Smu90, Ste94c, Wri92, Lie98]. Georges [Lev91b, Smi91d]. Georgia [Cra97]. Georgina [Bra97]. Gerald [Bar92c, Dou95, Ing94, Lai92]. Gerard [Goo95, Mah97, War93]. Gerber [And94]. Gerhard [Bra90, Hal96, LF93]. Germaine [Ste98]. German [Neu93, Nye93b, Pay95, Tho91a, Tho93, Tod97, Tra95, Wal99a, HK99, Kue99, Sto91]. Germany [All98b, Mor93, Wal94, Hey98, Kön96, Neu90, Zei95, Mor91c]. Gerstle [Scr91b]. Gerstner [Bro99b]. Gertjan [Haf97]. Geschichte [LF93, Lin93a, Reu93a, Tod99, Ver97, Web96, Har99d, Blu92]. geschiedenis [Hom94]. Geschlechterverhaltnis [Zac99]. Geschlechterverhaltnisse [Bro99a]. Gesellschaft [Lon93b, Fin98]. gesellschaftlichen [Lin93b]. Get [Sch99a]. Getting [Coo93c, Hil95, Sla98]. Ghost [Ada95b]. Gieck [Rob90a]. Gifted [Tei99]. Gijutsu [Has93]. Gill [Moh97a]. Gillespie [Sok93]. Gilman [Rue90]. Gilson [Buc97]. Gimbel [Mor91c]. Gimpel [Szo96]. Ginning [Cou94]. Giorgio [Lon93a]. Girls [Aik98, Coo92a]. Gisela [Blu92]. Gispen [Tho91a]. Gladys [Ran97]. Glass [Kid91, Her94]. Glenn [Hou97, Maz94]. Glennan [Kra96]. Glick [Hol91]. Glitters [Gor91]. Global [Bar92a, Meu98, Per99b]. Gloeilampenfabrieken [Blo91]. glorreiche [Lin93a]. Glory [Min97, Tra98a]. Goddard [Hil95, Hun95]. Goes [Bow98, Dia99]. Goethals [Sch94c]. Going [Ada95a]. Gold [New91, Was92]. Gold-dredging [New91]. Golden [Sti91, vK92a, Con92c]. Goldman [Coz93a]. Goldschmidt [Wal92]. Golin [Gro90a]. Golinski [Mau93]. Gollin [Dou91a]. Gomery [Ber94b]. Gone [Swe92]. Goodbye [Gre95a]. Goode [Por93]. Goodman [Tri90a]. Goodstein [Low93a]. Göran [Ols97]. Gorbachev [Jos95b]. Gordon [Lan95b, See98a]. Gorgas [Fri96, Col99]. Gothic [Bow99a]. Götlind [O’D94b]. Gottlieb [Rob92b]. Goubert [Bul91]. Gouk [Bla97b]. Government [Dia99, Dou91a, Hei91, Hig91, Maz90a, Nor90, Rob92b, See90, Les90b, Gus91]. Governmental [Mer90]. Grace [Sal92a]. Graeme [Sal95]. Graham [Fie94b, Hev95, Kry92c, Ryc93, Ste95, Tri94, Wer92, Ada95b, Bro95a, Ras92b]. Grammar [HK99, Hoc98]. Grand [Elv99]. Grande [Kah96]. Grande-Bretagne [Kah96]. Grasshopper [Rey98]. Grave [Dub93]. Gray [Gom97, Lea98, Sha90a]. Great [Ano99n, Bro95b, Bro95c, Con92b, Die97, Hol94, Lan99, Les99, Lin97b, McG97a, Sha96b, She94b, Smi95a, Smi97a, Sum91]. Greater [Ham91, McS98]. Greek [Rol92]. Greeks [Sim95a]. Green [Par92]. Greenberg [Kel96]. Greenhill [Gib91]. Greenhouse [Kra98]. Greg [Hou90]. Gregg [Smi93a]. Gregory [Bro97b, Lai92, Min95c, Tho94]. Griechische [Lon93b]. Grieves [Hig91]. Grint [Moh97a]. Griset [Lav93, Smi92a]. Groot [Haf97]. Gross [Jer94b]. Group [Kap92]. Grown [O’B91]. Growth [Cal98, Cla92, Gra91, Hil92b, Jef95, Kie93, Mer98, Rob92b, Wes93b, Nas90]. Guagnini [Kni95]. Guano [Sha96b]. Guaranteed [Wil92a]. Guarantees [Coo94]. Guardians [Bad94]. Guarding [Bad94]. Guidance [DeV92]. Guide [All90, Cer97, Cut95, Jac93a, Lar98, Wes99, Low93b]. Guild [Old97].

Guillaume [Smi90a]. Guillerme [Gei90, Kah96]. Guilmartin [Cap90b]. Gulag [Jos99]. Gulf [New94]. Gun [Gui91]. Gunboats [Dud92]. Gundovald [Hal98]. Gunfounding [Gui93a]. Gunnar [Bjö97]. Gunpowder [HV99, Mau90, Gui98b]. Guns [Bas92b, Wes95]. Gunther [Vie90]. Gurcke [Sta90a]. Gustav [Fin98]. Gusterson [Olw98]. Gustina [Lon93a]. Gutenberg [Ree99]. Gutmann [Haf90]. Gwenyth [Hil97]. Gwin [Gro90b]. Gwynn [Rei93a]. Gwynn-Jones [Rei93a]. Gyroscope [Mac93a].

H [Ang90, Bac90, Bel98a, Bra96, Bra91b, Bri93, Bro93b, Con91, Cut94a, Cut94b, Dav96, Daw91, DeV91, Eva92, Fit96, Ful95, Goo90, Gre91, Hac90b, Hac94b, Hig90, Hil91b, Hol94, Hom94, Hun95, Kay93, Kay91, Kir95, Kle95, LeC94, Led93, Lew90, Lin91, Lon96b, Mal91, Mar91, Nas90, Neu92, Per93, Sha95b, Sim94, Smu90, Sta94, Ste90, Sum96, Til94, Uss90, Vol93, Wor91, vK92a]. Hård [Coo95]. Haber [vL99, Bro93b]. Habitat [Pin92]. Hackbutt [Phi99]. Hacker [Hac90c, Sei96a]. Hackmann [Kid96]. Hafter [Til97]. Hagan [Gar92]. Hagia [Ter94]. Hakfoort [Kip97, vL90]. Hal [Hal94]. Hale [Rot94a, vK92a]. Half [Day99]. halft [Har99a]. Hall [Bla90b]. Hallion [Mor91d, New94]. Hambly [Bed90a]. Hamburg [CS90]. Hamden [Vic92]. Hamilton [Sta97b]. Hamlin [Bro92a]. Hammersley [Har91]. Hampton [Han93]. Hand [Cha95]. Handbook [Sta90a]. Hanford [And94]. Hankins [Bas97, Mor97]. Hannay [Nob97]. Hans [Brü94, Hei95a, Kid90, Rot97]. Hans-Liudger [Brü94]. Harbor [Hil92a]. Hard [Ale99b, LaL90, You93]. Hardison [Vol91a]. Hardware [Fie94a, Ras99]. Harley [Lin94a, The90, Wig96]. Harold [Bix98, Cla92, Hug93b]. Harris [Jan91, Edg98, Fer98, Lay93, Skr93, Sta97a]. Harrisburg [Ing94]. Harry [Mul95]. Harsh [Pfa90]. Hartman [Lup91]. Hartwell [Har98b]. Harvest [Lan99]. Harvesting [And99b, Han99, Ras92a]. Harvey [Bat98, Ser92, Sum96, Tan92]. Hashimoto [Sel97]. Hassenfratz [Bel98a]. Hatcher [Har95a]. Hath [Rol97]. Haudricourt [Lev91b]. Hauptinstrumentes [Dek98]. Haus [Fin98]. Hausman [Old98]. Haussmann [Rei96]. Hawthorne [Sok93]. Haynes [Jac93b, Bas96]. Hayward [Hig92]. Hazards [Sha94, Aik98]. Hazen [Sim94, Sim94]. Headrick [Fin93]. Health [Aik98, Bri93, Bul91, Dav93, McE95, SS99, Sim90, Ham92a]. Heart [Jef95]. Hearth [Was91]. Heartland [O’D96a]. Heat [Kay91, Lay92b, Pac95, Die97]. Heather [Jos95a, Whi92c]. Heating [Har94]. Heaven [Cor95]. Heaviside [Hon97]. Hebron [Win90]. Hechstetter [Har91]. Hecla [FO92]. Hedin [Gie96]. Hedy [Bry92]. Heerding [Blo91]. Hegmon [Sto97]. Heidegger [Gol91a]. Heijmans [Dav96]. Heilbron [Kri92, War92a]. Heims [Kap92]. Hein [LaF91]. Heiner [Jac91]. Heinrich [Kli97a, Tra98a]. Heinz [Nye95]. Heirs [Kra90]. Heisenberg [Wal99a]. Helden [Mor97]. Helen [Lan90b]. Hellemans [Mag94]. Heller [HV98]. Hellions [Zac98]. Hello [Fis92]. Helmuth [Bra90]. Henisch [Nye95, Nye95]. Henri [Lyn98, SB95, Smi90a].

Henri-Jean [Lyn98, SB95]. Henry [Fer95, Gor92b, Har96b, HV98, Isr94a, Joh97a, Lan94c, Lew97, Low96, Nye93b, Oli90, Sch94b, Tri93, Bra98, Dou99, Isr94b, Jac91, Zem92, Bau93, Jac91, O’D97]. Henrys [Mor99]. Herbert [Fri91b, Pur90]. Here [LMO97a]. Heritage [Bla92b, O’D94a]. Herken [Smi93a]. Herlihy [Fle92]. Herman [Por94]. Heroes [Bat90]. Herrick [Fro92b]. Herrigel [Tod97]. Herring [Coo91d]. Herrschaft [Brü94]. Herty [Ste98]. Hertz [Kli97a]. Hesse [End97]. Hevly [Wes93b]. Hewlett [Hac91a]. Heyden [Win98]. Heymann [Ver97]. Hezinger [Lin93a]. Hickman [Gol91a]. Hidage [BA90]. Hidden [Rob92b]. Hideo [Mul95]. Higgins [Moy96, Oli90]. High [Arn97, Bes95, Cha99a, Con92c, Han99, Hig98, Lev97b, Nor90, Szo96, Gui95b]. High-End [Con92c]. High-Frequency [Hig98]. High-Tech [Szo96]. High-Technology [Nor90]. High-Temperature [Cha99a]. High-Vacuum [Arn97]. High-Voltage [Lev97b]. Higher [Abb96]. Highway [Rot99a]. Highways [Cor95]. Hilde [LaF91]. Hill [And99a, Ble91, Gam92]. Hills [Cro99, Bow92, Gen92, Mue95]. Hilton [Sal92b]. Himmel [Web96]. Hindle [Dup92]. hip [Kal95]. Hippel [Isr90]. Hiroaki [Dal90]. Hiroshima [Pur90, Bat90, Her99b]. Hirsh [Chr91]. histoire [Coh90, Lav93, Coh90, Hal93a, Smi92a]. Histones [Fer95]. historia [Car99b]. Historian [Van96]. Historians [Lai98]. Historic [You93]. Historical [Asp93, Bat98, Bra91a, Bro97a, Buc92, Dav93, DeV92, Fel99, Gie96, Goo95, Gor91, Hil91b, Hoo99, Ing91, IR91, Ker99, Lip97, Mal91, Mar93a, O’M93, Olw96, Ree98, Rom97b, Sta90a, Thr91, Wil99a, Wer97]. historie [Hei99]. Histories [All98a, Lau95, ST98]. Historiographical [Lon96b, Rol93a]. Historiography [LMO97a, Mac92, Sin99a]. Historische [Lun90a]. History [Adr94, Ale99b, Ano90d, Ano91f, Ano92e, Ano93e, Ano95e, Ano96e, Ano97e, AC95, Bac90, Bal97, Bar90, Bed90b, Ber94b, Ber95a, Ber95b, Big95a, Bin91, Blo91, Bor97, Bow92, Bra98, Bre99, Bri90, Buc91, Cal98, Cam92, Cam93, Cer97, Cod92, Coh90, Con90, Con92a, Con95, Coo93b, Coo98d, Cow96, Cro94, Cut97, Cut99, Dav92b, Dor91, Dun93, Edm96, Eis93, Eva96, Eze90, Far96, Fel99, Fin95, Fit91, Fit97, Fle98, Ger96, Gol99, Goo90, Gor92b, Gra97b, Gra98a, Gro90b, Gui98b, Hac92a, Hac94c, Hal98, Ham92a, Har95a, Hat90, Hil92a, Hon98, Hor97a, Hyd99, Jac95a, Jar90, Joh99a, Jos95b, Kar99, Kay94, Kri92, Kun95, Kyn99, Lai96b, Lau99, Law91, Lay91, Lea99, LMO97b, LM91, Lev96, Lew96, Lin94a, Lon91a, Lon93a, Low93a, LAC^{+}90, LAC^{+}91, Low91, LAC^{+}92]. History [Low92, LAC^{+}93, LAC^{+}94, LCH^{+}95, LCH^{+}96b, LCH^{+}97, Low99a, Low99b, Mac93a, Mag94, Mah97, Mar91, Mau91, May98, Maz90b, McG97a, MM92, Mea98, Mer98, Mis97, Mor91d, Mue95, Mul96a, Nas90, Nye98b, Pet90, Pre98, Pur97, Qui95, Reu95, Rey94, Rol90, Rol97, Ron99, Rot93, Roz96, Rug96, Rum92, SB95, Scr91c, Scr95, Scr98b, See95b, See95c, See95a, Sei98, Sha90a, SS97, Sha95b, Sha94, She93, Sim95b, Sla92, Sla98, Smi95a, Smu94, Sok93, Ste92, Ste99, Str90b, Sul98b, Sum96, Svi93, Swe92, Swe93, Tal94, Tan92, The90, Tho97a, Vol91c, Vol94a, Vol94b, Was91, Wen98, Wig96, Wil99b, Wor98, Zel97, All99b, Car99b, Hei99, Kay99, Ale99b, Ano94e, Coa99, Hor92, Str90b, Ten92, Bas92a, Ben91a, Bro95a, Dou91b, Flo92, Fri92a, Gen92, Led93, McW95, Nel90]. History [Rhe98, Sma95, Ste90, Wil95b, Wor94, Yat96, Kay91, Kel99, Kir95, Kry92c, Lay92b, Lig97a, McG97b, Skr91, Ste95, Tri94, Wer92, Wil95c]. Hitchhiker [Cut95]. Hitler [Tho93]. Hochschule [Bra90]. Hoddeson [See95b]. Hodge [Lan94a]. Hoerr [Mar90]. Hoertz [Tas97]. Hoffman [Leo96]. Hogan [Jos95a]. Hoke [O’M93, Zem93]. Holding [Fis98]. Holes [Har95b]. Holifield [Hac98]. Holl [Hac91a]. Holland [Mis93]. Hollister [Kry92c, Ste95, Tri94, Wer92]. Hollister-Short [Kry92c, Ste95, Tri94, Wer92]. Holly [Wor93b]. Hollywood [Kra94a, Gom96]. Holmes [Ste98, Don91]. Holocaust [Ale99a]. Hölsken [Neu96b]. Holt [Mue90]. Holtzmann [Her98]. Holy [Lig97a]. Holzman [Mah97]. Homburg [Wac94]. Home [And94, Bra96, Day92, Gol99, Gol97, Mat99c, Nic99b, Pye92, Rot95, Was91, Thr91]. Homer [Cor97, Bla95]. Homes [Ber93]. Homestead [Reu93b]. Homework [Thr91]. Hongkong [Mor91b]. Honor [Fit95, Har98b, Kay91, Mar91, Rot93]. Hook [Tra97a]. Hoopes [Vas98]. Hoover [Smi95a]. Hope [O’D96b]. Horizon [Buc92]. Horizons [Day99]. Horror [Ald93]. Horse [Kin99, Lon97, Sum93, McS98]. Horse-Powered [Kin99]. Horses [Kin99]. Hoses [Win98]. Hosler [Chi97]. Hospital [Gam97]. Hot [Pos98b, Vol93]. Hounshell [Bir90]. House [Lee96, Lun92, Mor90, Whi93]. Household [Kea98]. Housewares [Nic99b]. Housewives [Hoy92]. Howard [Bla97a, Dou91b, End97, Hof92, Ler98, Log92, Lup91, Sla95]. Howe [Jer97]. Howell [Gam97, Sha95a]. Howse [Gin90]. Hoy [Mel96]. Hubble [Spi95]. Hubris [Hey98, McC95]. Hudson [Bra93, Rug96]. Hufbauer [Har92]. Hugh [Car98, Koh96, Olw98, Dou94]. Hughes [All98a, Ben92a, Ben92a, Mor91a, Sim95a]. Hugill [Gui95a]. Hull [Pry93]. Hult [O’D94a]. humaine [Lev91b]. Human [Fri98, Pfa91, Rey94]. Humanists [Lon95a]. Humans [Kat94b]. Humble [Oli98]. Hundley [She94b]. Hungary [Hor97a]. Hungry [Bel97]. Huning [Lue90]. Hunt [Bro93a]. Hunter [Coo93b, Smi91a, Fer90]. Hunting [Tra98a]. Huntington [Bau93]. Hurt [Fin95, Fin95]. Husbandry [Fle90]. Hüttenwerk [All96a]. Huxley [O’H98]. Hybrid [Fit93a, Sha91]. Hydraulic [Gui98a, Lig97b, Mil91a]. Hydraulics [Pfa90]. hydraulique [But92]. Hydro [Eva93a]. Hydrographic [Bur92]. hydropower [Kai98]. hygiene [Rei99b]. Hyldtoft [Hår98a, Myl97]. Hypertext [McM94a].

Ian [Low92, Pur97]. Ibironke [Kri97]. IBM [Edw96a, Kid94]. ICOHTEC [Ber95b, CS90, Hau90]. Idaho [Cor90]. Idea [Cru99, McG97b, Old98]. Ideal [Coo93a, Lon91b]. Ideas [Dav93, Edg98, Hea91b, Gor94]. Identity [Ler98, Pye92, Tap99]. Ideological [All96d]. Ideologies [Fer99, Hea90, Moo98]. Ideology [Chr95b, Kli97b, Sch94a]. Ido [Hon97]. IEEE [Ano90d, Ano91f, Ano92e, Ano93e, Ano94e, Ano95e, Ano96e, Ano97e]. Ignition [Hoo99]. Igor [Hal90]. Ihde [Gol91a, Mit92]. ihre [Lin93b]. II [Bar92c, Cha90, Har96b, Hou97, Lue90, Cap93, Kue99, Nah97, Rus99, Ber95a, Coz93b, Gor92a, Hac94b, Hig97b, Moy96, Zac98]. III [Ste94c, Sum96]. Illinois [Sha91]. Illusion [Neu96a]. Illusions [Kid90]. Illustrated [Cro94, Kid91, Tas97, Wor94]. Im [Fin98, All96a, All96c, Bra90, Die93, Hal96, Web96]. Image [Koc93, Nee96, Wos92, Cre99]. Images [Cod92, Joe90, Nye95, Pan95, Tob92]. Imagination [Pis96, Tho94, Bas97, Hac90b, Neu93, Reu95, Sla91]. Imagine [Hor99]. Imaging [Her98]. Immigrants [Haf99]. Impact [Dou91a, Kie92, Lau99, May94, Oga90]. Imperatives [Sch93b]. Imperfect [Sla95]. Imperial [Gra97b, Per99a]. Imperialism [HC98]. Important [Mag94]. Impractical [Tra98a]. Improved [Lai91]. Impurity [Bro92a]. in-dustri [Hår98a]. Inaugural [Por97]. Inca [Pet96]. Incorporated [Smi93b]. Incremental [Log92]. Incubator [Koc98]. Independence [Str91]. Independent [Egg94]. Indeterminacy [Scr95]. Index [Ano90e, Ano91b, Ano91g, Ano91j, Ano91k, Ano92b, Ano92f, Ano92k, Ano93b, Ano93f, Ano93j, Ano94b, Ano94f, Ano94k, Ano95b, Ano95f, Ano95k, Ano97b, Ano97f, Ano97k, Ano98d, Ano99a, Ano99b, Ano99l, Ano99o, Ano99p, LCH^{+}96a, LCH^{+}96c]. India [Par92, Set99, Ada98]. Indian [Hea92, Kal98, Sam99]. Indiana [Wal99b]. Indians [Was93]. Indica [Sam99]. Individualism [Ree98]. Indonesia [Moo98]. Indoors [Die97]. industri [Hår98a]. Industrial [Aik98, All97, Ben91c, Ben96, Bla92b, Bra92, Bra93, Buc94, Buc98a, Bul91, Car98, Con96b, Coo93b, Coo96, Dal90, Dou95, Gei94, Gib91, Gib93, Gib99, Gor96, Gra95, Hac93, Haf96, Har96c, Hoc98, Jac93a, Jar90, Jer94b, Jer94a, Kas99, Kni95, Lea98, Len91, Leu98, Lie92, Low97, Mal91, Mal98, McB92, McE95, Mis93, Moh97b, O’D96a, Ols97, Rab98, Rob90b, Rob94a, Ryc93, Scr91a, Scr98a, Smi90c, Tal94, Tod97, Tri90a, Twe92, Web90a, Car99b, Fer92, O’D96b, Lip98, Tod97]. Industrialisation [Mic94, Hår93a, Dav92a]. Industrialisering [Kai98]. Industrialisierung [Tod96]. Industrialising [Nye93a]. Industrialist [Dan93]. Industrialists [Car98, Pri98, WN98]. Industrialization [Ber95c, Bro93b, Coo90, Gli94, Gra98b, JJ98, Kai98, Lan95b, Mat99b, MS92, Nel90, Opi90, Smi93b, Voe90]. Industrializes [Ing94]. Industrializing [Lai92, Bro96]. industrie [Dav96, Hom94, Ros98, Hal93a]. Industriegesellschaft [Lun90a]. Industriële [vL90, Wac94]. Industrielle [Web90c, All96a, Die93]. Industrieller [All96c]. industrielles [Kah96]. industriellt [Car99b]. Industries [Ben91c, Gar90, yH98, Kra94a, Liu97, Min95c, Pay95, See95a]. industriforskare [Har99a]. Industry [Ald96, Ben91c, Bla90b, Bla96b, Cla90, Coo94, Dav92b, Don94, Edw96a, Eva96, Fri92a, Gre91, Haf90, Haf99, Har95a, Hei91, Hil99, Hor97b, HM96, Jar90, Jer93, Jer94a, Kat94a, Kid94, Kli90, Kön96, Lan95b, Lay93, Lea92, Lie90, Lin98, Mea98, Meu98, Mil94, Moe94, Mor99, New97, O’B91, O’D94a, Qui93, Roc94, Ryd98, Scr90, Ser94, Set99, Smi91b, Tan93, TJ96, Tra90, Tra95, Twe92, Voe90, Wes95, Wor93a, Ahv90, Ham91, Mol91, Wri96, Bin91, Bla92a, Blo90, Chr91, Cor97, Fro92b, Gro90b, Hig92, Jam90, Mar90, Sch92a, Tri93, Bro93b, Cas98a, Coo98c, Rol93b, Sch99b, Vol98a]. Industry-Based [Kön96]. Inequality [Liu97]. Infantry [Bac98]. Infertility [Lay97]. InfoCulture [Chr95a]. Information [Cap91, Cer98, Chr95a, Con96b, Gom96, Hei91, Hil91a, Joh91, Kal95, Kid90, Lip97, Lue90, Pfa91, Sch95, Sei98, Shi97, Ten92, Yat92, Lav93, Asp93, Fri92b, Coo98d, Rei95b]. infrastructures [Rei99b]. ingénieur [Lan95a]. Ingenieurausbildung [Bra90]. Ingenieure [Brü94]. Ingénieurs [Lan90a]. Ingenious [Zem93]. ingenjörskonst [Bjö97]. Ingenuity [Wal99b]. Ingham [Egg94]. Inglis [Mau91]. Initiative [Fri91b]. Injury [Bra91b]. Inland [Hof91]. innern [Bro99b]. Innovation [Ald97, Car99a, Cha99b, Cla93, Cut94a, HD90, Hol99, Hol93, Jam90, Kle99, Lin97b, Lip98, MKR93, Mul90a, Sha95c, Uss91, Vas98, Wil92b, Ros98, Isr94b, Isr90, Isr91, Hei91, Isr92a, Ros94b]. Innovations [CS90, GS98, Mar93a]. Innovators [Isr94b]. Inquiry [Ing99, Lon96b]. Insider [Jac91]. Instant [Vie90]. Institute [Low93a, Col90, Ser94]. Institution [Gra97b, Mil95, Roz96]. Institutional [Mea98]. Institutionalization [Bra97, Lun92]. Institutions [Fea99, Hac93, Hac94c, Hea91b, Maz93]. Instrument [Ben91c, Kid91, War91, War92b]. Instrument-Making [War92b]. Instrumental [Mit92]. Instrumentation [Hon98]. Instruments [Bed90a, Bed91, Eva93b, Eva96, Goo95, Hea91b, Kid91, Man99, Maz93, Ste96, War98b, Dic92, Bed90b, Mor97, Bas97]. Integrating [Str90b]. Intellectual [Lon91a, Ben92a]. Intelligence [Kid90, Fri98]. Intensive [Koc98, MKR93]. Inter [Coo92b]. Inter-war [Coo92b]. Interchangeable [Ald97]. Interconnection [Sin99c]. Interdisciplinary [Fle98]. Interest [Lew96]. Interface [Mit92]. Internal [Mue91b]. International [Ber95b, Fin95, Fin93, Gui98b, Lew93, Mai96, Mei98a, Nel90, Ols97, Ste99, Wan96, Hal93a, Hor97a, LM91, Pye92]. Interpretation [Coo95, Jud90]. Interpretations [Mah90, Miy94]. Interpreting [Cod92]. Intersection [ST98]. Interstate [Ald93]. Interwar [Hol99, McB97]. Introduccion [PO99]. Introduction [Eva93b, Fer92, Hac91c, Lav93, Mil95]. Introspection [Coo93a]. Introspective [Pfa97]. inventeurs [But92]. Inventing [Rei91, DeV92, Rob90b]. Invention [And99a, Bal98, Bas92b, Bri92, Coo91c, Daw96, Fri91a, Gou98, Gra95, Gra97b, HP91, Kid96, Kra94b, Lon91a, Mac91, Mac93b, McG95, Mor91a, Mul90a, Owe91b, Pot99, Ree99, Sch99a, Tym99, Wes93a, Cor91, Gor94]. Inventions [Coo91b, Doo91, Chr95a]. Inventiveness [CC95]. Inventor [Kli91, Mai96, Mar98]. Inventors [McG97a, Tho91b, Uss91]. investigation [Car99b]. Investigative [Don91]. Invisible [Bal98, Fin93, Maz93]. Ionic [Kra98]. Iosif [Bra91a]. Iowa [Hil91b]. Iraq [New94]. IRCAM [Bra97]. Ireland [War91]. Iron [Egg94, GK93, Jar95, Kal98, Kri99, Ryd98, Scr96, See95a, See98a, Sum93, Lew97, Mul94, Roz98]. Ironwork [Rey90a]. Irrigation [Gui98a]. Irwin [Sol98, Ber98]. Isern [Ras92a]. Isfahan [Ozd98]. Islamic [Lin94a]. Island [Gor97b, Les90a, Twe92]. Islands [Ell92a]. Isles [BP91]. Israel [Mer95, Gra95]. Issue [Ano98b, Ano98c, Ano99h, Ano99i, Ano99j, Ano99k, Cow93, Qui95, Bro97a]. Issues [Coo98d, Gor92a, Kid90, Gie96, Ham98b, Hil91a, Til94]. Italia [Man99]. Italian [Kin93, Car98]. Italy [Man99, Ste96]. Ithiel [Bar92a]. Ivan [Ste92].

J [Ang90, Bad91, Bas97, Bel98a, Ben92b, Ben92c, Ben96, Big91, Big95a, Bri92, Bro93a, Bro98b, Cla90, Con94b, Coo97, Dan93, Dou94, Dun93, Edg98, Edw96b, Fit93b, Fri92a, Gar92, Gar90, Gli94, Gui95a, Hac94a, Hei95a, Hil92a, Hur90, Hyd90, Jac91, Jac95a, Jam90, Kap92, Kid96, Kli90, Kon97, Kri92, Lay96, Lea92, Len91, Lin94a, Log95, Lon95a, Mah97, Mei90, Mil91b, Min97, Nas91, O’M90, Pet97, Pfa92, Por94, Rab96, Rei93b, Rey91, Rob94b, San92, See98b, Sim95a, Sim95b, Smi95a, Ste95, Swe93, The90, Tri94, War91, War92b, War92a, Was91, Web90a, Wig96, Wos94]. J. [Mau90, Kas99, Wil92b]. J.-L. [Mau90]. Jack [Hac91a, Rob90a]. Jackaway [Hil97]. Jackson [Bla97c, Moy96, Sum96, Nye98a, Rom97a]. Jacob [Bil97, Frä91, Tri93]. Jacobi [Kry92b]. Jacobson [Hea96]. Jacquelyn [Bla90b]. Jacques [vR90]. Jahre [dV95]. Jahren [All96a]. Jahrhundert [All96c, Har99d, Die93, Hal96, Lin93b]. Jakab [Cor91]. Jakle [Jac95b]. James [Ada93, Bro95c, Bry96, Buc92, Cer97, Cut93, Eis93, Ell92b, Joh92, Joh97a, Kid94, Lea92, Led93, Low96, Lun90b, McM94a, Ors97, Pry93, Rab96, Ros94a, Sei96b, Ser92, Smi94, Tho98, Wil93b, Zel97, Kry92c, Lon95a, Sla99, Ste95, Tri94, Wer92, Dan93, Lan90b]. Jan [Mau93, O’D94a, Sun97, Whi92a, Win98]. Jane [Har94, Min96]. Janet [Gom96]. January [Isr94a, Ano90j, Ano91m, Ano92m, Ano93n, Ano94m, Ano95m, Ano96j, Ano97m, Ano98f, Ano99q]. Japan [Coo90, Coo98b, Bac99, Col90, FF96, All97, Bla97b, Coo92a, Far96, Mul95, Rot99c]. Japanese [Jer97, FF96, Mat99b, MS92]. Jarrell [Pry93]. Jean [All96b, Bin93, Bul91, Gol92, Gui91, Lyn98, Pet96, SB95, Smi95b, Szo96, JJ98]. Jean-Antoine [JJ98]. Jean-Claude [All96b, Bin93]. Jean-Pierre [Bul91, Pet96, Smi95b]. Jed [Kli97a]. Jefferson [Dan92, Har99b]. Jeffrey [Bla96b, Ham95, Har96b, Koc98, Lew96, Oga90, Qui95, Rhe98, Sum93, Cha90]. Jenkins [Kli91, Sha95b]. Jennings [Whi92a]. Jeremy [Lea92]. Jerold [Sch91]. Jerry [Moy96]. Jersey [Nas90]. Jervis [Cal92a]. Jessica [Bra96]. Jewish [Ben92b]. Jim [Mah90]. Jimmy [Sha96b]. J’Nell [Yea90]. Jo [Kri98]. Joan [Ano90f, Ano91h, Ano92g, Ano93g, Ano94g, Ano95g, Ano96f, Ano97g, Ham98a, Kay93, Les93]. JoAnne [Lub91a, Mul95]. Job [Aik98, Kay93, Bix99]. Joel [Bol90, Fit93b, Gam97, Har96c, Hea91a, Qui95, Rom97b, Sha95a]. Johan [Fer92]. Johann [Ree99]. John [Arm96, Bac90, Bir90, Bri93, Buc94, Cap90b, Chr95b, DeV93, Dol97, Egg94, FO92, Gib93, Gie94, Gor98, Ham92a, Har95a, Hea91b, Hil91a, Hol94, Jac95b, Jer93, Joh99a, Kal95, Kay91, Kle95, Lay93, Lee96, Lie90, Lie93, Lin98, Mar90, Mar93a, Maz90a, McN90, McW95, Min97, Mor91c, Nye98a, Rei93b, Reu91, Rhe94, Rob93, Rol95a, Rom97a, Rug96, Ryc92, Sch93a, Sel90, Sha94, Sla92, Smi92b, Spe95, Sut90, Vas98, Was92, Win90, Wor93a, Col99, Fer98, Gol91a, Han99, Kas99, Lai97, Mul90b, Pol90, Cal92a, Kid92]. Johnson [Wor90]. Johnston [Les90b, Sum93, Far96]. Jon [Bat98, O’D96a, Owe91a, Per90]. Jonathan [Kat92, Pet92b, Wei94]. Jones [Web90a, Jar95, Rei93a]. Joppke [Wal94]. Jordan [Jud90, Rei96, Chr95b]. José [D’A97]. Josef [Web96]. Joseph [Bad91, Bla90a, Cla90, FG98, Huf92, Isr94a, Kle91, Lev91c, Mul96b, Sum93, Wes95, Dou99, Gam99, Mul96c]. Josiah [Fri96, Col99]. Journal [Qui95, Ben91a]. Journalistic [Kie90b]. Journals [Fri96]. journée [Lan98b]. Journey [Fie98, Gru90]. Jowitt [Gar90]. Joyce [Dan90, Sha95c, Sla99]. Jr [Cap90b, Kay91]. Jr. [Bir90, Bry95, Buc92, FG98, Hoc98, Hug96a, Kle95, Lew97, Mal91, Min97, Ryc93, Scr91a, Smi91a, Smi92b, Vol91a, Win98]. Judgment [Fer95]. Judith [Dav92b, Lev91c, Low93a, Pas96]. Judt [All98b]. Julie [Pan95, Ube94]. Julius [Dav96]. July [LM91, Ano90l, Ano91p, Ano92o, Ano93p, Ano94o, Ano95p, Ano96l, Ano97o, Ano98h, Ano99s, Ano99u, Ber95b]. June [Har93b, Kli91]. Jürgen [All96c]. Justinian [Ter94]. Justitia [Lea92]. Justus [Von99].

Kahn [Tho94, McM94a]. Kaijser [Gie96]. Kaiser [Jac91, Dia99]. Kamminga [Fer92]. Kane [Scr90]. Kansas [Hir96, Wil92b]. Kaplan [Øve97]. Karal [Bra96]. Karaoke [Per99b]. Karen [Cou94, Smi97a, Sta92b]. Karl [Har92, Lar98, Sta90a, Zak92]. Karlsson [Rey90a]. karriar [Har99a]. Kathleen [Ebn92, Fle90]. Katsh [Phi91]. Kauffman [Mul95]. Kaups [Jud90]. Kay [Bud94]. Keating [Hof92]. Keepers [Sim94]. Keeping [Rum92]. Kees [Tho91a]. Keith [Cor90, Eva93a, Hig91, Hig92, Jac95b, Kra96, Moh97a, Moy94, San90, Sta96a]. Keizô [Sel97]. Kelham [Twe92]. Keller [Lin97a]. Kelley [Hof91]. Kelly [Bac98, Gor92c, Mul95, Par94, Ber94a, Ros92]. Kelvin [Por91]. Kemp [Ada91]. Kenkyu [Has93]. Kenly [Bir90]. Kennedy [Eat90]. Kenneth [Gar92, Hig90, McM94b, Mue91a, Rob96a, San97, See98b]. Kennett [Omi92]. Kenyon [DeV93]. Kernforschung [Rüd92]. Kertz [Gis98]. Kevin [Cap91, Rot92]. Kevles [Her98]. Khayyam [Ozd98]. Kicherer [Pul92]. Kidwell [Sma95]. Kikaiseizu [Has93]. Kiley [Dav99]. Kill [Cro99]. Kimoto [Lin93b]. King [Cas98b, Ste94c, Gar91, Gol91b, Cor97]. Kingdom [Win90, Hay96]. Kingery [How98, Wil95c]. Kip [Har95b]. Kirkham [Kat97a]. Kitt [Fie94b]. Klaus [Die93]. Klein [Gor95b]. Kline [Chr93]. Kluger [Smi94]. Knights [Gol91b]. Knot [Gut95b]. Know [Flo92, Old92]. Know-How [Old92]. Knowing [Pin90, Edw97]. Knowledge [Ada98, Con99, Daw96, Fit97, Hår96a, Hea99b, Ler97, Lon91b, Ree98, Scr00, Sok93, Sta97b, Rei91, Nob97, Dan95, Joh91, Pan98a]. Known [Bed90c]. Knut [Qui98]. Köhle [Lin93a]. Kohler [Kay92]. Kohlstedt [Por93, Pye92]. Königreich [Bra90]. Konzepte [Buc97]. Koos [Hou90]. Korean [Hon98]. Kosaten [Has93]. Kotz [Hol90]. Kowaleski [Cla97a]. Kraft [Tho98]. Kramer [Hal96]. Kranzberg [Isr92a, Rol97, SMP^{+}96, Van96, Mar91]. Krarup [Kra94b]. Krause [Moy94, Reu93b]. Krematorien [All96b]. Krige [Dol97]. Krimsky [Ham98b]. kring [Kai98]. Krisciunas [Rot92]. Kriss [Han99]. Kristine [Hår93a, Voe90]. Kroes [Rob94a]. Kronick [Low93b]. Kroyer [WN98]. Kuchen [Lin93a]. Kuhn [Ful98, tZS90]. Kumar [Pfa98]. Kunst [dV95]. Kuromiya [Dal90].

L [All97, Bas97, Bla90b, Bow92, Cla90, Cla92, Col93, Con96a, Cor91, Coz93b, Coz93a, Cut93, Dav96, Dud92, Gen92, Gom97, Gor91, Hac90a, Hec92, Her93, Hil97, Hol97a, Kel96, Kri92, Lai93, Lev91a, Lon95b, Lon95a, Lup91, Mag95, Mic94, Mit91, Moh90, Mor93, Mue95, Old98, Pet97, Pin92, Rhe94, Rhe98, Rol90, Ryc93, Sch93a, Seg94, Smi91a, Ste95, Tri94, War92a, Wri94, Yea90]. L. [Mau90]. Labor [Bla90b, Day92, Gom97, Gro93, Hor97b, Kas99, Lie92, Nuw90, Sal92a, Scr91b, See90, Thr91, Whi98, Smi91b, Cam93]. Laboratoria [vL90]. Laboratories [Cla93, Wes91]. Laboratory [Gra95, Har93b, Hol93, How93, Kri92, Olw98, Sei96b, vk92b, Kri92, LaF91]. Labors [Rei96]. Laborsaving [Kli97b]. Labour [Ben91b, Bla92a, Cha98, Gar90, Hil99, Mic94, Pfa98, Roz90, HV98]. Lady [Dup92, Pur93a]. LaFollette [Hin94, Tob92]. Laithwaite [Bri90]. Lake [Dub93, Mou99]. Lakoff [Fri91b]. Lamandé [Lan91b]. Lambright [Joh97a]. Lamiral [Smi91d]. Lamoreaux [Lip97]. Lamp [Sun97, War98a]. lampan [Sun97]. Lancashire [Jer93]. Lancaster [Bro92b]. Lance [Kel96, Pur97, Ste92]. Land [Jud99, Mul96a, Sch91, Lig97a]. Landau [Dre97]. Landing [Con94a, LV95, Vin94]. Landmarks [War97, Mai98, Sma95]. Lands [Hug99, Spe95]. Landsat [Bar92b]. Landscape [Ber98, Lev97b, Mou99, Pfa92, You93, Sha95b]. Landscapes [Car96, LeC99]. Language [Bel91, Kid96, MB92, Tra97b, Mal91]. Languages [Her94]. Lanham [Sin96]. Lanka [Pfa90]. Lankton [Dub93]. Lanouette [Hew94]. Lansing [Pfa92]. Large [Jos95b, Min96, Pin90, San94, Scr98a, Wes93b]. Large-Scale [Jos95b, San94, Wes93b]. Larkin [Cal92a]. Larry [Bro98b, DeV91, Dub93, Gol91a, Hig90]. Lars [Cut97]. Lary [Miy94]. Laser [Les93, Mac93a]. Lasers [All90]. Lash [Sum93]. Lassell [War98b]. Last [Sin99c]. Late [Gui93a, Nye93a, Whi92c, Per99a]. Later [Lai93]. Latour [Rei91]. Latrobe [Har96b]. Laughlin [Bla95]. Laughlins [Jar95]. Launch [Nea98]. Laundrymen [Moh97b]. Launius [Bro97a]. Laura [Liu97]. Laurence [Jer94b]. L’Automobile [Lau98a]. Lavoisier [Rob95, Tra97b]. Law [Gor98, Haf92, Lub91b, Rob94b, Gie94, Kie92, Phi91, Kon98]. Lawal [Kri97]. Lawn [Sha95b]. Lawrence [Don91, Kri92, Les90a, Kri92]. Lawyers [Oli98]. Lay [Vol94b]. Layton [Kay91]. Lazonick [Lam93]. Leadership [Day99]. Learned [Sal95, Tho93]. Learning [vHT96, How98, Kid96, Wil95a]. Leary [Sti91]. Lebenslauf [Lin93a]. L’École [Lan98b]. Lee [Jam90, Lai92, LeC94, Liu97, Min95c, Omi92, Wor91, Edw96b, Lev91c]. Legacy [And94, Cow93, Gor98, Yea90, Har95b, Gui98a]. Legal [Har99b]. Legislator [Hac98]. Lehigh [Rey94]. Leiden [War98a, War99]. Leiss [Car91]. lektorer [Har99a]. LeMahieu [Moh90]. Lemonnier [Ell94a]. Lenard [Tal94]. Lenin [Jos95b]. Lennart [Rey90a]. Lenoir [War92b]. Lens [Eno98]. Leo [Gut92, Hew94]. Leon [Skr91]. Leonard [Cod91, War97]. Leonardo [Kel99, Ano90g, Ano91i, Ano92h, Ano93h, Ano94h, Ano95h, Ano96g, Ano97h, Lon97]. Leonhard [Ste94a]. Leroy [And99a]. Leslie [Gei94]. Less [Par97]. Lessoff [Rei96]. Lessons [Cla90]. Lester [Kie92, Pin91]. L’étonnant [Bel98a]. Letter [Cap90b]. Letters [Har91, Paq97]. Levant [Lig97b]. Lever [Hea91a]. Levers [All90]. Levidow [Cap91]. Levine [Log95]. Levinson [Ano90i, Ano91l, Ano92j, Ano93i, Ano94j, Ano95j, Ano96h, Ano97j, Pin90]. Levittown [Ber94a]. Lewchuk [Blo90]. Lewis [Ada96, Hol93, Sut90, Bla91, Gor96, Smu93, Web90a, Ben92a, Bia97, Hug90, Hug96a]. Liberal [Fie98]. Liberalism [Car98, Chr95b, Kel93c]. Liberalization [Bau99]. Liberty [Rue90]. Librarians [Kal95]. Libraries [Ber97a, End97]. Library [War97, Kuc94]. Licht [Bro96, Coo93c]. Lichtenstein [Pet90]. Lick [Lan90b, Rot92]. Lieberman [Tho97b]. Liebermann [Kra93]. Liebig [Von99]. Lienhard [Kay91]. Life [And99a, Ano90d, Ano91f, Ano92e, Ano93e, Ano94e, Ano95e, Ano96e, Ano97e, Bud94, Cas98b, Coo91a, Dav99, Dub93, Fos98, Gol91c, Gol91b, Har91, Hei95b, Jac96, Kay93, Kay94, Ler97, Lie91, Lie92, Lin97a, Lüt98, Mat92a, Mou99, Pot99, Rei96, RK97, Sam99, Sum96, Vol98b, Alb97, Cap95, Dav92a, Fis98, Gol92, Gor95b, Hug90, May92, Mei93, Bix98]. Lifeworld [Gol91a, Pin91]. Lifting [Mei99b]. Light [BA90, Gol97, Kip97, Opi90, Pre98, Sun97, Tri91]. Lightfoot [Wor95]. Lighting [Wor93a]. Lightning [Kry90, KK92, Sel99]. Ligonnière [Coh90]. Like [Chr99, Hyd90, Bla90b, Bix98]. Lilienthal [Fie98]. Lillian [See95b]. Lily [Bud94]. Limb [Ben94]. Lime [See95a]. Limit [Pet92a, Rot94b]. Limited [Maz90b]. Limits [Cha95, Moh95, Wil90, Wil93a]. Limon [Sla92]. lin [Sel92]. Linda [Ell93]. Lindberg [Kel93b]. Lindbergh [Rei95a]. Lindley [Kie92]. Lindroth [Frä91]. Line [Fis98, Pet90, Smi97b]. Linear [Bri90]. l’informatique [Coh90]. Ling [Big91]. Linkages [Tho91b]. Links [Les90b, Hig97a]. L’Invention [Lan95a]. Lipartito [Mue91a]. Lisa [Les93, Roz98]. Litai [Sut90]. Literacy [Hoc98]. Literary [Tas97]. Literature [Bow98, RK97, Sta92a, Bas96, Cro94, Lev91c, Kel96]. Little [Hal98]. Liu [Gut95b]. Liudger [Brü94]. Liverpool [Bla92b, Jer90b]. Lives [Big96, Coo96, Eam97, Hou90, Sha95a]. Livestock [Yea90]. Living [Mic94]. ljus [Sun97]. Llego [PO99]. Lloyd [Web90a]. Lloyd-Jones [Web90a]. Loading [Rol94]. Lobster [Lun90b]. Local [O’D96b, Per99b, Cla97a, Leo96]. Locomotive [Mal98, Mit95, Sch99b, vR90]. Loeb [Bix98, Maz94]. Logic [Cre99, San92]. Logistics [Kop91, Rol95a]. Lois [Gom97, Sha93]. London [Div95, Rhe94, Ano97i, Rob96b, Bow99b, Mic94]. Long [Bin91]. Longbow [Phi99]. Longitude [Lan93]. Look [CC95, Jac93b]. Loom [Con95]. López [Coo91a]. Lord [Por91, Lai92]. Loren [Ada95b, Bro95a, Ras92b]. Losing [Eva93b, Ryc93]. Loss [Hut92]. Lost [Rou96]. Lotchin [Cor93]. Loubère [Gut92]. Louis [Bla90a, Coo93b, Gut90, Lau98b, Mag95, Smi90b, Rei95a]. Love [Low97]. Loved [Lan90b]. Lovejoy [Wos91]. Lovell [Smi91c]. Low [War91]. Lowell [Jer94b, Ano95i]. Lower [Sha95c, Smi92b]. Lowlands [Mul96a]. Lowood [Nye93b]. Lubar [Wil95c, Chr95a]. Luccarelli [Bia97]. Lucic [Sta92b]. Lucien [Lau98a]. Lucius [Gol92]. Luckin [Coo92b]. Lucky [Dup92]. Luddities [Tan93]. Lumber [Cor90]. Lumières [Lan90a]. Luminaries [McG97a]. Luminiferous [Mac93a]. Lumley [Gol90]. Lunar [Swe92]. Luz [PO99]. Lyle [Bry95]. Lynn [Mor91b, Hea91b, Rol95a]. Lynwood [Coo93b]. Lyon [Kid90]. lysende [Myl97].

M [Adr94, Alb97, Ber94a, Chr95b, Dav93, Edm96, Eze90, Hac91a, Ham98a, Har98b, Har97, Hom94, Isr92a, Jan91, Kas99, Lan95b, Lew97, Lie93, Lip97, Lun90b, McG97b, McM94a, Mil91a, Mit96, Miy94, Phi91, Pla96, Por91, Qui95, Ros92, Sha96b, Sim94, Spe95, Sta97a, Ste92, Ste98, Sti91, Tan92, Til97, Tra97a, Uss93, Was93, Web90a]. Ma [Dav99]. Macaulay [All90]. Macdonald [Cur91, Goo90]. Machiavelli [Kel99]. Machine [Cor91, Gol90, Gra95, Gre99, Hoc98, Koc98, May92, May94, Mis93, Rab96, Rob93, Sca99, Sta90c, Sum97, dlP99, vHT96, vR90, McM94a, Sta92b, Chr95b, Sei98, Smi99]. Machine-Age [Chr95b]. Machinery [Coo91c, Mac93b, Nor90, Tan93, Old92]. Machines [Adl90, Ben96, Bix99, Cro94, Edw97, Kid90, Lau99, Lon95b, Pos98b, All90, Kat94b, Uss90, Coo95, Hea90, Ros98]. Macht [All96a, Rüd92]. Mack [Bar92b]. MacKenzie [DeV92, Edw97]. MacLeod [Gus91, Pye90, Rob90b]. Maddin [Wes90]. Maddison [Kid96]. Made [Hay90, Hew94, Lan94c, Seg90, Smu93, Mul96a]. Madison [Fit97, Ano92i, Sin95]. Madrid [Hau90]. Maffioli [Mac96]. Magic [Spe95]. Magnetic [Cla93]. Magnificent [Kel99]. Magnus [Fin98]. Maidens [Her99b]. Mailer [Sla97]. Maillart [Far99, Isr92b]. Main [Bro97b]. Maine [Hay90, Seg90, Eve92, Lun90b]. Maintenance [Rey99]. Maiolica [Kin93]. Major [McN90, Cut99]. Make [Lan98a]. Maker [Til90]. Makers [Eva96, Roc94]. Makes [Par97]. Making [Ben92c, Bla90b, Coo91b, Fit98, Gar92, Hor99, Hor94, Kal98, Kid91, Kli91, Lay96, Leo96, McG95, Nee96, Old97, Ors97, Pas96, Sin99c, Slo96, Vin97, War92b, Wil92a, Zak92, Rot95, Egg94, Gal92a, Goo95, Gus98, Hea99b, Kat94a, Pur90, Rab96, Rug96, Tob92, Wor98]. Malavis [Oli98]. Male [Cow96, Old97]. Mallet [Col99]. Malmö [Car99b]. Malmska [Car99b]. Malone [Lan95b, Was93]. Man [Cur91, Hew94, Kle91, Lan90b, O’B91, Ree99, Rot97, Swe92, Mul96a]. Man-Made [Mul96a]. Management [And99a, Ben91c, Big94, Con96b, Coo91c, Coo98d, Elv99, Joh97b, Lee97, Mac93b, Mul96a, Nel92, Pul92, See90, Uss93, Wor93a, Lub91a, Sul98a]. Managerial [Sch99b]. Managers [Cra97, Don94, Jos95a, Gro98]. Managing [Low97, Rog95, Jac97, Jer94a]. Manchester [Ham91, Mor99, Tra94, Web90a]. Manfrass [Die93]. Mangle [Con97]. manifattura [Sel92]. Manned [Daw91]. Manning [Bur92, McM94b]. Manufacture [Gut93, Lea92, Wal95, Zei95]. Manufactures [Zem93]. Manufacturing [Ald97, Gor97b, Mil93, Mit95, Scr94, Arm96, Gor95a, Bel98b, Liu97, Sok93]. Manuscripts [Lon93a]. Many [And99a, Blo91]. Manzini [Fri91a]. Mapping [Hod93]. Maps [Kon99, Poo99, Cod92, Kie90b, Kon97]. Marc [Ber95a, Bla96b, Isr94a, Kie90a]. Marcel [Bed90b, Hin94, Tob92]. March [Har93b]. Marchand [Cur91]. Marco [Rob95]. Marconi [Hon94]. Marcos [Gua98]. Marcus [Dou91b, App96]. Margaret [Big97, Lay97, Rob92b, Roz98, Sim94, Sta92a]. Margolis [Por94]. Margot [Wos91]. Mari [Eva96]. Maria [Rüd92]. Marian [Lüt98]. Marianne [Och92a]. Marie [Neu98]. Marika [Gie96]. Marine [Mat99b]. Mariner [Bed90c]. Maritime [Bac90, Jer90b]. Marius [Lau98b]. Mark [Abb96, Bia97, Bil97, Hun96, Jac91, Kel96, Kie90b, Nee96, Ter94, San94, Tri91]. Market [Lam93, Sei96b, Wil92a]. Marketing [Daw96, Lai91, Skr93]. Markets [Cla97a, Con94b, Liv92, Scr94]. Markley [Sin97]. Marks [Rho93]. Markt [Rüd92]. Marling [Bra96]. Marlou [Haf97]. Marrey [Dre95]. Marriage [Lau95]. Mars [Rol95a]. Marsh [Lay97]. Marshall [Was92, Bla97a, Cur91, Cur91]. Martha [Big96, Kuc94, Lai93]. Martijn [Rob94a]. Martin [Ada91, Hac92b, Nye92, Shi97, Swe93, Fis92, Hil92b, Lyn98, SB95]. Martinique [Paq97]. Marx [Adl90]. Mary [Fin95, Gre91, Kri98, Per93]. Maryanne [Cla97a]. Maryland [Sha93]. Maschine [dV95]. Maschinen [All96c]. Maschinenbau [All96c]. Maschinenlinien [dV95]. Maskelyne [Gin90]. Mass [Alv99, Bla95, Blu99, Hyd99, Lai91, Moh90, O’D97, Sch94d, See99b, Wil92a, Zei95]. Massachusetts [Ano95i, Jer94b, O’H96, Tra94]. Massenfertigung [All96c]. Massenmordes [All96b]. Master [Ben91c, Hac98, Scr96, Tri91]. Mastered [Hen99]. Mastering [Die97]. Masterpiece [Mor91b]. Matejka [Hou90]. Material [Cre99, Fer92, Fri91a, Hor92, How98, Kri98, Pan98a, Was91, Wil95c, Bla99a]. Materials [Fri91a, Wer97, Gor92a]. matériaux [Kah96]. Mathematical [Jes97, Ozd98]. Mathias [Gib93]. Matter [Rob95, Rol97, Bla99a]. Matters [Pos99]. Matthew [Gor95a, Her94]. Matthewman [Wor90]. Matthias [All98b, Ver97]. Matti [Jud90]. Matures [Ada97]. Mauer [Cap90b]. Maureen [GS98, Kea98]. Maury [Gor95b]. Mauskopf [Ful95]. Mauve [Tra90]. Max [Hig97b, Mis93]. Maxwell [Vol94b]. Maxwellians [Hon94, Bro93a]. Maya [Bed90a]. Mayr [And92]. Mazlish [Kat94b]. McBride [Smi95a]. McCaffray [Gra98b]. McCann [Zel97]. McChesney [Sch94d]. McClellan [Wil93b]. McClelland [Tho93]. McCurdy [Log92]. McDaniel [Oli90]. McFarland [Mor93, Con96a]. McGaw [Pas96]. McHugh [Bla90b]. McIvor [Gar90]. McKelvey [GS98]. McKiven [Lew97]. McLuhan [Cur91]. McNamara [All97]. McNeil [Low92, Pur97]. McPhee [Reu91]. McShane [Hug95]. Mead [Smi94]. Meaning [Cha93]. Meanings [Hir92, Sha95b]. Measure [Hea90, Por98]. Measurement [Aga98]. Meatpacking [Hor97b]. Mechanical [All99a, Cou94, Kel93a, Mai98, Pan98c, Has93]. Mechanick [War91]. Mechanics [Fer92, vHT96, Fer92]. Mechanik [Lon93b]. Mechanisierung [Mer97]. Mechanisms [Sul98a]. Mechanization [And99b, Rei99a, San90]. Mechanized [Cam95, Cum91]. Medal [Ano90g, Ano91i, Ano93h, Ano94h, Ano95h, Ano96g, Ano97h, Ano92h]. Media [Joh92, Oga90, Phi91, Sch94d, Wos91, Hil97]. Medical [Big95a, Her98, Jef95, Mar93b, Mel99, MKR93, Mar93a, Sha95a]. Medicine [Eva93b, Mar96, Pry93, Sha95a, Sta97b, How93, Coo97, Led93]. Medieval [Cla97a, Fle90, Fle92, Gui98a, Lon96a, Mat91, Mat92b, Mue90, O’D94b, Sun99, Tar95, The90, Tra97a, Whi92c, Mue99, DeV93, Par94, Rey90a, Whi92c]. meditaciones [MA97]. Mediterranean [Bac90, The90]. Medium [Cur91, Hyd99]. Medizin [Bro99a]. Meeting [Ano90h, Ano91c, Ano92i, Ano93l, Ano94i, Ano95i, Ano96b, Ano97i, Ano98e, Ano99m, LM91]. Meiji [Coo92a]. Meikle [Rhe98]. Meiksins [Cha98]. Mellon [Ser94]. Melosi [Nye92]. Melvin [Mar91, SMP^{+}96, Van96]. Melzer [McG97b, Mit96]. Members [Ano90d, Ano91f, Ano92e, Ano93e, Ano94e, Ano95e, Ano96e, Ano97e]. Membership [Rhe94]. Memex [McM94a, Rei95b]. Memoir [Ale99a, Jos99, Jak98]. Memoirs [Bad91, McC93, Hol97a]. Memorabilia [Har91]. Memorandum [Har99b]. Memoriam [SMP^{+}96]. Memory [O’D96b]. Men [Fan91, Gro90b, Hea90, Hor97b, Old92, Pos98b, Rol90, Roz98, Smu93, Til90, Uss90, Edg98, Gre91]. Menahem [Sch95]. Mendelian [Pal93]. Mendelsohn [Hac91b]. Menlo [Mil97]. Mennonite [Fis98]. Mental [Big94]. Menzies [yH98]. Merchant [Ste91a]. Merck [Car99a]. Mere [Pos99]. Meridiankreises [Dek98]. Merseyside [Jer90b]. Message [Sla98, Cur91]. Messenger [Cur91]. Metal [Gor97a]. Metallurgical [Chi97]. Metallurgy [Lon91b, Gor91]. Metals [Wes90]. Metalworkers [Jos95a, Len91]. Metalworking [Liu97, Kri97]. Metaphor [Lai92]. Metaphors [Lin97a]. Meteorological [Fel99]. Meteorology [Fel97, Fel99, Kie90a, Ell92b]. Meter [Daw96]. Method [How98, Ror91, Win98]. Methods [Bat98, Scr98b, Vol98a]. Metropolis [Con92b, Dan95, Twe92]. Metropolitan [She94a]. Metropolización [D’A97]. Metz [Lan98b, Ste92]. Meulen [Bil97, Tri93]. Mexico [Bro93b, Gre99, Chi97]. Meyer [Hor94, Pet90, Tri93]. Miasma [Dav93]. Michael [Ang90, Cha96, Coh90, Edw95, Fri92a, Gol91a, Gor91, Haf92, Hea90, Hea98, Hod93, Hou90, Hug96b, Ing91, Jer94a, Lee96, Len91, Leo92, Mei93, Ror91, Rüd92, Rum92, Sim95b, Ste90, Str90a, Wri94]. Michele [And94, Fis92]. Michelle [Sto97]. Michiel [Leo92]. Michigan [Sha95a, Ser96]. Microchip [Kid90]. Microelectronics [Pfa98]. Microfilm [Cha90]. Microphysics [Cre99]. Microscope [Kun95, Smi98, Bal98]. Microscopy [Lüt98, Ras99]. Microwave [Can95]. Mid [Cla96, Ler97, Voe90]. Mid-19th-Century [Ler97]. Mid-Century [Cla96]. Middle [Kai96, Rol95a, Rug96]. Middle-Western [Rug96]. Midwest [Hur90, O’D96a]. Might [Bix98, Mol91, Per93]. Migrate [Ras99]. Mikael [Coo95]. Mikroelektronik [Rüd92]. Milberry [Hig90]. Mildred [Gro90b]. Miles [Sum97, Cas98b, Sin99c]. Milieu [SD98, vR90]. Militant [Hor94, Woo92]. Military [BA90, Cap91, Dor91, Edg98, Gei94, Hac92a, Hac93, Hac94c, Hac94b, HD90, Hal98, Hol94, Hol99, Kuc94, Moy94, Mul96b, Par94, Tri93, vk92b, Dor91, Kuc94, Hac91b]. Military-Industrial-Academic [Gei94]. Mill [Bla90b, Moe94, O’D96b, TJ96, Bla90b, LeC94]. Millard [Isr91, Tho97a]. Millennium [Tei99]. Miller [Bla97c, Cro94, Hug90]. Millhands [Cra97]. Millikan [Low93a]. Millinery [Lew99]. Milling [Eve92, TJ96]. Mills [Cla95, Coo92a, Egg94, Gro90b, Jer94b, Lea99, Mue90, Mue99, Pet97, Rey99]. Millwrights [Kyn99]. Milton [Tra96]. Minami [Coo90]. Mind [Gol95, Kra94c, McG95, McM94a, Moh90, Pin90]. Minds [Ben99]. Mine [Ald95, LeC94, Was92]. Miner [San90, Hir96]. Miners [HM96]. Mines [Cla95, Mou99, Mue99, Ald95, Och92b, Str91, Dub93, Bel98a]. Mining [Gor97a, HM96, Lon91b, New91, Och92b, Was92, You93, FO92, San90]. Minute [Sal95]. Miron [Gra97a]. Mirror [Sim91]. Misa [Lay96]. Miscellany [Har99b]. Misconduct [Hin94]. Missile [Buc98b, DeV92, Hev95, Log95]. Missiles [Mul96b, Neu96b]. Missing [Scr00]. Mission [Ada96]. Missions [Swe92]. Mississippi [Ano99n]. Misunderstanding [Sol98]. Mitcham [Bec91, Lue90, Caw95, Cut98, Ile96]. Mitchell [San92, Wos94]. Mitsubishi [Mat99b]. Mixed [Sla95, Lan99]. Mixing [Ham95]. Mobile [Ber93]. Mobilitat [Tod99]. Mobilizing [Wal94]. Moby [Zie92]. Model [Con92c]. Modern [All96d, Bac99, Bal98, Bas92b, Bix99, Bla97b, Cla91, Cra97, Eis93, Far96, Ful95, Haf90, Jac91, Kat94a, Kea98, Kid92, Kie90a, Kon98, Kuc94, Lay96, Lie93, Lon95a, Lon96a, McG99, Ors97, Per91, Phi99, Pot99, Rey99, Skr93, Sta94, Tas97, Tho93, You95, dlP99, All99b]. Moderna [Sel92]. moderne [Lan95a, LF93, Lav93]. modernes [Dre95]. Modernisierung [Wen92]. Modernism [All96d, O’D97, Cla96]. Modernity [Gol91a, Kat92, Øve97, Sha95c]. Modernization [Mat99c, Nye92, Rob96a, Sel99, Whi92b]. Modernizing [Eva92]. Modest [Bow98]. Moffett [Hol95]. Mokyr [Har96c, Hea91a]. Molecular [Bud94, Bud96]. Mollenhoff [Cap90a]. Molloy [Smi97b]. Molson [Eze90]. Moltke [Gui93b]. Moment [Nye98b]. Momentum [HS96]. Money [Hol90]. Monika [Hun96]. Monitor [Fer90, Min95a]. Monitoring [MKR93]. Monmonier [Kie90b]. Monopoly [Mis90, Sin99c, Owe91b]. Monster [Bas92b]. Montana [New91]. Montgomery [Lea92]. Montmartre [Bel98a]. Monument [Lan90b, Lon97]. Moon [Log91]. Mooney [Hut92]. Moore [Big95a, Win90]. Moores [Bla97c]. Moorhouse [Vol93]. Morel [Hal93a]. Mori [Has93]. Morison [MMH96]. mörker [Sun97]. Morone [Cla90]. Morris [Mil94, Coo98b, Nas91]. Morris-Suzuki [Coo98b]. Morrison [War91]. Morrison-Low [War91]. Morro [Was92]. Morrow [Hol94]. Morse [Lai97, Mil91b]. Moshe [Adr94]. Mosque [Ozd98]. Most [Bar91b, Mag94, Mor90, Ste89]. Motel [Mei99a]. moteur [But92]. Motion [Gol95, Zem92]. Motor [Blo90, Cas98a, Cor92, Min95c, Olw96, Vol91c]. Motors [Bri90]. Mountain [Bla96a, Rey98]. Movable [Pal91]. Move [SD98]. Movement [Fit93b, Jud99, Pur93b, Con91]. Moves [Ras99]. Movie [Ber94b, Bro98a]. Moviemakers [Per91]. Movies [Bro97b]. Mr. [Dia99]. Mueller [McN90]. Mulford [Car99a]. Mulholland [Rot99a]. Mulhouse [Tra94]. Muliebria [Fle92]. Mulligan [Mal91]. Multidisciplinary [Low96]. Mumford [Ben92a, Bia97, Hug90, Hug96a, Bla91]. Munich [And92]. Muron [Lau98b]. Murray [Log91, Gru90]. Museum [Ale99b, And92, Bla92b, Bra98, Bra95, CF92, Coa99, De 99, Eze90, Fit93b, Gol90, Ham91, Hay90, Hor92, Jer90b, Kid91, LaF91, Lie92, Mat99a, Mat92a, May98, Mil91a, Mol91, NP94, O’M93, Per91, Pos98b, Rol98b, Roz96, Str90b, Ten92, Tur95, Vic92, Wri96, Yu99, Abb93, Arr94, Bat90, Bro95b, Div95, Isr94b, Lan94c, Lea99, Olw96, Ste94c, Twe92, Vol95, Wil90, Wil91, Zem92, ML98]. Museums [Gro93, Hay96, O’D96b, Skr91]. Music [Hor99, Mah90]. Music/Consuming [Hor99]. Musical [Bra97]. Musicians [Kra94a, Tho98]. Musings [Cow96]. Muskets [Ste94a]. Musschenbroek [War98a]. Musser [Mil93]. mutation [Lan91b]. muutos [Ahv90]. MX [Her94]. My [Lie91]. Myllyntaus [Ahv90, Nye93a]. Myron [Haf90]. Mystery [Tri91]. Mysticism [O’H98]. Mystique [Dup92]. Myth [Lam93].

N [Arm96, Blo91, Cam92, Edw96b, Egg94, Hal90, Huf92, Kid91, Kon98, Low97, Mer95, Sum93, Uss97]. Nabokov [Mat90]. Nagengast [Die97]. Naked [Her98]. Nancy [Moe94, Scr90]. Nantes [Lan91b]. Naomi [Lip97]. Narrative [Big96, Buc91, Law91, Scr91c, Mit91]. Narrow [Sal92b]. NASA [Bar91a, Bil98, Kay99, Joh97a, Nee96, Nea98, Vin93]. Nasaw [Ada95a]. Nash [Bar92c]. Nathan [Lay92a, Yat96]. Nation [Bro95c, Lay96, Neu93, Par97, Rei96, Rob92a, Swe93, Woo92]. National [Eze90, Hec94, Hun96, Kle99, Lar98, Mol91, Mul95, Pye92, Sha90a, Ten92, Wes91, Wil90, Ale99b, Bat90, Coa99, Hor92, Roz96, Str90b, Wri96, Yu99]. Nationalism [Edm96]. Nationalized [Hec92]. Nations [Kel93d]. Native [Mat90]. Natur [Brü94]. Natural [Con94b, Haf92, Kay92, Lau95, Lon95a, Por93, Roz96]. Nature [All99a, Bes95, Cha93, Hea99b, Kip97, Lon96a, Reu95, Sim95b, Ste91a, dlP99, Bla96a, Cha95, Reu91, Cha96, Con92b, Ell94b, Smi93b]. Naturvorstellungen [Brü94]. Naturwissenschaft [Bro99a]. natuurkunde [Dav96]. Natverksentreprenoren [Car99b]. Naval [Bur92, Gui91, Hol95, McB92, Owe91a, Sum91, Tan92, Fri99, McG99, Ols98, vk92b, Gar91]. Navigation [Lee97]. Navigational [Dic92]. navires [Fer99]. Navy [Dic92, Dud92, Dup92, Gar92, Tan92, Wal91]. Nazi [Wal99a, All96d, Tho90]. NCR [Kid94]. Near [Hei95a]. Nebeker [Fel97]. Necah [Wes91]. Neck [Sch94c]. Nederland [Hom94, vL90]. Needham [Mul96b, Mul96c]. Needless [Ald95]. Nef [Pol90]. Negev [Mer95]. Negotiating [Tra97b, Koc93]. Neighborhood [Leo96]. Neil [Fro93, Skr93]. Nelson [Pet90, Big94, Nuw90, Wor91]. Neonatal [MKR93]. Net [Sin99b]. Netherlands [Mul96a]. network [Car99b]. Networked [Bol90]. Networks [Car99a, Mah97]. neuen [Buc97]. Neufeld [Len91]. Neumann [Kid92].

Neuzeit [Har99d]. Nevada [Bla97a]. Never [Bin91]. Nevil [Gin90]. Neville [Gor97b]. Newberry [Kuc94]. Newborn [Koc98]. Newell [Gib91, Jac93b, O’B91]. Newlands [Hol97b]. Newman [Eam97]. News [Kie90b, Lai97, Sch95]. Newspapers [Hil97]. Newton [Cut94a, Mor93]. Next [Hof92]. Niagara [Ber98]. Niccolo [Kel99]. Nicholas [Gut95a, yH98, Jac95a, Lon95b, Oli98]. Nichols [Wil92b, Wor95]. Nick [Hol90]. Nickelodeon [Mil93]. Nicolás [Gli91]. Night [Opi90]. Nightmares [Sta96a]. Niklaus [Reu93a]. Nineteenth [Bro92a, Bro96, Gut93, Gut95a, Haf97, Kat92, Kin99, Lie93, Mac93b, Mar96, ML98, Opi90, Per93, San97, Tal94, Tho93, Ube94, Voe90, Wes95, vK92a, Dre99, Reu99b]. Nineteenth-Century [Gut95a, Kin99, Lie93, Mac93b, Per93, Tal94, Wes95, Dre99]. Nissen [Hei95a]. Nixon [Maz94]. No [HS96, Mul95, Pos99, Swe92, Fer99, Fit97, Has93, Qui95]. Noble [Pri95]. Noe [Rob96a]. noir [Smi95b]. Noll [Ell93]. Norbert [Buc97, Gor94]. Nordic [O’D94a, Gie96]. Norling [Roz98]. Norman [Arm96, Kon97, Mor91b, Was91, Wor94, Fri90]. Norms [Cla91]. Norrell [Dan93, Rey91]. Norris [She94b]. North [Bry96, Cod92, Fit98, Gei90, Lan99, Lan95b, MM92, Ras92a, Ste90, Whi92b, Wor93a, Cru99, Her99b, Jag99, Mac91, Rei99a]. Northeast [Cla95, O’D96b]. Northeastern [Con94b]. Northern [Hug99]. Northrop [Con94a, Vin94]. Norton [Alb97, Por91, Sim95b]. Norwegian [Qui98, Voe90, Wor98]. Note [Fri95, Kry90, Sta98]. Notebooks [Bro98b]. Notes [Ano98e, Ano99m, Lon91a, Sin95, Sla91]. Nouveau [Lev91a]. Novak [Hug96a]. Novelty [Bow99a, Pat95]. November [Ano92i, Fit97]. novembre [Lan98b]. nucléaire [Smi91d]. Nuclear [And94, Bad94, Car97, Cla90, DeV92, Gol93, Hac90a, Hac94a, Hac98, Ham98a, Hea96, Hec94, Her94, Hir96, Moh95, Sei96a, Shi97, Wal91, Wal94, Olw98]. Nucleus [Maz90b]. Nuevas [MA97]. Number [Ano90k, Ano90j, Ano90l, Ano90m, Ano91n, Ano91o, Ano91m, Ano91p, Ano91q, Ano92n, Ano92m, Ano92o, Ano92p, Ano93o, Ano93n, Ano93p, Ano93q, Ano94n, Ano94m, Ano94o, Ano94p, Ano95n, Ano95o, Ano95m, Ano95p, Ano95q, Ano96k, Ano96j, Ano96l, Ano96m, Ano97n, Ano97m, Ano97o, Ano97p, Ano98g, Ano98f, Ano98h, Ano98i, Ano99r, Ano99q, Ano99s, Ano99t]. Numbers [Alb97]. Numerical [Ben92c]. Nuremberg [Len91]. Nursery [Koc98]. Nyce [McM94a]. Nye [Hir92, Mei95]. Nyegaard [Qui98]. Nyström [O’D94a].

O [Cut93, Cut97, Kri97, Vol91a]. O-Rings [Cut93]. Oakland [Pos98b]. Oates [Cal92b]. Objectivity [Alb97]. Obscurity [Hon97]. Observations [BA90]. Observatory [Rot92, Lan90b]. Observer [Kat92]. Obsession [Rot97, Sha95b]. Obsolescence [McN90]. Occasion [Kry90]. Occupation [Col90]. Ocean [Roz96]. O’Connell [Rol90]. o’Cotton [Cou94]. October [Ano90h, Ano91c, Ano92i, Ano94i, Ano95i, Ano96b, Ano98e, Ano99m, Fit97, Ano90m, Ano91q, Ano92p, Ano93q, Ano94p, Ano95q, Ano96m, Ano97p, Ano98i, Ano99t]. Odds [Och92a]. Odyssey [Pur90]. Offensive [Hig97b]. offentlichen [Tod99]. Office [Bur92, Jan91, Kie93, Sta94, Sti98, Tan92, CK98, IR91]. Ogle [Kea98]. Ohanian [Moe94]. Ohio [Bla92b, Bro95c, Col93, Rob90a, Uss93]. Oil [Coa99, Lig97a, Nas90, Oli98, Tra95, Oli90]. Olaus [Bjö97]. Old [Dud92, Fis98, Gra97b, Mil91a, Tho91a, Wil93b, Win90]. Oldfield [Cor97]. Ole [Hår98a, Myl97]. Olive [Lig97a]. Oliver [Hon97, Jac93a]. Olivetti [Pul92]. Olivier [Gal92a]. Ollantaytambo [Pet96]. Olney [Lai93]. Olof [Bjö97]. Olsson [Cut97]. Olympus [Law92]. om [Car99b]. O’Malley [Rum92]. Omar [Ozd98]. Omissi [Ken92]. Once [Til90]. onder [Dav96]. onderwijs [Wac94]. One [Ano91j, Hew94, Min99]. Only [Kir95]. Ontario [Eva93a, Gib91]. Openness [Lon91b]. Opera [Fle92]. Operation [Kra92]. Operations [Hig97b, Joh97b]. Opinion [Rog95]. Opium [Lai96b]. opkomst [Wac94]. Oppenheimer [See95b]. Optical [Ada91, Fie94a]. Optics [Kip97]. Opting [Tra95]. Oral [Smi95a]. Orange [Lan98a]. Order [Ald96, Dou95, Fis98, Hac93, She94a, Tho91a, Ras92b]. Ordering [You95]. Orders [Eat90]. ordinateurs [Coh90]. Ordnance [Col99]. Ordway [Kra93, Neu94]. Oregon [LaL90]. Organic [May92, Sca99, Tra90]. Organization [BA90, Div94, Lam93, Lip97, Lun92]. Organizational [Ano98e, Ano99m, Law92, Sch99b]. Organizations [Moh95, Tho93]. Organized [Nuw90]. Organizing [Bro97a]. Oriental [War98a]. Origin [Bry96, Lon91a]. Origines [Coh90, vR90]. Origins [Ait94, Can95, Gro98, Hea96, Hon94, Hug99, Jer97, Kid92, Koc98, Lew94, MS92, Por93, Pur93a, Roc94, Sei93, Sta94, Wri94]. Ormiston [Mit91, Pin92]. Ormrod [Rot95]. Ornamental [KK92]. Ornstein [Dav96]. Orosz [Fit93b]. Orvell [Sum97]. Orville [Gol91c]. Osborne [Cha96]. Osborough [Kon98]. Osietzki [Rüd92]. Osiris [Les94]. Osiris. [Fit97, Mor97]. Osterbrock [Rot92, Rot94a]. Ostroff [Nye90]. Other [Cod91, Fle90, McG97a, Mor99, Rei95b]. Otis [Ryc93, Stu94]. O’Toole [Gal92b]. Ottawa [Sum91]. Otto [And92, Jam90]. Our [Cow96, Hou90, Per93, Tob92, Gor98]. Ourselves [Kel93d]. Outpost [Qui94]. Outrageous [Har95b]. Outstanding [Fit95]. Overseas [Ada96, Hea98, Sha96b]. Overview [Wri96]. Owen [All98a]. Own [Tob92]. Owning [Bla96a, Hil91a]. Oxford [Poo99]. Oxygen [MKR93]. oyobi [Has93].

P [Ben92a, Ben90, Bro98b, Dou91b, Edg98, Gra98b, Gut95b, Haf90, Ham98a, Hol93, Isr92b, Jac91, Jak98, Koc98, Lie90, Mar90, Mil95, Mor91a, Mor91d, Mor93, Mul96a, Nel92, New94, Pol91, Pry93, Rei96, Sla95, Tho98, vL90]. P. [WN98]. Paarlberg [Ras90]. Pacey [Bas92a]. Pacific [Jer97, O’B91, Lew90]. Pacing [Jef95]. Packard [Arm96]. Paid [Thr91]. Painter [Kas99]. Painterly [Kin93]. Painting [WN98]. Palace [Gut93, Whi90]. Paleolithic [Bar90]. Palestine [Ben92b]. Palladino [Sal92a]. Palmer [Hoy92]. Pamela [Bar92b, Gor92a]. Pamphlets [Sta92a]. Pan [Sim95a]. Panhandle [Mea98]. Panorama [Hyd99]. Papayanis [Gut95a]. Paper [Hil99, Moe94, Wri92]. Papermaking [Rey99, Gen92]. Papers [Cha90, Har93b, Har96b, HC98, Isr94a, Kli91, Lev97a, Mil97, Jan91, Sta97a]. papetière [Ros98]. papier [Ros98]. Para [PO99]. Paradigms [Fer95]. Paradise [Big97, LaL90, Slo95, Bla97a, Kel93a]. Paradoxes [Mac91]. Parascandola [Rei93b]. parcours [Bel98a]. Paris [Gut95a, LM91, Rei96, Haf99, Fro92a]. Parisian [Tal94]. Park [Dav92b, Mil97]. Parker [Cam92, HD90]. Part [Ano91j, Ano91k, Ano91n, Ano91o, Cha90]. Participation [Gol93]. Partners [Ser94, Kay92]. Partnership [Kra92]. Parts [Ald97, Hou97, Blo91]. Pasadena [Ano98e]. Pascal [Lav93, Smi92a]. Passage [Vol98b]. Passenger [Min95c]. Passfield [Rey90b]. Passing [Pos95]. Passion [Pis96]. Past [Lup91, Miy94, Pry93, Yat92, Bri92, Sin93]. Pasteur [Cha99b]. Pasteurization [Rei91]. Pastoral [Fle90]. Pat [Bra93, Hut92, Kat97a]. Pate [Yea90]. Patent [Coo91b, Coo91c, IR91, Jan91, Lub91b, Mac93b, Rob90b]. Patenting [Mac91]. Patents [Doo91, Lon91a, Owe91b, Uss91, Gra97b]. Paterson [Gro90a]. Path [Caw95, Cum91, Hug95]. Paths [Ing99]. Patient [Gam97]. Patrice [Lav93]. Patricia [Coo92a, Pet97, Til90]. Patrick [Hug96a, Lan95b, Was93]. Patriotism [Nye93b]. Patron [Ser94]. Patronage [Koh96]. patrons [Lin94b]. Patterns [DeV91, Moy94, Bla92a]. Pattillo [Oli90]. Paul [Cla91, Con90, Dav93, Gor97a, Gra95, Hal93a, Kop91, Lie91, McM94a, Mil91b, Pin90, Reu93b, See90, Sma95, Uss97]. Pauline [Pet92a]. Paulinyi [Tod96]. Paulo [Pri98]. Pauper [Rot94a]. Pavilions [Bod99]. Paws [All99a]. Pay [Lai93]. Payen [vR90]. Payson [Ano90a, Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Ano97a]. PCR [Gus98]. Peace [Hac91a, Pri98, Pur90, Hig91]. Peebles [Zie92]. Peenemunde [Buc98b, Jak98]. Peggy [Sma95]. Pehrson [Mah97]. Pencil [Gor92b]. Pendulum [Lan93]. Pendulums [Ste94a]. Penelope [Bla97b]. Pennsylvania [Sal92a, Gor97b]. Penslar [Ben92b]. Pentagon [Cer98]. People [Bla97a, Dou91a, Gar92, Hug99, Jam90, Kel95, Lie92, Mag94, Pac95, Sch92b, Abb93, Pfa91, All99a, Zel97]. Perception [Dav93, Nei98, Kop91]. Perceptions [Bes95]. Perestroika [Ger96]. Perfection [GS99]. Performance [Gui95b, Hil99, Kni95, Ube99]. Peril [Ald95, Ald99]. Period [Hol99, For93]. Periodicals [Low93b]. Periphery [Lig97b]. Perkins [Rue90, Tra90]. Perry [How93, Kie93]. Persistence [Bro92b]. Persistent [Bro97a]. Personal [Bra91b, Gre95a]. Personalities [Hyd90]. Personality [Ube99]. Personennahverkehrs [Tod99]. Perspective [Seg90, Lan91b, Cha98, Har96c, Mar93a, Mei98a, Ols97, Rom97b, Ryc92]. Perspectives [Asp93, Ber90, Bro97a, Edw95, Fit95, Gie96, Lip97, Rob94a, Sun97, Swe93, Thr91, Low96]. perspektiv [Sun97]. Persuasion [Øve97]. Perú [Gua98]. Pessimism [Nye97]. Pesticides [Gor98]. Pestilence [Dav93]. Peter [Asp93, Bad94, Big91, Bri92, Cha98, Coo91a, Cor91, Eva92, Gib93, Gol91b, Gui95a, Jes97, Low97, Mat90, Mul90a, Nas91, Neu93, Rob94a, Sim95b, Smi90b, Ste94b, Tri90b, War98a, Wes93b, dV95]. Peters [Smi90a]. Petersburg [Hur98, Jos95a, Kry90]. Petersen [Cor90, Vol95]. Peterson [Dan90]. Petroleum [Kra92]. Petroski [Fer95, Gor92b, Sch94b]. Petter [Qui98]. Pflicht [Tod99]. Phantom [Hou97]. Pharmaceutical [Lie90, Lin98, Qui98]. Pharmacology [Rei93b]. Philadelphia [Coo93c, Liv92, Rab96, Gib99, Ler97, Jac93a]. Philip [Blo91, Cur91, Hac90a, Liv92]. Philippe [Coh90]. Philosophers [Lon95a]. Philosophical [Con98, Cut98, Ree98]. Philosophically [Dur98]. Philosophy [And99a, Bal98, Bec91, Cru99, Cut98, Mit92, Caw95, Ile96, Lue90, Sel91]. Photo [Rob90a]. Photographic [Nye95, Sil93, Wos94]. Photographs [Sam99]. Photography [Bla99b, Bro98b, Joh99b, Luk99, Nye90, Pan95, Sam99]. Photojournalism [Wri94]. Photometry [Aga98]. Phyllis [Hoy92, Sch94c]. Physical [Col90]. Physicist [Pur90]. Physics [War99]. Physikalischen [Fin98]. Piano [Lin91]. Pianotyp [Zac99]. Piasecki [Mul90a]. Pickering [Con97]. Pickstone [Mar93a]. Picon [Lan90a, Lan95a, Lan98b]. picoteux [Fer99]. Pictorial [Sma95]. Picture [Bro98a]. Pictures [Wos92]. Picturing [Ree98, Sam99]. Pierre [Bul91, Ell94a, Lan91b, Lau98a, Paq97, Pet96, Smi95b]. Pierre-Lucien [Lau98a]. Pigeon [Cap93]. Pioneer [Cla97b, New91, Ros94a, Cal92a]. Pioneers [Arm96, Gol91b, Jud99, Edw96b, Nye90]. Pipe [Ham92b]. Pipe-and-Brick [Ham92b]. Pipeline [Ste94b, Coa99]. Pipelines [Con94b, Mea98]. Pirker [Web90c]. Pisani [Rob94b]. Pitch [Sim91]. Pittsburgh [Egg94]. Place [Ale99b, Rom97a, Sla92, Bla97a]. Placer [New91]. Places [You93]. Plagiarism [Hin94]. Plains [Han99, Ras92a]. Plan [Rob93]. Planet [Roz96, Ale99a]. Planetary [Huf92, Sul98b]. Planned [Wil92b]. Planning [Fie94b, Hur98, Ros93, Sch91, Seg94, Bra98, Bia97, Bro90, Gui93b, Sha96a]. Plant [Pal93]. Plantation [Lee96]. Planter [Paq97]. Plastic [Rhe98]. Platt [Cla92, Gor95a]. Pleasure [War92b]. Pleasures [Ber94b, Pfa97]. Plenty [Bod99]. Plow [Zel97]. Plumbing [Kea98]. Plunder [Mor91c]. PLUTO [Kra92]. Pocket [O’M93, Sch93b]. Poder [Gli91]. Poetics [Har97, San97, Sla99]. Poetry [Tas97]. Point [Bat90]. Points [Ada93]. Pol [Rei99a]. Pol-Cats [Rei99a]. Polaris [Hev95]. Poliak [Kuc94, Kuc94]. Policies [Maz90a]. Policy [Bla90a, Bud96, Coz93b, Gal92b, Ham98b, Hec94, Her93, Hof91, Mac92, Owe91a, Rob94b, Ryc93, SS99, Slo98, Cla90, Kel93c, Swe93]. Political [Car98, Hec94, Hof91, Mit96, Ryc92, Tra95]. Politics [Ben96, Bia97, Bla97c, Ble91, Bud96, Dal90, DB94, Fin93, Fri91b, Gol91a, Gra98b, Haf92, Ham95, Hol90, Hut92, JM97, Kon99, Leo92, Lev91a, Lew93, Lin98, Log92, May98, Nob97, Rei96, Reu93b, Rol98a, Scr91b, Sha96a, Sim90, Sti98, Sto99, Tod95, Tra91, Tri93, Uss97, Vol91b, Bar91a, Bur92, Oga90, Smi90b, Ber95c, Nee96].

Pollution [Bri97, New97, Qui95, Rom97b]. Polly [Rue90]. Polytechnical [Tho90]. polytechnische [Wac94]. Pont [Bir90, Kle91]. Ponts [Lan95a, Dre95]. Pool [Bar92a]. Popular [Neu93]. Popularization [Neu98]. Porcelain [Mar98]. Pork [Ell93]. Port [Fit90]. Portable [Mei93]. Porter [Alb97, Mil93]. Portraits [Sti91, Frä91, Neb90]. Positions [Ree98]. Possible [Isr94b, Sch93a]. Post [Ald96, CK98, Kie93, Tra97a, Wos94, Gra91, Gui95b, Mar91]. Post-Medieval [Tra97a]. Post-photographic [Wos94]. Post-Revolutionary [Ald96]. Postal [Lai97]. Postman [Fro93]. Postmodern [Rol98a, Sla97, Wos91]. Postmodernism [Nye97]. Postwar [Bes95, Hec94, Mor91c, Mul95, Ras99, Smi99, Wes91, Par97]. Potlatch [Cor90]. Potosí [Coo91a]. Pottery [Bla92a, Bla96b, Sto97, Vas98]. Poullaouen [Bri97]. Pounds [Was91]. Pouzet [Lau98a]. Power [Bin93, Chi97, Cla90, Con95, Coo93b, Coo92b, Daw96, Dou91a, Fit90, Gam95, Gol93, Gol97, Hat90, Hec92, Hor97b, Joh91, Ken92, Lea92, Lev97b, Liv92, Lub95, May94, New94, Per99a, Pfa92, Rho93, SB95, Shi97, Wil99b, Hyd90, Zem92, Coo93b, Gar91, Gar92, Sta90c, Tod97, Bow92, Coo90, Eva93a, Hig90, Mue95, Pan98b]. Powered [Kin99, Kir95]. Powerhouse [CF92]. Powering [Joh97a]. Practical [Con99, Gui91, Ler97, Mil95]. Practice [Con97, Hig97a, Koh96, Mar96, MM92, ML98, Lue90, Mal98]. Practices [Dur98, Hår96a, Her99a, Mel99, Rey99]. Pragmatic [Gol91a]. praktiska [Bjö97]. Pratiques [Fer99]. Pratt [Hig90, Kid90, Bla90a]. Praxis [Bra90]. Pre [Fer92]. Pre-industrial [Fer92]. Precision [Con96a, Eva96, FF96, Reu99a, War92b, Alb97]. Préhistoire [Coh90]. Prehistoric [The90]. Preindustrial [Ben94, Dav93, Kuc92, Til97]. Prelapsarian [GS99]. premiers [Coh90]. Preparedness [vk92b]. Preparing [Kel93d, Ler97]. Present [Ale99b, Lup91, May94, Bar90, Con90, Lay97, Rol95a, Ter94]. Presentation [Ber94b]. Presented [Goo95, HC98, Kid96]. Presidential [Smi93a]. Press [Bat98, Ell92a]. Pressac [All96b]. Pressure [Eva93b]. Preston [Jan91]. Presuppositions [Hår96a]. Preventing [Ald95]. Prices [Smi90b]. Priests [Pfa92]. Primary [Goo90]. Prince [Rot94a]. Princeton [Isr94a]. Print [Hea99b, Hin94]. Printed [Lie91]. Priscilla [Bin91]. Private [Lan94b, Ser94, Zem93]. privée [Lav93]. Privilege [Ben90]. Privileges [Gra97b]. Privy [Wor90]. Prize [Ano90b, Ano90d, Ano90f, Ano91d, Ano91f, Ano91h, Ano92c, Ano92e, Ano92g, Ano93c, Ano93e, Ano93g, Ano94c, Ano94e, Ano94g, Ano95c, Ano95e, Ano95g, Ano96c, Ano96e, Ano96f, Ano97c, Ano97e, Ano97g, Ano90a, Ano90i, Ano91a, Ano91l, Ano92a, Ano92j, Ano93a, Ano93i, Ano94a, Ano94j, Ano95a, Ano95j, Ano96a, Ano96h, Ano97a, Ano97j]. Probabilistic [Car97]. Problem [vHT96, Bor99]. Problems [Cut99, Fit97, Gro93, Ree98, Rei91, Sim95a]. Probst [Fie94b]. Proceedings [Fit97, Lup91]. Process [Ben91c, Ben91b, Ber97b, Cor91, Mer90, Moh97b, Per90, Ros94b, Roz90, vHT96, Her93]. Processing [CK98, Cer98]. Producers [Gra98b]. Product [Sch93b]. Production [Bla95, Col99, Cum91, Div94, Ham98a, Liv92, Moy94, O’D97, Pet97, Rey98, Scr94, Sta90b, Vol98a, Zei95, Hal93a, Alv99, See99b, Gor97a]. Productions [Big96]. Productivity [Hil99, Pfa91, Fit96]. Profession [Gol99, Hyd90, Lay91, Tho91a, Chr99]. Professional [Div94, Sim90, Ste98]. Professionalization [Tho93]. Professionals [Gus91]. Professions [Tho93]. Professor [Seg90]. Profit [Ben90, Nye93b, War92b]. Program [Str91, Lew96, Nas91]. Programmers [Pfa92]. Programs [SD98]. Progress [Bri92, Fan91, Rei96, Rum97, WN98, Coz93a, Hea91a, McG97b, Neu92]. Progressive [Uss93, Jac93b]. Progressivism [vk92b]. Project [Joh97b, Wal99a, Zie92, Cap93]. Projects [Jos95b]. prole [Sel92]. Prometheus [Rot97]. Promising [Rey91]. Promotion [Nye93b]. Prompt [Rot99c]. Property [Lon91a]. Propulsion [Hol93]. Propyläen [All99b]. Prospect [Sta98]. Prosper [Kra90]. Protect [Gor98]. Protest [Gre91]. Protzen [Pet96]. Proust [Mau90]. Prozesse [Tod96]. Prussia [Tod95, Sto99, Ber95c]. Prussian [Gui93b]. Pryor [Bac90]. Psychologists [Coo93a]. Psychology [Fri90, Lev91a]. Pt [yH98, Mul96b, tZS90]. Pts [Har96b]. Public [Ada95a, Alb97, Ben92a, Bla90a, Bra91b, Bri93, Car97, Dav93, Eis93, Gal92b, Gol93, Ham92a, Her94, Hof91, Kel93c, Lai98, Lew96, Mau93, McS98, Pre98, Rob94b, Rog95, Sch95, Seg94, Sha90b, Sol98, Tob92, Lav93, Kel95, Kel93c, Lan94b]. publique [Rei99b]. Publishing [Hin94]. Puebloan [Sto97]. Pugh [Edw96a]. Pullman [Wri92]. Pulos [Mei90]. Pulp [Moe94]. Pulse [Lan93]. Pump [Wor90]. Pumps [CK95, Ron99]. Punched [Nor90]. Pure [Oli98, Ser92]. Purge [Jos99]. Purity [Gal92b]. Purloined [Uss91]. Pursell [Pac95]. Pursuing [Doo91]. Pursuit [Alb97, Con96a, Mel96, Sha95c]. Putting [Gol90]. Puzzle [All96d]. Pyenson [Ada96]. Pyne [Dun93]. Pyramids [Jac95a].

Q [Bro98b]. Qanats [Lig97b]. Quadrant [Cha99b]. Quantification [Por98]. Quantifying [War92a]. Quantitative [CC95]. Quastler [Van97]. Que [PO99]. Quedarse [PO99]. Queisser [Kid90]. Quest [Bow99c, Coo93a, Min97]. Question [Lea98, Pal91, Smi91b]. Questions [Coo92b]. Quincentennial [Por94]. Quiroga [Coo91a].

R [Ada95b, All98a, Arm96, Bri90, Bro95a, Bro92b, Cap90a, Chr93, Coh90, Coo98c, Cor97, Coz93b, DeV93, Eam97, Ell92b, Ell93, End97, Eva93a, FG98, Fer98, Fin95, Fin93, Fri92b, Gar91, Goo90, Goo95, Gos90, Har98b, Her93, Hig97a, Jer90a, Kie93, Koh96, Lai97, Lay91, Lay93, Les90a, Lev97a, Lip97, Low93a, May94, McM94b, O’M93, Ras92b, Rol93b, See95a, Sei93, Smi94, Tal94, Thr91, Tra97a, Van97, Zem93, vL90]. R&D [Bir90]. rôle [Lan91b]. Rabinow [Cla91]. Race [Gre95b, Lew97, Log91, Paq97, Lew90, Log95, Daw91]. Racing [Gui95b]. Radar [Sul98b]. Radiation [Sei96a]. Radicalism [Fro92b]. Radio [Ait94, Hil97, Kra95, Mei93, Sch93b, Slo95, Slo96, Smu90, Tho94, Slo95, Smu93]. Radiology [Koc93]. Radium [Aik98]. Rae [McW95]. Rafael [Lig97a]. Raff [Lip97]. Rail [Ald99, Hil95, Jos96, Min95c, Sum93]. Railroad [Bro95b, Bro95c, Bry96, Dav99, Gam92, Gor95b, Kle95, Min95b, Rho93, Rob96a, Whi92b, Uss93, Min95b]. Railroads [Ber95c, Bro99b, Hui99, Ste89, Sto99, Uss91, Vol98b, Sal92b, Hil92b]. Rails [McS98]. Railway [Cru99]. Railways [Cru99, Gra91]. Rainbow [Roc94, Slo96]. Raising [Jac96]. Raitz [Lar98, Lar98]. Ralf [Neu96a]. Ralph [Gib91, Sha90a]. Ranabir [Pfa98]. Rand [Kid94]. Randall [Tan93, Tur95]. Randy [Kra93]. Ranger [Dav93]. Raphael [Mit91]. Rasmussen [Fit95]. Rassweiler [Hat90]. Rational [Con99, See99b]. Rationality [Ald97]. Rationalizing [Mel99, Bra97]. Rawling [McN93]. Ray [Arn97, Cas98b, Her99b, Kat97a, O’D97]. Raymond [Kra90, Tra95]. Re [Leo92, Ros92]. Re-defining [Leo92]. Re-examined [Ros92]. Reaction [Gol93]. Reactors [Car97, Ham98a, Hec94]. Readership [Lyn98]. Reading [Eis93, Kel95, You93]. Readings [Gor91, Bla97c]. Ready [Haf99]. Ready-To-Wear [Haf99]. Ready-To-Work [Haf99]. Real [Bla99b]. Realism [Mit92]. Realities [Fro92a, Sin97]. Reality [Por98, Seg94]. Reappraisals [Kel93b]. Rear [Min97]. Reardon [Hac92a]. Rebirth [Hil92b]. Rebuilding [Ber94a]. Recasting [Jar95]. Recharging [dlP99]. Recherches [Lev91b, Lan98b]. Reclaim [Rob94b]. Reclaiming [Hol97b]. Reclamation [Jac93b, Mul96a]. Reconfigured [Wos94]. Reconnaissance [Zie92]. Reconsidering [Mat99b]. Reconstruction [Mul95, Sol98, Tar95]. Record [Tho97a]. Recorded [Tho97a]. Recorder [Mor98]. Recording [Cla93]. Records [Har99b, IR91, Bro98b]. Recovery [Hol91]. Recreational [Rei99a]. Red [Rey98]. Redefinition [Jef95]. Redisplayed [Bra95]. Reduced [Gam95]. Reed [Lan94b, Sha96a, Ste98]. Reeling [tZS90]. Rees [Jos96]. Reese [Kli91]. Reference [DeV90]. Refiguring [Lin97a]. Refining [Mer90]. Reflections [MKR93]. Reform [Aik98, Big91, Øve97]. Refrigeration [Coo95]. Regard [Bra91a]. Regime [Bau99, Gra97b, Wil93b]. Region [Bia97, CC95, Rey98]. Regional [Cla97a, McS98, Mue91a]. Regions [Sal92a]. Regulate [Mea98]. Regulated [Con94b]. Regulation [Ait94, App96, Hac94a, Oli98, Lin98, Roz90, Fit96]. Regulatory [Bud96, Mea98, Slo96]. Rehn [Jam90]. Reich [Buc98b, Neu96b, Kel93d]. Reiches [Neu96a]. Reid [Col93, Fro92a]. Reign [Bla95]. Reinforced [Isr92b]. Reingold [Lay92a]. Reinterpreting [Kli97b]. Reintjes [Ben92c]. Reissuing [Doo91]. Rejection [Hil92b]. Relating [Jan91]. Relation [Moh97a]. Relations [Car97, Gom97, Lau95, McB92, Nic99a]. Relationship [Rob94a]. Relative [Lea92]. Relativity [Hol91]. Reliable [Bar91b, Con99, Scr00, Ste89]. Religion [Lyn98, O’D94b]. Relocation [Scr90, Moe94]. Remaking [Pri98, Sca99]. Remembered [Hou90]. Remington [Kid94]. Removing [Her99b, Adr94]. Renaissance [Ada93, HV99, Kin93, Lon95a]. Renewed [Lon93a]. Renneberg [Hun96]. Renzo [Sel92]. Reorganization [Col90, Sch94c]. Reparations [Mor91c]. Report [Gre94]. Representation [Kuc94, Lub95]. Representations [Bas96, Fro92a]. Representing [Pan98c]. Reprography [Ber97b]. Republic [Jam90, Smi98]. republicain [Bel98a]. Republican [Rob93]. Research [Bat98, Bil98, Cha99a, Col90, Dre99, Fit97, Hei99, Jar90, Joh97b, LMO97a, Lie90, Lip98, Mer90, ML98, Mor98, Nah97, Nas91, Pan98a, Reu99a, See93, See98b, Ser96, Smi90c, vk92b, Car99b, Moh97a, Tan92, Wes93b, Bec91, Cut98, Fle98]. Reshaping [Bar92c]. Residential [Mil91a, Wil92b]. Resistance [Shi97]. Resonant [AC95]. Resource [Mue99]. Response [Bar91b, Bul94, KM92, Law91, LV95, MB92, Pfa98]. Responsibility [Bra91b, Sim90]. Restoration [Sha96a]. Restored [Kel93a]. Rethinking [Gol99]. Retractable [Con94a, LV95, Vin94]. Retreat [McN90]. Retrospective [Gal92b, Str90a]. Review [BS94, Bla97c, Kat97b, KM92, Lay92b, Por97, Qui95, Abb96, Ada91, Ada93, Ada95b, Ada95a, Ada96, Ada98, Adr94, Aga98, Ahv90, Aik98, Alb97, Ald96, Ale99a, All98a, All90, All96a, All96c, All96b, All98b, All99a, All97, And94, And92, Ang90, App96, Arm96, Asp93, Bac90, Bac98, Bad91, Bad94, Bal98, Bal97, Bal99, Bar90, Bar91a, Bar92a, Bar92b, Bar92c, Bas92a, Bas96, Bas97, Bat98, Bau99, Bau93, Bec91, Bed90a, Bed90c, Bed90b, Bed91, Bel98a, Bel98b, Ben92a, Ben92b, Ben90, Ben91a, Ben91b, Ben92c, Ben96, Ber93, Ber94a, Ber94b, Ber95a, Ber97a, Ber95c, Ber98, Ber90, Bia97, Big91, Big94, Big95a, Big96, Big97, Bil97, Bil98, Bin91, Bin93, Bir90, Bix98, Bix99, Bjö97, Bla96a, Bla97a, Bla90a, Bla97b, Bla90b, Bla92a, Bla96b, Bla99a, Ble91, Blo91]. review [Blo90, Blu92, Bog98, Bol90, Bor97, Bow92, Bow99b, Bra96, Bra91a, Bra92, Bra93, Bra91b, Bra90, Bra97, Bri93, Bri92, Bri90, Bro92a, Bro93a, Bro97a, Bro95a, Bro97b, Bro98a, Bro92b, Bro93b, Bro95c, Bro96, Bro98b, Bro90, Brü94, Bry92, Bry95, Bry96, Buc92, Buc94, Buc98a, Buc97, Buc98b, Bud94, Bud96, Bul94, Bul91, Bur92, But92, Cal98, Cal92a, Cal92b, Cam92, Cam93, Cam95, Cap90a, Cap90b, Cap95, Cap91, Car98, Car91, Cas98a, Cas98b, Caw95, Cer97, Cer98, Cha90, Cha91, Cha93, Cha95, Cha98, Cha96, Chi97, Chr95a, Chr99, Chr91, Chr93, Chr95b, Cla96, Cla90, Cla92, Cla97a, Cla91, Cla97b, Cod91, Cod92, Coh90, Col93, Con90, Con91, Con92a, Con92b, Con94b, Con96a, Con96b, Con97, Con98, Coo91a, Coo97, Coo90, Coo92a]. review [Coo93c, Coo93b, Coo95, Coo96, Coo98a, Coo98b, Coo91d, Coo92b, Coo98c, Coo98d, Cor91, Cor92, Cor95, Cor90, Cor93, Cor97, Cou94, Coz93b, Coz93a, Cra97, Cre99, Cro94, Cum91, Cur91, Cut93, Cut94a, Cut94b, Cut97, Cut98, Cut99, D’A97, Dal90, Dan90, Dan92, Dan93, Dan95, Dav96, Dav92a, Dav92b, Dav93, Daw91, Day99, DeV92, DeV91, DeV93, Dek98, Die93, Die97, Dol97, Don91, Dou91a, Dou91b, Dou95, Dre95, Dre97, Dub93, Dud92, Dun93, Dup92, Dur98, Eam97, Eat90, Ebn92, Edg98, Edm96, Edw95, Edw96a, Edw96b, Edw97, Egg94, Eis93, Ell94a, Ell92b, Ell93, Ell94b, End97, Eva92, Eva93a, Eva96, Eze90, Far96, FG98, Fel97, Fer90, Fer92, Fer95, Fie94b, Fie98, Fin98, Fin95, Fin93, Fis92, Fis98, Fit93b, Fit90, Fit95, Fit96]. review [Fit97, Fit98, Fle90, Fle92, Fle98, Flo92, FO92, For97, Fos98, Frä91, Fri92a, Fri90, Fri91a, Fri92b, Fri98, Fri91b, Fri96, Fro92a, Fro92b, Fro93, Ful95, Gal92a, GS98, Gal92b, Gam97, Gar91, Gar92, Gar90, Gei90, Gei94, Gen92, Gib91, Gib93, Gie96, Gie94, Gin90, Gis98, Gli91, Gli94, Gol92, Gol93, Gol95, Gol91a, Gol90, Gol91c, Gol91b, Gom96, Gom97, Goo90, Goo95, Gor91, Gor92a, Gor92b, Gor95a, Gor95b, Gor96, Gor97a, Gor94, Gor98, Gos90, Gra95, Gra97a, Gra91, Gra97b, Gra98b, Gra98a, Gre91, Gro90a, Gro90b, Gro98, Gru90, Gua98, Gui98a, Gui91, Gui93a, Gui93b, Gui95b, Gui95a, Gui98b, Gus98, Gus91, Gut90, Gut92, Gut93, Gut95a, Gut95b, Hår93a, Hår96a, Hår98a, Hac90a, Hac90b, Hac91a, Hac91b, Hac92a, Hac92b, Hac94a, Hac94b]. review [Hac98, Haf90, Haf92, Haf96, Haf97, Hal94, HD90, Hal93a, Hal96, Hal98, Hal90, Ham98a, Ham92a, Ham95, Ham98b, Han93, Har94, Har98b, Har97, Har91, Har93b, Har95a, Har96c, Har96b, Har92, Har95b, Has93, Hat90, Hea90, Hea91a, Hea91b, Hea92, Hea96, Hea98, Hea99a, Hea99b, Hec92, Hei99, Hei95a, Hei95b, Hei91, HV98, Her93, Her94, Her98, Hes99, Hev95, Hew94, Hig97a, Hig90, Hig91, Hig92, Hig97b, Hig98, Hig96, Hil91a, Hil97, Hil91b, Hil92a, Hil92b, Hil95, Hin94, Hir92, Hir96, Hoc98, Hod93, Hof91, Hof92, Hol90, Hol94, Hol95, Hol97a, Hol93, Hol91, Hol97b, Hom94, Hon97, Hon98, Hor94, Hou97, Hou90, How98, How93, Hoy92, yH98, Huf92, Hug90, Hug96a, Hug93b, Hug95, Hug96b, HC98, Hun96, Hur98, Hur90, Hur99, Hut92, Hyd90, Hyd99, Ile96]. review [Ing91, Ing94, Isr90, Isr91, Isr92a, Isr92b, Isr94a, Jac91, Jac93a, Jac95b, Jac95a, Jac96, Jac97, Jak98, Jam90, Jan91, Jar90, Jer90a, Jer93, Jer94b, Jer94a, Jer97, Jes97, Joh91, Joh92, Joh97a, Joh99a, Jos95a, Jos96, Jud90, Kah96, Kai98, Kal95, Kal98, Kap92, Kat92, Kat94b, Kat94a, Kat97a, Kay92, Kay93, Kay94, Kay91, Kea98, Kei94, Kel93a, Kel93b, Kel96, Kel99, Kel93c, Kel93d, Kel95, Ken92, Kid90, Kid91, Kid92, Kid94, Kid96, Kie90a, Kie90b, Kie92, Kie93, Kin99, Kip97, Kir95, Kle91, Kle95, Kle99, Kli90, Kli91, Kli97a, Kni95, Koc98, Koh96, Kon97, Kon98, Kop91, Kra93, Kra95, Kra98, Kra90, Kra94c, Kri92, Kri97, Kri98, Kry92a, Kry92b, Kry92c, Kuc92, Kuc94, Kyn99, LaF91, LaL90, Lai91, Lai92, Lai93, Lai96b, Lai96a, Lai97]. review [Lam93, Lan94a, Lan91a, Lan93, Lan98a, Lan99, Lan90a, Lan91b, Lan94b, Lan95a, Lan98b, Lan90b, Lan95b, LF93, Lar91, Lar98, Lau98a, Lau98b, Lav93, Lay97, Lay91, Lay92a, Lay93, Lay96, LeC94, LeC99, Lea92, Lea98, Led91, Led93, Lee96, Lee97, Len91, Leo92, Leo96, Ler98, Les90a, Les90b, Les93, Les94, Les99, Leu98, Lev91a, Lev91b, Lev91c, Lev97a, Lew96, Lew90, Lew93, Lew97, Lie90, Lie91, Lie93, Lie98, Lig97a, Lin97a, Lin98, Lin94a, Lin93b, Lin93a, Lin94b, Lin97b, Lin91, Lip97, Lip98, Liu97, Liv92, Log91, Log92, Log95, Lon93a, Lon93b, Lon95b, Lon95a, Lon96b, Lon96a, Lon97, Low93a, Low96, Low97, Low91, Low92, Low93b, Lub91a, Lue90, Lun90a, Lun92, Lun90b, Lup91, Lüt98, Lyn98, Mac93b, Mac96, Mac92, Mag94, Mag95, Mah90, Mah97, Mai96]. review [Mai98, Mal91, Mal98, Mar90, Mar96, Mar93a, Mar91, Mat90, Mau91, Mau93, May91, May92, May94, Maz90a, Maz90b, Maz93, Maz94, McC93, McG97b, McM94a, McM94b, McN90, McN93, McS98, McW95, MM92, Mea98, Mei90, Mei93, Mei95, Mei96, Mei98a, Mei98b, Mei99b, Mel96, Mer97, Mer98, Mer95, Mer96, Meu98, Mic94, Mil94, Mil93, Mil97, Mil91a, Mil95, Mil91b, Min96, Min97, Min95c, Min95b, Mis90, Mis93, Mis97, Mit91, Mit92, Mit96, Mit97, MA97, Miy94, Moe94, Moh90, Moh94, Moh95, Moh97a, MM93, Mor91a, Mor90, Mor91b, Mor91c, ML98, Mor91d, Mor93, Mor97, Mou99, Moy94, Moy96, Mue91a, Mue90, Mue95, Mul96a, Mul90a, Mul94, Mul95, Mul96b, Myl97, Nas90, Nas91, Nea98, Neb90, Nee96, Nei98, Nel90, Nel92, Neu92, Neu93, Neu94, Neu96a, Neu96b]. review [New91, New94, Nob97, Nuw90, Nye90, Nye92, Nye93a, Nye93b, Nye95, Nye97, Nye98a, O’B91, O’D94a, O’D94b, O’D96a, O’D97, O’H96, O’H98, O’M90, O’M93, Och92a, Oga90, Old92, Old98, Oli90, Oli98, Ols97, Olw98, Omi92, Opi90, Ors97, Øve97, Owe91a, Pac95, Pan98b, Pan98a, Pan95, Paq97, Par94, Pas96, Pat95, Pay95, Per90, Per93, Pet92a, Pet96, Pet98, Pet90, Pet92b, Pet97, Pfa91, Pfa92, Pfa98, Pfa97, Phi91, Phi93, Pin90, Pin91, Pin92, Pis96, Pla96, Pol91, Por91, Por93, Por94, Por98, Pre97, Pri95, Pri98, Pry93, Pul92, Pur90, Pur97, Pye90, Pye92, Qui98, Qui94, Rab96, Rab98, Ran97, Ras90, Ras92a, Ras92b, Rau98, Ree98, Rei93a, Rei96, Rei91, Rei93b, Rei95b, Reu91, Reu93a, Reu95, Reu93b, Rey90a, Rey90b, Rey91, Rey94, Rey98]. review [Rhe94, Rhe98, Rho93, Rob90a, Rob92a, Rob93, Rob96a, Rob95, Rob90b, Rob92b, Rob94b, Rob94a, Roc94, Rol90, Rol95a, Rol98a, Rol93b, Rom97a, Rom97b, Ror91, Ros92, Ros93, Ros98, Ros94a, Ros94b, Rot92, Rot93, Rot94a, Rot97, Rot95, Rot99c, Rou96, Roz90, Roz98, Rüd92, Rue90, Rug96, Rum92, Rum97, Ryc92, Ryc93, Sal92a, Sal92b, Sal95, Sam99, San92, San94, San97, San90, Sch93a, Sch92a, Sch92b, SB95, Sch99b, Sch94b, Sch91, Sch94c, Sch94d, Sch95, Scr90, Scr91a, Scr91b, Scr96, Scr98a, Scr98b, See90, See95b, See95a, See98a, See98b, Seg94, Sei93, Sei96b, Sei96a, Sei98, Sel90, Sel91, Sel92, Sel97, Ser92, Sha90a, Sha96a, SS97, Sha95b, Sha95a, Sha91, Sha93, Sha95c, Sha94, Sha96b, She94a, She94b, She93, Shi97, Sim90, Sim94]. review [Sim95a, Sim95b, Sin93, Sin96, Sin97, Skr91, Skr93, Sla91, Sla92, Sla95, Sla97, Sla98, Slo98, Sma95, SD98, Smi90b, Smi90a, Smi91c, Smi91a, Smi91d, Smi92b, Smi92a, Smi93a, Smi93b, Smi94, Smi95a, Smi95b, Smi97a, Smi97b, Smi98, Smi99, Smu90, Smu93, Sok93, Sol98, Spe95, Spi95, Sta94, Sta92a, Sta96a, Sta97a, Sta97b, Sta90a, Sta90c, Sta90b, Sta92b, Ste92, Ste95, Ste98, Ste91a, Ste94b, Ste90, Ste96, Sti98, Sti91, Sto97, Str90a, Sul98a, Sul98b, Sum91, Sum93, Sum96, Sum97, tZS90, Sun97, Sut90, Svi93, Swe92, Swe93, Szo96, Tal94, Tan92, Tan93, Tas97, Ter94, The90, Tho91a, Tho93, Tho94, Tho97a, Tho97b, Tho98, Thr91, Til94, Til90, Til97, Tob92, Tod95, Tod96, Tod97, Tod99, Tra97a, Tra95, Tra97b, Tra98b, Tri90a, Tri91, Tri90b]. review [Tri93, Tri94, Tur95, Ube94, Ube99, Uss90, Uss93, Uss97, Van97, Vas98, Ver97, Vie90, Vin93, Voe90, Vol91a, Vol91c, Vol91b, Vol93, Vol94a, Vol94b, Vol98a, Vol98b, Wac94, Wal91, Wal92, Wal94, Wal99a, Wan96, War91, War92b, War92a, War93, War97, War98b, War98a, Was91, Was93, Was92, Way90, Web90a, Web90b, Web90c, Web96, Wei94, Wen92, Wen98, Wer92, Wer97, Wes90, Wes91, Wes93b, Wes95, Whi92b, Whi92a, Whi93, Whi98, Whi92c, Wig96, Wil92a, Wil95a, Wil93a, Wil95b, Wil95c, Wil92b, Wil93b, Win98, Win90, Woo92, Wor98, Wor91, Wor93a, Wor93b, Wor94, Wor95, Wos91, Wos92, Wos94, Wri94, Yat92, Yat96, Yea90, You93, You95, Zac98, Zak92, Zel97, Zem93, Zie92, dV95, vK92a, vL90, vR90, vL99]. Reviewer [Bow98]. Reviewing [Pre98]. Revisited [Bal99, Hon94, Lun92]. Revolution [Big94, Bra92, Coo90, Dal90, Dor91, Fit96, GA98, Gut92, HD90, Har96c, How93, Jos95a, Lon95a, Lon96b, Par92, Rob90b, Rol93a, San94, Ste94a, Tho98, Bra93, Coo96, Eam97, Jes97, Kas99, Kel93b, Len91, Web90a, Web90c]. Revolutionary [Ald96, Eat90, War92b]. Revolutions [Ste91a, Kah96, Lav93, Gib93]. Revolver [Bel98b]. Revue [Fer99]. Rey [Was92]. Rezun [Gra97a]. Rice [Car96]. Richard [Bla92a, Bow92, Bra91b, Chr91, Gen92, Hac91a, Hac98, Har97, Lai91, Lai97, Lan90b, Mit97, Mor91d, Mue90, Mue95, New94, Pry93, Sin96, Sok93, Tho97b, You93]. Richardson [Goo90]. Riches [Hea91a]. Richmond [Lie92, Mat92a]. Rick [Bra92, Lin97b]. Rickover [Hew94, Wal91]. Rider [War92a]. Ridges [Hal94]. Riding [Whi90]. Right [Kay93]. Riken [Col90]. Rikoon [Hur90]. Rims [Hal94]. Rings [Cut93]. Rip [Jac97, Mac92]. Rise [Ada95a, Bel97, Ber93, Bla90a, Bol90, Bud94, Cas98a, Dia99, Dou99, Dre97, Gor96, Hac93, HD90, Lub91a, Mer96, O’D96a, Och92b, Per93, Pur93b, Ros94b, Scr98a, Smi92b, Sti98, Vie90, Whi93, Zem93]. Rising [Ano99n, Cal92b]. Risk [And99a, Car97, Cut94a]. Risky [Nea98]. Risø [Hei99]. Risøs [Hei99]. Ritchey [Rot94a]. Rites [Olw98]. Ritter [Bro99b]. Rittman [Mer90]. Rivalry [Bur92]. Rivals [Wal92]. River [Col93, Kel99, Reu99a, Smi92b, Sca99]. Rivers [Rob93, Sum96]. Rivulet [Eve92]. Road [Bro95c, Hil95, Lai96a, Low97, Sin95, Bro95b, Lar98, Lar98, Rho93]. Roads [Han93, Vol94b]. Roadside [Whi92a]. Robbins [LaL90]. Robert [Ald96, Bar91a, Bas97, Ben96, Bul94, Cla96, Coh90, Col93, Dan93, Ebn92, Fit97, Gra91, Gui95b, Gut90, Har93b, Hec92, Hof91, Kay92, Kay94, Kel93b, Kel93d, Kie90a, Kni95, Kri92, Kuc92, Lan95b, Lea98, Les90b, Lew90, Mar91, Mat90, Maz90b, Maz93, Mil97, Mil91a, Min97, Neb90, Oli90, Pat95, Pis96, Rei91, Rey90b, Rey91, Rob92b, Rol90, Rot93, Rum97, San94, Sch94d, Scr98b, See98a, She94a, She93, Sim94, Sin97, Smi91a, Ter94, Tri91, Wes90, Far99, Hun95, Isr92b, Sha93]. Robin [Mul96b, War92a]. Robins [Ste94a, Cap91]. Robinson [Ano90f, Ano91h, Ano92g, Ano93g, Ano94g, Ano95g, Ano96f, Ano97g]. Robots [Bla95]. Rock [Ale99b, Ler98, Mah90]. Rockefeller [Bud94]. Rocket [Buc98b]. Rocketry [Neu90, vK92a]. Rockets [vK92a, vK92a]. Rockies [And99b]. Rockwell [Hew94]. Rod [Kry90, Vol93]. Roderick [Pye92]. Rodger [Low96]. Rodney [Ham98a]. Rods [KK92, Pos98b]. Roell [Lin91]. Roger [Bro97a, Cor93, Eat90, Ell93, Ham98b, Kei94, Web90a]. Rogge [Jac96]. Rolando [See95a]. Role [Bra92, Con99, Sim94]. Roles [Don94]. Rolf [Blu92]. Rolle [Neu96a]. Roller [TJ96]. Rolt [Lev97a]. Rolv [Qui98]. Roman [Lan94a]. Romans [Sim95a]. Romantic [Kas99]. Rommelspacher [Web96]. Ron [Cro94, Sch92b, Adr94]. Ronald [Chr93, Gom97, Kel95, Rob92a, Sim90]. rond [Hom94]. Rondo [Bar90]. Ronell [MM93]. Ronner [Lay97]. Roos [Bal97]. Root [Rei91]. Root-Bernstein [Rei91]. Roots [Ell94b, Her99b, JJ98]. Rosalind [Moh97a, Sla91]. Rose [Ano90i, Ano91l, Ano92j, Ano93i, Ano94j, Ano95j, Ano96h, Ano97j]. Rosen [Hof92, Qui95]. Rosenberg [Har93b, Mil97, Yat96]. Rosenzweig [Skr91]. Roslynn [Bas96]. Ross [Cum91, Rey94, Wil95a, Wil93a]. Rosslyn [Ano94i]. Rossum [LF93]. Rotary [Tri90b]. Rotary-Wing [Tri90b]. Rotenberg [She94a]. Roth [Gor95a]. Rothenberg [Isr94a, Ber95a, Cha95]. Rothman [Coo97, Hal94]. Rouny [Hol97a]. Routes [Div95]. Rowley [Hol97b]. Roy [Gus91, Pye90, Skr91]. Royal [Ken92, Kid91, Lar91, Pet92b]. Rubber [Nas91, Nuw90]. Ruby [All96c]. Rudbeck [Bjö97]. Rudbecks [Bjö97]. Rudder [Dot99]. Rudi [Cha91, McW95]. Rudolf [Dav92a]. Rudolph [FG98]. Ruhr [Web96]. Ruhrbergbaus [Mer97]. Ruhrgebiet [Web96]. Rule [Ada98]. Run [Con96b]. Rural [Bor97, Eva93a, Eve92, KP96, Lan99, RK97]. Rush [Sha96b]. Russell [Tra98b]. Russia [Bro95a, Coo98c, Gra97b, Gra98b, McC93, Rho93, Wei94, Wes95, Zie92]. Russian [Gra97b, Hal90]. Russo [Dol97]. Rüstungspolitik [Neu96a]. Rutland [Wes93a]. Ruza [Rot95]. Rybczynski [Mor90]. Rydell [Rum97]. Ryoshin [Coo90].

S [Ada93, And94, Buc94, Dav96, For93, Ful98, Gom97, Hal98, Har95b, Jac91, Jer90a, Kel93b, Kra94c, Lai91, Mac96, Maz94, Mer96, Mil93, Mor91b, Mul96b, Nye98a, Roc94, Roz98, Svi93, Vol94a, Wes91, Wil92b, WN98, Ter94]. Sabbagh [Zak92]. Sabbatini [Sel92]. Saberes [Gua98]. Sacramento [Hof91, Ano90h]. Sacred [Chi97]. Sadahiko [Has93]. Safer [Sha94]. Safety [McE95, Moh95, Sei96a, Sim90, Whi98, Whi98]. Saga [Lan90b, Wor93b]. Sagan [Moh95]. Sahara [Kri97]. Saint [Les90a, Wil93b]. Sala [D’A97]. Sales [Gol97]. Salesman [Lai96a]. Saliga [Pet92a]. Salinenwesen [Mul90a]. Sally [Fie94b, Fit96, Por93, Pye92, Hac90c]. Salmon [O’B91]. Salmon-Canning [O’B91]. Sam [Neu92]. Samaddar [Pfa98]. samfundets [Hei99]. Sampler [Rob96b]. Samuel [Hac94a, Lee97, Pfa97, Ano90i, Ano91l, Ano92j, Ano93i, Ano94j, Ano95j, Ano96h, Ano97j, Buc99, Con95, Mil91b]. San [Vie90]. Sanderson [Cur91]. Sandia [Wes91]. Sanford [Fri91b, Hur90]. Sangallo [Lon95b]. Sangwan [Hea92]. Sanitarians [Ham92a]. São [Pri98]. Sapolsky [Tan92]. Sarah [Dre97, Fri96]. Sassower [Mit91]. Satellite [Bar92b, Cod92, Bul94]. Satisfaction [Wil92a]. Satpal [Hea92]. Saving [Gos90]. Savings [CK98]. savoir [Frä91]. Sawislak [Smi97a]. Saxby [Yat92]. Saxon [BA90, Wal95]. Say [Lev96]. Scaglia [Lon93a]. Scale [Jos95b, San94, Wes93b, Scr91a]. Scamehorn [LeC94]. Scandinavian [Hår93a]. Schaaf [Bro98b]. Schabel [Neu96a]. Schachterle [Kel96]. Schaffer [Bro90]. Scharff [Cor92]. Schedules [She94a]. Schedvin [Jar90]. scheikunde [Hom94]. Schemes [Pfa90]. Scheppler [Lew90]. Schiffer [Edw95, Hug96b, Mei93]. Schivelbusch [Opi90]. Schizophrenia [MM93]. Schlieffen [Gui93b]. Schlumberger [Con96b]. Schmandt [Phi93]. Schmandt-Besserat [Phi93]. Schnitter [Jac95a, Reu93a]. Schoenfeld [Hig97b]. Schofield [Rot93]. Scholars [Hac92a, Lon95a]. School [Low93a, Och92b]. Schorzman [Mil95]. Schroeder [Sha95b]. Schrover [Haf97]. Schubert [Mer95]. Schüler [Wen92]. Schultz [Ros93]. Schürmann [Lon93b]. Schurr [Neu92]. Schwantes [Qui94]. Schwarz [Hal96, Leo92, Mic94]. Schweiz [Reu93a]. Science [Ada91, Ada97, Ada98, Alb97, Bal97, Bar91a, Ber95a, Bix99, Bre99, Bro92a, Bur92, Cha99b, Cha96, Con98, Cow96, Coz93b, Cro94, Cut94a, Dan95, Dur98, Fin95, Gei94, Ger96, Gor98, Ham91, Ham98b, Har92, Hea90, HV98, Isr92a, Jar90, Kay92, Kle99, Koc93, Koh96, Kön96, Kra93, Kri99, Lau95, Les90b, Les94, Lev91c, Lew96, Low91, Mac96, Mat99a, Mer96, Mit91, Mit92, ML98, Mul95, Nye90, Nye93b, Pal93, Pan98b, Pan98a, Pla96, Por93, Pye90, Pye92, Ree98, Rob94a, Rol95b, Rot93, See93, Sel97, Sha90b, Sin99a, Sla92, Smi93a, Smu94, Sol98, Sta97b, Ste91a, Ste94c, Tei99, Tob92, War92b, Wes93b, WH96, Wil93a, Wil93b, Wri96, Yu99, vk92b, Lev91b, Bra95, Gam95, Lie93, Bjö97, Cal92b, Con97, Dan92, Fit97, Hol91, Hug93b, Kry92a]. Science [Lun92, Maz93, Och92a, Pri95, Ree98, Aik98, Ald96, Bal99, Bir90, Bro95a, Con96b, Coz93a, Goo90, Gra97a, Hac91b, Hac94b, Hea92, Hig96, yH98, Hun96, Lar91, Lay92a, Leu98, Lin98, Lon96a, Low96, Mar96, Mau93, Mor91c, Mul95, Mul96b, O’H98, Pry93, Ras92b, Sim95b, Sin93, Ste96, tZS90, Tan92]. Science-Based [Kön96]. Science-Technology [Rob94a]. Sciences [Ada96, Dan95, Ful95, Hac94b, Lar91, Rhe94, Lav93, Sim95b, Kay93, Sim95b, Lan91b]. Scientia [Pry93]. Scientific [Ada98, Bed90b, Bed91, Big94, Daw91, Eam97, Fit98, Goo95, Hei91, Hil91a, Hin94, Jar90, Jes97, Kel93b, Kid91, Kli97a, Lon96b, Nel92, Nye93b, Pye90, Pye92, Rei91, Reu99a, Rot99b, San94, Sha95a, Smi90c, Sta92a, Tra97a, War91, War92b, Low93b]. Scientification [Coo95]. scientifique [Lan91b]. Scientist [Bas96, Dou99]. Scientists [Bad91, Her93, Kay92, Lie90, Wan96]. scienze [War93]. Sclove [Mit97]. Scope [Scr91a]. Scotland [Kid91, Kid91]. Scott [Moh95, Rei91, Sha95b, Smi97b]. Scranton [Liv92]. Scrapheap [Jer93]. Screen [Gom96]. Sculle [Jac95b]. SDI [Sei93, Smi93a]. Sea [Bed90c, Cod92, Gar91, Gar92, Hof91]. Seaborg [Maz94]. Seafaring [Roz98]. Seaman [Gin90]. Seamless [Kun95]. Search [Gal92b, Jud99, Oli90, Rom97b]. Seas [Gib99]. secoli [Man99]. Second [Bil97, Lan99, Lea92, Mor91b]. Secrecy [Rei95b, Rol92]. Secret [Hea96, Hig98]. Secrets [Lon96a]. Sectional [Rob96a]. Sector [Zem93]. Securing [Ser92]. Security [War92b]. See [Ols98, Sul98b]. Seeber [Gom97]. Seeds [Por94]. Seeking [Smi91b]. Seelye [Rob93]. Segal [Dou91b, Sla95]. Segrè [Gol95, Gol95]. Seidel [Kri92]. sein [Fin98]. seit [Wen92]. Seizukyoiku [Has93]. Seizure [She94a]. Selected [Fit97, Bed90b]. Selection [McB92, Vin94]. Selections [Pry93]. Selective [Ber95a, Cha90, Jac93a]. Self [Moo98]. Self-Sufficiency [Moo98]. Selling [Kra95, Slo98]. Semyonovich [Kry92b]. Sendzimir [Hei95b, Hei95b]. Sense [Rom97a, Wor98]. September [Hau90]. ser [Les94]. Serge [Lin94b]. Sericulture [MS92]. Series [Har96b, Gra98a]. Servant [Ben91c]. Servants [Hoy92]. Service [Gol97, Gre95a, Gut95a, Kel93c, Sin99c, Jac93b]. Servitude [Bin93]. ses [Lin94b]. Settings [Bix99]. Settlement [Ben92b]. Setzmaschinenentwicklung [Zac99]. Seurat [Ada91]. Seven [Cor97, Gib99]. Seventeenth [Coo91a, Coo98b, Low93b, Lüt98, Ste96]. Seventeenth-Century [Coo91a, Ste96]. Seventy [Gui91]. Seventy-Four [Gui91]. Sewermen [Fro92a]. Sewers [Fro92a, Ham92b, Sul98a]. Sex [Old98]. Sexual [Ble91, Hor97b]. Seymour [Ful95]. Sforza [Lon97]. Shadow [Tod95]. Shadows [Hew94]. Shallat [Mer96]. Shallow [Kub90]. Shallow-Draft [Kub90]. Shanghai [Jam90]. Shapes [Pur99]. Shapin [Mis97]. Shaping [Edw96a, Har94, Hey98, McM94b, Rei93b, Sch92b, Slo95, Smi98, Vin94, Wil92b, Gie94, Jar90, Mac93b]. Shared [Ber94b]. Sharon [Sta94]. Sharp [Car99a]. Shattering [Hut92]. Shaun [Bla97c]. Shaw [Dan90, Rob92a]. Shear [And99b]. Shearer [Coo98c]. Sheeler [Sta92b]. Sheffield [Twe92]. Sheldon [Ham98b]. Sheridan [Sin93]. Shield [Ber90, Fri91b]. Shields [Abb96]. Shifting [Bes95]. Shifts [HS96]. Shines [Bin91]. Shinso [Has93]. Ship [Gui91, Rey90b, Ron99]. Shipboard [DeV90]. Shipbuilders [Kuc92]. Shipbuilding [Gib99]. Shipping [HC98]. Ships [Gib99, Kei94]. Shipwrights [Kyn99]. Shirt [Sch93b]. Shirt-Pocket [Sch93b]. Shneller [Min97]. Shoe [Cum91, Gre91]. Shoemaking [Tho91b]. Shop [Gra95]. Shops [Hou90]. Shore [Ste90]. Short [Bro95a, Gen92, Kry92c, Led93, Ste95, Tri94, Wer92]. Shortland [Sim95b]. Shoshana [Sta90c]. SHOT [See95c]. Shoulders [LMO97a]. Shovel [Smi94]. Show [Bro98a]. Shropshire [Buc98a]. Shrum [Hig96]. Shrunk [Eve92]. Shuttle [Cap90b]. Siberian [Bal99, Rho93]. Sibylle [Pul92]. Sicherheit [Lun90a]. Side [Gre94]. Sidney [Adr94, Tra96]. Siècle [Lan90a, Lev91a, But92, Lav93, vR90]. siècles [Dre95, Smi92a, Hea99a]. Siege [Tar95]. Sieges [Mul96b]. Siemens [Mai96]. Siglos [Gli91]. Sign [War98a]. Signal [Sla98]. Signals [Mat92a]. Significance [Gam92, Yat92]. Signs [Hey98]. Sikorsky [Hal90]. Silicon [Neb90]. Silk [Gro90a]. Silver [Gom96, Coo91a]. Silverman [Bas97, Lev91a]. Silvio [Dan92, Lan93, Ste96]. Simon [Dan95, Hyd90, Jam90, Wor95]. Simple [Hyd90]. since [Big94, Buc92, Coz93b, Gui95a, Hal94, Har92, Nel92, Sta97b]. Singing [Per99b]. Single [Lai96c]. Singleton [Jer93]. Sink [Rom97b]. Sir [Bas92b, Hig91]. Site [And94]. Sixteenth [Buc92, Sel90]. Sixteenth-Century [Sel90]. Sizemore [Wor95]. Skaggs [Sha96b]. Ski [Rei99a]. Ski-Dogs [Rei99a]. Skill [Adl90, Ryd98]. Skilled [Bla96b, Len91]. Skin [Les90a]. Skinner [Cap95, Cap93, Low97]. Skulking [Was93]. Skulptur [dV95]. Sky [Mor91d, Mor93, Pet92a, Rot92, Rot94b, Sch92a, Slo96, BS94]. Skylight [Vol91a]. Skyscraper [Dre97, Mor91b, Zak92]. Skyscrapers [Pet92a]. Slack [Dav93]. Slade [Lev91c]. Sladovich [Bry92]. Slater [Con95, Neb90]. Slave [Scr96]. Slavery [Hen99, Paq97, Lee96]. Slaves [Don94]. Slick [Cas98b]. Slide [Bla96a]. Sloan [Mil95]. Sloops [Dud92]. Sloss [Gor96]. Slotten [Koh96]. Slyck [Ber97a]. Small [Bel98b, Bro98a, Gor95a, Mil91a, WH96, Wes95]. Small-Arms [Bel98b]. Small-Town [Bro98a]. Smaller [Rob96b]. Smart [Sta90c]. Smelting [GK93]. Smil [Wil95b]. Smiles [Buc99]. Smith [Bir90, Eis93, Por91, Bar91a, Cha98, Coz93b, Gol92, Her93, Hug93a, Jam90, Kat94a, Kei94, Mis90]. Smithsonian [Chr95a, Mil95, Roz96, Sma95, Mol91]. Smoke [Mat92a]. Smoldering [Smi97a]. Smulyan [Kra95]. Snapper [Hil91a]. Snapshots [Fit97]. Snelders [Hom94]. Snowmobiling [Rei99a]. sobre [MA97]. Sochor [Lew93]. Social [Abb96, Ald96, Arn97, Bar92b, Big91, Bla92a, Bry92, Chr95b, Cla91, Coo91c, Coz93a, Cut98, Hår96a, Hac94c, Ham98b, Har94, Hir92, KP96, Kli97b, Lai96b, Leo92, Ler97, Ler98, Mar93b, Mis97, Per90, Pri98, Ras92b, Ros94b, Sam99, Slo95, Sto97, Str90b, Tal94, Vol94a, Wil99b]. Socialism [Hun96]. Socialist [Chr93]. Socializing [Scr98a]. Societal [Fea99]. Societies [Nye93b, Lin94a, Vas98, Wig96]. Society [Bow99b, Buc94, Coo98c, Ell94b, Gie94, Her99a, Hig96, Jac93a, Jac97, Kid90, Lup91, Mah90, Pan98a, Pfa90, Por98, Pur99, Rhe94, Sch92b, Scr94, Sha93, Sla91, Tho91a, Vol91b, Hei99, Fel99, Pur93a, Tho90, Cap91, Joh92, Kon97, Pfa98, Cha91]. Sociology [DeV92, Joe90, McG95]. Sociotechnical [Gie94, Mei96]. Software [Fie94a, Ras99]. Sogner [Qui98]. Sola [Bar92a]. Solar [Har92]. Soldiers [Hac92a]. Solidarity [Gut95b]. Solingen [Wan96]. Solomon [Lun92]. Solving [Rei91]. som [Har99a]. Some [BA90, Bla99a, Bow98, Cut97, RK97]. Sonata [Ozd98]. Sonenscher [Haf92]. Sonnemann [Blu92]. Sons [Tho97b]. Soo [Rey90b]. Soot [See95a]. Sophia [Ter94]. Sound [Hor99, Per91, Tho94, Tho98, Vin97, Tho97a]. Sounds [Chi97]. Source [IR91]. Sources [Goo90, Isr90, Tod97]. South [Cra97, Dan93, Kri97, Lin94a, Mue91a, Sha95c, CC95, Col99, Hen99, Bri93]. Southeast [Wig96]. Southern [Bau93, Bla90b, Bro97b, Gra98b, Gro90b, Lup91, Set99]. Southwest [Win90, Eva92, Pet97, Sto97]. Southwestern [Rob96a]. souverainété [Hea99a]. Soviet [Ada95b, Bro95a, Bul94, Jos96, Ran97, Ras92b, Jos95b]. sozialer [Tod99]. Space [Ale99a, Bar91a, Bat90, Bro97a, Day99, Dol97, Fri91b, Gru90, Hac94b, Hor97b, Joh99a, Kra93, Log92, Log95, Mac92, Nee96, Swe93, Wil90, Wor98, Kay99, vK92a, Mac92, Huf92]. Spacefaring [Swe93]. Spaceflight [Neu90]. Spaceship [Cro94]. Spain [Gui98a, Hol91]. spanabhebender [All96c]. Spanish [Ile96, Sel90]. Sparks [Hoo99]. Speaking [Ile96]. Spears [Lai96a]. Special [Qui95]. Specialists [Gus91]. Specialty [SS97]. Spectrum [Ait94]. Speech [MM93]. Spence [New91]. Spies [Hea96]. Spiller [Bel98b]. Spinardi [Hev95]. Spinners [Jer90a]. Spinning [tZS90]. Spirit [Coo95, Rei95a, War92a, Wes93a]. Spiritual [Hig97a]. Sponsored [Ser96]. Spot [Adr94]. Spotless [Lan98a]. Spreading [Lai97]. Sputnik [Bul94]. Sputniks [Mac92]. Spying [Hea96]. Sr [Win98]. Sri [Pfa90]. SS [All96d]. SST [Rei95a]. St [Hur98, Jos95a, Kry90, Rei95a, Was92]. staatlichen [Lon93b]. Stadt [Tod99]. Stafford [Pla96]. Stage [Tho98]. Stahlindustrie [All96a]. Staircase [Sha94]. Staiti [Mil91b]. Stalin [Coo98c, Hor94, McC93, Jos99, Dal90]. Stalinism [Jos96]. Stalking [Hig97b]. Stand [Leo92, LMO97a]. Standard [Nas90]. Standardizing [Coo93a]. Stanford [See99a, Gei94]. Stanley [Hug93a, Lai92, Ros93]. Starkey [Eam97]. Starlight [Aga98]. Stars [Gom97, Hac90b, Lan90b]. State [Coo98c, Hac93, Hay90, HP91, Jos95a, Kai96, Mea98, Mer90, Moy94, Sal92a, Wan96, Fro92b]. States [Ber95a, Ber95c, Cla96, Coo93b, Fit96, Gro98, Han93, Hoy92, Mac92, Reu99a, Rob92a, Sei96b, Sha90b, Skr91, Str90a, Fan91, Hey98, KP96, Nic99a, Ols98, Pur93b, Sch94a, Zei95, Bad94, Ber94b, Cum91, Hoc98, Slo98, Van97, Wal94, Kyn99]. Statesman [Dan92, Hac98]. Station [Hir96, Jac95b, Joh99a, Log92, Rol94, Isr94b]. Stationary [Bow92]. Stealing [Hin94]. Steam [Bow92, Dud92, Sch99b, Stu94]. Steamboat [Wor93b]. Steamers [Kub90]. Stech [Way90]. Steel [Ald99, Egg94, Gra98b, Jar95, Kal98, Lay96, Lew97, Mar90, Mul94, Pay95, See90, Twe92, Kle95, Reu93b, Hei95b]. Steichen [Bla99b]. Stein [Gut90]. Steinberg [Bla96a, Smi93b]. Steiner [Hod93]. Steinmetz [Chr93]. Steinway [Tho97b]. Stellung [Lin93b]. Sten [Frä91]. Stennis [Joh99a]. Stephanie [Mor91b]. Stephen [Bro93b, Con96a, Dun93, Hår96a, Hil95, Hor94, Les90a, Mar91, Mor93, Nel92, Pet90, Pfa92, Til94, Vol91b, Yat92]. Stereoscope [Sil93]. Stern [Bla96b]. Steve [Gro90a, Kap92]. Steven [Chr95a, Coz93a, Hei91, Mis97, Rho93, Wil95c]. Stevens [Hoc98, She93]. Stevi [Bla97c]. Stimson [Bed90c]. Stine [Qui95, Ham95, Lew96]. Stockholm [Arr94]. Stockyards [Yea90]. Stokes [Kra90, Tra95]. Stone [Wal99b, Kel93c]. Stopes [Neu98]. Stopped [Mis93]. Storia [War93]. Storm [New94]. stormaktstiden [Bjö97]. Storrow [For93]. Story [Gus98, Lai91, McS98, Sch92a, Cap90a, Hig90, Mei99b]. Strahan [Moy96]. Strange [Rus99, Wil93a]. Strangelove [Bas96]. Stranger [O’M90]. Strasser [Wil92a]. Strategic [Fri91b, McB92, McB97, Sch92a]. Strategy [Bir90]. Stratosphere [Daw91]. Stream [Eve92, Mer96]. streams [Kai98]. Street [Bro97b, Gra91]. Streetcar [Smi97b]. Streeter [Slo98]. Strike [Gro90a, Mor91d]. Striking [Hon98]. stroke [Nah97]. Strom [Sta94]. Structuration [Bor99]. Structure [Ler97, Moe94, San94, Tri91, Web90a]. Structures [Wan96, Mer96]. Struggle [Gut95b, Hil95]. strumenti [Man99, War93]. STS [Cut95]. Stuart [Big95a, Gei94]. Studebaker [Fos98]. Studien [Lon93b]. Studier [Kai98]. Studies [Bac90, Chr99, Con98, Cut94a, Flo92, Gie94, Has93, Her99a, Mor99, Pan98a, Pin92, Rhe94, Sha94, Sim90, Kai98, Ben91b, Bla97c, How98, Gra98a, San97]. Studio [Smi99, Tho98]. Study [Bar91a, BP91, Div94, DB94, For93, Hur90, Law92, Mal98, May92, Mil91a, Och92b, Par92, Por91, Pul92, Smi90b, Tra97a, Wal99a, Wil99a, Wor93a]. Stuff [Con99, Scr00]. Stuhlinger [Neu94]. Stunde [LF93]. Style [Sto97, Whi90, Lay92a, Lev91a]. Styles [Hey98]. Styrelsen [Car99b]. Subject [Ano91k, Ano92k, Ano93j, Ano94k, Ano95k, Ano97k, Ano99o, Ano99p, Coo93a, LCH^{+}96c]. Sublime [Sla97, Mei95]. Subrata [Pre97]. Substitution [Day92]. Suburban [Ste90]. Suburbia [Ros92]. Success [Car97, Jag99, Kai96, Kun95]. Suède [Frä91]. Suellen [Mel96]. Sufficiency [Moo98]. Sugar [Alv99, Gut90, Alv99, Paq97]. Sullivan [Hig90]. Sumida [Owe91a]. Summary [Hil91b]. Summer [Pos98a]. Summerton [Har94, Min96]. Sumter [Fer90]. Sun [Bin91, Hal93b, Har92]. Sunbelt [Ebn92]. Suomen [Ahv90]. Super [Sel90]. Superconductivity [Cha99a]. Superintendency [Jan91]. Superior [Dub93, Mou99, Rot97]. Superiority [Mor93, Slo96]. Superlative [Wor91]. Superweapon [Hac90b]. suppl [Fit97]. Supplement [Ano92l, Ano93m, Ano94l, Ano95o, Ano95l, Ano96i, Ano97l, Ano99u]. Supply [Gal92b, Sch94c, Lan94a]. Supplying [Ham98a]. Support [Mel99]. Suppressing [Cla93]. Supremacy [Jer97, Owe91a]. Surgery [Koc93]. Surrender [Fro93]. Surveillance [Hea96]. Survey [Bed90c, Bur92, Cut97, Koh96]. Surveys [Kli97b]. Survival [Moe94]. Survivals [Gib91]. Surviving [Bed90c, McN93]. Susan [Bud96, Coo91d, Gra98b, Kra95, Lie93, Maz93, Rot95, Wil92a]. Suspension [Smi90a]. Sustainable [Jud99]. Suzuki [Coo98b]. svensk [Kai98, Bjö97]. Svenska [Har99a]. Swann [Lie90]. Swanson [Whi92c]. Sweat [See95a]. Sweatshops [Ale99b]. Sweden [Ano93l, Bau99, Lar91, Ols97, O’D94b, Rey90a]. Swedish [Kai98, Lar91, Car99b, Kai96, Ols98, Ryd98]. Swiderski [Har97]. Sword [Ber90, Wal95]. Sylvia [Swe93, Vin93]. Symbol [BS93, Jag99]. Symbolism [Kri99]. Symposium [Ber95b, CS90, Hau90, Hor97a, Lup91]. Synthesis [Nye98b, Smi91b]. Synthetic [Nas91, Roc94, Str91, Str90a]. System [Bla90b, Cam95, Col99, Gre95a, Gre95b, Gro98, HS96, Lai97, Lub91a, Mue91a, Rob90b, Sch94c, Smi90b, Zem93, Bar92b, Har94, Hea96, Sin99c]. Systems [Gie96, Ham92b, Joh97b, Lee97, Lev97b, Scr94, Fis92, Hou97, Min96, Rau98]. Szilard [Hew94]. Szostak [Bra92, Lin97b].

T [Gor97a, Kay91, Kid91, Maz94, Nas91, O’M90]. T. [Kra96]. Tabbi [Sla97]. Taboo [Gar99]. Tacit [Hår96a]. Tactics [Bac98, HV99, Was93]. Tad [Hei95b]. Tadaaki [Lin93b]. Taiwan [Yu99]. Takeoff [Moo98]. Taking [Sta96b, Cor92, Hug96b]. Talbot [Bro98b]. Tale [Dot99, Mis93]. talets [Har99a]. Talking [Pur90, Bix99]. Tallapoosa [Sum96]. Tamara [Way90]. Tanks [Edg98]. Tapestry [Liv92]. Tapia [Gli91]. Target [Scr00]. Tarr [Bol90, Qui95, Qui95, Rom97b]. Tartakow [Ano90i, Ano91l, Ano92j, Ano93i, Ano94j, Ano95j, Ano96h, Ano97j]. tarv [Hei99]. Task [Moh97b]. Taste [Gut93]. Tastes [Skr93]. Tata [Kal98]. Tatarewicz [Huf92]. Taylor [Big96, Min99, Big94]. Taylored [Big96]. Taylorism [Nel92]. Teaching [Cut97, ML98]. Teaford [O’D96a]. Tech [Szo96]. Technical [Bix99, Bra91a, Cha98, Edw97, Hac93, Hei91, Hil91a, Hoc98, Kyn99, Low93b, Min96, Old92, Old97, Rau98, Ryd98, Sch94a, See95b, Slo96, SD98, Sta92a, Car99b, Sta97b, Vol91b]. Technicalities [Pos99]. Techniek [vL90]. Technik [All96b, Dek98, Bro99a, Dek98, Lun90a, Web90c].

Technikdebatte [Wen92]. Technikgeschichte [All99b]. Technikkritik [Wen92]. Technikwissenschaften [Buc97, Blu92]. technique [Hal93a]. Techniques [Hei95a, Mel99, Lan91b, Smi92a, Lev91b, Kat92]. Technische [Bra90, Bra90, Gis98]. technischem [Tod99]. Technischen [Gis98, Lon93b]. Techno [Sla99]. Techno-Poetics [Sla99]. Technocracy [Ben92b, Bix98]. Technohistory [Fle98]. Technological [Ahv90, Arn97, CK98, Cha91, Cha99b, Coo98b, Coo91d, Dot99, Fan91, Fit90, For93, Gro93, Haf97, Hea91a, Hur90, Kai96, KP96, Kra94a, LeC99, Ler97, Log92, McG97a, Mei95, Mor91a, Neu92, Par92, Per90, Pin90, Sel99, Str90b, Tho91b, Tod95, Tra95, Wal95, Wen98, Wil95a, WN98, Woo92, Wos92, Edw95, Jam90, Lin97b, Ols97, Rob94a, Scr98b]. Technologie [Die93, Lev91b, Lin93b, Hea99a]. Technologies [Jos95b, Oga90, Pfa92, Ras99, Til94, Ran97, Bar92a, Sch92b]. Technology [Abb96, Ada95b, Ada97, Ald99, All96d, Alv99, Ano93k, Bac90, BA90, Bar91a, Bau99, Bec91, Ben96, Ber95a, Ber95b, Ber98, Bes95, Big91, Blo90, Bor99, Bow99b, Buc91, Car96, Car91, Caw95, Cer98, Cha95, Chi97, Con92c, Coo98d, Cow93, Cow96, Coz93a, Cut97, Cut98, Dav92b, Dic92, Ell93, Ell94b, Fer92, Fin95, Fis92, Fit91, Fit97, Fle98, Fri91b, Fri99, Gam95, Ger96, Gie94, Gol91a, Goo90, GK93, Gra97a, Gre95b, Gre99, Gui98a, Gui95b, Gui95a, Hac91b, Hac94c, Hal93b, HK99, Hay96, Hea90, Hea92, HV98, HV99, Her99a, Hey98, Hig97a, Hig96, Hil99, Hor97a, Hor99, yH98, Huf92, HC98, Hun96, Ile96, Ing99, Jac97, Jef95, Jer97, Joe90, JM97, Joh97b, Kal95, Kat97b, Ker99, Kli97b, Kni95, Koc98, KRT90, Kri99, Kun95, Lau95]. Technology [Law91, Lay93, Leo92, LMO97b, Leu98, LM91, Lev96, Lev91c, Lie93, Lig97a, Lig97b, Low96, LAC^{+}90, LAC^{+}91, LAC^{+}92, LAC^{+}93, LAC^{+}94, LCH^{+}95, LCH^{+}96b, LCH^{+}97, Low99a, Low99b, Lue90, Lup91, Mag94, Mar93b, Mar91, Mat99c, Mau91, May92, May94, McB92, McB97, McN93, Min95a, Moh94, Moh97a, MM93, Mor91c, Mue99, Mul95, Mul96b, Myl91, Nea98, Neu90, Neu98, Nor90, Nye93a, Nye98b, Old98, Owe91a, Pan98a, Par97, Pas96, Pay95, Per99a, Per90, Per99b, Pfa90, Pfa91, Pfa98, Phi99, Pry93, Pur93b, Pur99, Qui95, Rab98, Rei95a, RK97, Rob94a, Rob96b, Rol92, Rol93a, Rol95b, Rol97, Rol93b, Rot95, Roz90, San94, Sch93a, Scr90, Scr91c, Scr95, See95c, See99b, Set99, Shi97, Sin96, Sin99a, Sla91, Sla95, Sla97, Slo95, Smi91b, Smu94, Sta97b, Ste94b, ST98]. Technology [Sti98, Sto91, Str90a, tZS90, Svi93, TJ96, Tap99, Tei99, Tho90, Ube94, Van96, Vas98, Vin94, Voe90, Vol91a, War97, Was93, Wen98, Wes95, Whi98, Wil93a, Yat96, Yu99, All99b, Has93, Cla90, Fit97, Gol91a, Mat99a, Rot94b, Bac98, Bal97, Bec91, Ben92c, Ber90, Cha93, Coh90, Cut94a, Cut98, Edw96a, Ell94a, Fro93, Har97, Hev95, Hir92, Hol93, Isr92a, Kel96, Kry92c, Lee97, Low93a, Low92, Mag94, May91, McS98, Mit91, Mit92, Mit97, Mue95, Nye90, Pac95, Par94, Pin91, Pin92, Pur97, Rot97, Scr98b, Sel91, Sol98, Ste95, Tri94, Wal91, Wer92, Wer97, All98b, Ano90e, Ano91g, Ano92f, Ano93f, Ano94f, Ano95f, Ano97f, Bas92a, Ben91b, Bol90, Chr91, Con98, Cut94b, Cut99, Dou91b, Gam97, Gol91a, Gui98b, Hår93a, Hea98, Hes99, Lea92]. Technology [Mit96, Nob97, Nye97, O’D94a, O’D94b, Pre97, Rey90b]. Technology/Building [Gie94]. Technopoly [Fro93]. Technoscientific [Coo93a]. Técnica [D’A97, MA97, Gli91]. tecnología [Gua98]. Tectonic [San97]. Tedlow [Lai91]. Teich [Cut94b]. teknisk [Car99b]. tekniska [Bjö97]. Teknologinen [Ahv90]. Teknologiske [Hår98a]. Telecommunications [Bar92a, Bau99, Fin93, Kel93c, Hea99a, Sin99c, Sch94d]. Telegraph [Kie92, Nic99a, Sch95, Sel99]. Telegrapher [Dav99]. Telegraphy [Fie94a, Gra95, Hon94]. Telephone [Fis92, Fis98, MM93, Mue91a, Sin99c, Vol94a]. Telephony [Can95]. Telescope [Bar91a]. Telescopes [Rot94a]. Television [Ste99, Wes99, Cla97b]. Tell [Sal95]. Teller [Ber90]. Telling [Wos92]. Temin [Smi90b, Asp93]. Temperature [Cha99a]. Templer [Sha94]. Ten [Sin99c]. Tennessee [Ham95, Rol93b]. Tennessee-Tombigbee [Ham95]. teollisuudessa [Ahv90]. Terence [Dav93]. Terri [Mil95]. Territorial [Laz99]. Terror [Gre99]. Terry [Jud90, Kat94a, Rei93a]. Tessa [Coo98b]. Tessie [Gut95b]. Testing [Don94, Sei96a]. Tetsuro [Owe91a]. Texas [Oli90, Oli98, Har99c]. Textbooks [Nye98b]. Textile [Bla90b, Con95, Gar90, Gro90b, Lea99, Scr90, Scr91b, tZS90, Voe90]. Textiles [Liv92, All97, Scr90]. Texture [Lan95b]. Teyler [ML98]. Thackray [Fit97, Les94]. Thailand [Edm96]. Thames [Bow99b]. Their [Dou91a, Ing91, Moh97b, Nye90, Old97, Ron99, Sin97, Tho93, War98b]. Them [Vol94b]. Theme [Cow93]. Themes [Ada91, Ber95b, Scr98b]. Theo [Web90c]. Theodore [Alb97, Bla96a, Hew94, Smi93b]. Theories [Sha94]. Theory [Ada93, Bat98, Blu99, Bor99, Buc91, Hår96a, How98, Kra98, Law91, Lue90, Mei96, Moh97a, Nel92, Ror91, Scr91c, Tri90a, Vin97]. There [Hil95, Mar98]. Thing [Sch99a]. Things [All90, Bla99a, How98, Ols98, Pas96, Pos99, Wil95c, Fri90, Sch92b, Sch94b, Web90b]. Thinking [BA90, Luk99, WH96, Kid90, Caw95, Chr99]. Third [Hig96, HC98, Neu96b, Hor97a]. Thirst [She94b, Rob92b]. Thirteenth [Kel93a]. Thomas [All98a, All96a, Bas97, Bro92b, Bur92, Cha90, Coo96, Gos90, Har93b, Hol91, Kli91, Lay96, Mac93b, Mil97, Mor91a, Mor97, Ras92a, Slo98, Web96, Ben96, Coo91c, Ful98, IR91, Cha90, Dan92, Nye92]. Thompson [Hil91b, Leo92, Min95c]. Thomson [Ros94b, Cum91, Gli94, Wil95a]. Thornburg [Whi93]. Thorne [Har95b]. Thornton [Jan91, Sta97a]. Thought [Edg98, Fri98, Sch99a]. Thousand [Bas92a]. Thousand-Year [Bas92a]. Thread [Coo92a]. Three [Gru90, Joh97b, McG99]. Threshers [Ras92a]. Threshing [Ras92a, Hur90]. Thrills [Dav92b]. Thrower [Kon97]. Thumb [Car91]. Tibetan [Pal91]. Tide [Ano99n]. Tidewater [Wor93b]. Tiffany [See90]. Til [Hei99]. till [Car99b]. Tilting [Mol95b]. Time [Arr94, Con97, Gol90, Har95b, Hen99, Lan91a, Lan93, O’M93, Ryc93, Sla92, Ste89, Tur95, You95, Bar91b, Ste89, Rom97a, Rum92, Sal95, All98a, O’M90, She94a]. Time-Machine [Gol90]. Timekeepers [FF96]. Times [Bar90, Cas98b, Coo91a, Gol91b, LaL90, Lay97, Maz94]. Timetables [Mag94]. Timo [Ahv90, Nye93a]. Timothy [Lai96a, Pin91]. Tinctoria [Adr94]. Tire [Fri92a]. Titanic [Coo91d]. Title [Ano91j]. Todd [Mer96]. Together [Con92a]. Tokugawa [FF96]. Tom [Cas98b, Gol91c, Pfa91, Smi90a, Smu93]. Tombigbee [Ham95]. Tool [Coo98a, Jag99]. Tools [Kay93, LMO97b, Hea91b]. Tore [Lar91, Lun92, War92a]. Torpedoes [Zac98]. toscano [Sel92]. tot [Hom94]. Touch [Eva93b]. Tourism [Ber98]. Tower [Wil91]. Town [Bro98a, Buc94, Cor90, Har94, Wor93a]. Towns [O’D96b, Big97]. Toxic [Gor98]. Toyota [Vol91c, Vol91c]. Toys [Old97]. Tracks [Dou95]. Traction [Tar95]. Tractors [Fit91]. Trade [Bla96b, Cla97a, Gam95, Gui95a, Jer97, War91, Whi90, Moy94, Tra97a]. Trades [Lew99, Haf92]. Trading [Tas97]. Tradition [Daw99, Hår96a, Sel97, Smi90c, War97, Fie94b, Mit96]. Traditional [Fer92, Hur90, Lin94a, Wig96]. Traditions [McG99, NP94]. Tragedy [Hac90a]. Trailer [Whi93]. Trails [McS98]. Train [Ald93, Min95c]. Training [For93]. Trans [Rho93, Ste94b, Coa99]. Trans-Alaska [Ste94b, Coa99]. Trans-Siberian [Rho93]. Transatlantic [Mei98b, Tra94]. transatlantiques [Hea99a]. Transcontinental [Cru99]. Transcutaneous [MKR93]. Transfer [All98b, Car96, For93, Hår93a, Kay91, Mat99b, Myl91, Nye93a, Phi99, Ras99, Lay92b]. Transformation [Chr91, Coo98b, Daw99, Hac92b, HM96, Lan99, Lub91b, Moh97b, Phi91, See99a, TJ96, War92b]. Transformations [Mis90, Ros98]. Transformed [Nel92, Vol98a, Vol98b]. Transforming [Fie94b, Gam97, Cer98, Coo96, Rei96]. Transit [McS98]. Transition [All97, Bla96b, Rey90b, Scr90]. Transitions [Kel93c, Smi90a]. Transmission [Coo93b, Lev97b]. Transport [Div95, Gra98a, Jos96, Sum93, Ang90]. Transportation [Bra92, Buc92, Mil91a, Hil91b]. Transsexualism [Old98]. Trapped [Sin99b]. Trautmann [All98a]. Travail [Sim95a]. Traveler [Sin95]. Traveling [Lai96a]. Travis [Roc94]. Treadwheels [Mat92b]. Treatise [Gui91]. Treatises [Kuc94]. Trebuchet [Tar95]. Trelease [Whi92b]. Trench [McN93]. trente [Smi91d]. Trenton [Bla96b]. Trevor [Lan94a]. Tribe [Rhe94]. Trident [Hev95]. triere [Fer99]. Trimble [Hol95]. Trinder [Buc98a]. Triumphant [Hil92b]. Trolley [Smi97b]. Troubled [Maz94]. Truman [Rot99c]. Trust [Alb97]. Truth [Wos94, For97, Mis97]. Tsar [Wes95]. Tsarist [Gra98b]. Tsurumi [Coo92a]. Tube [Arn97, Mau91]. Tuberculosis [Far96]. Tufte [Fri92b]. Tula [Bra91a]. Tuning [Con92c]. Tunnel [Sta96a, Rog95, Tra91]. Tunnels [Jak98]. Tupolev [Jos99]. Turbine [Mat99b, Nah97]. Turin [Kuc94]. Turner [Kas99, War92b, Bed91, Hår96a, Kid96, War93, Goo95]. Turning [Smi94]. tussen [Dav96]. TVA [Seg94]. Twentieth [Bla90a, Coo91d, Gam97, Gro98, Gut92, Haf97, Her98, Kat97a, Kay93, LeC94, Lie90, Lin97a, Ryc92, San97, Sch99b, Ube99, Vol91a, Nic99b, Tap99]. Twentieth-Century [Gro98, Kay93, Lie90, Lin97a, Ryc92, Sch99b, Nic99b, Tap99]. Twenty [Coo98b, Hor97a, Kyn99, Lew96]. Twenty-First [Coo98b, Kyn99]. Twenty-third [Hor97a]. Two [Aga98, Ano91k, Bel91, Lev91c, Man99, MB92, Nah97, O’H96, Bro90]. Two-stroke [Nah97]. Tycho [War98b]. Type [Pal91]. Typewriter [Sta94]. Typography [Tas97].

U [Hig97b, Mer96, Pol90]. U-Boat [Hig97b]. U.S [Bix99, Bla90a, Cla90, Fri92a, Hac92a, Hac94b, Hev95, Hig98, Jan91, Koh96, Mer95, Sch91, Sch94c, Sch94d, Scr90, Sla98]. U.S. [Dic92, Gre99, Jac93b, McB92, Bur92]. U.S.-Mexico [Gre99]. Uhren [LF93]. Ulrich [Pay95]. Ultimate [Nye98a, Rom97b]. Ultra [Rei95b]. Ultrasound [Koc93]. Umble [Fis98]. Umwälzung [Lin93b]. Umwelt [Web96]. Underdevelopment [Bro93b]. Undergraduate [Cut97]. Underground [Ham91, Sla91]. Understand [Dur98]. undervisning [Bjö97]. Underwood [Ebn92]. Unemployment [Fan91, McB97]. Unfinished [Gor95b]. Unforgiving [Sal95]. Unglued [Ran97]. Uniform [Col99]. Unintended [Sin99b]. Uninventive [CC95]. Union [Bro95a, Vol98b, Ada95b]. Unionism [Hor94]. United [Bad94, Ber94b, Ber95a, Ber95c, Cla96, Coo93b, Cum91, Fit96, Gro98, Han93, Hoc98, Hoy92, Kyn99, Mac92, Rob92a, Sei96b, Skr91, Slo98, Str90a, Van97, Wal94, Fan91, Hay96, Hey98, KP96, Nic99a, Ols98, Pur93b, Reu99a, Sch94a, Sha90b, Zei95]. Universal [Bel91, MB92, Sin99c]. Universität [Gis98]. université [Frä91]. universiteit [Dav96]. Universities [Cut97, Div94, Les90b, Pal93]. University [Rey94, See99a, Ser96, See98b]. Unmaking [Hor94]. Unpatentable [Coo94]. Unreliable [Con99, Rey99, Scr00]. Unseen [Sul98b]. unserer [Lin93a]. Unstable [McB97]. Untimely [Bel98b]. Upper [Bla92b]. Uppsala [Ano93l]. Urban [Ebn92, Gei90, Hur98, JM97, Kuc94, Leo96, Qui94, Qui95, Rom97b, Ros93, Sul98a, Whi92c]. Urbanization [Vie90]. USA [Cut97]. USAAF [Hig97b]. Use [Asp93, Bro97a, Cla95, Coo98a, Eno98, Mue99, Ree98, Rot99c, Wes90, vHT96]. Used [Vol94b, Win98]. Useful [Jac95a, Sch94b]. Users [Bor99, KP96, War98b]. Uses [Hac92a, Kay94]. Usher [Ano90a, Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Ano97a]. Using [Fle98]. USS [Min95a]. USSR [Ger96, Gra97a]. Utilities [Mea98]. Utility [Chr91, HS96, SD98]. Utrecht [Dav96]. utredningen [Car99b]. Utter [Rot99c]. utveckling [Car99b]. utvecklingsarbete [Car99b]. Uwe [Mer97].

V [Blo91, Fit95, Kli91, Kry92b, Lon95a, Mar93a, Neu96b, Nye92, You93, Neu96b, Neu96b]. V-1 [Neu96b]. V-2 [Neu96b]. V-Missiles [Neu96b]. Vaccine [Car99a]. Vaclav [Wil95b]. Vacuum [Arn97]. Valentine [Lie92, Mat92a]. Valenze [Haf96]. Valley [Bla97a, Hof91, Bla92b]. Value [Hil91a]. Values [Alb97, Buc99, Chu99]. Values-Driven [Chu99]. Vance [Bry96, Buc92]. Vancouver [NP94]. Vanda [Hei95b]. Vander [Bil97, Tri93].

Vanderbilt [Cas99]. Vandiver [Gor92a]. Vanguard [Kei94]. Vannevar [Bas99, McM94a, Rei95b]. Vardalas [Arm96]. Variation [Vin94]. Variation-Selection [Vin94]. Vasa [Arr94]. vattenkraftutbyggnad [Kai98]. Veblen [Big96]. Vehicle [Blo90]. Vehicles [McS98, Vol94b]. Velho [Was92]. Ven [Mul96a]. Venetian [Kuc92]. Vengeance [Hac94b]. Venture [Bel98b]. Verkehrs [Har99d]. Verkehrskonzepte [Har99d]. verksamhet [Bjö97]. Vermont [See95a, Wes93a, Rol94]. Vernacular [Sha95b]. Vernon [Kid90]. Versailles [Laz99]. Versatile [LMO97b]. Vertical [Ell92a, Mil91a]. Vetenskap [Har99a]. via [Tra94]. Vickers [O’H96]. Victor [Ell94b, See95a]. Victorian [Ang90, Bow99b, Kie93, Tra98b, Buc99, Web90b]. Victorians [Bri92]. Videohistory [Mil95]. vie [Lav93]. Vienna [She94a]. Viest [Ste92]. View [Bat90, Lan95b, LMO97a, Twe92, Jac91]. Viewing [Bar92b]. Views [Mat92b]. Vilhelm [Kie90a]. ville [Kah96]. Vincent [Tra97b]. Vincenti [Flo92]. Vinci [Ano92h, Kel99, Lon97, Wes93a, Ano90g, Ano91i, Ano93h, Ano94h, Ano95h, Ano96g, Ano97h]. Viola [Por94]. Virginia [Ano94i, Ano96b, Cor92, Hol93, Rob96a, Sha95b, Lie92, Mat92a]. Virilio [Kop91]. Virtual [Bow98, Sin97]. Virtue [Wal95]. Virtuoso [Rhe94]. Vision [Bjö97, Bud94, Kat92, Seg94]. Visions [Sta96a, Cor91, Hur98]. Visits [BP91]. Visser [vL90]. Visual [All90, Nei98, Pla96, Sum97, Ube94, Wos94]. Visuality [Ree98]. Visualization [Luk99, Sam99]. Vital [Hig97a, Hil92b, May92]. Vitamania [Bog98]. Vitamins [Bog98]. Vivian [Hil91a]. Vlach [Lee96]. Vogel [Mil91a]. Voices [Per93, Rol98b]. Vol [Coo93b, Cut98, Fit97, Hår98a, Har96b, yH98, Isr94a, Kli91, Lan91b, Lin94a, Mil97, Mor97, Mul96b, Pan98a, Qui95, She93, Sta97a, The90, War93, Bec91, Blo91, Bry95, Cam92, Dud92, Har93b, Har95a, Les94, Lon95b, Phi93, Sha94, tZS90, Wig96]. Volker [All96c, Tod96]. Voltage [Lev97b]. Volti [Cha91, McW95]. Volume [Ano90e, Ano90k, Ano90j, Ano90l, Ano90m, Ano91g, Ano91n, Ano91o, Ano91m, Ano91p, Ano91q, Ano92f, Ano92n, Ano92m, Ano92o, Ano92p, Ano92l, Ano93f, Ano93o, Ano93n, Ano93p, Ano93q, Ano93m, Ano94f, Ano94n, Ano94m, Ano94o, Ano94p, Ano94l, Ano95f, Ano95n, Ano95o, Ano95m, Ano95p, Ano95q, Ano95l, Ano96k, Ano96j, Ano96l, Ano96m, Ano96i, Ano97f, Ano97n, Ano97m, Ano97o, Ano97p, Ano97l, Ano98d, Ano98g, Ano98f, Ano98h, Ano98i, Ano99l, Ano99r, Ano99q, Ano99s, Ano99t, Ano99u, Ber95a, Fel99]. Vom [Gis98, Lin93b, Zac99]. Voyage [Bow98, Tra94]. Vrooman [Uss93]. vs [Bur92]. Vulgar [War91].

W [Bar91a, Ben90, Cap90a, Cap95, Cer97, Col99, Cor93, Cut94a, Dav96, Dre97, Edw96a, Eva96, Gar91, Gei94, Goo95, Gor95a, Gor96, Hac98, Hal96, Hil91a, Hoc98, How98, Iri97, Jam90, Joh92, Joh97a, Kas99, Kid94, Kid96, Kon97, Kon98, Kri92, Lai92, Lev91c, Mar96, MM92, Oli90, Pye92, Rey90b, Rey94, Rob96a, Ror91, Ros94b, Rum97, Sal92b, Sch94c, Sch94d, Sei96b, SS97, She93, Sta92a, Tri90b, Vas98, Whi92b, Wil95c, Wor95, vL90]. Waffentechnik [Hal96]. Wagenhoff [All96c, Tod96]. Wages [Haf92, Scr90].

Wagner [Mul94]. während [Lin93b]. Wakefield [Kir95]. Walked [Kin99]. Walker [Cap90b, Hac94a, Hun96]. Wall [Kle91, Nas90]. Waller [Bro97b]. Wallis [Ber93]. Walter [Bro96, Coo93c, Flo92, Gis98, Hyd90, Lin93a, Seg94]. Walton [Gut93]. Wanda [Lay97]. Waning [Szo96]. War [And94, Bac90, Bar92c, Cap93, Coz93b, DeV91, Gei94, Gui93b, Hac91a, Hac94b, Hal98, Ham92b, Hea91b, Hig91, Hig97b, Hil97, Hol94, Kön96, Kue99, Min95a, Moy96, Nah97, Omi92, Rol90, Rol92, Rol93a, Rol95b, Rol98a, Rus99, Sal92a, Uss97, Was93, Zac98, Zei95, Zie94, Coo92b, Bil97, Blu99, See99a, Hac92b, Jer97, McN90, New94, Olw98, Pol91, Hac90b, Kop91]. Ware [Rey98]. Warfare [Bac98, Bil97, Cor93, Hea91b, HV99, McN93, Rol95a, Sch92a]. Wargo [Gor98]. Waring [Nel92]. Warps [Har95b]. Warren [Skr91]. Wars [Smi97b, Jer94a, Moh90, Spi95]. Wartime [Kra92]. Was [Mar98, Wil91]. Washday [Par97]. Washington [Pos98a, Rei96, Dia99]. Wasko [Gom96]. Wasserbaus [Reu93a]. Wastage [Fan91]. Waste [Low97]. Watch [Rum92]. Watches [O’M93]. Water [Bro92a, Bul91, Gal92b, Gei90, Hon98, Kin99, Lan94a, Mer96, Mul96a, Nye98a, Rob92b, Rob94b, She94b, Smi94, Ste93, Win98, Hof92]. Water-Clocks [Hon98]. Waterfront [Fit90]. Waters [Ham95, Mac96, Smi93b]. Waterway [Sha90a, Ham95]. Waterways [Sha90b]. Waterwheels [Cla95]. Waves [Can95, Kli97a]. Way [All90, Jes97, Kai96, Mat92a, Min99, Ols98, Was93, Joh99a]. Wayman [Gor91]. Wayne [Blo90, Fit95, Hod93, Sch92a]. Ways [Vol94b]. Weakness [Lai96c]. Wealth [Reu95, Qui98]. Weapon [Bra91a, Fin93, Hig98]. Weapons [Bad94, Gar99, Hac94c, HV99, Olw98, Phi99, Pur90, Rol90, Sei96a, Zie94, Moh95]. Wear [Haf99]. Weather [Fel97, Kie90a, Wil93a]. Weaver [Gut95b]. Web [Kun95]. Webb [Joh97a]. Weber [Lun90a, Lun90a]. Weberian [Coo95]. Wechselverhatnisses [Dek98]. Wedge [Ell92a]. Weg [Gam99, Pet98]. Wegener [Jak98]. Weil [Hil91a]. Weimar [Neu90]. Weinstein [Pri98]. Weir [Kas99]. Welfare [Kai96, Cor93]. Well [Ras90]. Well-Fed [Ras90]. Wells [Bil97]. Wellsprings [Bla97b]. Welskopp [All96a]. Weltkrieges [All96c]. Weltraum [Neu94]. Wende [Lin93b]. Wendy [Øve97]. Wengenroth [Pay95]. Went [See90]. Were [Lig99, Ols98]. Werkbund [Blu99]. Werkplaatsen [vL90]. Werkstattpraxis [All96c]. Werkzeugmaschinen [All96c]. Werner [Mai96]. Wernher [Neu94]. Wesley [Hig96, Mor93]. West [Bar92c, Chi97, Fit90, HD90, Hol97b, Rob94b, Tra95, Wor93a, Ker99, Rot94b, Sto91, Con92b, Jac91, Nye98a, Hur99, Les99]. Western [Ada91, Bas96, Cam92, Coo97, FO92, Gut95b, Hea90, Mit96, Pis96, Por98, Pri95, Roc94, Rol95a, Rug96, Spe95, War97, Ada97, HM96, Ing99, Mar98, LeC99]. Westman [Kel93b]. Westrum [Sch92b]. Wetenschap [vL90, Dav96]. Wheat [Han99]. Wheel [Cor92, Hyd90, Wri92, Ber93]. Wheelbarrow [Lew94, Mat91]. Wheels [Wor94]. Where [Hor97b, Bin91, Sch91, Swe92]. Whipp [Bla92a]. Whiskey [Cal98]. White [Fie94b, Kle95, Rey98, Lie93, Pac95]. Whitehead [Tho94]. Whitney [Gut93, Hig90, Vic92, Kel93a]. Who [Gol91b, JM97, Lan90b, Smu93, Gre94].

Whose [Bra98]. Wider [Cam92]. Wiebe [Gie94, Mei96]. Wiener [Gor94]. Wiggins [Fri96]. Wight [Bla97a]. Wilbur [Gol91c]. Wild [Hol90]. Wildcatters [Cas98b]. Wilderness [Hur99]. Wildlife [Gos90]. Wilfried [Mai96]. Will [Hei95b, Hor97b, Per93]. Willard [Uss93]. William [Bau93, Ber98, Cam92, Car91, Con91, Con92b, Far96, Fit90, Hew94, Hil91b, Hol95, Hol97b, Jan91, Kid92, LaL90, Lam93, Lie93, Lon96a, Mal91, Mei98a, Min95b, San92, Sta97a, Sti91, Swe92, Vas98, War98b, Wil92b, Wos94, Bas92b, Ste89]. Williams [Coh90, Eva96, How93, Ing91, Mor91b, Ors97, Sla91]. Williot [Smi95b]. Willis [Ben91b]. Wilson [Sut90, Con91, Kli90, Sta96a, Wor93a]. Wind [Hey98, Jak98, Kar99, Mue95, Tri91]. Windenergienutzung [Ver97]. Windmill [Mue95]. Windmills [Mol95b]. Wine [Gut92]. Winecraft [Gou98]. Wing [Tri90b]. Wingless [Mei99b]. Wings [Pis96]. Winning [Fri99]. Winpenny [Bro92b]. Winslow [Min99]. Winter [Rei99a, vK92a, vK92a]. Wire [Mor98]. Wireless [Hon94, Tho94]. Wires [Sch95, Sel99]. Wirtschaftswunder [Sto91]. Wisconsin [Ano92i, Fit97, Hor94]. Wise [Alb97, Por91]. Wissenschaft [Bra90, Lin93b, Lon93b, Rüd92]. Without [Lai96c, Bar92a, Hea96, Pri95]. Witkin [McC93, McC93]. Witold [Mor90]. Witzell [Jam90]. Wohl [Pis96]. Wolf [Mar90, Hir96]. Wolfgang [Opi90]. Wolfhard [Lun90a]. Wolfsonian [Por97]. Woman [Haf96]. Women [Aik98, Coo92a, Coo96, Cor92, Fle92, Gol99, Gre95b, Gre91, Kli97b, Lig99, McG97a, Och92a, Old92, Pri95, RK97, Roz98, Til90, Til97, Pur93a, Ble91, Haf97, Moh94]. Won [Moy96]. Wood [Kle95, Lie93]. Wood-Car [Kle95]. Wooden [Roz98, Sch94a]. Woodhouse [Cla90]. Woodward [Lin94a, The90, Wig96]. Woodworking [Coo91c]. Woolfolk [Fri96]. Woollen [Tan93]. Word [Sin96]. Words [Tas97, Wal95, Wos92]. Work [All90, Bix99, Bla92a, Ble91, Coo93c, Coo96, Dub93, Fit93a, Fle92, Gre91, Gre99, Haf99, Hon97, Hor97b, HM96, Kay93, Kel93d, Kra94a, Mel99, O’H96, Pos99, Rhe94, See99b, Sta94, Sta90c, Til90, Vol91b, Whi98, Car99b, Led91, Pet90, Abb96, Cam93, Coo98a, Fit90, Haf92]. Worker [Fit90, Pfa98]. Workers [Cam95, Cha98, Dal90, Dan90, Gro98, Gut95a, Haf97, Kuc92, Nuw90, Sim90, Smi97b, Vol91b]. Workforce [Jar95]. Working [Fro92b, Kal98, Lie92, McE95, Pri98, Scr91b]. Working-Class [Fro92b, Scr91b]. Workingman [Big97]. Workingwomen [Per93]. Workplace [Kuc92, Per90]. Works [Har91, Mal98, Sha90b, Wor95]. Workshop [Har93b, Meu98, War98a, Jac93a]. Workshops [Kid91]. Worktime [Nel90]. World [All90, Bal98, Bar90, Bas92a, Bel97, Bil97, Blu99, Cam92, Con98, Coz93b, Cut93, Eat90, Hac94b, Hig97b, Hig96, Hor92, Jac93a, Jer90b, Moh97b, Moy96, Pan95, Per99b, Pol91, Pri95, Rot99b, Rou96, Roz98, Ryc92, Sha93, Ste93, Uss97, Vol94b, Wil95b, Zac98, Bla90b, Cap93, Kön96, Kue99, Nah97, Rot92, Rus99, Zie94, Cam92, Ful95, Mor90, Ras90, Szo96, Bar92c, Gui95a, Rum97]. Worlds [All99a, And99a, Cap91, Til94]. Worley [Wil92b]. Worster [Reu95]. Worth [Yea90]. Wosk [Ube94]. Wright [Cor91, Lan90b, Bud96, Sch93a, Gol91c]. Writing [Con95, Hei95a, Phi93, Sil99, Sla97, Wac99, SB95, Sla92, SD98]. Writings [Lon91b]. Wrong [See90]. Wrought [Rol97]. Wrubel [Ham98b]. Wunderwaffen [Neu96a]. Württemberg [Bra90]. WWII [Dup92]. Wynn [Rue90]. Wynne [Sol98].

X [Hac98, Arn97, Her99b]. X-Ray [Arn97]. XIX [But92, Lav93, Smi92a, vR90]. XIXe [Hea99a]. Xue [Sut90]. XVI [Gli91]. XVII [Gli91]. XX [Lav93, Smi92a]. XXe [Hea99a]. Xylography [Pal91].

Yankees [Zem93]. Yard [Sha95b]. Yaron [Nye97]. Yaroshinskaya [For97]. Yarotsky [Kry92b]. Yates [Lub91a, Mul96b, Rey94]. Yavetz [Hon97]. Year [Bas92a, Fer90]. Years [Chu99, Lai96a, Lew96, Rei93a, See95b, See95a, Ste92, Ste99, Uss90, Bil98, Hal90, Isr94a, Mil97, Vol91c]. Yellow [Bri93]. Yonder [Hol90]. York [Con92a, Dan95, Dre97, Hil92a, Sul98a, Fri91b, Haf99, Pur90]. Yoshikawa [Mul95]. Young [Edm96, Ser92]. Younger [Har91, Lon95b]. Yves [Die93].

Z [Kli97a]. Zaheer [Ada98]. Zanella [Cut94a]. Zara [McC93]. Zboray [Kel95]. Zeilensetzmaschine [Zac99]. Zeitordnung [LF93]. Zenzen [Ham98a]. Zeppelin [De 99]. Ziegler [Hea96, Lin93a]. Zimmerman [Fis98, Gol91a, Sch94c, Sum91]. Zionism [Ben92b]. Zipper [Pat95]. Zonderman [Cam95]. Zuboff [Sta90c]. Zukunft [Har99d]. zum [All96c, Har99d, Lin93b, Tod96]. Zunz [Gal92a]. Zur [Lin93a, Gam99, Gis98, Lon93b, Wen92, Zac99]. Zweckbronner [Bra90]. zwischen [Bra90, Buc97, Dek98, Tod99]. Zworykin [Cla97b].

References

Abbate:1993:PCC

[Abb93] Janet Abbate. “People and computers” at the Computer Museum. Technology and Culture, 34(3):665–669, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

         https://muse.jhu.edu/pub/1/article/888136/pdf.

Abbate:1996:WTH

[Abb96] Janet Abbate. Work and Technology in Higher Education: The Social Construction of Academic Computing ed. by Mark A. Shields (review). Technology and Culture, 37(3):653–655, July 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

         URL https://muse.jhu.edu/pub/1/article/887463/pdf.

Arns:1995:RCH

[AC95] Robert G. Arns and Bret E. Crawford. Resonant cavities in the history of architectural acoustics. Technology and Culture, 36(1): 104–135, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888006/pdf.

Adams:1991:SAO

[Ada91] Nicholas Adams. The Science of Art: Optical Themes in Western Art from Brunelleschi to Seurat by Martin Kemp (review). Technology and Culture, 32(3):599–601, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888699/pdf.

Adams:1993:DPE

[Ada93] Nicholas Adams. Distance Points: Essays in Theory and Renaissance Art and Architecture by James S. Ackerman (review). Technology and Culture, 34(1):138–140, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888464/pdf.

Adams:1995:GRF

[Ada95a] Judith Adams. Going Out: The Rise and Fall of Public Amusements by David Nasaw (review). Technology and Culture, 36(2):425–426, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887903/pdf.

Adams:1995:GEE

[Ada95b] Mark B. Adams. The Ghost of the Executed Engineer: Technology and the Fall of the Soviet Union by Loren R. Graham (review). Technology and Culture, 36(1):168–170, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888008/pdf.

Adas:1996:CME

[Ada96] Michael Adas. Civilizing Mission: Exact Sciences and French Overseas Expansion, 1830–1940 by Lewis Pyenson (review). Technology and Culture, 37(1):190–192, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887555/pdf.

Adas:1997:FMT

[Ada97] Michael Adas. A field matures: Technology, science, and Western colonialism. Technology and Culture, 38(2):478–487, April 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887332/pdf.

Adas:1998:SES

[Ada98] Michael Adas. The Science of Empire: Scientific Knowledge, Civilization, and Colonial Rule in India by Zaheer Baber (review). Technology and Culture, 39(2):354–355, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887160/pdf.

Adler:1990:MMS

[Adl90] Paul S. Adler. Marx, machines, and skill. Technology and Culture, 31(4):780–812, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901650/pdf.

Adrosko:1994:BTA

[Adr94] Rita J. Adrosko. Bibliotheca Tinctoria: Annotated Catalog of the Sidney M. Edelstein Collection in the History of Bleaching, Dyeing, Finishing and Spot Removing by Moshe Ron (review). Technology and Culture, 35(1):181–183, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887964/pdf.

Agar:1998:MST

[Aga98] Jon Agar. The Measurement of Starlight: Two Centuries of Astronomical Photometry (review). Technology and Culture, 39(4): 764–766, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33246.

Ahvenainen:1990:TMS

[Ahv90] Jorma Ahvenainen. Teknologinen muutos Suomen teollisuudessa 1885–1920 (Technological change in Finnish industry 1885–1920) by Timo Myllyntaus, et al. (review). Technology and Culture, 31(2):328– 329, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888901/pdf.

Aiken:1998:HJI

[Aik98] Katherine G. Aiken. Hazards of the Job: From Industrial Disease to Environmental Health Science, and: Radium Girls, Women and Industrial Health Reform: 1910–1935 (review). Technology and Culture, 39(4):794–797, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33257.

Aitken:1994:ASO

[Ait94] Hugh G. J. Aitken. Allocating the spectrum: The origins of radio regulation. Technology and Culture, 35(4):686–716, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887719/pdf.

Alborn:1997:TNP

[Alb97] Timothy L. Alborn. Trust in Numbers: The Pursuit of Objectivity in Science and Public Life by Theodore M. Porter, and: The Values of Precision ed. by M. Norton Wise (review). Technology and Culture, 38(1):259–261, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887390/pdf.

Aldrich:1993:CCH

[Ald93] Mark Aldrich. Combating the collision horror: The Interstate Commerce Commission and automatic train control, 1900–1939. Technology and Culture, 34(1):49–77, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888457/pdf.

Aldrich:1995:PNP

[Ald95] Mark Aldrich. Preventing “The needless peril of the coal mine”: The Bureau of Mines and the campaign against coal mine explosions, 1910–1940. Technology and Culture, 36(3):483–518, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887788/pdf.

Alder:1996:SIS

[Ald96] Ken Alder. Science, Industry, and the Social Order in Post-Revolutionary France by Robert Fox (review). Technology and Culture, 37(4):836–838, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887411/pdf.

Alder:1997:IAE

[Ald97] Ken Alder. Innovation and amnesia: Engineering rationality and the fate of interchangeable parts manufacturing in France. Technology and Culture, 38(2):273–311, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887323/pdf.

Aldrich:1999:PBR

[Ald99] Mark Aldrich. “The peril of the broken rail”: The carriers, the steel companies, and rail technology, 1900–1945. Technology and Culture, 40(2):263–291, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33326.

Alexander:1999:PDM

[Ale99a] Jennifer Karns Alexander. Planet Dora: a Memoir of the Holocaust and the Birth of the Space Age (review). Technology and Culture, 40(4):908–910, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33424.

Alexander:1999:BRH

[Ale99b] Mary Alexander. Between a rock and a hard place: a history of American sweatshops, 1820–present, at the National Museum of American History. Technology and Culture, 40(4):861–865, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33425.

Allen:1990:WTW

[All90] Frederick Allen. The Way Things Work: From Levers to Lasers, Cars to Computers — a Visual Guide to the World of Machines by David Macaulay (review). Technology and Culture, 31(4):911–912, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901685/pdf.

Allen:1996:AMI

[All96a] Michael Allen. Arbeit und Macht im Hüttenwerk: Arbeits- und industrielle Beziehungen in der deutschen und amerikanischen Eisenund Stahlindustrie von den 1860er bis zu den 1930er Jahren by Thomas Welskopp (review). Technology and Culture, 37(1):187–188, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887553/pdf.

Allen:1996:KAT

[All96b] Michael Allen. Die Krematorien von Auschwitz: Die Technik des Massenmordes by Jean-Claude Pressac (review). Technology and Culture, 37(1):198–199, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887559/pdf.

Allen:1996:IMI

[All96c] Michael Allen. Industrieller Maschinenbau im 19. Jahrhundert: Werkstattpraxis und Entwicklung spanabhebender Werkzeugmaschinen im deutschen Maschinenbau, 1870–1914 by Volker Benad-Wagenhoff, and: Maschinen für die Massenfertigung: Die Entwicklung der Drehautomaten bis zum Ende des Ersten Weltkrieges by Jürgen Ruby (review). Technology and Culture, 37(4):848–851, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887418/pdf.

Allen:1996:PNM

[All96d] Michael Allen. The puzzle of Nazi modernism: Modern technology and ideological consensus in an SS factory at Auschwitz. Technology and Culture, 37(3):527–571, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887433/pdf.

Allinson:1997:TIT

[All97] Gary D. Allinson. Textiles and Industrial Transition in Japan by Dennis L. McNamara (review). Technology and Culture, 38 (1):267–269, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887394/pdf.

Allan:1998:THE

[All98a] Stuart Allan. Time: Histories and Ethnologies ed. by Diane Owen Hughes and Thomas R. Trautmann (review). Technology and Culture, 39(1):183–185, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887208/pdf.

Allen:1998:TTG

[All98b] Michael Allen. Technology Transfer Out of Germany after 1945 ed. by Matthias Judt and Burghard Ciesla (review). Technology and Culture, 39(1):130–132, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887178/pdf.

Allen:1999:CPC

[All99a] Frederick Allen. Cats’ Paws and Catapults: Mechanical Worlds of Nature and People (review). Technology and Culture, 40(3):652–654, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33411.

Allen:1999:HME

[All99b] Michael Thad Allen. A history of modern European technology: the Propyläen technikgeschichte. Technology and Culture, 40(3):599– 606, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33415.

Alvarez:1999:CSA

[Alv99] Alejandro García Alvarez. Cuban sugar in the Age of Mass Production: Technology and the economics of the Sugar Central, 1899– 1929. Technology and Culture, 40(3):675–677, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33389.

Anderson:1992:DMM

[And92] R. G. W. Anderson. The Deutsches Museum, Munich by Otto Mayr, et al. (review). Technology and Culture, 33(1):210–212, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888383/pdf.

Anders:1994:HFC

[And94] Roger M. Anders. On the Home Front: The Cold War Legacy of the Hanford Nuclear Site by Michele S. Gerber (review). Technology and Culture, 35(3):631–633, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887864/pdf.

Anderson:1999:MWI

[And99a] John David Anderson. So many worlds: Invention, management, philosophy, and risk in the life of Leroy Hill. Technology and Culture, 40(2):438–439, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33348.

Andrews:1999:BFS

[And99b] Ian A. Andrews. Broadaxe to flying shear: The mechanization of forest harvesting East of the Rockies. Technology and Culture, 40 (1):166–167, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33311.

Angus:1990:TVB

[Ang90] James T. Angus. Transport in Victorian Britain ed. by Michael J. Freeman, Derek H. Aldcroft (review). Technology and Culture, 31(3):531–534, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888857/pdf.

Anonymous:1990:APU

[Ano90a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 31(3):475, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888838/pdf.

Anonymous:1990:DP

[Ano90b] Anonymous. The Dexter prize. Technology and Culture, 31(3):476– 478, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888839/pdf.

Anonymous:1990:DA

[Ano90c] Anonymous. The dibner award. Technology and Culture, 31(3): 469, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888834/pdf.

Anonymous:1990:ILM

[Ano90d] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 31(3):470–471, July 1990. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888835/pdf.

Anonymous:1990:IVT

[Ano90e] Anonymous. Index to volume 31 (1990) of Technology and Culture. Technology and Culture, 31(4):930–945, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901690/pdf.

Anonymous:1990:JCR

[Ano90f] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 31(3):473–474, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888837/pdf.

Anonymous:1990:LVM

[Ano90g] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 31(3):479–487, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888840/pdf.

Anonymous:1990:SCM

[Ano90h] Anonymous. The Sacramento, California, Meeting, October 12–15, 1989. Technology and Culture, 31(3):488–495, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888841/pdf.

Anonymous:1990:SER

[Ano90i] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 31(3):472, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888836/pdf.

Anonymous:1990:VNJa

[Ano90j] Anonymous. Volume 31, number 1, January 1990. Technology and Culture, 31(1):??, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1990:VNA

[Ano90k] Anonymous. Volume 31, number 2, April 1990. Technology and Culture, 31(2):??, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1990:VNJb

[Ano90l] Anonymous. Volume 31, number 3, July 1990. Technology and Culture, 31(3):??, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1990:VNO

[Ano90m] Anonymous. Volume 31, number 4, October 1990. Technology and Culture, 31(4):??, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1991:APU

[Ano91a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 32(3):574–575, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888693/pdf.

Anonymous:1991:AI

[Ano91b] Anonymous. Author index (1989). Technology and Culture, 32 (3):777–835, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895778/pdf.

Anonymous:1991:CMO

[Ano91c] Anonymous. The Cleveland Meeting, October 18–21, 1990. Technology and Culture, 32(3):587–593, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888696/pdf.

Anonymous:1991:DP

[Ano91d] Anonymous. The Dexter prize. Technology and Culture, 32(3):575– 578, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888694/pdf.

Anonymous:1991:DA

[Ano91e] Anonymous. The dibner award. Technology and Culture, 32(3):571– 572, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888689/pdf.

Anonymous:1991:ILM

[Ano91f] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 32(3):572–573, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888690/pdf.

Anonymous:1991:IVT

[Ano91g] Anonymous. Index to volume 32 (1991) of Technology and Culture. Technology and Culture, 32(4):1153–1168, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888684/pdf.

Anonymous:1991:JCR

[Ano91h] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 32(3):573–574, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888692/pdf.

Anonymous:1991:LVM

[Ano91i] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 32(3):579–586, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888695/pdf.

Anonymous:1991:POA

[Ano91j] Anonymous. Part one: Author-title index. Technology and Culture, 32(2 (part 2)):1–245, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888735/pdf.

Anonymous:1991:PTS

[Ano91k] Anonymous. Part two: Subject index. Technology and Culture, 32(2 (part 2)):247–403, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888736/pdf.

Anonymous:1991:SER

[Ano91l] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 32(3):573, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888691/pdf.

Anonymous:1991:VNJa

[Ano91m] Anonymous. Volume 32, number 1, January 1991. Technology and Culture, 32(1):??, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1991:VNAa

[Ano91n] Anonymous. Volume 32, number 2, April 1991, part 1. Technology and Culture, 32(2 (part 1)):??, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1991:VNAb

[Ano91o] Anonymous. Volume 32, number 2, April 1991, part 2. Technology and Culture, 32(2 (part 2)):??, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1991:VNJb

[Ano91p] Anonymous. Volume 32, number 3, July 1991. Technology and Culture, 32(3):??, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1991:VNO

[Ano91q] Anonymous. Volume 32, number 4, October 1991. Technology and Culture, 32(4):??, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1992:APU

[Ano92a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 33(3):574–576, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888250/pdf.

Anonymous:1992:AI

[Ano92b] Anonymous. Author index (1990). Technology and Culture, 33 (Supplement):139–169, ???? 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888034/pdf.

Anonymous:1992:DP

[Ano92c] Anonymous. The Dexter prize. Technology and Culture, 33(3):576– 578, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888251/pdf.

Anonymous:1992:DA

[Ano92d] Anonymous. The dibner award. Technology and Culture, 33(3): 571, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888246/pdf.

Anonymous:1992:ILM

[Ano92e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 33(3):571–572, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888247/pdf.

Anonymous:1992:IVT

[Ano92f] Anonymous. Index to volume 33 (1992) of Technology and Culture. Technology and Culture, 33(4):863–878, October 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888121/pdf.

Anonymous:1992:JCR

[Ano92g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 33(3):573–574, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888249/pdf.

Anonymous:1992:LVM

[Ano92h] Anonymous. The Leonardo Da Vinci Medal. Technology and Culture, 33(3):579–584, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888252/pdf.

Anonymous:1992:MWM

[Ano92i] Anonymous. The Madison, Wisconsin, Meeting, October 30– November 3, 1991. Technology and Culture, 33(3):585–590, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888253/
pdf.

Anonymous:1992:SER

[Ano92j] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 33(3):572–573, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888248/pdf.

Anonymous:1992:SI

[Ano92k] Anonymous. Subject index (1990). Technology and Culture, 33 (Supplement):170–211, ???? 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888035/pdf.

Anonymous:1992:VS

[Ano92l] Anonymous. Volume 33, 1992 supplement. Technology and Culture, 33(Supplement):??, ???? 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1992:VNJa

[Ano92m] Anonymous. Volume 33, number 1, January 1992. Technology and Culture, 33(1):??, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1992:VNA

[Ano92n] Anonymous. Volume 33, number 2, April 1992. Technology and Culture, 33(2):??, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1992:VNJb

[Ano92o] Anonymous. Volume 33, number 3, July 1992. Technology and Culture, 33(3):??, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1992:VNO

[Ano92p] Anonymous. Volume 33, number 4, October 1992. Technology and Culture, 33(4):??, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1993:APU

[Ano93a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 34(3):642–643, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888132/pdf.

Anonymous:1993:AI

[Ano93b] Anonymous. Author index (1991). Technology and Culture, 34 (Supplement):136–163, ???? 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888000/pdf.

Anonymous:1993:DP

[Ano93c] Anonymous. The Dexter prize. Technology and Culture, 34(3):643– 644, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888133/pdf.

Anonymous:1993:DA

[Ano93d] Anonymous. The dibner award. Technology and Culture, 34(3):638– 640, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888128.

Anonymous:1993:ILM

[Ano93e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 34(3):640, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888129/pdf.

Anonymous:1993:IVT

[Ano93f] Anonymous. Index to volume 34 (1993) of Technology and Culture. Technology and Culture, 34(4):977–991, October 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888075/pdf.

Anonymous:1993:JCR

[Ano93g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 34(3):641–642, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888131/pdf.

Anonymous:1993:LVM

[Ano93h] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 34(3):645–652, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888134/pdf.

Anonymous:1993:SER

[Ano93i] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 34(3):641, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888130/pdf.

Anonymous:1993:SI

[Ano93j] Anonymous. Subject index (1991). Technology and Culture, 34 (Supplement):164–209, ???? 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888001/pdf.

Anonymous:1993:TCE

[Ano93k] Anonymous. Technology and culture editorship. Technology and Culture, 34(2):472–473, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888454/pdf.

Anonymous:1993:USM

[Ano93l] Anonymous. The Uppsala, Sweden, Meeting, August 16–20, 1992. Technology and Culture, 34(3):653–664, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888135/pdf.

Anonymous:1993:VS

[Ano93m] Anonymous. Volume 34, 1993 supplement. Technology and Culture, 34(Supplement):??, ???? 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1993:VNJa

[Ano93n] Anonymous. Volume 34, number 1, January 1993. Technology and Culture, 34(1):??, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1993:VNA

[Ano93o] Anonymous. Volume 34, number 2, April 1993. Technology and Culture, 34(2):??, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1993:VNJb

[Ano93p] Anonymous. Volume 34, number 3, July 1993. Technology and Culture, 34(3):??, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1993:VNO

[Ano93q] Anonymous. Volume 34, number 4, October 1993. Technology and Culture, 34(4):??, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1994:APU

[Ano94a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 35(3):584–585, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887844/pdf.

Anonymous:1994:AI

[Ano94b] Anonymous. Author index (1992). Technology and Culture, 35 (Supplement):151–181, ???? 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887703/pdf.

Anonymous:1994:DP

[Ano94c] Anonymous. The Dexter prize. Technology and Culture, 35(3):585– 586, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887845/pdf.

Anonymous:1994:DA

[Ano94d] Anonymous. The dibner award. Technology and Culture, 35(3): 581, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887840/pdf.

Anonymous:1994:ILM

[Ano94e] Anonymous. The IEEE Life Members Prize in Electrical History. Technology and Culture, 35(3):581–582, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887841/pdf.

Anonymous:1994:IVT

[Ano94f] Anonymous. Index to volume 35 (1994) of Technology and Culture. Technology and Culture, 35(4):913–925, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887750/pdf.

Anonymous:1994:JCR

[Ano94g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 35(3):583–584, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887843/pdf.

Anonymous:1994:LVM

[Ano94h] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 35(3):587–596, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887846/pdf.

Anonymous:1994:RVM

[Ano94i] Anonymous. The Rosslyn, Virginia, Meeting, October 14–17, 1993. Technology and Culture, 35(3):597–604, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887847/pdf.

Anonymous:1994:SER

[Ano94j] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 35(3):582, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887842/pdf.

Anonymous:1994:SI

[Ano94k] Anonymous. Subject index (1992). Technology and Culture, 35 (Supplement):182–232, ???? 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887704/pdf.

Anonymous:1994:VS

[Ano94l] Anonymous. Volume 35, 1994 supplement. Technology and Culture, 35(Supplement):??, ???? 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1994:VNJa

[Ano94m] Anonymous. Volume 35, number 1, January 1994. Technology and Culture, 35(1):??, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1994:VNA

[Ano94n] Anonymous. Volume 35, number 2, April 1994. Technology and Culture, 35(2):??, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1994:VNJb

[Ano94o] Anonymous. Volume 35, number 3, July 1994. Technology and Culture, 35(3):??, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1994:VNO

[Ano94p] Anonymous. Volume 35, number 4, October 1994. Technology and Culture, 35(4):??, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:APU

[Ano95a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 36(3):654–655, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887801/pdf.

Anonymous:1995:AI

[Ano95b] Anonymous. Author index (1993). Technology and Culture, 36 (Supplement):147–175, ???? 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887715/pdf.

Anonymous:1995:DP

[Ano95c] Anonymous. The Dexter prize. Technology and Culture, 36(3):655– 656, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887802/pdf.

Anonymous:1995:DA

[Ano95d] Anonymous. The dibner award. Technology and Culture, 36(3): 651, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887797/pdf.

Anonymous:1995:ILM

[Ano95e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 36(3):651–652, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887798/pdf.

Anonymous:1995:IVT

[Ano95f] Anonymous. Index to volume 36 (1995) of Technology and Culture. Technology and Culture, 36(4):1063–1077, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887786/pdf.

Anonymous:1995:JCR

[Ano95g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 36(3):653–654, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887800/pdf.

Anonymous:1995:LVM

[Ano95h] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 36(3):657–666, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887803/pdf.

Anonymous:1995:LMM

[Ano95i] Anonymous. The Lowell, Massachusetts, Meeting, October 6–9, 1994. Technology and Culture, 36(3):667–675, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887804/pdf.

Anonymous:1995:SER

[Ano95j] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 36(3):652–653, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887799/pdf.

Anonymous:1995:SI

[Ano95k] Anonymous. Subject index (1993). Technology and Culture, 36 (Supplement):176–230, ???? 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887716/pdf.

Anonymous:1995:VS

[Ano95l] Anonymous. Volume 36, 1995 supplement. Technology and Culture, 36(Supplement):??, ???? 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:VNJa

[Ano95m] Anonymous. Volume 36, number 1, January 1995. Technology and Culture, 36(1):??, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:VNAa

[Ano95n] Anonymous. Volume 36, number 2, April 1995. Technology and Culture, 36(2):??, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:VNAb

[Ano95o] Anonymous. Volume 36, number 2, April 1995 supplement. Technology and Culture, 36(2S):??, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:VNJb

[Ano95p] Anonymous. Volume 36, number 3, July 1995. Technology and Culture, 36(3):??, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1995:VNO

[Ano95q] Anonymous. Volume 36, number 4, October 1995. Technology and Culture, 36(4):??, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1996:APU

[Ano96a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 37(3):587–588, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887439/pdf.

Anonymous:1996:CVM

[Ano96b] Anonymous. The Charlottesville, Virginia, Meeting, October 18–22, 1995. Technology and Culture, 37(3):596–616, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887442/pdf.

Anonymous:1996:DP

[Ano96c] Anonymous. The Dexter prize. Technology and Culture, 37(3):588– 589, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887440/pdf.

Anonymous:1996:DA

[Ano96d] Anonymous. The dibner award. Technology and Culture, 37(3): 583, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887435/pdf.

Anonymous:1996:ILM

[Ano96e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 37(3):584, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887436/pdf.

Anonymous:1996:JCR

[Ano96f] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 37(3):586, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887438/pdf.

Anonymous:1996:LVM

[Ano96g] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 37(3):590–595, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887441/pdf.

Anonymous:1996:SER

[Ano96h] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 37(3):585, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887437/pdf.

Anonymous:1996:VS

[Ano96i] Anonymous. Volume 37, 1996 supplement. Technology and Culture, 37(Supplement):??, ???? 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1996:VNJa

[Ano96j] Anonymous. Volume 37, number 1, January 1996. Technology and Culture, 37(1):??, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1996:VNA

[Ano96k] Anonymous. Volume 37, number 2, April 1996. Technology and Culture, 37(2):??, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1996:VNJb

[Ano96l] Anonymous. Volume 37, number 3, July 1996. Technology and Culture, 37(3):??, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1996:VNO

[Ano96m] Anonymous. Volume 37, number 4, October 1996. Technology and Culture, 37(4):??, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1997:APU

[Ano97a] Anonymous. The Abbot Payson Usher Prize. Technology and Culture, 38(3):719, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887278/pdf.

Anonymous:1997:AI

[Ano97b] Anonymous. Author index (1995). Technology and Culture, 38 (Supplement):168–205, ???? 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887210/pdf.

Anonymous:1997:DP

[Ano97c] Anonymous. The Dexter prize. Technology and Culture, 38(3):720– 721, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887279/pdf.

Anonymous:1997:DA

[Ano97d] Anonymous. The dibner award. Technology and Culture, 38(3):714– 715, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887274/pdf.

Anonymous:1997:ILM

[Ano97e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 38(3):715–717, July 1997. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887275/pdf.

Anonymous:1997:IVT

[Ano97f] Anonymous. Index to volume 38 (1997) of Technology and Culture. Technology and Culture, 38(4):1031–1047, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887266/pdf.

Anonymous:1997:JCR

[Ano97g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 38(3):717–718, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887277/pdf.

Anonymous:1997:LVM

[Ano97h] Anonymous. The Leonardo da Vinci medal. Technology and Culture, 38(3):721–725, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887280/pdf.

Anonymous:1997:LMA

[Ano97i] Anonymous. The London Meeting, August 1–4, 1996. Technology and Culture, 38(3):726–743, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887281/pdf.

Anonymous:1997:SER

[Ano97j] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 38(3):717, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887276/pdf.

Anonymous:1997:SI

[Ano97k] Anonymous. Subject index (1995). Technology and Culture, 38 (Supplement):206–265, ???? 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887211/pdf.

Anonymous:1997:VS

[Ano97l] Anonymous. Volume 38, 1997 supplement. Technology and Culture, 38(Supplement):??, ???? 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1997:VNJa

[Ano97m] Anonymous. Volume 38, number 1, January 1997. Technology and Culture, 38(1):??, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1997:VNA

[Ano97n] Anonymous. Volume 38, number 2, April 1997. Technology and Culture, 38(2):??, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1997:VNJb

[Ano97o] Anonymous. Volume 38, number 3, July 1997. Technology and Culture, 38(3):??, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1997:VNO

[Ano97p] Anonymous. Volume 38, number 4, October 1997. Technology and Culture, 38(4):??, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1998:A

[Ano98a] Anonymous. Awards. Technology and Culture, 39(3):512–525, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33223.

Anonymous:1998:Ia

[Ano98b] Anonymous. In this issue. Technology and Culture, 39(3):i–ii, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33214.

Anonymous:1998:Ib

[Ano98c] Anonymous. In this issue. Technology and Culture, 39(4):ix, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33268.

Anonymous:1998:IV

[Ano98d] Anonymous. Index, volume 39 (1998). Technology and Culture, 39 (4):829–848, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33272.

Anonymous:1998:ONP

[Ano98e] Anonymous. Organizational notes: The Pasadena Meeting, October 16–19, 1997. Technology and Culture, 39(3):526–537, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33219.

Anonymous:1998:VNJa

[Ano98f] Anonymous. Volume 39, number 1, January 1998. Technology and Culture, 39(1):??, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1998:VNA

[Ano98g] Anonymous. Volume 39, number 2, April 1998. Technology and Culture, 39(2):??, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1998:VNJb

[Ano98h] Anonymous. Volume 39, number 3, July 1998. Technology and Culture, 39(3):??, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1998:VNO

[Ano98i] Anonymous. Volume 39, number 4, October 1998. Technology and Culture, 39(4):??, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1999:AIa

[Ano99a] Anonymous. Author index (1996–1997). Technology and Culture, 40(3):190–215, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/168997.

Anonymous:1999:AIb

[Ano99b] Anonymous. Author index (1996–1997). Technology and Culture, 40 (Supplement):190–215, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/38023.

Anonymous:1999:Aa

[Ano99c] Anonymous. Awards. Technology and Culture, 40(3):623–637, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/169001.

Anonymous:1999:Ab

[Ano99d] Anonymous. Awards. Technology and Culture, 40(Supplement):623– 637, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33373.

Anonymous:1999:Ba

[Ano99e] Anonymous. Bibliography. Technology and Culture, 40(3):5–189, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
168999.

Anonymous:1999:Bb

[Ano99f] Anonymous. Bibliography. Technology and Culture, 40(Supplement): 5–189, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
38025.

Anonymous:1999:E

[Ano99g] Anonymous. Erratum. Technology and Culture, 40(1):207, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33317.

Anonymous:1999:Ia

[Ano99h] Anonymous. In this issue. Technology and Culture, 40(1):ix, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33325.

Anonymous:1999:Ib

[Ano99i] Anonymous. In this issue. Technology and Culture, 40(2):ix–x, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33369.

Anonymous:1999:Ic

[Ano99j] Anonymous. In this issue. Technology and Culture, 40(3):ix, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33416.

Anonymous:1999:Id

[Ano99k] Anonymous. In this issue. Technology and Culture, 40(4):ix–x, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33459.

Anonymous:1999:IV

[Ano99l] Anonymous. Index, volume 40 (1999). Technology and Culture, 40 (4):933–954, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33441.

Anonymous:1999:ONB

[Ano99m] Anonymous. Organizational notes: The Baltimore Meeting, 15-18 October 1998. Technology and Culture, 40(3):638–647, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33417.

Anonymous:1999:RTG

[Ano99n] Anonymous. Rising tide: The great Mississippi flood of 1927 and how it changed America. Technology and Culture, 40(2):375–387, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33366.

Anonymous:1999:SIa

[Ano99o] Anonymous. Subject index (1996–1997). Technology and Culture, 40(3):216–251, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/169000.

Anonymous:1999:SIb

[Ano99p] Anonymous. Subject index (1996–1997). Technology and Culture, 40 (Supplement):216–251, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/38026.

Anonymous:1999:VNJa

[Ano99q] Anonymous. Volume 40, number 1, January 1999. Technology and Culture, 40(1):??, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1999:VNA

[Ano99r] Anonymous. Volume 40, number 2, April 1999. Technology and Culture, 40(2):??, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1999:VNJb

[Ano99s] Anonymous. Volume 40, number 3, July 1999. Technology and Culture, 40(3):??, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1999:VNO

[Ano99t] Anonymous. Volume 40, number 4, October 1999. Technology and Culture, 40(4):??, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1999:VSJ

[Ano99u] Anonymous. Volume 40, supplement, July 1999. Technology and Culture, 40(Supplement):??, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Apple:1996:CBF

[App96] Rima D. Apple. Cancer from Beef: DES, Federal Food Regulation, and Consumer Confidence by Alan Marcus (review). Technology and Culture, 37(2):396–397, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887502/pdf.

Armstrong:1996:FPP

[Arm96] Christopher Armstrong. Ferranti-Packard: Pioneers in Canadian Electrical Manufacturing by Norman R. Ball, John N. Vardalas (review). Technology and Culture, 37(1):201–202, January 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887561/pdf.

Arns:1997:HVX

[Arn97] Robert G. Arns. The high-vacuum X-ray tube: Technological change in social context. Technology and Culture, 38(4):852–890, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887213/pdf.

Arrison:1994:TCC

[Arr94] John G. Arrison. Time capsule from the 17th Century: Stockholm’s Vasa Museum. Technology and Culture, 35(1):158–167, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887958/pdf.

Aspray:1993:IBE

[Asp93] William Aspray. Inside the Business Enterprise: Historical Perspectives on the Use of Information ed. by Peter Temin (review). Technology and Culture, 34(3):707–708, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888157/pdf.

Atkinson:1999:CEE

[Atk99] George Atkinson. Civil engineers and engineering in Britain, 1600– 1830. Technology and Culture, 40(1):148–150, January 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33312.

Bachrach:1990:MTG

[BA90] Bernard S. Bachrach and Rutherford Aris. Military technology and garrison organization: Some observations on Anglo-Saxon military thinking in light of the Burghal Hidage. Technology and Culture, 31(1):1–17, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888916/pdf.

Bachrach:1990:GTW

[Bac90] Bernard S. Bachrach. Geography, Technology, and War: Studies in the Maritime History of the Mediterranean, 649–1571 by John H. Pryor (review). Technology and Culture, 31(1):183–185, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888957/pdf.

Bachrach:1998:IWE

[Bac98] Bernard S. Bachrach. Infantry Warfare in the Early Fourteenth Century: Discipline, Tactics, and Technology by Kelly DeVries (review). Technology and Culture, 39(2):362–363, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887164/pdf.

Bacay:1999:EEM

[Bac99] Virginia S. Bacay. The economic emergence of modern Japan. Technology and Culture, 40(1):169–171, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33315.

Badash:1991:ABS

[Bad91] Lawrence Badash. Atomic Bomb Scientists: Memoirs, 1939–1945 ed. by Joseph J. Ermenc (review). Technology and Culture, 32(2 (part 1)):448–450, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888777/pdf.

Badash:1994:GGC

[Bad94] Lawrence Badash. Guarding the Guardians: Civilian Control of Nuclear Weapons in the United States by Peter Douglas Feaver (review). Technology and Culture, 35(1):216–217, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887983/pdf.

Ball:1997:BHC

[Bal97] Norman R. Ball. A Bibliography of the History of Canadian Science and Technology by Arnold E. Roos (review). Technology and Culture, 38(4):1013–1014, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887259/pdf.

Baldwin:1998:IWE

[Bal98] Martha Baldwin. The Invisible World: Early Modern Philosophy and the Invention of the Microscope (review). Technology and Culture, 39(3):560–562, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33210.

Balzer:1999:NAR

[Bal99] Harley D. Balzer. New Atlantis Revisited: Akademgorodok, the Siberian City of Science (review). Technology and Culture, 40 (4):915–917, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33423.

Barbezat:1990:CEH

[Bar90] Daniel Barbezat. A Concise Economic History of the World: From Paleolithic Times to the Present by Rondo Cameron (review). Technology and Culture, 31(4):864–865, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901657/pdf.

Barke:1991:STS

[Bar91a] Richard P. Barke. The Space Telescope: a Study of NASA, Science, Technology, and Politics by Robert W. Smith (review). Technology and Culture, 32(3):637–639, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888723/pdf.

Bartky:1991:CRM

[Bar91b] Ian R. Bartky. Comment and response on “The Most Reliable Time”. Technology and Culture, 32(1):183–185, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888825/pdf. See [Ste89].

Barry:1992:TBT

[Bar92a] Andrew Barry. Technologies without Boundaries: On Telecommunications in a Global Age by Ithiel de Sola Pool (review). Technology and Culture, 33(2):399–400, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888331/pdf.

Barry:1992:VES

[Bar92b] Andrew Barry. Viewing the Earth: The Social Construction of the Landsat Satellite System by Pamela E. Mack (review). Technology and Culture, 33(3):647–648, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888288/pdf.

Barth:1992:WWI

[Bar92c] Gunther Barth. World War II and the West: Reshaping the Economy by Gerald D. Nash (review). Technology and Culture, 33(2):397– 399, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888330/pdf.

Basalla:1992:TWC

[Bas92a] George Basalla. Technology in World Civilization: a Thousand-Year History by Arnold Pacey (review). Technology and Culture, 33(2):347–348, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888298/pdf.

Bastable:1992:BMG

[Bas92b] Marshall J. Bastable. From breechloaders to monster guns: Sir William Armstrong and the invention of modern artillery, 1854– 1880. Technology and Culture, 33(2):213–247, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888291/pdf.

Basalla:1996:FSR

[Bas96] George Basalla. From Faust to Strangelove: Representations of the Scientist in Western Literature by Roslynn D. Haynes (review). Technology and Culture, 37(3):623–624, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887446/pdf.

Basalla:1997:IIE

[Bas97] George Basalla. Instruments and the Imagination ed. by Thomas L. Hankins and Robert J. Silverman (review). Technology and Culture, 38(1):233–235, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887375/pdf.

Bassett:1999:EFV

[Bas99] Ross Knox Bassett. Endless Frontier: Vannevar Bush, Engineer of the American Century. Technology and Culture, 40(3):685–686, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33413.

Batzli:1990:HHN

[Bat90] Samuel A. Batzli. From heroes to Hiroshima: The National Air and Space Museum adjusts its point of view. Technology and Culture, 31(4):830–837, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901653/pdf.

Batty:1998:DHR

[Bat98] David Batty. Databases in Historical Research: Theory, Methods, and Applications by Charles Harvey and Jon Press (review). Technology and Culture, 39(1):144–146, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887186/pdf.

Baur:1993:HHC

[Bau93] John E. Baur. Henry E. Huntington and the Creation of Southern California by William B. Friedricks (review). Technology and Culture, 34(2):447–448, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888439/pdf.

Bauer:1999:LTS

[Bau99] Johannes M. Bauer. The Liberalization of Telecommunications in Sweden: Technology and Regime Change from the 1960s to 1993 (review). Technology and Culture, 40(4):926–927, October 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33426.

Beckwith:1991:RPT

[Bec91] Guy V. Beckwith. Research in Philosophy and Technology, vol. 9: Ethics and Technology ed. by Carl Mitcham, Frederick Ferré (review). Technology and Culture, 32(4):1137–1140, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888675/pdf.

Bedini:1990:DIM

[Bed90a] Silvio A. Bedini. Drawing Instruments 1580–1980 by Maya Hambly (review). Technology and Culture, 31(3):548–550, July 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888866/pdf.

Bedini:1990:SCH

[Bed90b] Silvio A. Bedini. Selected, Contributions to the History of Cartography and Scientific Instruments by Marcel Destombes (review). Technology and Culture, 31(3):550–551, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888867/pdf.

Bedini:1990:MAS

[Bed90c] Silvio A. Bedini. The Mariner’s Astrolabe: a Survey of Known, Surviving Sea Astrolabes by Alan Stimson (review). Technology and Culture, 31(4):867–869, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901659/pdf.

Bedini:1991:ESI

[Bed91] Silvio A. Bedini. Early Scientific Instruments: Europe 1400–1800 by Anthony Turner (review). Technology and Culture, 32(1):169–170, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888817/pdf.

Belofsky:1991:EDU

[Bel91] Harold Belofsky. Engineering drawing — a universal language in two dialects. Technology and Culture, 32(1):23–46, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888784/pdf. See

comment and response [MB92].

Belasco:1997:ABH

[Bel97] Warren Belasco. Algae burgers for a hungry world? The rise and fall of Chlorella cuisine. Technology and Culture, 38(3):608–634, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887269/pdf.

Belhoste:1998:LPD

[Bel98a] Bruno Belhoste. L’étonnant parcours du republicain J. H. Hassenfratz (1755–1827): Du faubourg Montmartre au Corps des Mines (review). Technology and Culture, 39(4):766–767, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33245.

Bellesiles:1998:CBS

[Bel98b] Michael A. Bellesiles. Colonel Burton’s Spiller & Burr Revolver: an Untimely Venture in Confederate Small-Arms Manufacturing (review). Technology and Culture, 39(4):780–782, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33252.

Benhamou:1990:PPB

[Ben90] Reed Benhamou. Privilege and Profit: a Business Family in Eighteenth-Century France by P. W. Bamford (review). Technology and Culture, 31(3):528–530, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888855/pdf.

Benhamou:1991:JDH

[Ben91a] Reed Benhamou. Journal of Design History ed. by Christopher Bailey (review). Technology and Culture, 32(1):135–136, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888797/pdf.

Bennett:1991:TLP

[Ben91b] Stuart Bennett. Technology and the Labour Process: Australasian Case Studies ed. by Evan Willis (review). Technology and Culture, 32(2 (part 1)):429–430, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888765/pdf.

Bennett:1991:IIM

[Ben91c] Stuart Bennett. “The industrial instrument — master of industry, servant of management”: Automatic control in the process industries, 1900–1940. Technology and Culture, 32(1):69–81, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888786/pdf.

Bender:1992:LMP

[Ben92a] Thomas Bender. Lewis Mumford: Public Intellectual ed. by Thomas P. Hughes and Agatha C. Hughes (review). Technology and Culture, 33(2):388–389, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888325/pdf.

Benfey:1992:ZTE

[Ben92b] Otto Theodor Benfey. Zionism, and Technocracy: The Engineering of Jewish Settlement in Palestine, 1870–1918 by Derek J. Penslar (review). Technology and Culture, 33(4):830–831, October 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888104/pdf.

Bennett:1992:NCM

[Ben92c] Stuart Bennett. Numerical Control: Making a New Technology by J. Francis Reintjes (review). Technology and Culture, 33(4):841–842, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888111/pdf.

Benhamou:1994:ALP

[Ben94] Reed Benhamou. The artificial limb in preindustrial France. Technology and Culture, 35(4):835–845, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887723/pdf.

Bennett:1996:WMC

[Ben96] Stuart Bennett. What Machines Can’t Do: Politics and Technology in the Industrial Enterprise by Robert J. Thomas (review). Technology and Culture, 37(1):207–208, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887564/pdf.

Bennett:1999:BED

[Ben99] S. Bennett. Brainchildren: Essays on designing minds. Technology and Culture, 40(3):720–722, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33387.

Bernstein:1990:BST

[Ber90] Barton J. Bernstein. Better a Shield than a Sword: Perspectives on Defense and Technology by Edward Teller (review). Technology and Culture, 31(4):846–861, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901655/pdf.

Berger:1993:WER

[Ber93] Michael L. Berger. Wheel Estate: The Rise and Decline of Mobile Homes by Allan D. Wallis (review). Technology and Culture, 34(2):452–453, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888442/pdf.

Berger:1994:EAD

[Ber94a] Molly W. Berger. Expanding the American Dream: Building and Rebuilding Levittown by Barbara M. Kelly (review). Technology and Culture, 35(4):878–879, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887733/pdf.

Berger:1994:SPH

[Ber94b] Molly W. Berger. Shared Pleasures: a History of Movie Presentation in the United States by Douglas Gomery (review). Technology and Culture, 35(1):199–201, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887974/pdf.

Berger:1995:HST

[Ber95a] Molly W. Berger. The History of Science and Technology in the United States: a Critical and Selective Bibliography — Volume II by Marc Rothenberg (review). Technology and Culture, 36(3):731– 732, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887832/pdf.

Berger:1995:ITH

[Ber95b] Molly W. Berger. International themes in the history of technology — ICOHTEC symposium, bath, July–August 1994. Technology and Culture, 36(3):676–679, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887805/pdf.

Berk:1995:PIE

[Ber95c] Gerald Berk. Politics and Industrialization: Early Railroads in the United States and Prussia by Colleen A. Dunlavy (review). Technology and Culture, 36(3):703–705, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887817/pdf.

Berger:1997:FAC

[Ber97a] Molly W. Berger. Free to All: Carnegie Libraries and American Culture, 1890–1920 by Abigail A. Van Slyck (review). Technology and Culture, 38(2):521–523, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887347/pdf.

Berner:1997:ACP

[Ber97b] R. Thomas Berner. The ancient Chinese process of reprography. Technology and Culture, 38(2):424–431, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887327/pdf.

Berkowitz:1998:NNT

[Ber98] Michael A. Berkowitz. The New Niagara: Tourism, Technology, and the Landscape of Niagara Falls, 1776–1917 by William Irwin (review). Technology and Culture, 39(1):171–173, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887202/pdf.

Bess:1995:EAS

[Bes95] Michael D. Bess. Ecology and artifice: Shifting perceptions of nature and high technology in postwar France. Technology and Culture, 36 (4):830–862, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887755/pdf.

Bianco:1997:LME

[Bia97] Martha J. Bianco. Lewis Mumford and the Ecological Region: The Politics of Planning by Mark Luccarelli (review). Technology and Culture, 38(4):990–992, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887243/pdf.

Biggs:1991:AAT

[Big91] Lindy Biggs. America and the Automobile: Technology, Reform and Social Change, 1893–1923 by Peter J. Ling (review). Technology and Culture, 32(3):627–628, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888716/pdf.

Biggs:1994:MRS

[Big94] Lindy Biggs. A Mental Revolution: Scientific Management since Taylor ed. by Daniel Nelson (review). Technology and Culture, 35(2):427–428, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887940/pdf.

Biggs:1995:CAH

[Big95a] Lesley Biggs. Chiropractic in America: The History of a Medical Alternative by J. Stuart Moore (review). Technology and Culture, 36(2):440–442, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887912/pdf.

Biggs:1995:EF

[Big95b] Lindy Biggs. The engineered factory. Technology and Culture, 36 (2S):S174–S188, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887713/pdf.

Biggs:1996:TLN

[Big96] Lindy Biggs. Taylored Lives: Narrative Productions in the Age of Taylor, Veblen, and Ford by Martha Banta (review). Technology and Culture, 37(4):846–848, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887417/pdf.

Biggs:1997:BWP

[Big97] Lindy Biggs. Building the Workingman’s Paradise: The Design of American Company Towns by Margaret Crawford (review). Technology and Culture, 38(4):988–990, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887242/pdf.

Bilstein:1997:BBJ

[Bil97] Roger E. Bilstein. Building the B-29 by Jacob Vander Meulen, and: Courage and Air Warfare: The Allied Aircrew Experience in the Second World War by Mark K. Wells (review). Technology and Culture, 38(3):788–789, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887306/pdf.

Bilstein:1998:FDY

[Bil98] Roger E. Bilstein. Flights of Discovery: 50 years at the NASA Dryden Flight Research Center (review). Technology and Culture, 39(3):588–589, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33208.

Bindocci:1991:WSN

[Bin91] Gay Bindocci. Where the Sun Never Shines: a History of America’s Bloody Coal Industry by Priscilla Long (review). Technology and Culture, 32(3):610–611, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888706/pdf.

Bindocci:1993:SPE

[Bin93] Cynthia Gay Bindocci. In the Servitude of Power: Energy and Civilisation through the Ages by Jean-Claude Debeir, et al. (review). Technology and Culture, 34(3):708–709, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888158/pdf.

Birr:1990:SCS

[Bir90] Kendall Birr. Science and Corporate Strategy: Du Pont and R&D, 1902–1980 by David A. Hounshell, John Kenly Smith, Jr. (review). Technology and Culture, 31(2):338–340, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888907/pdf.

Bix:1998:LTW

[Bix98] Amy Sue Bix. Life in a Technocracy: What It Might Be Like by Harold Loeb (review). Technology and Culture, 39(1):126–128, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887176/pdf.

Bix:1999:TAM

[Bix99] Amy Sue Bix. Talking About Machines: an Ethnography of a Modern Job, and: Between Craft and Science: Technical Work in U.S. Settings (review). Technology and Culture, 40(4):883–885, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33427.

Bjorck:1997:AVO

[Bjö97] Henrik Björck. The Atlantic Vision: Olaus Rudbeck and Baroque Science by Gunnar Eriksson, and: Svensk ingenjörskonst under stormaktstiden: Olof Rudbecks tekniska undervisning och praktiska verksamhet by Per Dahl (review). Technology and Culture, 38(1):246–248, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887382/pdf.

Blackford:1990:RCC

[Bla90a] Mansel G. Blackford. The Rise of the Corporate Commonwealth: U.S. Business and Public Policy in the Twentieth Century by Louis Galambos, Joseph Pratt (review). Technology and Culture, 31 (1):132–133, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888927/pdf.

Blackwelder:1990:MFL

[Bla90b] Julia Kirk Blackwelder. Mill Family: The Labor System in the Southern Cotton Textile Industry, 1880–1915 by Cathy L. McHugh, and: Like a Family: The Making of a Southern Cotton Mill World by Jacquelyn Hall, et al. (review). Technology and Culture, 31(2):325– 327, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888900/pdf.

Blake:1991:LM

[Bla91] Casey Blake. Lewis Mumford (1895–1990). Technology and Culture, 32(1):187–190, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888827/pdf.

Blaszczyk:1992:PLW

[Bla92a] Regina Lee Blaszczyk. Patterns of Labour: Work and Social Change in the Pottery Industry by Richard Whipp (review). Technology and Culture, 33(3):611–613, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888266/pdf.

Blaszczyk:1992:UOV

[Bla92b] Regina Lee Blaszczyk. The Upper Ohio Valley’s industrial heritage: The East Liverpool Museum of Ceramics. Technology and Culture, 33(4):788–796, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888083/pdf.

Blaszczyk:1995:RRH

[Bla95] Regina Lee Blaszczyk. “Reign of the Robots”: The Homer Laughlin China Company and flexible mass production. Technology and Culture, 36(4):863–911, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887756/pdf.

Black:1996:SMF

[Bla96a] Michael Black. Slide Mountain, or the Folly of Owning Nature by Theodore Steinberg (review). Technology and Culture, 37(3):628– 629, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887449/pdf.

Blaszczyk:1996:PIT

[Bla96b] Regina Lee Blaszczyk. The Pottery Industry of Trenton: a Skilled Trade in Transition, 1850–1929 by Marc Jeffrey Stern (review). Technology and Culture, 37(3):637–639, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887454/pdf.

Black:1997:PVN

[Bla97a] Brian Black. Paradise Valley, Nevada: The People and Buildings of an American Place by Howard Wight Marshall (review). Technology and Culture, 38(4):995–996, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887247/pdf.

Blackford:1997:WAC

[Bla97b] Mansel G. Blackford. Wellsprings of Achievement: Cultural and Economic Dynamics in Early Modern England and Japan ed. by Penelope Gouk (review). Technology and Culture, 38(1):248–249, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887383/pdf.

Blaszczyk:1997:ACR

[Bla97c] Regina Lee Blaszczyk. Acknowledging Consumption: a Review of New Studies ed. by Daniel Miller, and: The Politics of Domestic Consumption: Critical Readings ed. by Stevi Jackson and Shaun Moores (review). Technology and Culture, 38(2):493–497, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887334/pdf.

Blaszczyk:1999:MCW

[Bla99a] Regina Lee Blaszczyk. Material Cultures: Why Some Things Matter (review). Technology and Culture, 40(4):874–875, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33429.

Blaszczyk:1999:RFE

[Bla99b] Regina Lee Blaszczyk. Real fantasies: Edward Steichen’s advertising photography. Technology and Culture, 40(3):705–707, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33392.

Blewett:1991:WWS

[Ble91] Mary H. Blewett. Women, Work, and Sexual Politics in Eighteenth-Century England by Bridget Hill (review). Technology and Culture, 32(2 (part 1)):424–426, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888762/pdf.

Bloomfield:1990:ATB

[Blo90] Gerald T. Bloomfield. American Technology and the British Motor Vehicle Industry by Wayne Lewchuk (review). Technology and Culture, 31(1):168–169, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888948/pdf.

Bloemen:1991:HVP

[Blo91] Erik Bloemen. The History of N. V. Philips’ Gloeilampenfabrieken, vol. 2: a Company of Many Parts by A. Heerding (review). Technology and Culture, 32(4):1110–1111, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888659/pdf.

Blumtritt:1992:GTE

[Blu92] Oskar Blumtritt. Geschichte der Technikwissenschaften ed. by Gisela Buchheim, Rolf Sonnemann (review). Technology and Culture, 33 (1):154–156, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888349/pdf.

Blum:1999:WDT

[Blu99] Martina Blum. The Werkbund: Design theory and mass culture before the First World War. Technology and Culture, 40(2):446–448, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33356.

Bodker:1999:PPE

[Bod99] Henrik Bodker. Pavilions of plenty: Exhibiting American culture abroad in the 1950s. Technology and Culture, 40(1):182–183, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33286.

Bogin:1998:VVA

[Bog98] Mary Bogin. Vitamania: Vitamins in American Culture (review). Technology and Culture, 39(4):804–806, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33225.

Bolton:1990:TRN

[Bol90] Roger Bolton. Technology and the Rise of the Networked City in Europe and America ed. by Joel A. Tarr, Gabriel Dupuy (review). Technology and Culture, 31(2):299–301, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888884/pdf.

Borish:1997:BCH

[Bor97] Linda J. Borish. Born in the Country: a History of Rural America by David B. Danbom (review). Technology and Culture, 38(3):799– 801, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887312/pdf.

Borg:1999:CPE

[Bor99] Kevin L. Borg. The “Chauffeur Problem” in the Early Auto Era: Structuration theory and the users of technology. Technology and Culture, 40(4):797–832, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33428.

Bowditch:1992:PSH

[Bow92] John Bowditch. Power from Steam: a History of the Stationary Steam Engine by Richard L. Hills (review). Technology and Culture, 33(4):821–823, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888098/pdf.

Bowker:1998:MRG

[Bow98] Geoffrey C. Bowker. Modest reviewer goes on virtual voyage: Some recent literature of cyberspace. Technology and Culture, 39(3):499– 511, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33213.

Bowditch:1999:GN

[Bow99a] John Bowditch. A Gothic novelty. Technology and Culture, 40 (1):99–101, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33273.

Bowers:1999:TEE

[Bow99b] Brian Bowers. The Thames Embankment: Environment, Technology, and Society in Victorian London (review). Technology and Culture, 40(4):894–895, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33430.

Bowers:1999:FQ

[Bow99c] Brian Bowers. The fusion quest. Technology and Culture, 40(2):416– 417, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33337.

Bradley:1991:CDS

[BP91] Margaret Bradley and Fernand Perrin. Charles Dupin’s study visits to the British Isles, 1816–1824. Technology and Culture, 32(1):47–68, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888785/pdf.

Braun:1990:TBZ

[Bra90] Hans-Joachim Braun. Technische Bildung zwischen Wissenschaft und Praxis: Die Technische Hochschule Braunschweig 1862–1914 by Helmuth Albrecht, and: Ingenieurausbildung im Königreich Württemberg by Gerhard Zweckbronner (review). Technology and Culture, 31(3):536–537, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888859/pdf.

Bradley:1991:DTW

[Bra91a] Joseph Bradley. Description of the Tula Weapon Factory in Regard to Historical and Technical Aspects by Iosif Gamel (review). Technology and Culture, 32(2 (part 1)):442–443, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888773/pdf.

Brannigan:1991:EAP

[Bra91b] Vincent M. Brannigan. Environmental Accidents: Personal Injury and Public Responsibility by Richard H. Gaskins (review). Technology and Culture, 32(4):1145–1146, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888679/pdf.

Bradley:1992:RTI

[Bra92] Margaret Bradley. The Role of Transportation in the Industrial Revolution: a Comparison of England and France by Rick Szostak (review). Technology and Culture, 33(4):816–817, October 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888095/pdf.

Bradley:1993:IRP

[Bra93] Margaret Bradley. The Industrial Revolution by Pat Hudson (review). Technology and Culture, 34(3):683–684, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888144/pdf.

Braun:1995:SMA

[Bra95] Hans-Joachim Braun. The Science Museum’s Aeronautics Gallery redisplayed. Technology and Culture, 36(3):625–629, July 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887792/pdf.

Braden:1996:AAH

[Bra96] Donna R. Braden. The Arts in the American Home, 1890–1930 ed. by Jessica H. Foy, Karal Ann Marling (review). Technology and Culture, 37(2):375–377, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887491/pdf.

Braun:1997:RCI

[Bra97] Hans-Joachim Braun. Rationalizing Culture: IRCAM, Boulez, and the Institutionalization of the Musical Avant-Garde by Georgina Born (review). Technology and Culture, 38(3):803–804, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887314/pdf.

Braden:1998:WHI

[Bra98] Donna R. Braden. Whose history is it? Planning Henry Ford Museum’s Clockwork Exhibit. Technology and Culture, 39(3):489–498, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33211.

Brentjes:1999:EHA

[Bre99] Sonja Brentjes. Encyclopedia of the history of Arabic science. Technology and Culture, 40(2):399–401, April 1999. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33352.

Brittain:1990:HLE

[Bri90] James E. Brittain. A History of Linear Electric Motors by Eric R. Laithwaite (review). Technology and Culture, 31(2):337–338, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888906/pdf.

Briggs:1992:IPV

[Bri92] Asa Briggs. The Invention of Progress: The Victorians and the Past by Peter J. Bowler (review). Technology and Culture, 33(1):163–164, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888354/pdf.

Brieger:1993:YFP

[Bri93] Gert H. Brieger. Yellow Fever and Public Health in the New South by John H. Ellis (review). Technology and Culture, 34(4):928–929, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888047/pdf.

Briggs:1997:PP

[Bri97] J. Morton Briggs. Pollution in poullaouen. Technology and Culture, 38(3):635–654, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887270/pdf.

Brownell:1990:TCA

[Bro90] Blaine A. Brownell. Two Centuries of American Planning ed. by Daniel Schaffer (review). Technology and Culture, 31(2):301–302, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888885/pdf.

Brock:1992:SIW

[Bro92a] William H. Brock. A Science of Impurity: Water Analysis in Nineteenth Century Britain by Christopher Hamlin (review). Technology and Culture, 33(1):164–165, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888355/pdf.

Brown:1992:BBA

[Bro92b] John K. Brown. Bending Is Not Breaking: Adaptation and Persistence among 19th Century Lancaster Artisans by Thomas R. Winpenny (review). Technology and Culture, 33(2):377–378, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888317/pdf.

Bromberg:1993:MBJ

[Bro93a] Joan Lisa Bromberg. The Maxwellians by Bruce J. Hunt (review). Technology and Culture, 34(1):186–187, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888491/pdf.

Brown:1993:IUI

[Bro93b] Jonathan C. Brown. Industry and Underdevelopment: The Industrialization of Mexico, 1890–1940 by Stephen H. Haber (review). Technology and Culture, 34(1):179–180, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888487/pdf.

Brooks:1995:SRS

[Bro95a] Nathan M. Brooks. Science in Russia and the Soviet Union: a Short History by Loren R. Graham (review). Technology and Culture, 36(3):725–727, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887829/pdf.

Brown:1995:AGR

[Bro95b] John K. Brown. “America’s Great Road” at the B&O Railroad Museum. Technology and Culture, 36(3):641–645, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887795/pdf.

Brown:1995:GRB

[Bro95c] John K. Brown. The Great Road: The Building of the Baltimore & Ohio, the Nation’s First Railroad, 1828–1853 by James D. Dilts (review). Technology and Culture, 36(1):189–191, January 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888020/pdf.

Brown:1996:IAN

[Bro96] John K. Brown. Industrializing America: The Nineteenth Century by Walter Licht (review). Technology and Culture, 37(4):838–840, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887412/pdf.

Bromberg:1997:OUS

[Bro97a] Joan L. Bromberg. Organizing for the Use of Space: Historical Perspectives on a Persistent Issue ed. by Roger D. Launius (review). Technology and Culture, 38(4):973–974, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887229/pdf.

Broussard:1997:MSA

[Bro97b] Joyce L. Broussard. Main Street Amusements: Movies and Commercial Entertainment in a Southern City, 1896–1930 by Gregory A. Waller (review). Technology and Culture, 38(3):771–773, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887297/pdf.

Broussard:1998:PSS

[Bro98a] Joyce L. Broussard. At the Picture Show: Small-Town Audiences and the Creation of Movie Fan Culture (review). Technology and Culture, 39(4):803–804, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33241.

Brown:1998:RDP

[Bro98b] Elspeth Brown. Records of the Dawn of Photography: Talbot’s Notebooks P and Q by Larry J. Schaaf (review). Technology and Culture, 39(1):153–155, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887191/pdf.

Brouwer:1999:GMN

[Bro99a] Christien Brouwer. Geschlechterverhaltnisse in Medizin, Naturwissenschaft und Technik. Technology and Culture, 40(2):433–434, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33346.

Brown:1999:EAR

[Bro99b] John K. Brown. Early American railroads: Franz Anton Ritter von Gerstner’s Die innern Communicationen (1842–1843). Technology and Culture, 40(1):142–144, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33283.

Bruggemeier:1994:HNN

[Brü94] Franz-Josef Brüggemeier. Herrschaft über die Natur? Naturvorstellungen deutscher Ingenieure, 1871–1914 by Hans-Liudger Dienel (review). Technology and Culture, 35(4):902–904, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887747/pdf.

Bryant:1992:ESE

[Bry92] Lynwood Bryant. Engineering as a Social Enterprise ed. by Hedy E. Sladovich (review). Technology and Culture, 33(4):846–848, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888114/pdf.

Bryant:1995:DEV

[Bry95] Lynwood Bryant. Diesel’s Engine, vol. 1: From Conception to 1918 by C. Lyle Cummins, Jr. (review). Technology and Culture, 36(3):694–696, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887812/pdf.

Bryant:1996:NAR

[Bry96] Keith L. Bryant, Jr. The North American Railroad: Its Origin, Evolution, and Geography by James E. Vance, Jr. (review). Technology and Culture, 37(4):843–845, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887415/pdf.

Bryden:1993:AAS

[BS93] D. J. Bryden and D. L. Simms. Archimedes as an advertising symbol. Technology and Culture, 34(2):387–391, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888414/pdf.

Biddle:1994:CRB

[BS94] Wayne Biddle and Eric Schatzberg. Comment on the review of Barons of the Sky. Technology and Culture, 35(2):450–451, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

887952/pdf. See [Sch92a].

Buchanan:1991:TNH

[Buc91] R. A. Buchanan. Theory and narrative in the history of technology. Technology and Culture, 32(2 (part 1)):365–376, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888742/pdf.

Buchanan:1992:CHH

[Buc92] R. Angus Buchanan. Capturing the Horizon: The Historical Geography of Transportation since the Sixteenth Century by James E. Vance, Jr. (review). Technology and Culture, 33(2):348–349, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888299/pdf.

Buchanan:1994:CTA

[Buc94] R. Angus Buchanan. The Company Town: Architecture and Society in the Early Industrial Age ed. by John S. Garner (review). Technology and Culture, 35(3):619–620, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887856/pdf.

Buchhaupt:1997:KEE

[Buc97] Siegfried Buchhaupt. Konzepte von Elektrizitätsversorgung und Elektrizitätswirtschaft: Die Entstehung eines neuen Fachgebietes der Technikwissenschaften zwischen 1880 und 1945 by Norbert Gilson (review). Technology and Culture, 38(3):779–781, July 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887301/pdf.

Buchanan:1998:IAS

[Buc98a] R. Angus Buchanan. The Industrial Archaeology of Shropshire by Barrie Trinder (review). Technology and Culture, 39(2):319–321, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887142/pdf.

Bucholz:1998:RRP

[Buc98b] Arden Bucholz. The Rocket and the Reich: Peenemunde and the Coming of the Ballistic Missile Era (review). Technology and Culture, 39(3):574–577, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33197.

Buchanan:1999:SSC

[Buc99] R. A. Buchanan. Samuel Smiles and the construction of Victorian values. Technology and Culture, 40(1):147–148, January 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33289.

Bud:1994:MVL

[Bud94] Robert Bud. The Molecular Vision of Life: Caltech, the Rockefeller Foundation, and the Rise of the New Biology by Lily E. Kay (review). Technology and Culture, 35(4):883–885, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887736/pdf.

Bud:1996:MPD

[Bud96] Robert Bud. Molecular Politics: Developing American and British Regulatory Policy for Genetic Engineering, 1972–1982 by Susan Wright (review). Technology and Culture, 37(2):397–399, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887503/pdf.

Bullough:1991:CWA

[Bul91] Vern L. Bullough. The Conquest of Water: The Advent of Health in the Industrial Age by Jean-Pierre Goubert (review). Technology and Culture, 32(2 (part 1)):401–403, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888749/pdf.

Bulkeley:1994:SCE

[Bul94] Rip Bulkeley. The “Sputnik” Challenge: Eisenhower’s Response to the Soviet Satellite by Robert A. Divine (review). Technology and Culture, 35(1):214–215, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887982/pdf.

Burstyn:1992:UCS

[Bur92] Harold L. Burstyn. U.S. Coast Survey vs. Naval Hydrographic Office: a 19th-Century Rivalry in Science and Politics by Thomas G. Manning (review). Technology and Culture, 33(2):369–370, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888312/pdf.

Butrica:1992:MHF

[But92] Andrew J. Butrica. Le moteur hydraulique en France au XIX^{e} siècle: Concepteurs, inventeurs et constructeurs by Bruno Belhoste, et al. (review). Technology and Culture, 33(2):356–358, April 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888304/pdf.

Calhoun:1992:JBJ

[Cal92a] Daniel Calhoun. John B. Jervis: an American Engineering Pioneer by F. Daniel Larkin (review). Technology and Culture, 33(2):360– 361, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888306/pdf.

Callenbach:1992:ERE

[Cal92b] Ernest Callenbach. Earth Rising: Ecological Belief in an Age of Science by David Oates (review). Technology and Culture, 33(2): 387, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888324/pdf.

Caldwell:1998:HDC

[Cal98] Peter Caldwell. The History of the Distillers Company, 1877– 1939: Diversification and Growth in Whiskey and Chemicals (review). Technology and Culture, 39(3):573–574, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33209.

Cameron:1992:EAW

[Cam92] Rondo Cameron. Europe, America, and the Wider World: Essays on the Economic History of Western Capitalism, vol. 2: America and the Wider World by William N. Parker (review). Technology and Culture, 33(4):805–806, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888089/pdf.

Cameron:1993:WET

[Cam93] Ardis Cameron. Work Engendered: Toward a New History of American Labor ed. by Ava Baron (review). Technology and Culture, 34(3):705–707, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888156/pdf.

Cameron:1995:AAN

[Cam95] Ardis Cameron. Aspirations and Anxieties: New England Workers and the Mechanized Factory System, 1815–1850 by David A. Zonderman (review). Technology and Culture, 36(2):410–412, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887894/pdf.

Cantelon:1995:OMT

[Can95] Philip L. Cantelon. The origins of microwave telephony — waves of change. Technology and Culture, 36(3):560–582, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887790/pdf.

Caporael:1990:AFF

[Cap90a] Linnda Caporael. Atanasoff: Forgotten Father of the Computer by Clark R. Mollenhoff, and: The First Electronic Computer: The Atanasoff Story by Alice R. Burks, Arthur W. Burks (review). Technology and Culture, 31(1):199–201, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888966/pdf.

Capshew:1990:SCA

[Cap90b] James H. Capshew. A Shuttle Chronology 1964–1973: Abstract Concepts to Letter Contracts by John F. Guilmartin, Jr. and John Walker Mauer (review). Technology and Culture, 31(4):899–900, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901679/pdf.

Caputi:1991:CWM

[Cap91] Jane Caputi. Cyborg Worlds: The Military Information Society ed. by Les Levidow, Kevin Robins (review). Technology and Culture, 32 (4):1131–1132, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888672/pdf.

Capshew:1993:EBP

[Cap93] James H. Capshew. Engineering behavior: Project Pigeon, World War II, and the conditioning of B. F. Skinner. Technology and Culture, 34(4):835–857, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888041/pdf.

Capshew:1995:BFS

[Cap95] James H. Capshew. B. F. Skinner: a Life by Daniel W. Bjork (review). Technology and Culture, 36(2):427–428, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887904/pdf.

Carpenter:1991:UTT

[Car91] Stanley R. Carpenter. Under Technology’s Thumb by William Leiss (review). Technology and Culture, 32(4):1143–1144, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888678/pdf.

Carney:1996:LTT

[Car96] Judith Carney. Landscapes of technology transfer: Rice cultivation and African continuities. Technology and Culture, 37(1):5–35, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887535/pdf.

Carlisle:1997:PRA

[Car97] Rodney P. Carlisle. Probabilistic risk assessment in nuclear reactors: Engineering success, public relations failure. Technology and Culture, 38(4):920–941, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887215/pdf.

Cardoza:1998:IIL

[Car98] Anthony L. Cardoza. Italian Industrialists from Liberalism to Fascism: The Political Development of the Industrial Bourgeoisie, 1906– 34 by Franklin Hugh Adler (review). Technology and Culture, 39 (2):348–350, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887157/pdf.

Carlson:1999:NIV

[Car99a] W. Bernard Carlson. Networks of innovation: Vaccine development at Merck, Sharp & Dohme, and Mulford, 1895–1995. Technology and Culture, 40(3):683–684, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33390.

Carlsson:1999:NHO

[Car99b] Anders Carlsson. Natverksentreprenoren: En historia om teknisk forskning och industriellt utvecklingsarbete fran den Malmska utredningen till Styrelsen for teknisk utveckling. (Swedish) [The network entrepreneur: a history of technical research and industrial development work from the Malmö investigation to the Board for Technical Development]. Technology and Culture, 40(2):397–399, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33362.

Casey:1998:RDB

[Cas98a] Robert Casey. The Rise and Decline of the British Motor Industry (review). Technology and Culture, 39(4):789–791, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33234.

Castaneda:1998:KWL

[Cas98b] Christopher J. Castaneda. “King of the Wildcatters”: The Life and Times of Tom Slick, 1883–1930 by Ray Miles (review). Technology and Culture, 39(2):311–313, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887138/pdf.

Casey:1999:VC

[Cas99] Robert Casey. The Vanderbilt cup, 1908. Technology and Culture, 40(2):359–362, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33364.

Caws:1995:TTT

[Caw95] Peter Caws. Thinking through Technology: The Path between Engineering and Philosophy by Carl Mitcham (review). Technology and Culture, 36(4):1060–1061, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887785/pdf.

Carlton:1995:USQ

[CC95] David L. Carlton and Peter A. Coclanis. The Uninventive South? A quantitative look at region and American inventiveness. Technology and Culture, 36(2):302–326, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887882/pdf.

Ceruzzi:1997:BGH

[Cer97] Paul Ceruzzi. A Bibliographic Guide to the History of Computer Applications, 1950–1990 by James W. Cortada (review). Technology and Culture, 38(4):1012–1013, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887258/pdf.

Ceruzzi:1998:TCT

[Cer98] Paul E. Ceruzzi. Transforming Computer Technology: Information Processing for the Pentagon, 1962–1986 (review). Technology and Culture, 39(3):596–598, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33198.

Chambers:1992:AMP

[CF92] Wade Chambers and Rachel Faggetter. Australia’s museum powerhouse. Technology and Culture, 33(3):548–559, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888243/pdf.

Channell:1990:TEP

[Cha90] David F. Channell. Thomas A. Edison Papers. A Selective Microfilm Edition: Part II (1879–1886) ed. by Thomas E. Jeffrey (review). Technology and Culture, 31(2):332–333, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888903/pdf.

Channell:1991:STC

[Cha91] David Channell. Society and Technological Change by Rudi Volti (review). Technology and Culture, 32(4):1150–1151, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888682/pdf.

Channell:1993:ANE

[Cha93] David Channell. Action into Nature: an Essay on the Meaning of Technology by Barry Cooper (review). Technology and Culture, 34(3):711–713, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888159/pdf.

Channell:1995:HET

[Cha95] David Channell. Hand’s End: Technology and the Limits of Nature by David Rothenberg (review). Technology and Culture, 36(2):438– 440, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887911/pdf.

Charette:1996:NES

[Cha96] François Charette. Nature, the Exotic, and the Science of French Colonialism by Michael A. Osborne (review). Technology and Culture, 37(1):192–194, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887556/pdf.

Channell:1998:ELT

[Cha98] David F. Channell. Engineering Labour: Technical Workers in Comparative Perspective by Peter Meiksins and Chris Smith (review). Technology and Culture, 39(1):149–151, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887189/pdf.

Channell:1999:ABE

[Cha99a] David F. Channell. After the breakthrough:the emergence of hightemperature superconductivity as a research field. Technology and Culture, 40(1):195–197, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33300.

Channell:1999:PQB

[Cha99b] David F. Channell. Pasteur’s quadrant: Basic science and technological innovation. Technology and Culture, 40(2):390–392, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33360.

Childs:1997:SCP

[Chi97] S. Terry Childs. The Sounds and Colors of Power: The Sacred Metallurgical Technology of Ancient West Mexico by Dorothy Hosler (review). Technology and Culture, 38(2):503–505, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887338/pdf.

Christie:1991:TTA

[Chr91] Jean Christie. Technology and Transformation in the American Electric Utility Industry by Richard F. Hirsh (review). Technology and Culture, 32(4):1126–1127, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888669/pdf.

Christie:1993:SES

[Chr93] Jean Christie. Steinmetz: Engineer and Socialist by Ronald R. Kline (review). Technology and Culture, 34(3):699–701, July 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888153/pdf.

Christiansen:1995:ISB

[Chr95a] Donald Christiansen. InfoCulture: The Smithsonian Book of Information Age Inventions by Steven Lubar (review). Technology and Culture, 36(3):719–721, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887826/pdf.

Christie:1995:MAI

[Chr95b] Jean Christie. Machine-Age Ideology: Social Engineering and American Liberalism, 1911–1939 by John M. Jordan (review). Technology and Culture, 36(2):423–424, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887902/pdf.

Christiansen:1999:TLE

[Chr99] Donald Christiansen. Thinking Like an Engineer: Studies in the Ethics of a Profession (review). Technology and Culture, 40(4): 881–883, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33431.

Churella:1999:ECY

[Chu99] Albert J. Churella. The engine that could: 75 years of valuesdriven change at Cummins Engine Company. Technology and Culture, 40(3):681–683, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33385.

Collins:1995:AAP

[CK95] H. M. Collins and M. Kusch. Automating air pumps: an empirical and conceptual analysis. Technology and Culture, 36(4):802–829, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887754/pdf.

Campbell-Kelly:1998:DPT

[CK98] Martin Campbell-Kelly. Data processing and technological change: The post Office savings bank, 1861–1930. Technology and Culture, 39(1):1–32, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887169/pdf.

Clark:1990:CIN

[Cla90] John G. Clark. Collapse of an Industry: Nuclear Power and the Contradictions of U.S. Policy by John L. Campbell, and: The Demise of Nuclear Energy? Lessons for Democratic Control of Technology by Joseph G. Morone, Edward J. Woodhouse (review). Technology and Culture, 31(2):346–349, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888912/pdf.

Clarke:1991:FMN

[Cla91] David Clarke. French Modern: Norms and Forms of the Social Environment by Paul Rabinow (review). Technology and Culture, 32(1):137, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888798/pdf.

Clark:1992:ECE

[Cla92] John G. Clark. The Electric City: Energy and the Growth of the Chicago Area, 1880–1930 by Harold L. Platt (review). Technology and Culture, 33(3):593–595, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888255/pdf.

Clark:1993:SIB

[Cla93] Mark Clark. Suppressing innovation: Bell laboratories and magnetic recording. Technology and Culture, 34(3):516–538, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888123/pdf.

Clavering:1995:CMN

[Cla95] Eric Clavering. The coal mills of Northeast England: The use of waterwheels for draining coal mines, 1600–1750. Technology and Culture, 36(2):211–241, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887879/pdf.

Clancey:1996:MMC

[Cla96] Gregory Clancey. Modernism, at Mid-Century: The Architecture of the United States Air Force Academy ed. by Robert Bruegmann (review). Technology and Culture, 37(3):646–647, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887459/pdf.

Clark:1997:LMR

[Cla97a] Elaine Clark. Local Markets and Regional Trade in Medieval Exeter by Maryanne Kowaleski (review). Technology and Culture, 38(2):510–512, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887341/pdf.

Claxton:1997:ZPT

[Cla97b] Robert H. Claxton. Zworykin, Pioneer of Television by Albert Abramson (review). Technology and Culture, 38(3):789–791, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887307/
pdf.

Coates:1999:OAB

[Coa99] Peter A. Coates. Oil from the Arctic: Building the Trans-Alaska Pipeline, at the National Museum of American History. Technology and Culture, 40(2):369–374, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33367.

Cody:1991:GCO

[Cod91] Jeffrey Cody. Gateway Cities and Other Essays by Leonard K. Eaton (review). Technology and Culture, 32(2 (part 1)):417–419, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888758/pdf.

Cody:1992:SCS

[Cod92] Jeffrey Cody. From, Sea Charts to Satellite Images: Interpreting North American History through Maps ed. by David Buisseret (review). Technology and Culture, 33(2):368–369, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888311/pdf.

Cohen:1990:HCT

[Coh90] I. Bernard Cohen. A History of Computing Technology by Michael R. Williams, and: Préhistoire et histoire des ordinateurs: Origines du calcul aux premiers calculateurs électroniques by Robert Ligonnière, and: Histoire de l’informatique by Philippe Breton (review). Technology and Culture, 31(4):907–911, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901684/pdf.

Coleman:1990:RRJ

[Col90] Samuel K. Coleman. Riken from 1945 to 1948: The reorganization of Japan’s Physical and Chemical Research Institute under the American occupation. Technology and Culture, 31(2):228–250, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888876/pdf.

Colten:1993:ARO

[Col93] Craig E. Colten. Always a River: The Ohio River and the American Experience ed. by Robert L. Reid (review). Technology and Culture, 34(1):153–155, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888473/pdf.

Collins:1999:SSJ

[Col99] Steven G. Collins. System in the South: John W. Mallet, Josiah Gorgas, and uniform production at the Confederate Ordnance Department. Technology and Culture, 40(3):517–544, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33414.

Condit:1990:CED

[Con90] Carl W. Condit. Cities and Economic Development: From the Dawn of History to the Present by Paul Bairoch (review). Technology and Culture, 31(2):295–297, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888882/pdf.

Condit:1991:CBM

[Con91] Carl W. Condit. The City Beautiful Movement by William H. Wilson (review). Technology and Culture, 32(2 (part 1)):419–421, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888759/pdf.

Condit:1992:ATH

[Con92a] Carl W. Condit. Alone Together: a History of New York’s Early Apartments by Elizabeth Cromley (review). Technology and Culture, 33(1):193–194, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888373/pdf.

Condit:1992:NMC

[Con92b] Carl W. Condit. Nature’s Metropolis: Chicago and the Great West by William Cronon (review). Technology and Culture, 33(3):591– 593, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888254/pdf.

Connell:1992:FTG

[Con92c] Joseph O’ Connell. The fine-tuning of a Golden ear: High-end audio and the evolutionary model of technology. Technology and Culture, 33(1):1–37, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888336/pdf.

Constant:1994:CRA

[Con94a] Edward W. Constant. Comment on The Retractable Airplane Landing Gear and the “Northrop Anomaly”. Technology and Culture, 35(2):447–449, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887951. See [Vin94].

Constant:1994:REN

[Con94b] Edward W. Constant II. Regulated Enterprises: Natural Gas Pipelines and Northeastern Markets, 1938–1954 by Christopher J. Castaneda (review). Technology and Culture, 35(4):879–881, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887734/pdf.

Conrad:1995:DBS

[Con95] James L. Conrad, Jr. “Drive that branch”: Samuel Slater, the power loom, and the writing of America’s textile history. Technology and Culture, 36(1):1–28, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888002/pdf.

Constant:1996:APP

[Con96a] Ed Constant. America’s Pursuit of Precision Bombing, 1910–1945 by Stephen L. McFarland (review). Technology and Culture, 37 (2):387–389, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887497/pdf.

Constant:1996:SRI

[Con96b] Edward W. Constant. Science on the Run: Information Management and Industrial Geophysics at Schlumberger, 1920–1940 by Geoffrey C. Bowker (review). Technology and Culture, 37(1):202–205, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887562/pdf.

Constant:1997:MPT

[Con97] Ed Constant. The Mangle of Practice: Time, Agency, and Science by Andrew Pickering (review). Technology and Culture, 38(3):815– 817, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887321/pdf.

Constant:1998:AWP

[Con98] Edward W. Constant. In and About the World: Philosophical Studies of Science and Technology (review). Technology and Culture, 39(3):540–541, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33200.

Constant:1999:RKU

[Con99] Edward W. Constant. Reliable knowledge and unreliable stuff: On the practical role of rational beliefs. Technology and Culture, 40(2):324–357, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/ 1/article/33365. See comment [Scr00].

Cooper:1990:PRI

[Coo90] Gail Cooper. Power Revolution in the Industrialization of Japan: 1885–1940 by Ryoshin Minami (review). Technology and Culture, 31(2):330–331, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888902/pdf.

Cook:1991:SES

[Coo91a] Noble David Cook. Silver and Entrepreneurship in Seventeenth-Century Potosí: The Life and Times of Antonio López de Quiroga by Peter Bakewell (review). Technology and Culture, 32(2 (part 1)):438– 440, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888771/pdf.

Cooper:1991:MIP

[Coo91b] Carolyn C. Cooper. Making inventions patent. Technology and Culture, 32(4):837–845, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888644/pdf.

Cooper:1991:SCI

[Coo91c] Carolyn C. Cooper. Social construction of invention through patent management: Thomas Blanchard’s woodworking machinery. Technology and Culture, 32(4):960–998, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888649/pdf.

Coopersmith:1991:TCA

[Coo91d] Jonathan Coopersmith. From the Titanic to the Challenger: an Annotated Bibliography on Technological Failures of the Twentieth Century by Susan Davis Herring (review). Technology and Culture, 32(1):152, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888807/pdf.

Cooper:1992:FGW

[Coo92a] Gail Cooper. Factory Girls: Women in the Thread Mills of Meiji Japan by E. Patricia Tsurumi (review). Technology and Culture, 33 (4):810–811, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888092/pdf.

Coopersmith:1992:QPE

[Coo92b] Jonathan Coopersmith. Questions of Power: Electricity and Environment in Inter-war Britain by Bill Luckin (review). Technology and Culture, 33(2):393–395, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888328/pdf.

Coon:1993:SSE

[Coo93a] Deborah J. Coon. Standardizing the subject: Experimental psychologists, introspection, and the quest for a technoscientific ideal. Technology and Culture, 34(4):757–783, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888038/pdf.

Cooper:1993:HIP

[Coo93b] Carolyn C. Cooper. A History of Industrial Power in the United States, 1780–1930. Vol. 3: The Transmission of Power by Louis C. Hunter, Lynwood Bryant (review). Technology and Culture, 34 (1):162–163, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888478/pdf.

Cooper:1993:GWP

[Coo93c] Patricia Cooper. Getting Work: Philadelphia, 1840–1950 by Walter Licht (review). Technology and Culture, 34(4):949–951, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888060/pdf.

Cooper:1994:CDE

[Coo94] Gail Cooper. Custom design, engineering guarantees, and unpatentable data: The air conditioning industry, 1902–1935. Technology and Culture, 35(3):506–536, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887836/pdf.

Cooper:1995:MFS

[Coo95] Gail Cooper. Machines Are Frozen Spirit: The Scientification of Refrigeration and Brewing in the 19th Century — a Weberian Interpretation by Mikael Hård (review). Technology and Culture, 36 (4):1036–1038, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887772/pdf.

Cooper:1996:TWW

[Coo96] Carolyn Cooper. Transforming Women’s Work: New England Lives in the Industrial Revolution by Thomas Dublin (review). Technology and Culture, 37(2):354–356, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887480/pdf.

Cook:1997:MWC

[Coo97] Harold J. Cook. Medicine and Western Civilization ed. by David J. Rothman, et al. (review). Technology and Culture, 38(1):237–238, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887377/pdf.

Cooper:1998:CTU

[Coo98a] Carolyn C. Cooper. Cognition and Tool Use: The Blacksmith at Work (review). Technology and Culture, 39(4):759–760, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33248.

Cooper:1998:TTJ

[Coo98b] Gail A. Cooper. The Technological Transformation of Japan from the Seventeenth to the Twenty-First Century by Tessa Morris-Suzuki (review). Technology and Culture, 39(1):161–162, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887195/pdf.

Coopersmith:1998:ISS

[Coo98c] Jonathan Coopersmith. Industry, State, and Society in Stalin’s Russia, 1926–1934 by David R. Shearer (review). Technology and Culture, 39(2):346–348, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887156/pdf.

Coopey:1998:ITB

[Coo98d] R. Coopey. Information Technology as Business History: Issues in the History and Management of Computers (review). Technology and Culture, 39(3):595–596, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33184.

Cornford:1990:CTP

[Cor90] Daniel A. Cornford. Company Town: Potlatch, Idaho, and the Potlatch Lumber Company by Keith C. Petersen (review). Technology and Culture, 31(1):161–163, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888944/pdf.

Corn:1991:VFM

[Cor91] Joseph J. Corn. Visions of a Flying Machine: The Wright Brothers and the Process of Invention by Peter L. Jakab (review). Technology and Culture, 32(3):628–630, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888717/pdf.

Corn:1992:TWW

[Cor92] Joseph J. Corn. Taking the Wheel: Women and the Coming of the Motor Age by Virginia Scharff (review). Technology and Culture, 33(3):598–600, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888258/pdf.

Cornford:1993:FCW

[Cor93] Daniel Cornford. Fortress California, 1910–1961: From Warfare to Welfare by Roger W. Lotchin (review). Technology and Culture, 34(2):445–447, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888438/pdf.

Corn:1995:HHA

[Cor95] Joseph J. Corn. Highways to Heaven: The Auto Biography of America by Christopher Finch (review). Technology and Culture, 36 (1):192–194, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888022/pdf.

Cortada:1997:KSD

[Cor97] James Cortada. King of the Seven Dwarfs: General Electric’s Ambiguous Challenge to the Computer Industry by Homer R. Oldfield (review). Technology and Culture, 38(4):1010–1012, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887257/pdf.

Cousins:1994:BOM

[Cou94] Peter H. Cousins. Bale o’Cotton: The Mechanical Art of Cotton Ginning by Karen G. Britton (review). Technology and Culture, 35(3):621–623, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887858/pdf.

Cowan:1993:DLT

[Cow93] Ruth Schwartz Cowan. Descartes’s legacy: a theme issue on biomedical and behavioral technology. Technology and Culture, 34 (4):721–728, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888036/pdf.

Cowan:1996:TSF

[Cow96] Ruth Schwartz Cowan. Technology is to science as female is to male: Musings on the history and character of our discipline. Technology and Culture, 37(3):572–582, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887434/pdf.

Cozzens:1993:STS

[Coz93a] Susan Cozzens. Science, Technology, and Social Progress ed. by Steven L. Goldman (review). Technology and Culture, 34(2):455– 456, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888444/pdf.

Cozzens:1993:ASP

[Coz93b] Susan E. Cozzens. American Science Policy since World War II by Bruce L. R. Smith (review). Technology and Culture, 34(1):191–192, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888494/pdf.

Crawford:1997:CMS

[Cra97] Margaret Crawford. Creating the Modern South: Millhands and Managers in Dalton, Georgia, 1884–1984 by Douglas Flamming (review). Technology and Culture, 38(1):253–255, January 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887386/pdf.

Crease:1999:ILM

[Cre99] Robert P. Crease. Image and Logic: a Material Culture of Microphysics (review). Technology and Culture, 40(4):924–925, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33433.

Crouch:1994:DMI

[Cro94] Tom D. Crouch. The Dream Machines: an Illustrated History of the Spaceship in Art, Science and Literature by Ron Miller (review). Technology and Culture, 35(4):886–887, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887738/pdf.

Crouch:1999:KDH

[Cro99] Tom D. Crouch. Kill Devil Hills, 17 December 1903. Technology and Culture, 40(3):595–598, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33394.

Cruikshank:1999:PRT

[Cru99] Ken Cruikshank. The philosophy of railways: The transcontinental railway idea in British North America. Technology and Culture, 40(2):427–429, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33350.

Cooper:1990:FII

[CS90] Gail Cooper and Bruce Sinclair. Failed innovations — ICOHTEC Symposium, Hamburg, August 1989. Technology and Culture, 31 (3):496–499, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888842/pdf.

Cumbler:1991:PMS

[Cum91] John T. Cumbler. The Path to Mechanized Shoe Production in the United States by Ross Thomson (review). Technology and Culture, 32(4):1113–1114, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888661/pdf.

Curtis:1991:MMM

[Cur91] James M. Curtis. Marshall McLuhan: The Man and His Message ed. by George Sanderson, Frank Macdonald, and: Marshall McLuhan: The Medium and the Messenger by Philip Marchand (review). Technology and Culture, 32(4):1141–1143, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888677/pdf.

Cutcliffe:1993:FBE

[Cut93] Stephen H. Cutcliffe. Flying Buttresses, Entropy, and O-Rings: The World of an Engineer by James L. Adams (review). Technology and Culture, 34(2):454–455, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888443/pdf.

Cutcliffe:1994:DIR

[Cut94a] Stephen H. Cutcliffe. Discovery, Innovation, and Risk: Case Studies in Science and Technology by Newton H. Copp and Andrew W. Zanella (review). Technology and Culture, 35(3):645–646, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887872/pdf.

Cutcliffe:1994:TFE

[Cut94b] Stephen H. Cutcliffe. Technology and the Future ed. by Albert H. Teich (review). Technology and Culture, 35(4):904–905, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887748/pdf.

Cutcliffe:1995:HGS

[Cut95] Stephen H. Cutcliffe. A hitchhiker’s guide to STS. Technology and Culture, 36(4):1015–1020, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887762/pdf.

Cutcliffe:1997:UTH

[Cut97] Stephen H. Cutcliffe. Undergraduate Teaching of the History of Technology: a Survey of the Teaching at Some Universities in the USA in 1993 by Lars O. Olsson (review). Technology and Culture, 38 (1):271–272, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887396/pdf.

Cutcliffe:1998:RPT

[Cut98] Stephen H. Cutcliffe. Research in Philosophy and Technology. Vol. 15, Social and Philosophical Constructions of Technology ed. by Carl Mitcham (review). Technology and Culture, 39(1):122–123, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887174/pdf.

Cutcliffe:1999:MPH

[Cut99] Stephen H. Cutcliffe. Major Problems in the History of American Technology (review). Technology and Culture, 40(3):648–650, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33405.

DAmbrosio:1997:CTM

[D’A97] Ubiratan D’Ambrosio. Ciencia y Técnica en la Metropolización de América by José Sala Catalá (review). Technology and Culture, 38 (1):244–246, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887381/pdf.

Dale:1990:SIR

[Dal90] Patrick Dale. Stalin’s Industrial Revolution: Politics and Workers, 1928–1932 by Hiroaki Kuromiya (review). Technology and Culture, 31(3):539–541, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888861/pdf.

Daniels:1990:AAW

[Dan90] George H. Daniels. American Automobile Workers, 1900–1933 by Joyce Shaw Peterson (review). Technology and Culture, 31 (1):166–168, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888947/pdf.

Daniels:1992:TJS

[Dan92] George H. Daniels. Thomas Jefferson: Statesman of Science by Silvio A. Bedini (review). Technology and Culture, 33(1):160–161, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888352/pdf.

Daniels:1993:JBA

[Dan93] George H. Daniels. James Bowron: The Autobiography of a New South Industrialist by Robert J. Norrell (review). Technology and Culture, 34(1):157–159, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888475/pdf.

Daniels:1995:KCS

[Dan95] George H. Daniels. Knowledge, Culture, and Science in the Metropolis: The New York Academy of Sciences, 1817–1970 by Simon Baatz (review). Technology and Culture, 36(1):204–206, January 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888029/pdf.

Davies:1992:IEL

[Dav92a] Alun C. Davies. Industrialisation and Everyday Life by Rudolf Braun (review). Technology and Culture, 33(3):605–607, July 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888263/pdf.

Davis:1992:AAP

[Dav92b] Susan G. Davis. The American Amusement Park Industry: a History of Technology and Thrills by Judith A. Adams (review). Technology and Culture, 33(4):797–798, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888084/pdf.

Davis:1993:MDP

[Dav93] Audrey B. Davis. Miasma and Disease: Public Health and the Environment in the Preindustrial Age by Carlo M. Cipolla, and: Epidemics and Ideas: Essays on the Historical Perception of Pestilence ed. by Terence Ranger, Paul Slack (review). Technology and Culture, 34(4):926–928, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888046/pdf.

Davids:1996:WTU

[Dav96] Karel Davids. Wetenschap tussen universiteit en industrie: De experimentele natuurkunde in Utrecht onder W. H. Julius en L. S. Ornstein, 1896–1940 by H. G. Heijmans (review). Technology and Culture, 37(2):383–385, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887495/pdf.

Davidson:1999:MKL

[Dav99] Janet F. Davidson. Ma kiley: The life of a railroad telegrapher. Technology and Culture, 40(3):696–697, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33391.

Dawson:1991:RSM

[Daw91] Virginia P. Dawson. Race to the Stratosphere: Manned Scientific Ballooning in America by David H. DeVorkin (review). Technology and Culture, 32(3):633–634, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888720/pdf.

Dawson:1996:KPG

[Daw96] Virginia P. Dawson. Knowledge is power: E. G. Bailey and the invention and marketing of the Bailey Boiler Meter. Technology and Culture, 37(3):493–526, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887432/pdf.

Dawson:1999:BTT

[Daw99] Virginia P. Dawson. BFGoodrich: Tradition and transformation, 1870–1995. Technology and Culture, 40(2):436–437, April 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33330.

Day:1992:CLS

[Day92] Tanis Day. Capital — labor substitution in the home. Technology and Culture, 33(2):302–327, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888294/pdf.

Day:1999:BHH

[Day99] Dwayne A. Day. Beyond Horizons: a Half Century of Air Force Space Leadership (review). Technology and Culture, 40(4):919–921, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33434.

Doig:1994:AFB

[DB94] Jameson W. Doig and David P. Billington. Ammann’s first bridge: a study in engineering, politics, and entrepreneurial behavior. Technology and Culture, 35(3):537–570, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887837/pdf.

DeSyon:1999:ZMF

[De 99] Guillaume De Syon. The Zeppelin Museum in Friedrichshafen. Technology and Culture, 40(1):114–119, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33324.

Dekker:1998:EMG

[Dek98] Elly Dekker. Die Entwicklung de Meridiankreises, 1700–1850: Genesis eines astronomischen Hauptinstrumentes unter Berucksichtigung des Wechselverhatnisses zwischen Astronomie, Astro-Technik und Technik (review). Technology and Culture, 39(3):559–560, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33193.

DeVries:1990:RSA

[DeV90] Kelly R. DeVries. A 1445 reference to shipboard artillery. Technology and Culture, 31(4):818–829, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901652/pdf.

DeVries:1991:PWT

[DeV91] Kelly DeVries. The Patterns of War through the Eighteenth Century by Larry H. Addington (review). Technology and Culture, 32(3):646– 648, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888727/pdf.

DeVorkin:1992:IAH

[DeV92] David DeVorkin. Inventing Accuracy: a Historical Sociology of Nuclear Missile Guidance by Donald MacKenzie (review). Technology and Culture, 33(3):649–652, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888289/pdf.

DeVries:1993:MFJ

[DeV93] Kelly DeVries. Medieval Fortifications by John R. Kenyon (review). Technology and Culture, 34(3):691–692, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888148/pdf.

Dias:1999:MKG

[Dia99] Ric Dias. Mr. Kaiser goes to Washington: The rise of a government entrepreneur. Technology and Culture, 40(2):439–441, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33327.

Dick:1992:CNT

[Dic92] Steven J. Dick. Centralizing navigational technology in America: The U.S. Navy’s Depot of Charts and Instruments, 1830–1842. Technology and Culture, 33(3):467–509, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888241/pdf.

Dienel:1993:FDF

[Die93] Hans-Liudger Dienel. Frankreich und Deutschland: Forschung, Technologie und industrielle Entwicklung im 19. und 20. Jahrhundert ed. by Yves Cohen, Klaus Manfrass (review). Technology and Culture, 34(2):430–432, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888429/pdf.

Dienel:1997:HCM

[Die97] Hans-Liudger Dienel. Heat and Cold: Mastering the Great Indoors by Barry Donaldson, Bernard Nagengast (review). Technology and Culture, 38(4):1000–1002, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887252/pdf.

Divall:1994:EDP

[Div94] Colin Divall. Education for design and production: Professional organization, employers, and the study of chemical engineering in British universities, 1922–1976. Technology and Culture, 35(2):258– 288, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887918/pdf.

Divall:1995:CRN

[Div95] Colin Divall. Changing routes? The New London Transport Museum. Technology and Culture, 36(3):630–635, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887793/pdf.

delaPena:1999:RFC

[dlP99] Carolyn Thomas de la Pena. Recharging at the fordyce: Confronting the machine and nature in the modern bath. Technology and Culture, 40(4):746–769, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33463.

Dolman:1997:ESJ

[Dol97] Everett Carl Dolman. Europe in Space, 1960–1973 by John Krige, Arturo Russo (review). Technology and Culture, 38(3):792–793, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887308/pdf.

Donovan:1991:ECC

[Don91] Arthur Donovan. Eighteenth-Century Chemistry as an Investigative Enterprise by Frederic Lawrence Holmes (review). Technology and Culture, 32(4):1106–1108, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888657/pdf.

Donnelly:1994:CMT

[Don94] James Donnelly. Consultants, managers, testing slaves: Changing roles for chemists in the British alkali industry, 1850–1920. Technology and Culture, 35(1):100–128, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887956/pdf.

Dood:1991:PEI

[Doo91] Kendall J. Dood. Pursuing the essence of inventions: Reissuing patents in the 19th century. Technology and Culture, 32(4):999–1017, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888650/pdf.

Dorn:1991:MRM

[Dor91] Harold Dorn. The “Military revolution”: Military history or history of Europe? Technology and Culture, 32(3):656–658, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888732/pdf.

Dotson:1999:DRT

[Dot99] John E. Dotson. The development of the rudder: a technological tale. Technology and Culture, 40(1):137–139, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33297.

Douglas:1991:IAP

[Dou91a] Deborah G. Douglas. The Impact of Air Power on the British People and Their Government, 1909–14 by Alfred Gollin (review). Technology and Culture, 32(1):163–166, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888814/pdf.

Douglas:1991:TAB

[Dou91b] Susan Douglas. Technology in America: a Brief History by Alan I. Marcus, Howard P. Segal (review). Technology and Culture, 32 (2 (part 1)):397, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888746/pdf.

Douglas:1994:HGJ

[Dou94] Susan J. Douglas. Hugh G. J. Aitken (1922–1994). Technology and Culture, 35(4):906–911, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887749/pdf.

Douglas:1995:ATC

[Dou95] George H. Douglas. Alternative Tracks: The Constitution of American Industrial Order, 1865–1917 by Gerald Berk (review). Technology and Culture, 36(3):705–707, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887818/pdf.

Douglas:1999:JHR

[Dou99] Alan Douglas. Joseph Henry: The rise of an American scientist. Technology and Culture, 40(1):162–164, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33298.

Dreicer:1995:PMS

[Dre95] Gregory K. Dreicer. Les ponts modernes: 18^{e} –19^{e} siècles by Bernard Marrey (review). Technology and Culture, 36(1):180–181, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888015/pdf.

Dreicer:1997:RNY

[Dre97] Gregory K. Dreicer. The Rise of the New York Skyscraper, 1865– 1913 by Sarah Bradford Landau, Carl W. Condit (review). Technology and Culture, 38(3):767–769, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887295/pdf.

Dreicer:1999:CBE

[Dre99] Gregory K. Dreicer. Constructing a bridge: an exploration of engineering culture, design, and research in Nineteenth-Century France and America. Technology and Culture, 40(2):425–427, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33344.

Dublin:1993:CGL

[Dub93] Thomas Dublin. Cradle to Grave: Life, Work, and Death at the Lake Superior Copper Mines by Larry Lankton (review). Technology and Culture, 34(1):160–161, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888477/pdf.

Dudley:1992:OSN

[Dud92] William S. Dudley. The Old Steam Navy, vol. 1: Frigates, Sloops, and Gunboats, 1815–1885 by Donald L. Canney (review). Technology and Culture, 33(2):371–372, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888313/pdf.

Dunlap:1993:BBF

[Dun93] Thomas R. Dunlap. Burning Bush: a Fire History of Australia by Stephen J. Pyne (review). Technology and Culture, 34(2):420–422, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888423/pdf.

Dupree:1992:LLN

[Dup92] A. Hunter Dupree. Lucky Lady and the Navy Mystique: The Chenango in WWII by Brooke Hindle (review). Technology and Culture, 33(3):634–636, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888280/pdf.

Durbin:1998:ESH

[Dur98] Paul T. Durbin. Engaging Science: How to Understand Its Practices Philosophically (review). Technology and Culture, 39(3):546–547, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33202.

deVries:1995:KMJ

[dV95] Marc J. de Vries. Kunst und Maschine: 500 Jahre Maschinenlinien in Bild und Skulptur by Peter Friess (review). Technology and Culture, 36(3):697–699, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887814/pdf.

Eamon:1997:GFL

[Eam97] William Eamon. Gehennical Fire: The Lives of George Starkey, an American Alchemist in the Scientific Revolution by William R. Newman (review). Technology and Culture, 38(3):758–760, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887290/pdf.

Eaton:1990:OFA

[Eat90] Leonard K. Eaton. Orders from France: The Americans and the French in a Revolutionary World, 1780–1820 by Roger G. Kennedy (review). Technology and Culture, 31(4):872–873, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901662/pdf.

Ebner:1992:ESC

[Ebn92] Michael H. Ebner. Essays on Sunbelt Cities and Recent Urban America ed. by Robert B. Fairbanks, Kathleen Underwood (review). Technology and Culture, 33(1):192–193, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888372/pdf.

Edgerton:1998:MIT

[Edg98] David Edgerton. Men, Ideas, and Tanks: British Military Thought and Armoured Forces, 1903–1939 by J. P. Harris (review). Technology and Culture, 39(2):364–365, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887165/pdf.

Edmonds:1996:ANH

[Edm96] Leigh Edmonds. Aerial Nationalism: a History of Aviation in Thailand by Edward M. Young (review). Technology and Culture, 37(3):644–646, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887458/pdf.

Edwards:1995:TPB

[Edw95] Paul N. Edwards. Technological Perspectives on Behavioral Change by Michael B. Schiffer (review). Technology and Culture, 36(3):721– 723, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887827/pdf.

Edwards:1996:BIS

[Edw96a] Paul N. Edwards. Building IBM: Shaping an Industry and Its Technology by Emerson W. Pugh (review). Technology and Culture, 37(2):385–387, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887496/pdf.

Edwards:1996:CPJ

[Edw96b] Paul N. Edwards. Computer Pioneers by J. A. N. Lee (review). Technology and Culture, 37(4):860–861, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887424/pdf.

Edwards:1997:KME

[Edw97] Paul N. Edwards. Knowing Machines: Essays on Technical Change by Donald MacKenzie (review). Technology and Culture, 38(4): 962–965, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887222/pdf.

Eggert:1994:MIS

[Egg94] Gerald G. Eggert. Making Iron and Steel: Independent Mills in Pittsburgh, 1820–1920 by John N. Ingham (review). Technology and Culture, 35(1):196–197, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887972/pdf.

Eisenstein:1993:PEH

[Eis93] Elizabeth L. Eisenstein. In the Public Eye: a History of Reading in Modern France, 1800–1940 by James Smith Allen (review). Technology and Culture, 34(1):140–142, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888465/pdf.

Ellen:1992:VWP

[Ell92a] Roy Ellen. A vertical wedge Press from the banda islands. Technology and Culture, 33(1):122–131, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888340/pdf.

Elliott:1992:MAJ

[Ell92b] Clark A. Elliott. Meteorology in America, 1800–1870 by James R. Fleming (review). Technology and Culture, 33(1):167–168, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888357/pdf.

Elliott:1993:TPB

[Ell93] Derek W. Elliott. The Technology Pork Barrel by Linda R. Cohen, Roger G. Noll (review). Technology and Culture, 34(2):465– 467, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888451/pdf.

Ellen:1994:EAT

[Ell94a] Roy Ellen. Elements for an Anthropology of Technology by Pierre Lemonnier (review). Technology and Culture, 35(3):652–654, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887876/
pdf.

Ellis:1994:NTS

[Ell94b] Jeffrey Ellis. Nature, Technology, and Society: Cultural Roots of the Current Environmental Crisis by Victor Ferkiss (review). Technology and Culture, 35(3):650–652, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887875/pdf.

Elvin:1999:CAD

[Elv99] Mark Elvin. Controlling from afar: The Daoguang Emperor’s management of the Grand Canal Crisis, 1824–1826. Technology and Culture, 40(3):664–666, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33393.

Endelman:1997:FLE

[End97] Judith E. Endelman. Future Libraries ed. by R. Howard Bloch, Carla Hesse (review). Technology and Culture, 38(3):817–818, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887322/pdf.

Enoch:1998:EEL

[Eno98] Jay M. Enoch. The enigma of early lens use. Technology and Culture, 39(2):273–291, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887130/pdf.

Evans:1992:CBM

[Eva92] Elizabeth S. Evans. The Corncribs of Buzet: Modernizing Agriculture in the French Southwest by Peter H. Amann (review). Technology and Culture, 33(3):615–616, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888268/pdf.

Evans:1993:PCO

[Eva93a] Gail Evans. Power at Cost: Ontario Hydro and Rural Electrification, 1911–1958 by Keith R. Fleming (review). Technology and Culture, 34(2):436–438, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888433/pdf.

Evans:1993:LTC

[Eva93b] Hughes Evans. Losing touch: The controversy over the introduction of blood pressure instruments into medicine. Technology and Culture, 34(4):784–807, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888039/pdf.

Evans:1996:PMH

[Eva96] Chris J. Evans. The Precision Makers: a History of the Instruments Industry in Britain and France, 1870–1939 by Mari E. W. Williams (review). Technology and Culture, 37(1):188–190, January 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887554/pdf.

Eves:1992:SCR

[Eve92] Jamie H. Eves. “Shrunk to a comparative rivulet”: Deforestation, stream flow, and rural milling in 19th-Century Maine. Technology and Culture, 33(1):38–65, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888337/pdf.

Ezell:1990:CNA

[Eze90] Linda N. Ezell. Canada’s National Aviation Museum: Its History and Collections by K. M. Molson (review). Technology and Culture, 31(3):559–560, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888873/pdf.

Fano:1991:WMT

[Fan91] Ester Fano. A “Wastage of Men”: Technological progress and unemployment in the United States. Technology and Culture, 32(2 (part 1)):264–292, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888738/pdf.

Farris:1996:MEH

[Far96] W. Wayne Farris. The Modern Epidemic: a History of Tuberculosis in Japan by William Johnston (review). Technology and Culture, 37(4):856–858, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887422/pdf.

Farnham:1999:RMB

[Far99] Jonathan E. Farnham. Robert maillart: Builder, designer, and artist. Technology and Culture, 40(1):154–156, January 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33276.

Fear:1999:FFF

[Fea99] Jeff Fear. Forcing the factory of the future: Cybernation and societal institutions. Technology and Culture, 40(1):160–162, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33290.

Feldman:1997:CWM

[Fel97] Theodore S. Feldman. Calculating the Weather: Meteorology in the 20th Century by Frederik Nebeker (review). Technology and Culture, 38(3):786–788, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887305/pdf.

Feldman:1999:HEM

[Fel99] Theodore Sherman Feldman. Historical essays on meteorology, 1919– 1995: The diamond anniversary history volume of the American Meteorological Society. Technology and Culture, 40(1):164–165, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33279.

Ferguson:1990:YMD

[Fer90] Eugene S. Ferguson. A Year on a Monitor and the Destruction of Fort Sumter by Alvah F. Hunter (review). Technology and Culture, 31(1):185–187, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888958/pdf.

Ferguson:1992:MPI

[Fer92] Eugene S. Ferguson. Mechanics of Pre-industrial Technology: an Introduction to the Mechanics of Ancient and Traditional Material Culture by Brian Cotterell, Johan Kamminga (review). Technology and Culture, 33(3):603–604, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888261/pdf.

Ferguson:1995:DPC

[Fer95] Eugene S. Ferguson. Design Paradigms: Case Histones of Error and Judgment in Engineering by Henry Petroski (review). Technology and Culture, 36(3):696–697, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887813/pdf.

Ferguson:1998:JRH

[Fer98] Eugene Ferguson. John R. Harris (1923–1997). Technology and Culture, 39(2):368–371, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887167/pdf.

Ferreiro:1999:BIP

[Fer99] Larrie D. Ferreiro. BTechnologies, ideologies, pratiques: Revue d’anthropologie des connaissances 13, no. 1 (1998). concevoir et construire les navires: de la triere au picoteux. Technology and Culture, 40(3):690–692, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33403.

Fernandez:1996:PTT

[FF96] M. P. Fernandez and P. C. Fernandez. Precision timekeepers of tokugawa Japan and the evolution of the Japanese domestic clock. Technology and Culture, 37(2):221–248, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887467/pdf.

Fehner:1998:EAB

[FG98] Terrence R. Fehner and F. G. Gosling. Energy: an Annotated Bibliography by Joseph R. Rudolph Jr. (review). Technology and Culture, 39(1):146–148, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887187/pdf.

Field:1994:FOT

[Fie94a] Alexander J. Field. French optical telegraphy, 1793–1855: Hardware, software, administration. Technology and Culture, 35(2):315– 347, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887920/pdf.

Field:1994:APG

[Fie94b] Cynthia R. Field. Architecture and Planning of Graham, Anderson, Probst and White, 1912–1936: Transforming Tradition by Sally A. Kitt Chappell (review). Technology and Culture, 35(1):203–205, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887976/pdf.

Field:1998:DLJ

[Fie98] Gregory Field. David E. Lilienthal: The Journey of an American Liberal (review). Technology and Culture, 39(4):813–814, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33251.

Finn:1993:IWT

[Fin93] Bernard S. Finn. The Invisible Weapon: Telecommunications and International Politics, 1851–1945 by Daniel R. Headrick (review). Technology and Culture, 34(1):180–182, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888488/pdf.

Finlay:1995:HAS

[Fin95] Mark R. Finlay. The History of Agricultural Science and Technology: an International Annotated Bibliography by R. Douglas Hurt and Mary Ellen Hurt (review). Technology and Culture, 36(2):442– 443, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887913/pdf.

Finkelstein:1998:GMS

[Fin98] Gabriel Ward Finkelstein. Gustav Magnus und sein Haus: Im Auftrag der Deutschen Physikalischen Gesellschaft (review). Technology and Culture, 39(3):568–569, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33194.

Finn:1999:EE

[Fin99] Bernard S. Finn. Exhibiting electricity. Technology and Culture, 40(1):184–185, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33275.

Fischer:1992:HCG

[Fis92] Claude S. Fischer. “Hello, Central?” Gender, Technology, and Culture in the Formation of Telephone Systems by Michèle Martin (review). Technology and Culture, 33(3):597–598, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888257/pdf.

Fischer:1998:HLT

[Fis98] Claude S. Fischer. Holding the Line: The Telephone in Old Order Mennonite and Amish Life by Diane Zimmerman Umble (review). Technology and Culture, 39(1):164–166, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887198/pdf.

Fitzgerald:1990:WWW

[Fit90] Donald Fitzgerald. Work on the Waterfront: Worker Power and Technological Change in a West Coast Port by William Finlay (review). Technology and Culture, 31(1):154–156, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888939/pdf.

Fitzgerald:1991:BTH

[Fit91] Deborah Fitzgerald. Beyond tractors: The history of technology in American agriculture. Technology and Culture, 32(1):114–126, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888792/pdf.

Fitzgerald:1993:FDH

[Fit93a] Deborah Fitzgerald. Farmers deskilled: Hybrid corn and farmers’ work. Technology and Culture, 34(2):324–343, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888412/pdf.

FitzGibbon:1993:CCM

[Fit93b] Michael FitzGibbon. Curators and Culture: The Museum Movement in America, 1740–1870 by Joel J. Orosz (review). Technology and Culture, 34(2):412–413, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888418/pdf.

Fitzgerald:1995:OHF

[Fit95] Deborah Fitzgerald. Outstanding in His Field: Perspectives on American Agriculture in Honor of Wayne D. Rasmussen ed. by Frederick V. Carstensen, et al. (review). Technology and Culture, 36 (4):1041–1042, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887774/pdf.

Fitzgerald:1996:RRU

[Fit96] Deborah Fitzgerald. Regulation and the Revolution in United States Farm Productivity by Sally H. Clarke (review). Technology and Culture, 37(4):851–853, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887419/pdf.

Fitzgerald:1997:OVC

[Fit97] Deborah Fitzgerald. Osiris. Vol. 10, Constructing Knowledge in the History of Science ed. by Arnold Thackray, and: Technology and Culture. Vol. 36, no. 2 (suppl.), Snapshots of a Discipline: Selected Proceedings, Conference on Critical Problems and Research Frontiers in the History of Technology, Madison, Wisconsin, October 30–November 3, 1991 ed. by Robert Friedel (review). Technology and Culture, 38(3):744–746, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887282/pdf.

Fitzgerald:1998:EEA

[Fit98] Deborah Kay Fitzgerald. Entomology, Ecology and Agriculture: The Making of Scientific Careers in North America, 1885–1985 (review). Technology and Culture, 39(3):584–585, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33199.

Fleming:1990:OEP

[Fle90] Robin Fleming. The Other Economy: Pastoral Husbandry on a Medieval Estate by Kathleen Biddick (review). Technology and Culture, 31(3):512–513, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888845/pdf.

Fleming:1992:OMW

[Fle92] Robin Fleming. Opera Muliebria: Women and Work in Medieval Europe by David Herlihy (review). Technology and Culture, 33(2):350– 351, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888300/pdf.

Fletcher:1998:TUH

[Fle98] Amy L. Fletcher. Technohistory: Using the History of American Technology in Interdisciplinary Research (review). Technology and Culture, 39(4):823–825, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33244.

Florman:1992:WEK

[Flo92] Samuel C. Florman. What Engineers Know and How They Know It: Analytical Studies from Aeronautical History by Walter G. Vincenti (review). Technology and Culture, 33(1):140–142, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888342/pdf.

Forster:1992:HCW

[FO92] Emy Lou Forster and Kathleen Ochs. Hecla: a Century of Western Mining by John Fahey (review). Technology and Culture, 33 (4):823–824, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888099/pdf.

Ford:1993:CSC

[For93] Peter A. Ford. Charles S. Storrow, civil engineer: a case study of European training and technological transfer in the Antebellum Period. Technology and Culture, 34(2):271–299, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888410/pdf.

Fortun:1997:CFT

[For97] Kim Fortun. Chernobyl: The Forbidden Truth by Alla Yaroshinskaya, et al. (review). Technology and Culture, 38(2):529–531, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887351/pdf.

Foster:1998:SLD

[Fos98] Mark S. Foster. Studebaker: The Life and Death of an American Corporation (review). Technology and Culture, 39(4):786–787, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33236.

Frangsmyr:1991:CDS

[Frä91] Tore Frängsmyr. Les chemins du savoir en Suède: De la fondation de l’université d’Upsal à Jacob Berzelius — études et portraits by Sten Lindroth (review). Technology and Culture, 32(1):173–174, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888820/pdf.

Friedel:1990:PET

[Fri90] Robert Friedel. The Psychology of Everyday Things by Donald A. Norman (review). Technology and Culture, 31(3):554–555, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888870/pdf.

Friedel:1991:MIM

[Fri91a] Robert Friedel. The Material of Invention: Materials and Design by Ezio Manzini (review). Technology and Culture, 32(1):133–135, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888796/pdf.

Fries:1991:SST

[Fri91b] Sylvia Doughty Fries. A Shield in Space? Technology, Politics, and the Strategic Defense Initiative by Sanford Lakoff, Herbert F. York (review). Technology and Culture, 32(3):641–643, July 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888725/pdf.

Fricke:1992:UTI

[Fri92a] Ernest B. Fricke. The U.S. Tire Industry: a History by Michael J. French (review). Technology and Culture, 33(4):825–826, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888100/pdf.

Friedel:1992:EIE

[Fri92b] Robert Friedel. Envisioning Information by Edward R. Tufte (review). Technology and Culture, 33(1):143–147, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888344/pdf.

Friedel:1995:EN

[Fri95] Robert Friedel. Editorial note. Technology and Culture, 36(2S):S1– S2, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887705/pdf.

Fries:1996:JJG

[Fri96] Russell I. Fries. The Journals of Josiah Gorgas, 1857–1878 ed. by Sarah Woolfolk Wiggins (review). Technology and Culture, 37 (4):845–846, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887416/pdf.

Friedman:1998:ATC

[Fri98] Edward A. Friedman. After Thought: The Computer Challenge to Human Intelligence (review). Technology and Culture, 39(4): 822–823, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33229.

Fries:1999:WEN

[Fri99] Russell I. Fries. The winning edge: Naval technology in action, 1939– 1945. Technology and Culture, 40(3):692–693, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33402.

Frost:1992:PSS

[Fro92a] Robert Frost. Paris Sewers and Sewermen: Realities and Representations by Donald Reid (review). Technology and Culture, 33(3):617– 618, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888269/pdf.

Frost:1992:SCW

[Fro92b] Robert L. Frost. State Capitalism and Working-Class Radicalism in the French Aircraft Industry by Herrick Chapman (review). Technology and Culture, 33(2):381–383, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888321/pdf.

Frost:1993:TSC

[Fro93] Robert L. Frost. Technopoly: The Surrender of Culture to Technology by Neil Postman (review). Technology and Culture, 34(3):714– 715, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888161/pdf.

Fullmer:1995:CSM

[Ful95] June Z. Fullmer. Chemical Sciences in the Modern World ed. by Seymour H. Mauskopf (review). Technology and Culture, 36(3):727– 729, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887830/pdf.

Fullmer:1998:TK

[Ful98] June Z. Fullmer. Thomas S. Kuhn (1922–1996). Technology and Culture, 39(2):372–377, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887168/pdf.

Gillispie:1998:ER

[GA98] Charles Coulston Gillispie and Ken Alder. Engineering the revolution. Technology and Culture, 39(4):733–754, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33264.

Galambos:1992:MAC

[Gal92a] Louis Galambos. Making America Corporate, 1870–1920 by Olivier Zunz (review). Technology and Culture, 33(2):362–364, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888308/pdf.

Galishoff:1992:SPR

[Gal92b] Stuart Galishoff. The Search for Purity: a Retrospective Policy Analysis of the Decision to Chlorinate Cincinnati’s Public Water Supply, 1890–1920 by Colleen K. O’Toole (review). Technology and Culture, 33(3):618–619, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888270/pdf.

Gamst:1992:CSA

[Gam92] Frederick C. Gamst. The context and significance of America’s first railroad, on Boston’s Beacon Hill. Technology and Culture, 33(1):66–100, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888338/pdf.

Gamber:1995:RSG

[Gam95] Wendy Gamber. “Reduced to Science”: Gender, technology, and power in the American dressmaking trade, 1860–1910. Technology and Culture, 36(3):455–482, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887787/pdf.

Gamble:1997:THT

[Gam97] Vanessa Northington Gamble. Technology in the Hospital: Transforming Patient Care in the Early Twentieth Century by Joel D. Howell (review). Technology and Culture, 38(3):770–771, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887296/pdf.

Gamst:1999:BWE

[Gam99] Frederick C. Gamst. Bayerns weg zur eisenbahn: Joseph von baader und die fruhzeit der eisenbahn in bayern 1800 bis 1835. Technology and Culture, 40(3):667–669, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33388.

Garside:1990:ELE

[Gar90] W. R. Garside. Employers and Labour in the English Textile Industries, 1850–1939 ed. by J. A. Jowitt, A. J. McIvor (review). Technology and Culture, 31(3):525–527, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888853/pdf.

GarciayRobertson:1991:NEA

[Gar91] Rodrigo Garcia y Robertson. Naval Engineering and American Sea Power ed. by R. W. King (review). Technology and Culture, 32 (1):156–158, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888810/pdf.

GarciayRobertson:1992:PNM

[Gar92] Rodrigo Garcia y Robertson. This People’s Navy: The Making of American Sea Power by Kenneth J. Hagan (review). Technology and Culture, 33(3):633–634, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888279/pdf.

Garrett:1999:CWT

[Gar99] Stephen A. Garrett. The chemical weapons taboo. Technology and Culture, 40(1):194–195, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33307.

Geiger:1990:AWU

[Gei90] Reed Geiger. The Age of Water: The Urban Environment in the North of France, A.D. 300–1800 by André E. Guillerme (review). Technology and Culture, 31(2):297–299, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888883/pdf.

Geiger:1994:CWA

[Gei94] Roger L. Geiger. The Cold War and American Science: The Military-Industrial-Academic Complex at MIT and Stanford by Stuart W. Leslie (review). Technology and Culture, 35(3):629–631, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887863/pdf.

Genz:1992:PBS

[Gen92] Marcella Genz. Papermaking in Britain, 1488–1988: a Short History by Richard L. Hills (review). Technology and Culture, 33(4):814–815, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888094/pdf.

Gerovitch:1996:PHT

[Ger96] Slava Gerovitch. Perestroika of the history of technology and science in the USSR: Changes in the discourse. Technology and Culture, 37(1):102–137, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887538/pdf.

Gibson:1991:SAI

[Gib91] Jane Mork Gibson. Survivals: Aspects of Industrial Archaeology in Ontario by Dianne Newell, Ralph Greenhill (review). Technology and Culture, 32(2 (part 1)):414–415, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888756/pdf.

Gibson:1993:FIR

[Gib93] Jane Mork Gibson. The First Industrial Revolutions ed. by Peter Mathias, John A. Davis (review). Technology and Culture, 34(3):681– 683, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888143/pdf.

Gibson:1999:SSS

[Gib99] Jane Mork Gibson. Ships for the Seven Seas: Philadelphia shipbuilding in the Age of Industrial Capitalism. Technology and Culture, 40 (1):141–142, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33288.

Gieryn:1994:STB

[Gie94] Thomas F. Gieryn. Shaping Technology/Building Society: Studies in Sociotechnical Change ed. by Wiebe E. Bijker, John Law (review). Technology and Culture, 35(2):438–441, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887947/pdf.

Giebelhaus:1996:NES

[Gie96] August W. Giebelhaus. Nordic Energy Systems: Historical Perspectives and Current Issues ed. by Arne Kaijser and Marika Hedin (review). Technology and Culture, 37(3):648–649, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887460/pdf.

Gingerich:1990:NMS

[Gin90] Owen Gingerich. Nevil Maskelyne: The Seaman’s Astronomer by Derek Howse (review). Technology and Culture, 31(4):869–870, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901660/pdf.

Gispen:1998:TUB

[Gis98] Kees Gispen. Technische Universität Braunschweig: Vom Collegium Carolinum zur Technischen Universität, 1745–1995 ed. by Walter Kertz (review). Technology and Culture, 39(2):340–343, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887153/pdf.

Gordon:1993:ATC

[GK93] Robert B. Gordon and David J. Killick. Adaptation of technology to culture and environment: Bloomery iron smelting in America and Africa. Technology and Culture, 34(2):243–270, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888409/pdf.

Glick:1991:TPC

[Gli91] Thomas F. Glick. Técnica y Poder en Castilla durante los Siglos XVI y XVII by Nicolás García Tapia (review). Technology and Culture, 32(2 (part 1)):409–410, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888753/pdf.

Glick:1994:DIC

[Gli94] Thomas F. Glick. A Distinctive Industrialization: Cotton in Barcelona, 1728–1832 by J. K. J. Thomson (review). Technology and Culture, 35(2):407–408, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887928.

Goler:1990:MTM

[Gol90] Robert I. Goler. The Museum. Time-Machine: Putting Cultures On Display ed. by Robert Lumley (review). Technology and Culture, 31(3):557–559, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888872/pdf.

Goldman:1991:TLG

[Gol91a] Steven L. Goldman. Technology and the Lifeworld: From Garden to Earth by Don Ihde, and: Heidegger’s Confrontation with Modernity: Technology, Politics, and Art by Michael E. Zimmerman, and: John Dewey’s Pragmatic Technology by Larry A. Hickman (review). Technology and Culture, 32(4):1135–1137, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888674/pdf.

Gollin:1991:KAL

[Gol91b] Alfred Gollin. Knights of the Air: The Life and Times of the Extraordinary Pioneers Who First Built British Aeroplanes by Peter King (review). Technology and Culture, 32(3):630–631, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888718/pdf.

Gollin:1991:BBL

[Gol91c] Alfred Gollin. The Bishop’s Boys: a Life of Wilbur and Orville Wright by Tom Crouch (review). Technology and Culture, 32(1): 166–167, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888815/pdf.

Goldberg:1992:LDC

[Gol92] Stanley Goldberg. Lucius D. Clay: an American Life by Jean Edward Smith (review). Technology and Culture, 33(1):175–177, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888362/pdf.

Goldberg:1993:CRE

[Gol93] Stanley Goldberg. Chain Reaction: Expert Debate and Public Participation in American Commercial Nuclear Power, 1945–1975 by Brian Balogh (review). Technology and Culture, 34(1):189–191, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888493/pdf.

Goldberg:1995:MAM

[Gol95] Stanley Goldberg. A Mind Always in Motion: The Autobiography of Emilio Segrè by Emilio Segrè (review). Technology and Culture, 36(1):206–208, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888030/pdf.

Goldstein:1997:SSH

[Gol97] Carolyn M. Goldstein. From service to sales: Home economics in light and power, 1920–1940. Technology and Culture, 38(1):121–152, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887370/pdf.

Golden:1999:RHE

[Gol99] Janet Lynne Golden. Rethinking home economics: Women and the history of a profession. Technology and Culture, 40(3):713–715, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33407.

Gomery:1996:HIA

[Gom96] Douglas Gomery. Hollywood, in the Information Age: Beyond the Silver Screen by Janet Wasko (review). Technology and Culture, 37(3):655–656, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887464/pdf.

Gomery:1997:USE

[Gom97] Douglas Gomery. Under the Stars: Essays on Labor Relations in Arts and Entertainment ed. by Lois S. Gray, Ronald L. Seeber (review). Technology and Culture, 38(4):986–988, October 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887240/pdf.

Good:1990:STC

[Goo90] Gregory A. Good. Science and Technology in Canadian History: a Bibliography of Primary Sources to 1914 comp. by R. Alan Richardson, Bertrum H. MacDonald (review). Technology and Culture, 31(2):356–357, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888920/pdf.

Good:1995:MIC

[Goo95] Gregory A. Good. Making Instruments Count: Essays on Historical Scientific Instruments Presented to Gerard L’Estrange Turner ed. by R. G. W. Anderson (review). Technology and Culture, 36(1):201–202, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888027/pdf.

Gordon:1991:AGR

[Gor91] Robert B. Gordon. All That Glitters: Readings in Historical Metallurgy ed. by Michael L. Wayman (review). Technology and Culture, 32(2 (part 1)):436–437, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888769/pdf.

Gordon:1992:MIA

[Gor92a] Robert B. Gordon. Materials Issues in Art and Archaeology II ed. by Pamela B. Vandiver, et al. (review). Technology and Culture, 33(3):602–603, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888260/pdf.

Gordon:1992:PHD

[Gor92b] Robert B. Gordon. The Pencil: a History of Design and Circumstance by Henry Petroski (review). Technology and Culture, 33 (1):142–143, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888343/pdf.

Gordon:1992:KC

[Gor92c] Robert B. Gordon. The “Kelly” converter. Technology and Culture, 33(4):769–779, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888081/pdf.

Gorman:1994:ICF

[Gor94] Michael Gorman. Invention: The Care and Feeding of Ideas by Norbert Wiener (review). Technology and Culture, 35(3):647–648, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887873/pdf.

Gordon:1995:PBC

[Gor95a] Robert B. Gordon. Platt Brothers and Company: Small Business in American Manufacturing by Matthew W. Roth (review). Technology and Culture, 36(2):412–413, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887895/pdf.

Gordon:1995:UBR

[Gor95b] Sarah Gordon. Unfinished Business: The Railroad in American Life by Maury Klein (review). Technology and Culture, 36(2):416–418, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887898/pdf.

Gordon:1996:SFR

[Gor96] Robert B. Gordon. Sloss Furnaces and the Rise of the Birmingham District: an Industrial Epic by W. David Lewis (review). Technology and Culture, 37(2):365–366, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887486/pdf.

Gordon:1997:EMM

[Gor97a] Robert B. Gordon. Early Metal Mining and Production by Paul T. Craddock (review). Technology and Culture, 38(3):751–753, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887286/pdf.

Gorman:1997:MBC

[Gor97b] Hugh Gorman. Manufacturing brownfields: The case of Neville Island, Pennsylvania. Technology and Culture, 38(3):539–574, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887267/
pdf.

Gorman:1998:OCT

[Gor98] Hugh Gorman. Our Children’s Toxic Legacy: How Science and Law Fail to Protect Us from Pesticides by John Wargo (review). Technology and Culture, 39(1):175–177, January 1998. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887204/pdf.

Gossel:1990:SAW

[Gos90] Patricia Peck Gossel. Saving America’s Wildlife by Thomas R. Dunlap (review). Technology and Culture, 31(4):887–888, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901671/pdf.

Gough:1998:WCC

[Gou98] J. B. Gough. Winecraft and chemistry in 18th-Century France: Chaptal and the invention of chaptalization. Technology and Culture, 39(1):74–104, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887171/pdf.

Grant:1991:SRG

[Gra91] H. Roger Grant. Street Railways and the Growth of Los Angeles by Robert C. Post (review). Technology and Culture, 32(2 (part 1)):421– 423, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888760/pdf.

Graham:1995:MSI

[Gra95] Margaret B. W. Graham. From, Machine Shop to Industrial Laboratory: Telegraphy and the Changing Context of American Invention, 1830–1920 by Paul Israel (review). Technology and Culture, 36(2):413–415, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887896/pdf.

Graham:1997:STE

[Gra97a] Loren Graham. Science, Technology, and Ecopolitics in the USSR by Miron Rezun (review). Technology and Culture, 38(3):804–806, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887315/pdf.

Grant:1997:PIR

[Gra97b] Jonathan Grant. Patents in Imperial Russia: a History of the Russian Institution of Invention Privileges under the Old Regime by Anneli Aer (review). Technology and Culture, 38(2):517–519, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887345/pdf.

Grant:1998:STH

[Gra98a] H. Roger Grant. Studies in Transport History Series (review). Technology and Culture, 39(3):549–550, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33178.

Grant:1998:PIT

[Gra98b] Jonathan Grant. The Politics of Industrialization in Tsarist Russia: The Association of Southern Coal and Steel Producers, 1874– 1914 by Susan P. McCaffray (review). Technology and Culture, 39 (1):159–160, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887194/pdf.

Greenberg:1991:MWW

[Gre91] Brian Greenberg. Men, Women, and Work: Class, Gender, and Protest in the New England Shoe Industry, 1780–1910 by Mary H. Blewett (review). Technology and Culture, 32(2 (part 1)):426–427, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888763/pdf.

Greenberg:1994:WCR

[Gre94] Brian Greenberg. “Who’s in charge”? A report from the “Dark Side”. Technology and Culture, 35(3):575–580, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887839/pdf.

Green:1995:GCA

[Gre95a] Venus Green. Goodbye central: Automation and the decline of “Personal service” in the Bell system, 1878–1921. Technology and Culture, 36(4):912–949, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887757/pdf.

Green:1995:RTA

[Gre95b] Venus Green. Race and technology: African American women in the Bell system, 1945–1980. Technology and Culture, 36(2S):S101– S144, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887711/pdf.

Greenberg:1999:TMT

[Gre99] Brian Greenberg. The terror of the machine: Technology, work, gender, and ecology on the U.S.-Mexico border. Technology and Culture, 40(1):158–160, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33303.

Gross:1990:FBP

[Gro90a] Laurence Gross. The Fragile Bridge: Paterson Silk Strike, 1913 by Steve Golin (review). Technology and Culture, 31(1):158–159, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888942/pdf.

Gross:1990:MMH

[Gro90b] Laurence Gross. The Men and the Mills: a History of the Southern Textile Industry by Mildred Gwin Andrews (review). Technology and Culture, 31(2):324–325, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888899/pdf.

Gross:1993:PEL

[Gro93] Laurence F. Gross. Problems in exhibiting labor in museums and a technological fix. Technology and Culture, 34(2):392–400, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888415/pdf.

Gross:1998:MWO

[Gro98] Laurence F. Gross. Managers and Workers: Origins of the Twentieth-Century Factory System in the United States, 1880–1920 (review). Technology and Culture, 39(4):778–780, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33250.

Gruen:1990:JSF

[Gru90] Adam L. Gruen. Journey into Space: The First Three Decades of Space Exploration by Bruce Murray (review). Technology and Culture, 31(4):897–898, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/901678/pdf.

Galambos:1998:EIB

[GS98] Louis Galambos and Jeffrey L. Sturchio. Evolutionary Innovations: The Business of Biotechnology by Maureen D. McKelvey (review). Technology and Culture, 39(2):332–334, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887149/pdf.

Grote:1999:PPB

[GS99] Jim Grote and P. Hans Sun. Prelapsarian perfection and blurred distinctions. Technology and Culture, 40(4):866–871, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33435.

Guajardo:1998:SAC

[Gua98] Guillermo Guajardo. Saberes andinos: Ciencia y tecnología en Bolivia, Ecuador, y Perú ed. by Marcos Cueto (review). Technology and Culture, 39(1):181–183, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887207/pdf.

Guilmartin:1991:SFG

[Gui91] John F. Guilmartin. The Seventy-Four Gun Ship: a Practical Treatise on the Art of Naval Architecture by Jean Boudriot (review). Technology and Culture, 32(2 (part 1)):440–441, April 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888772/pdf.

Guilmartin:1993:AGC

[Gui93a] J. F. Guilmartin. The Art of Gunfounding: The Casting of Bronze Cannon in the Late 18th Century ed. by Carel de Beer (review). Technology and Culture, 34(3):693–694, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888149/pdf.

Guilmartin:1993:MSP

[Gui93b] John F. Guilmartin. Moltke, Schlieffen, and Prussian War Planning by Arden Bucholz (review). Technology and Culture, 34(4):955–957, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888064/pdf.

Guilmartin:1995:WTS

[Gui95a] J. F. Guilmartin. World, Trade since 1431: Geography, Technology, and Capitalism by Peter J. Hugill (review). Technology and Culture, 36(1):170–172, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888009/pdf.

Guilmartin:1995:HPC

[Gui95b] John F. Guilmartin, Jr. High Performance: The Culture and Technology of Drag Racing, 1950–1990 by Robert C. Post (review). Technology and Culture, 36(4):1024–1026, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887765/pdf.

Guichand:1998:IHT

[Gui98a] Pierre Guichand. Irrigation and Hydraulic Technology: Medieval Spain and Its Legacy (review). Technology and Culture, 39(3):564– 566, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33188.

Guilmartin:1998:GHI

[Gui98b] John Francis Guilmartin. Gunpowder: The History of an International Technology (review). Technology and Culture, 39(3):547–549, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33181.

Guston:1991:GES

[Gus91] David H. Guston. Government and Expertise: Specialists, Administrators and Professionals, 1860–1919 ed. by Roy MacLeod (review). Technology and Culture, 32(3):606–607, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888703/pdf.

Gusterson:1998:MPS

[Gus98] Hugh Gusterson. Making PCR: a Story of Biotechnology (review). Technology and Culture, 39(4):792–794, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33253.

Guthrie:1990:FSB

[Gut90] Christopher Esco Guthrie. The French Sugar Business in the Eighteenth Century by Robert Louis Stein (review). Technology and Culture, 31(1):174–177, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888952/pdf.

Guthrie:1992:WRF

[Gut92] Christopher E. Guthrie. The Wine Revolution in France: The Twentieth Century by Leo Loubère (review). Technology and Culture, 33(2):402–404, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888333/pdf.

Guthrie:1993:FCP

[Gut93] Christopher E. Guthrie. France at the Crystal Palace: Bourgeois Taste and Artisan Manufacture in the Nineteenth Century by Whitney Walton (review). Technology and Culture, 34(4):942–944, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888056/pdf.

Guthrie:1995:CNC

[Gut95a] Christopher E. Guthrie. The Coachmen of Nineteenth-Century Paris: Service Workers and Class Consciousness by Nicholas Papayanis (review). Technology and Culture, 36(1):183–185, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888017/pdf.

Guthrie:1995:WKC

[Gut95b] Christopher E. Guthrie. The Weaver’s Knot: The Contradictions of Class Struggle and Family Solidarity in Western France, 1750– 1914 by Tessie P. Liu (review). Technology and Culture, 36(3):699– 701, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887815/pdf.

Hacker:1990:FAN

[Hac90a] Barton C. Hacker. Fallout: an American Nuclear Tragedy by Philip L. Fradkin (review). Technology and Culture, 31(2):344–346, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888911/pdf.

Hacker:1990:WSS

[Hac90b] Barton C. Hacker. War Stars: The Superweapon and the American Imagination by H. Bruce Franklin (review). Technology and Culture, 31(1):188–190, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888960/pdf.

Hacker:1990:SH

[Hac90c] Barton C. Hacker. Sally Hacker (1936–1988). Technology and Culture, 31(4):925–929, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/901689/pdf.

Hacker:1991:APW

[Hac91a] Barton C. Hacker. Atoms for Peace and War 1953–1961: Eisenhower and the Atomic Energy Commission by Richard G. Hewlett, Jack M. Holl (review). Technology and Culture, 32(2 (part 1)):450– 452, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888778/pdf.

Hacker:1991:STM

[Hac91b] Barton C. Hacker. Science, Technology and the Military ed. by Everett Mendelsohn, et al. (review). Technology and Culture, 32(3):643– 646, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888726/pdf.

Hacker:1991:I

[Hac91c] Barton C. Hacker. Introduction. Technology and Culture, 32(2 (part 2)):iii–iv, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888734/pdf.

Hacker:1992:SSU

[Hac92a] Barton C. Hacker. Soldiers and Scholars: The U.S. Army and the Uses of Military History, 1865–1920 by Carol Reardon (review). Technology and Culture, 33(4):832–833, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888105/pdf.

Hacker:1992:TWM

[Hac92b] Barton C. Hacker. The Transformation of War by Martin van Creveld (review). Technology and Culture, 33(3):630–631, July 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888277/pdf.

Hacker:1993:ENO

[Hac93] Barton C. Hacker. Engineering a new order: Military institutions, technical education, and the rise of the industrial state. Technology and Culture, 34(1):1–27, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888455/pdf.

Hacker:1994:CAN

[Hac94a] Barton Hacker. Containing the Atom: Nuclear Regulation in a Changing Environment, 1963–1971 by J. Samuel Walker (review). Technology and Culture, 35(2):444–446, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887950/pdf.

Hacker:1994:SVH

[Hac94b] Barton C. Hacker. Science with a Vengeance: How the Military Created the U.S. Space Sciences after World War II by David H. De-Vorkin (review). Technology and Culture, 35(3):628–629, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887862/pdf.

Hacker:1994:MIW

[Hac94c] Barton C. Hacker. Military institutions, weapons, and social change: Toward a new history of military technology. Technology and Culture, 35(4):768–834, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887722/pdf.

Hacker:1998:CHM

[Hac98] Barton C. Hacker. Chet Holifield: Master Legislator and Nuclear Statesman by Richard W. Dyke and Francis X. Gannon (review). Technology and Culture, 39(1):132–133, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887179/pdf.

Hafter:1990:TME

[Haf90] Daryl Hafter. Toward the Modern Economy: Early Industry in Europe, 1500–1800 by Myron P. Gutmann (review). Technology and Culture, 31(3):517–518, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888848/pdf.

Hafter:1992:WWN

[Haf92] Daryl M. Hafter. Work and Wages: Natural Law, Politics and the Eighteenth-Century French Trades by Michael Sonenscher (review). Technology and Culture, 33(3):609–610, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888265/pdf.

Hafter:1996:FIW

[Haf96] Daryl M. Hafter. The First Industrial Woman by Deborah Valenze (review). Technology and Culture, 37(4):834–836, October 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887410/pdf.

Hafter:1997:WWT

[Haf97] Daryl M. Hafter. Women Workers and Technological Change in Europe in the Nineteenth and Twentieth Centuries ed. by Gertjan de Groot, Marlou Schrover (review). Technology and Culture, 38(3):783–784, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887303/pdf.

Hafter:1999:RWR

[Haf99] Daryl M. Hafter. Ready-to-wear and ready-to-work: a century of industry and immigrants in Paris and New York. Technology and Culture, 40(1):177–178, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33284.

Hallion:1990:ISR

[Hal90] Richard P. Hallion. Igor Sikorsky: The Russian Years by K. N. Finne (review). Technology and Culture, 31(4):892–893, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901674/pdf.

Hall:1993:HTP

[Hal93a] Marie Boas Hall. Histoire technique de la production d’aluminium: Les apports français au développement international d’une industrie ed. by Paul Morel (review). Technology and Culture, 34(4):944–945, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888057/pdf.

Hall:1993:ATC

[Hal93b] Phyllis A. Hall. The appreciation of technology in Campanella’s “The City of the Sun”. Technology and Culture, 34(3):613–628, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888126/pdf.

Hales:1994:RRA

[Hal94] Peter B. Hales. On Rims and Ridges: The Los Alamos Area since 1880 by Hal K. Rothman (review). Technology and Culture, 35 (2):432–433, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887943.

Hall:1996:BSC

[Hal96] Bert Hall. Berthold Schwarz: Chemie und Waffentechnik im 15. Jahrhundert by Gerhard W. Kramer (review). Technology and Culture, 37(2):342–344, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887473/pdf.

Hall:1998:ALW

[Hal98] Bert Hall. The Anatomy of a Little War: a Diplomatic and Military History of the Gundovald Affair, 568–586 by Bernard S. Bachrach (review). Technology and Culture, 39(1):135–137, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887181/pdf.

Hamlin:1991:UMG

[Ham91] Christopher Hamlin. “Underground Manchester” at the Greater Manchester Museum of Science and Industry. Technology and Culture, 32(1):97–101, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888789/pdf.

Hamlin:1992:SHA

[Ham92a] Christopher Hamlin. The Sanitarians: a History of American Public Health by John Duffy (review). Technology and Culture, 33 (4):826–827, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888101/pdf.

Hamlin:1992:ECE

[Ham92b] Christopher Hamlin. Edwin Chadwick and the engineers, 1842–1854: Systems and antisystems in the pipe-and-brick sewers war. Technology and Culture, 33(4):680–709, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888077/pdf.

Hamlin:1995:MWE

[Ham95] Christopher Hamlin. Mixing the Waters: Environment, Politics, and the Building of the Tennessee-Tombigbee Waterway by Jeffrey K. Stine (review). Technology and Culture, 36(3):687–690, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887809/pdf.

Hamblin:1998:SNA

[Ham98a] Jacob D. Hamblin. Supplying the Nuclear Arsenal: American Production Reactors, 1942–1992 by Rodney P. Carlisle and Joan M. Zenzen (review). Technology and Culture, 39(2):365–367, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887166/pdf.

Hamlin:1998:ABE

[Ham98b] Christopher Hamlin. Agricultural Biotechnology and the Environment: Science, Policy, and Social Issues by Sheldon Krimsky and Roger Wrubel (review). Technology and Culture, 39(2):334–336, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887150/pdf.

Hansen:1993:FAC

[Han93] David M. Hansen. Fortress America: The Corps of Engineers, Hampton Roads, and United States Coastal Defense by David A. Clary (review). Technology and Culture, 34(1):150–151, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888471/pdf.

Hansen:1999:HHP

[Han99] James E. Hansen. Harvesting the high plains: John Kriss and the business of wheat farming, 1920–1950. Technology and Culture, 40(2):422–423, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33347.

Harris:1991:DHY

[Har91] J. R. Harris. Daniel Hechstetter the Younger: Memorabilia and Letters 1600–1639 — Copper Works and Life in Cumbria ed. by George Hammersley (review). Technology and Culture, 32(2 (part 1)):437– 438, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888770/pdf.

Harwit:1992:ESS

[Har92] Martin Harwit. Exploring the Sun: Solar Science since Galileo by Karl Hufbauer (review). Technology and Culture, 33(3):644–645, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888286/pdf.

Haard:1993:TTS

[Hår93a] Mikael Hård. Technology Transfer and Scandinavian Industrialisation ed. by Kristine Bruland (review). Technology and Culture, 34(3):687–688, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888145/pdf.

Harris:1993:PTE

[Har93b] C. M. Harris. The Papers of Thomas A. Edison, vol. 2: From Workshop to Laboratory, June 1873–March 1876 ed. by Robert A. Rosenberg, et al. (review). Technology and Culture, 34(3):696–698, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888151/pdf.

Hard:1994:DHC

[Har94] Mikael Hard. District Heating Comes to Town: The Social Shaping of an Energy System by Jane Summerton (review). Technology and Culture, 35(2):415–416, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887933/pdf.

Harris:1995:HBC

[Har95a] J. R. Harris. The History of the British Coal Industry, vol. 1: Before 1700: Towards the Age of Coal by John Hatcher (review). Technology and Culture, 36(2):399–401, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887888/pdf.

Harwit:1995:BHT

[Har95b] Martin Harwit. Black Holes and Time Warps: Einstein’s Outrageous Legacy by Kip S. Thorne (review). Technology and Culture, 36(1):208–209, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888031/pdf.

Haard:1996:STP

[Hår96a] Mikael Hård. The Social Theory of Practices: Tradition, Tacit Knowledge, and Presuppositions by Stephen Turner (review). Technology and Culture, 37(3):652–653, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887462/pdf.

Harris:1996:PBH

[Har96b] C. M. Harris. The Papers of Benjamin Henry Latrobe. Series II: The Architectural and Engineering Drawings. Vol. 2: The Architectural Drawings of Benjamin Henry Latrobe. Pts. 1 and 2 ed. by Jeffrey A. Cohen and Charles E. Brownell (review). Technology and Culture, 37(3):629–631, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887450/pdf.

Harris:1996:BIR

[Har96c] J. R. Harris. The British Industrial Revolution: an Economic Perspective ed. by Joel Mokyr (review). Technology and Culture, 37 (1):180–181, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887549/pdf.

Harper:1997:EAP

[Har97] Douglas Harper. Eldoret: an African Poetics of Technology by Richard M. Swiderski (review). Technology and Culture, 38(2):505– 508, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887339/pdf.

Haard:1998:TFD

[Hår98a] Mikael Hård. Teknologiske forandringer i dansk industri 1870–1896. Vol. 4, Dansk in-dustri efter 1870 by Ole Hyldtoft (review). Technology and Culture, 39(2):343–344, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887154/pdf.

Harley:1998:CCE

[Har98b] C. Knick Harley. Capitalism in Context: Essays on Economic Development and Cultural Change in Honor of R. M. Hartwell (review). Technology and Culture, 39(3):543–544, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33191.

Hard:1999:VOK

[Har99a] Mikael Hard. Vetenskap och karriar: Svenska fysiker som lektorer, akademiker och industriforskare under 1900-talets forsta halft. Technology and Culture, 40(1):197–199, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33292.

Harris:1999:JMB

[Har99b] Charles M. Harris. Jefferson’s memorandum books: Accounts, with legal records and miscellany, 1767–1826. Technology and Culture, 40(3):662–664, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33377.

Hart:1999:TCD

[Har99c] Stephen S. Hart. The Texas city disaster, 1947. Technology and Culture, 40(1):189–190, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33313.

Hartmann:1999:GZV

[Har99d] Rudi Hartmann. Geschichte der Zukunft des Verkehrs: Verkehrskonzepte von der Fruhen Neuzeit bis zum 21. Jahrhundert. Technology and Culture, 40(2):429–431, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33335.

Hashimoto:1993:BGN

[Has93] Takehiko Hashimoto. Bunha to Gijutsu no Kosaten: Kikaiseizu oyobi Seizukyoiku no Shinso no Kenkyu (Culture and technology at crossroads: Studies on mechanical drawing and drawing education) by Sadahiko Mori (review). Technology and Culture, 34(3):688– 689, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888146/pdf.

Hatch:1990:GPH

[Hat90] John Hatch. The Generation of Power: The History of Dneprostroi by Anne D. Rassweiler (review). Technology and Culture, 31(3):541– 542, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888862/pdf.

Hauck:1990:CED

[Hau90] George F. W. Hauck. Civil engineering during the enlightenment — ICOHTEC Symposium, Madrid, September 5–9, 1988. Technology and Culture, 31(1):114–120, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888924/pdf.

Hay:1990:MMM

[Hay90] Duncan Hay. “Made in Maine” at the Maine State Museum, Augusta. Technology and Culture, 31(3):456–462, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888832/pdf.

Hayward:1996:TMU

[Hay96] Alexander Hayward. Technology museums in the United Kingdom. Technology and Culture, 37(1):138–146, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887539/pdf.

Hugill:1998:STI

[HC98] Peter J. Hugill and British-Dutch Maritime History Conference. Shipping, Technology, and Imperialism: Papers Presented to the Third British–Dutch Maritime History Conference (review). Technology and Culture, 39(4):799–801, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33247.

Hall:1990:MRM

[HD90] Bert S. Hall and Kelly R. DeVries. The Military Revolution: Military Innovation and the Rise of the West, 1500–1800 by Geoffrey Parker (review). Technology and Culture, 31(3):500–507, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888843/pdf.

Headrick:1990:MMM

[Hea90] Daniel R. Headrick. Machines as the Measure of Men: Science, Technology, and Ideologies of Western Dominance by Michael Adas (review). Technology and Culture, 31(4):862–863, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901656/pdf.

Headrick:1991:LRT

[Hea91a] Daniel R. Headrick. The Lever of Riches: Technological Creativity and Economic Progress by Joel Mokyr (review). Technology and Culture, 32(4):1105–1106, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888656/pdf.

Headrick:1991:TWI

[Hea91b] Daniel R. Headrick. Tools of War: Instruments, Ideas, and Institutions of Warfare, 1445–1871 ed. by John A. Lynn (review). Technology and Culture, 32(3):648–650, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888728/pdf.

Headrick:1992:STC

[Hea92] Daniel R. Headrick. Science, Technology and Colonisation: an Indian Experience, 1757–1857 by Satpal Sangwan (review). Technology and Culture, 33(4):801–803, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888087/pdf.

Headrick:1996:SSO

[Hea96] Daniel R. Headrick. Spying without Spies: Origins of America’s Secret Nuclear Surveillance System by Charles A. Ziegler, David Jacobson (review). Technology and Culture, 37(2):391–392, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887499/pdf.

Headrick:1998:TEO

[Hea98] Daniel R. Headrick. Technology and European Overseas Enterprise: Diffusion, Adaption, and Adoption ed. by Michael Adas (review). Technology and Culture, 39(2):352–354, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887159/pdf.

Headrick:1999:ETS

[Hea99a] Daniel R. Headrick. Entreprise, technologie et souverainété: Les télécommunications transatlantiques de la France, XIXe–XXe siècles (review). Technology and Culture, 40(1):167–169, January 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33285.

Headrick:1999:NBP

[Hea99b] Daniel R. Headrick. The Nature of the Book: Print and Knowledge in the Making (review). Technology and Culture, 40(4):872–873, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33436.

Hecht:1992:ACN

[Hec92] Gabrielle Hecht. Alternating Currents: Nationalized Power in France, 1946–1970 by Robert L. Frost (review). Technology and Culture, 33(4):808–810, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888091/pdf.

Hecht:1994:PDN

[Hec94] Gabrielle Hecht. Political designs: Nuclear reactors and national policy in postwar France. Technology and Culture, 35(4):657–685, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887718/pdf.

Heinz:1991:ITT

[Hei91] Lisa C. Heinz. Innovation through Technical and Scientific Information: Government and Industry Cooperation by Steven Ballard, et al. (review). Technology and Culture, 32(4):1128–1129, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888670/pdf.

Heimpel:1995:ABW

[Hei95a] Wolfgang Heimpel. Archaic Bookkeeping: Writing and Techniques of Economic Administration in the Ancient Near East by Hans J. Nissen, et al. (review). Technology and Culture, 36(1):197–199, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888025/pdf.

Heims:1995:SWL

[Hei95b] Steve Joshua Heims. Steel Will: The Life of Tad Sendzimir by Vanda Sendzimir (review). Technology and Culture, 36(1):187–189, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888019/pdf.

Heide:1999:TST

[Hei99] Lars Heide. Til samfundets tarv: Forskningscenter Risøs historie (In the advantage of society: the history of Research Centre Risø) (review). Technology and Culture, 40(4):914–915, October 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33437.

Hensley:1999:MCT

[Hen99] Paul B. Hensley. Mastered by the clock: Time, slavery, and freedom in the American South. Technology and Culture, 40(2):406–408, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33358.

Herken:1993:ASP

[Her93] Gregg Herken. The Advisers: Scientists in the Policy Process by Bruce L. R. Smith (review). Technology and Culture, 34(2):464– 465, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888450/pdf.

Herken:1994:CAM

[Her94] Gregg Herken. Citizens against the MX: Public Languages in the Nuclear Age by Matthew Glass (review). Technology and Culture, 35(3):633–635, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887865/pdf.

Herzig:1998:NBM

[Her98] Rebecca Herzig. Naked to the Bone: Medical Imaging in the Twentieth Century by Bettyann Holtzmann Kevles (review). Technology and Culture, 39(2):330–332, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887148/pdf.

Herzig:1999:GPF

[Her99a] Rebecca M. Herzig. Gendered practices: Feminist studies of technology and society. Technology and Culture, 40(3):655–657, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33379.

Herzig:1999:RRN

[Her99b] Rebecca M. Herzig. Removing roots: “North American Hiroshima maidens” and the X ray. Technology and Culture, 40(4):723–745, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33438.

Hessler:1999:HHG

[Hes99] Martina Hessler. His and Hers: Gender, Consumption, and Technology (review). Technology and Culture, 40(4):875–878, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33439.

Hevly:1995:PTD

[Hev95] Bruce Hevly. From Polaris to Trident: The Development of U.S. Fleet Ballistic Missile Technology by Graham Spinardi (review). Technology and Culture, 36(2):430–432, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887906/pdf.

Hewlett:1994:GSB

[Hew94] Richard G. Hewlett. Genius in the Shadows: a Biography of Leo Szilard — the Man behind the Bomb by William Lanouette, and: The Rickover Effect: How One Man Made a Difference by Theodore Rockwell (review). Technology and Culture, 35(2):442–444, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887949.

Heymann:1998:SHS

[Hey98] Matthias Heymann. Signs of hubris: The shaping of wind technology styles in Germany, Denmark, and the United States, 1940– 1990. Technology and Culture, 39(4):641–670, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33265.

Higham:1990:PPW

[Hig90] Robin Higham. Power: The Pratt & Whitney Canada Story by Kenneth H. Sullivan and Larry Milberry (review). Technology and Culture, 31(4):896–897, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901677/pdf.

Higham:1991:SEG

[Hig91] Robin Higham. Sir Eric Geddes: Business and Government in War and Peace by Keith Grieves (review). Technology and Culture, 32(1):161, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888812/pdf.

Higham:1992:BAI

[Hig92] Robin Higham. The British Aircraft Industry by Keith Hayward (review). Technology and Culture, 33(2):380–381, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888320/pdf.

Higman:1996:STS

[Hig96] B. W. Higman. Science, Technology, and Society in the Third World: an Annotated Bibliography ed. by Wesley Shrum, et al. (review). Technology and Culture, 37(4):861–863, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887425/pdf.

Higgs:1997:PTE

[Hig97a] Eric Higgs. The Practice of Technology: Exploring Technology, Ecophilosophy, and Spiritual Disciplines for Vital Links by Alan R. Drengson (review). Technology and Culture, 38(3):811–812, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887319/pdf.

Higham:1997:SUB

[Hig97b] Robin Higham. Stalking the U-Boat: USAAF Offensive Antisubmarine Operations in World War II by Max Schoenfeld (review). Technology and Culture, 38(1):262–263, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887391/pdf.

Higham:1998:SWU

[Hig98] Robin D. S. Higham. Secret Weapon: U.S. High-Frequency Direction Finding in the Battle of the Atlantic (review). Technology and Culture, 39(4):808–809, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33263.

Hilgartner:1991:OST

[Hil91a] Stephen Hilgartner. Owning Scientific and Technical Information: Value and Ethical Issues ed. by Vivian Weil, John W. Snapper (review). Technology and Culture, 32(4):1129–1131, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888671/pdf.

Hilton:1991:TIH

[Hil91b] George W. Hilton. Transportation in Iowa: a Historical Summary by William H. Thompson (review). Technology and Culture, 32(3): 626, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888715/pdf.

Hilton:1992:BHF

[Hil92a] George W. Hilton. Over and Back: The History of Ferryboats in New York Harbor by Brian J. Cudahy (review). Technology and Culture, 33(2):373–375, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888315/pdf.

Hilton:1992:RTG

[Hil92b] George W. Hilton. Railroads Triumphant: The Growth, Rejection, and Rebirth of a Vital American Force by Albro Martin (review). Technology and Culture, 33(4):820–821, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888097/pdf.

Hilton:1995:GTE

[Hil95] George W. Hilton. Getting There: The Epic Struggle between Road and Rail in the American Century by Stephen B. Goddard (review). Technology and Culture, 36(3):708–710, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887820/pdf.

Hilmes:1997:MWR

[Hil97] Michele Hilmes. Media at War: Radio’s Challenge to the Newspapers, 1924–1939 by Gwenyth L. Jackaway (review). Technology and Culture, 38(2):525–521, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887349/pdf.

Hills:1999:PPP

[Hil99] Richard Leslie Hills. Productivity and performance in the paper industry: Labour, capital, and technology in Britain and America, 1860–1914. Technology and Culture, 40(3):679–681, July 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33397.

Hine:1994:SPF

[Hin94] Christine M. Hine. Stealing into Print: Fraud, Plagiarism, and Misconduct in Scientific Publishing by Marcel C. LaFollette (review). Technology and Culture, 35(3):643–645, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887871/pdf.

Hirsh:1992:EAS

[Hir92] Richard F. Hirsh. Electrifying America: Social Meanings of a New Technology by David E. Nye (review). Technology and Culture, 33(3):595–596, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888256/pdf.

Hirsh:1996:WCS

[Hir96] Richard F. Hirsh. Wolf Creek Station: Kansas Gas and Electric Company in the Nuclear Era by Craig Miner (review). Technology and Culture, 37(2):394–395, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887501/pdf.

Hard:1999:GTG

[HK99] Mikael Hard and Andreas Knie. The grammar of technology: German and French diesel engineering, 1920–1940. Technology and Culture, 40(1):26–46, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33318.

Hovis:1996:MET

[HM96] Logan Hovis and Jeremy Mouat. Miners, engineers, and the transformation of work in the Western mining industry, 1880–1930. Technology and Culture, 37(3):429–456, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887430/pdf.

Hochfelder:1998:GMT

[Hoc98] David Hochfelder. The Grammar of the Machine: Technical Literacy and Early Industrial Expansion in the United States by Edward W. Stevens Jr. (review). Technology and Culture, 39(2):299–300, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887132/pdf.

Hodson:1993:MAC

[Hod93] Joel Hodson. Mapping American Culture ed. by Wayne Franklin, Michael Steiner (review). Technology and Culture, 34(4):953–954, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888062/pdf.

Hoffman:1991:BIS

[Hof91] Abraham Hoffman. Battling the Inland Sea: American Political Culture, Public Policy, and the Sacramento Valley, 1850–1986 by Robert Kelley (review). Technology and Culture, 32(2 (part 1)):403– 404, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888750/pdf.

Hoffman:1992:WCN

[Hof92] Abraham Hoffman. Water and the City: The Next Century ed. by Howard Rosen, Ann D. Keating (review). Technology and Culture, 33(4):828, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888102/pdf.

Holley:1990:WBY

[Hol90] I. B. Holley. Wild Blue Yonder: Money, Politics, and the B-1 Bomber by Nick Kotz (review). Technology and Culture, 31(1):187–188, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888959/pdf.

Holton:1991:ESR

[Hol91] Gerald Holton. Einstein in Spain: Relativity and the Recovery of Science by Thomas F. Glick (review). Technology and Culture, 32 (1):179–182, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888824/pdf.

Holmfeld:1993:EIL

[Hol93] John D. Holmfeld. Engines and Innovation: Lewis Laboratory and American Propulsion Technology by Virginia P. Dawson (review). Technology and Culture, 34(4):968–969, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888071/pdf.

Holley:1994:GWA

[Hol94] I. B. Holley, Jr. The Great War in the Air: Military Aviation from 1909 to 1921 by John H. Morrow, Jr. (review). Technology and Culture, 35(3):626–628, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887861/pdf.

Holley:1995:AWM

[Hol95] I. B. Holley, Jr. Admiral William A. Moffett: Architect of Naval Aviation by William F. Trimble (review). Technology and Culture, 36(2):428–430, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887905/pdf.

Holley:1997:EMR

[Hol97a] I. B. Holley, Jr. Engineer Memoirs by E. L. Rouny (review). Technology and Culture, 38(1):263–265, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887392/pdf.

Holweger:1997:RAW

[Hol97b] Jennifer Holweger. Reclaiming the Arid West: The Career of Francis G. Newlands by William D. Rowley (review). Technology and Culture, 38(4):997–998, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887249/pdf.

Holley:1999:MII

[Hol99] I. B. Holley. Military innovation in the interwar period. Technology and Culture, 40(1):190–192, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33299.

Homburg:1994:GVS

[Hom94] Ernst Homburg. De geschiedenis van de scheikunde in Nederland: Van alchemie tot chemie en chemische industrie rond 1900 by H. A. M. Snelders (review). Technology and Culture, 35(3):613–614, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887853/pdf.

Hong:1994:MMO

[Hon94] Sungook Hong. Marconi and the Maxwellians: The origins of wireless telegraphy revisited. Technology and Culture, 35(4):717–749, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887720/pdf.

Hong:1997:OEW

[Hon97] Sungook Hong. From, Obscurity to Enigma: The Work of Oliver Heaviside, 1872–1889 by Ido Yavetz (review). Technology and Culture, 38(3):765–767, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887294/pdf.

Hong:1998:KWC

[Hon98] Sungook Hong. Korean Water-Clocks: “Chagyongnu,” the Striking Clepsydra, and the History of Control and Instrumentation Engineering (review). Technology and Culture, 39(3):553–555, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33192.

Hooley:1999:SFI

[Hoo99] Ray E. Hooley. Sparks and flames: Ignition in engines — an historical approach. Technology and Culture, 40(3):687–688, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33395.

Horrigan:1992:MWN

[Hor92] Brian Horrigan. “A material world” at the National Museum of American History. Technology and Culture, 33(1):132–139, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888341/pdf.

Horowitz:1994:SWM

[Hor94] Roger Horowitz. “Stalin over Wisconsin”: The Making and Unmaking of Militant Unionism, 1900–1950 by Stephen Meyer (review). Technology and Culture, 35(3):623–624, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887859/pdf.

Horning:1997:TTS

[Hor97a] Susan Schmidt Horning. Twenty-third symposium of the International Committee for the History of Technology, Budapest, Hungary, August 7–11, 1996. Technology and Culture, 38(2):488–492, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887333/pdf.

Horowitz:1997:WMW

[Hor97b] Roger Horowitz. “Where men will not work”: Gender, power, space, and the sexual division of labor in America’s meatpacking industry, 1890–1990. Technology and Culture, 38(1):187–213, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887372/pdf.

Horning:1999:SYC

[Hor99] Susan Schmidt Horning. Any sound you can imagine: Making music/consuming technology. Technology and Culture, 40(3):709–711, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33409.

Houshower:1990:BCS

[Hou90] Hans Houshower. Bloomington’s C&A Shops: Our Lives Remembered ed. by Michael G. Matejka, Greg Koos (review). Technology and Culture, 31(2):321–322, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888897/pdf.

Houchin:1997:EFP

[Hou97] Roy F. Houchin II. Engineering the F-4 Phantom II: Parts into Systems by Glenn E. Bugos (review). Technology and Culture, 38 (4):1017–1020, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887262/pdf.

Howell:1993:LRM

[How93] Joel D. Howell. The Laboratory Revolution in Medicine ed. by Andrew Cunningham, Perry Williams (review). Technology and Culture, 34(4):924–926, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888045/pdf.

Howe:1998:LTM

[How98] Barbara J. Howe. Learning from Things: Method and Theory of Material Culture Studies ed. by W. David Kingery (review). Technology and Culture, 39(2):317–319, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887141/pdf.

Hoy:1992:DDH

[Hoy92] Suellen Hoy. Domesticity and Dirt: Housewives and Domestic Servants in the United States, 1920–1945 by Phyllis Palmer (review). Technology and Culture, 33(1):149–150, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888346/pdf.

Hilaire-Perez:1991:ISC

[HP91] Liliane Hilaire-Pérez. Invention and the state in 18th-Century France. Technology and Culture, 32(4):911–931, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888647/pdf.

Hirsh:1996:MSA

[HS96] Richard F. Hirsh and Adam H. Serchuk. Momentum shifts in the American electric utility system: Catastrophic change — or no change at all? Technology and Culture, 37(2):280–311, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887469/pdf.

Hufbauer:1992:STP

[Huf92] Karl Hufbauer. Space Technology and Planetary Astronomy by Joseph N. Tatarewicz (review). Technology and Culture, 33(3):646– 647, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888287/pdf.

Hughes:1990:LML

[Hug90] Thomas Hughes. Lewis Mumford: a Life by Donald Miller (review). Technology and Culture, 31(4):913–915, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901686/pdf.

Hughes:1993:CSS

[Hug93a] Thomas P. Hughes. Cyril Stanley Smith (1903–1992). Technology and Culture, 34(3):716–719, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888162/pdf.

Hugill:1993:GSH

[Hug93b] Peter J. Hugill. The Geography of Science by Harold Dorn (review). Technology and Culture, 34(1):135–136, January 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888462/pdf.

Hugill:1995:APA

[Hug95] Peter J. Hugill. Down the Asphalt Path: The Automobile and the American City by Clay McShane (review). Technology and Culture, 36(2):420–421, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887900/pdf.

Hughes:1996:LMP

[Hug96a] Agatha C. Hughes. Lewis Mumford and Patrick Geddes: The Correspondence ed. by Frank G. Novak, Jr. (review). Technology and Culture, 37(4):828–830, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887407/pdf.

Hugill:1996:TCE

[Hug96b] Peter J. Hugill. Taking Charge: The Electric Automobile in America by Michael Brian Schiffer (review). Technology and Culture, 37(2):379–381, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887493/pdf.

Hughes:1999:CLC

[Hug99] J. Donald Hughes. Common lands, common people: The origins of conservation in Northern New England. Technology and Culture, 40(1):187–188, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33291.

Huibregtse:1999:AR

[Hui99] Jon R. Huibregtse. American railroads. Technology and Culture, 40(3):666–667, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33408.

Hunley:1995:ERH

[Hun95] J. D. Hunley. The enigma of Robert H. Goddard. Technology and Culture, 36(2):327–350, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887883/pdf.

Hunley:1996:STN

[Hun96] J. D. Hunley. Science, Technology and National Socialism ed. by Monika Renneberg, Mark Walker (review). Technology and Culture, 37(1):196–197, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887558/pdf.

Hurt:1990:TMS

[Hur90] R. Douglas Hurt. Threshing in the Midwest, 1820–1940: a Study of Traditional Culture and Technological Change by J. Sanford Rikoon (review). Technology and Culture, 31(1):178–180, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888954/pdf.

Hurley:1998:VEE

[Hur98] Andrew Hurley. Visions of Eden: Environmentalism, Urban Planning, and City Building in St. Petersburg, Florida, 1900–1995 (review). Technology and Culture, 39(4):806–808, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33259.

Hurt:1999:CWA

[Hur99] R. Douglas Hurt. The Culture of Wilderness: Agriculture as Colonization in the American West (review). Technology and Culture, 40(2):420–421, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33343.

Hutchins:1992:SFP

[Hut92] Louis P. Hutchins. Shattering: Food, Politics, and the Loss of Genetic Diversity by Cary Fowler and Pat Mooney (review). Technology and Culture, 33(2):404–405, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888334/pdf.

Henninger-Voss:1998:LST

[HV98] M. Henninger-Voss. Labour, Science, and Technology in France, 1500–1620 by Henry Heller (review). Technology and Culture, 39 (2):322–324, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887144/pdf.

Henninger-Voss:1999:WWR

[HV99] Mary J. Henninger-Voss. Weapons and warfare in Renaissance Europe: Gunpowder, technology, and tactics. Technology and Culture, 40(1):134–136, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33287.

Hyde:1990:AAC

[Hyd90] Charles K. Hyde. The Art of American Car Design: The Profession and Personalities — “Not Simple like Simon” by C. Edson Armi, and: Power behind the Wheel: Creativity and the Evolution of the Automobile by Walter J. Boyne (review). Technology and Culture, 31(2):314–315, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888893/pdf.

Hyde:1999:PHM

[Hyd99] Ralph Hyde. The Panorama: History of a Mass Medium (review). Technology and Culture, 40(4):900–901, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33440.

Ilerbaig:1996:PTS

[Ile96] Juan Ilerbaig. Philosophy of Technology in Spanish Speaking Countries ed. by Carl Mitcham (review). Technology and Culture, 37 (1):178–179, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887548/pdf.

Ingerson:1991:ATF

[Ing91] Alice E. Ingerson. Americans and Their Forests: a Historical Geography by Michael Williams (review). Technology and Culture, 32(2 (part 1)):399–401, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888748/pdf.

Ingham:1994:HIC

[Ing94] John N. Ingham. Harrisburg Industrializes: The Coming of Factories to an American Community by Gerald G. Eggert (review). Technology and Culture, 35(3):620–621, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887857/pdf.

Ingold:1999:PFA

[Ing99] Tim Ingold. Paths of fire: an anthropologist’s inquiry into Western technology. Technology and Culture, 40(1):130–132, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33274.

Israel:1991:POR

[IR91] Paul Israel and Robert Rosenberg. Patent Office records as a historical source: The case of Thomas Edison. Technology and Culture, 32 (4):1094–1101, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888654/pdf.

Irish:1997:CWC

[Iri97] Sharon Irish. Carl W. Condit (1914–1997). Technology and Culture, 38(4):1026–1030, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887265/pdf.

Israel:1990:SIE

[Isr90] Paul Israel. The Sources of Innovation by Eric von Hippel (review). Technology and Culture, 31(2):333–334, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888904/pdf.

Israel:1991:EBI

[Isr91] Paul Israel. Edison and the Business of Innovation by Andre Millard (review). Technology and Culture, 32(4):1115–1117, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888663/pdf.

Israel:1992:ICB

[Isr92a] Paul Israel. Innovation at the Crossroads between Science and Technology ed. by M. Kranzberg, et al. (review). Technology and Culture, 33(1):181–182, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888365/pdf.

Israel:1992:RMA

[Isr92b] Paul Israel. Robert Maillart and the Art of Reinforced Concrete by David P. Billington (review). Technology and Culture, 33(3):621– 622, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888272/pdf.

Israel:1994:PJH

[Isr94a] Paul B. Israel. The Papers of Joseph Henry. Vol. 6: January 1844– December 1846 — the Princeton Years ed. by Marc Rothenberg, et al. (review). Technology and Culture, 35(2):416–418, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887934/pdf.

Israel:1994:EIP

[Isr94b] Paul B. Israel. Enthusiasts and innovators: “Possible Dreams” and the “Innovation Station” at the Henry Ford Museum. Technology and Culture, 35(2):396–401, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887924/pdf.

Jackson:1991:HJK

[Jac91] Donald C. Jackson. Henry J. Kaiser: Builder in the Modern American West by Mark S. Foster, and: Henry J. Kaiser — American Empire Builder: an Insider’s View by Albert P. Heiner (review). Technology and Culture, 32(3):612–614, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888708/pdf.

Jackson:1993:WWS

[Jac93a] Donald C. Jackson. Workshop of the World: a Selective Guide to the Industrial Archeology of Philadelphia by the Oliver Evans Chapter of the Society for Industrial Archeology (review). Technology and Culture, 34(1):159–160, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888476/pdf.

Jackson:1993:EPE

[Jac93b] Donald C. Jackson. Engineering in the Progressive Era: a new look at Frederick Haynes Newell and the U.S. Reclamation Service. Technology and Culture, 34(3):539–574, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888124/pdf.

Jackson:1995:HDU

[Jac95a] Donald C. Jackson. A History of Dams: The Useful Pyramids by Nicholas J. Schnitter (review). Technology and Culture, 36 (1):179–180, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888014/pdf.

Jackson:1995:GSA

[Jac95b] Donald C. Jackson. The Gas Station in America by John A. Jakle, Keith A. Sculle (review). Technology and Culture, 36(4):1044–1045, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887776/pdf.

Jackson:1996:RAD

[Jac96] Donald C. Jackson. Raising Arizona’s Dams: Daily Life, Danger, and Discrimination in the Dam Construction Camps of Central Arizona, 1890s–1940s by A. E. Rogge, et al. (review). Technology and Culture, 37(2):358–360, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887482/pdf.

Jacobson:1997:MTS

[Jac97] Charles Jacobson. Managing Technology in Society: The Approach of Constructive Technology Assessment ed. by Arie Rip, et al. (review). Technology and Culture, 38(3):813–814, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887320/pdf.

Jager:1999:TSS

[Jag99] Ronald Jager. Tool and symbol: The success of the double-bitted axe in North America. Technology and Culture, 40(4):833–860, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33442.

Jakab:1998:PWT

[Jak98] Peter L. Jakab. The Peenemünde Wind Tunnels: a Memoir by Peter P. Wegener (review). Technology and Culture, 39(1):128–130, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887177/pdf.

James:1990:TIC

[Jam90] Dilmus D. James. Technological Innovation in China: The Case of Shanghai’s Electronics Industry by Denis Fred Simon, Detlef Rehn, and: Closing the Gap: Computer Development in the People’s Republic of China by Otto W. Witzell and J. K. Lee Smith (review). Technology and Culture, 31(4):905–907, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901683/pdf.

Janssen:1991:PRA

[Jan91] Barbara S. Janssen. Papers Relating to the Administration of the U.S. Patent Office during the Superintendency of William Thornton, 1802–1828 ed. by C. M. Harris, Daniel Preston (review). Technology and Culture, 32(4):1111–1113, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888660/pdf.

Jarrell:1990:SSI

[Jar90] Richard A. Jarrell. Shaping Science and Industry: a History of Australia’s Council for Scientific and Industrial Research, 1926–49 by Carl B. Schedvin (review). Technology and Culture, 31(2):353–355, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888917/pdf.

Jardini:1995:ISR

[Jar95] David Jardini. From iron to steel: The recasting of the Jones and laughlins workforce between 1885 and 1896. Technology and Culture, 36(2):271–301, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887881/pdf.

Jeffrey:1995:PHG

[Jef95] Kirk Jeffrey. Pacing the heart: Growth and redefinition of a medical technology, 1952–1975. Technology and Culture, 36(3):583–624, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887791/pdf.

Jeremy:1990:ASF

[Jer90a] David J. Jeremy. The Arkwrights: Spinners of Fortune by R. S. Fitton (review). Technology and Culture, 31(3):522–524, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888851/pdf.

Jeremy:1990:ENW

[Jer90b] David J. Jeremy. The “Emigrants to a New World” Gallery at the Merseyside Maritime Museum, Liverpool. Technology and Culture, 31(2):278–283, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888880/pdf.

Jeremy:1993:LSC

[Jer93] David J. Jeremy. Lancashire on the Scrapheap: The Cotton Industry, 1945–1970 by John Singleton (review). Technology and Culture, 34 (1):175–177, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888485/pdf.

Jeremy:1994:MID

[Jer94a] David J. Jeremy. Managing Industrial Decline: Entrepreneurship in the British Coal Industry between the Wars by Michael Dintenfass (review). Technology and Culture, 35(3):624–626, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887860/pdf.

Jeremy:1994:CID

[Jer94b] David J. Jeremy. The Course of Industrial Decline: The Boott Cotton Mills of Lowell, Massachusetts, 1835–1955 by Laurence F. Gross (review). Technology and Culture, 35(4):868–869, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887727/pdf.

Jeremy:1997:OJT

[Jer97] David J. Jeremy. The Origins of Japanese Trade Supremacy: Development and Technology in Asia from 1540 to the Pacific War by Christopher Howe (review). Technology and Culture, 38(2):501– 503, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887337/pdf.

Jesseph:1997:DEM

[Jes97] Douglas M. Jesseph. Discipline and Experience: The Mathematical Way in the Scientific Revolution by Peter Dear (review). Technology and Culture, 38(3):756–758, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887289/pdf.

Jeff:1998:JAC

[JJ98] Ph. D. Horn Jeff and Margaret C. Jacob. Jean-Antoine Chaptal and the cultural roots of French industrialization. Technology and Culture, 39(4):671–698, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33266.

Johnson-McGrath:1997:WBB

[JM97] Julie Johnson-McGrath. Who built the built environment? Artifacts, politics, and urban technology. Technology and Culture, 38(3):690–696, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887272/pdf.

Joerges:1990:ITS

[Joe90] Bernward Joerges. Images of technology in sociology: Computer as butterfly and bat. Technology and Culture, 31(2):203–227, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888875/pdf.

John:1991:KPD

[Joh91] Richard R. John. Knowledge Is Power: The Diffusion of Information in Early America, 1700–1865 by Richard D. Brown (review). Technology and Culture, 32(3):617–619, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888710/pdf.

John:1992:CCE

[Joh92] Richard R. John. Communication as Culture: Essays on Media and Society by James W. Carey (review). Technology and Culture, 33 (1):200–202, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888377/pdf.

Johnson:1997:PAJ

[Joh97a] Stephen B. Johnson. Powering Apollo: James E. Webb of NASA by Henry W. Lambright (review). Technology and Culture, 38(3):793– 795, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887309/pdf.

Johnson:1997:TAB

[Joh97b] Stephen B. Johnson. Three approaches to big technology: Operations research, systems engineering, and project management. Technology and Culture, 38(4):891–919, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887214/pdf.

Johnson:1999:WSS

[Joh99a] Stephen B. Johnson. Way Station to Space: a History of the John C. Stennis Space Center (review). Technology and Culture, 40(4):922–924, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33443.

Johnston:1999:BDC

[Joh99b] Patricia A. Johnston. Burning with desire: The conception of photography. Technology and Culture, 40(3):702–705, July 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33376.

Josephson:1995:FRM

[Jos95a] Paul R. Josephson. Forging Revolution: Metalworkers, Managers, and the State in St. Petersburg, 1890–1914 by Heather Hogan (review). Technology and Culture, 36(3):701–702, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887816/pdf.

Josephson:1995:PCS

[Jos95b] Paul R. Josephson. “Projects of the Century” in Soviet history: Large-scale technologies from Lenin to Gorbachev. Technology and Culture, 36(3):519–559, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887789/pdf.

Josephson:1996:SSR

[Jos96] Paul Josephson. Stalinism, and Soviet Rail Transport, 1928–1941 by E. A. Rees (review). Technology and Culture, 37(3):641–642, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887456/pdf.

Josephson:1999:SAG

[Jos99] Paul R. Josephson. Stalin’s aviation gulag: a memoir of Andrei Tupolev and the Purge Era. Technology and Culture, 40(1):150–152, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33294.

Judd:1990:ABF

[Jud90] Richard W. Judd. The American Backwoods Frontier: an Ethnic and Ecological Interpretation by Terry G. Jordan and Matti Kaups (review). Technology and Culture, 31(4):873–875, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901663/pdf.

Judd:1999:NPB

[Jud99] Richard William Judd. New pioneers: The back-to-the-land movement and the search for a sustainable future. Technology and Culture, 40(2):423–425, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33340.

Kahlow:1996:BVR

[Kah96] Andreas Kahlow. Bâtir la ville: Révolutions industrielles dans les matériaux de construction, France–Grande-Bretagne, 1760–1840 by André Guillerme (review). Technology and Culture, 37(3):631–633, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887451/pdf.

Kaiserfeld:1996:CSW

[Kai96] Thomas Kaiserfeld. Computerizing the Swedish welfare state: The middle way of technological success and failure. Technology and Culture, 37(2):249–279, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887468/pdf.

Kaijser:1998:IAA

[Kai98] Arne Kaijser. Industrialisering av älvar: Studier kring svensk vattenkraftutbyggnad 1900–1918 (Swedish) [Industrialization of streams: studies on Swedish hydropower expansion 1900–1918] (review). Technology and Culture, 39(3):569–571, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33196.

Kalfatovic:1995:CAI

[Kal95] Martin R. Kalfatovic. Critical Approaches to Information Technology in Librarians hip: Foundations and Applications ed. by John Buschman (review). Technology and Culture, 36(1):199–200, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888026/pdf.

Kalia:1998:MIW

[Kal98] Ravi Kalia. The Making of the Indian Working Class: a Case of the Tata Iron and Steel Company, 1880–1946 (review). Technology and Culture, 39(4):782–784, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33227.

Kaplan:1992:CGS

[Kap92] Bonnie Kaplan. The Cybernetics Group by Steve J. Heims (review). Technology and Culture, 33(4):843–844, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888112/pdf.

Karnoe:1999:WEA

[Kar99] Peter Karnøe. Wind energy in America: a history. Technology and Culture, 40(2):417–419, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33353.

Kasson:1999:BRJ

[Kas99] John F. Kasson. J. M. W. Turner: Romantic Painter of the Industrial Revolution, and: John Ferguson Weir: The Labor of Art. Technology and Culture, 40(3):697–700, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33404.

Katz:1992:TOV

[Kat92] Barry Katz. Techniques of the Observer: On Vision and Modernity in the Nineteenth Century by Jonathan Crary (review). Technology and Culture, 33(1):206–207, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888380/pdf.

Katz:1994:MMI

[Kat94a] Barry M. Katz. Making the Modern: Industry, Art, and Design in America by Terry Smith (review). Technology and Culture, 35(3):642–643, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887870/pdf.

Katz:1994:FDC

[Kat94b] Barry M. Katz. The Fourth Discontinuity: The Co-Evolution of Humans and Machines by Bruce Mazlish (review). Technology and Culture, 35(4):888–889, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887739/pdf.

Katz:1997:CRE

[Kat97a] Barry M. Katz. Charles and Ray Eames: Designers of the Twentieth Century by Pat Kirkham (review). Technology and Culture, 38(4):954–955, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887217/pdf.

Katz:1997:RET

[Kat97b] Barry M. Katz. Review essay: Technology and design — a new agenda. Technology and Culture, 38(2):452–466, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887329/pdf.

Kays:1991:HHT

[Kay91] William M. Kays. History of Heat Transfer: Essays in Honor of the 50th Anniversary of the ASME Heat Transfer Division ed. by Edwin T. Layton, Jr., John H. Lienhard (review). Technology and Culture, 32(1):144–145, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888803/pdf. See comment [Lay92b].

Kay:1992:PSF

[Kay92] Lily E. Kay. Partners in Science: Foundations and Natural Scientists, 1900–1945 by Robert E. Kohler (review). Technology and Culture, 33(3):638–639, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888282/pdf.

Kay:1993:RTJ

[Kay93] Lily E. Kay. The Right Tools for the Job: At Work in Twentieth-Century Life Sciences ed. by Adele E. Clarke, Joan H. Fujimura (review). Technology and Culture, 34(4):964–966, October 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888069/pdf.

Kay:1994:ULH

[Kay94] Lily E. Kay. The Uses of Life: a History of Biotechnology by Robert Bud (review). Technology and Culture, 35(3):648–650, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887874/pdf.

Kay:1999:NSH

[Kay99] W. D. Kay. NASA and space history. Technology and Culture, 40 (1):120–127, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33316.

Keating:1998:AMC

[Kea98] Ann Durkin Keating. All the Modern Conveniences: American Household Plumbing, 1840–1890 by Maureen Ogle (review). Technology and Culture, 39(1):166–167, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887199/pdf.

Keith:1994:VES

[Kei94] Donald H. Keith. Vanguard of Empire: Ships of Exploration in the Age of Columbus by Roger C. Smith (review). Technology and Culture, 35(4):899–901, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887745/pdf.

Keller:1993:PRM

[Kel93a] Alex Keller. Paradise Restored: The Mechanical Arts from Antiquity through the Thirteenth Century by Elspeth Whitney (review). Technology and Culture, 34(1):136–138, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888463/pdf.

Keller:1993:RSR

[Kel93b] Alex Keller. Reappraisals of the Scientific Revolution ed. by David C. Lindberg, Robert S. Westman (review). Technology and Culture, 34(3):677–679, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888141/pdf.

Kellogg:1993:PSL

[Kel93c] Nelson R. Kellogg. Public Service Liberalism: Telecommunications and Transitions in Public Policy by Alan Stone (review). Technology and Culture, 34(1):182–184, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888489/pdf.

Kellogg:1993:WNP

[Kel93d] Nelson R. Kellogg. The Work of Nations: Preparing Ourselves for 21st-Century Capitalism by Robert B. Reich (review). Technology and Culture, 34(2):469–471, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888453/pdf.

Kelly:1995:FPA

[Kel95] Catherine E. Kelly. A Fictive People: Antebellum Economic Development and the American Reading Public by Ronald Zboray (review). Technology and Culture, 36(4):1049–1051, October 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887779/pdf.

Keller:1996:LTE

[Kel96] Alex Keller. Literature and Technology ed. by Mark L. Greenberg and Lance Schachterle (review). Technology and Culture, 37(3):624– 626, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887447/pdf.

Keller:1999:FRL

[Kel99] Alex Keller. Fortune is a River: Leonardo da Vinci and Niccolo Machiavelli’s Magnificent Dream to Change the Course of Florentine History (review). Technology and Culture, 40(4):878–879, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33444.

Kennett:1992:APC

[Ken92] Lee Kennett. Air Power and Colonial Control: The Royal Air Force, 1919–1939 by David E. Omissi (review). Technology and Culture, 33(3):632–633, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888278/pdf.

Kerr:1999:TWH

[Ker99] K. Austin Kerr. Technology & the West: a historical anthology from technology and culture. Technology and Culture, 40(2):388–390, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33333.

Kidwell:1990:ISI

[Kid90] Peggy Kidwell. The Information Society: Issues and Illusions by David Lyon, and: Thinking Machines: The Evolution of Artificial Intelligence by Vernon Pratt, and: The Conquest of the Microchip by Hans Queisser (review). Technology and Culture, 31(1):201–202, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888967/pdf.

Kidwell:1991:BGS

[Kid91] Peggy Aldrich Kidwell. Brass & Glass: Scientific Instrument Making Workshops in Scotland as Illustrated by Instruments from the Arthur Frank Collection at the Royal Museum of Scotland by T. N. Clarke, et al. (review). Technology and Culture, 32(1):171–172, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888818/pdf.

Kidwell:1992:JNO

[Kid92] Peggy Aldrich Kidwell. John von Neumann and the Origins of Modern Computing by William Aspray (review). Technology and Culture, 33(1):174–175, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888361/pdf.

Kidwell:1994:BCI

[Kid94] Peggy Aldrich Kidwell. Before the Computer: IBM, NCR, Burroughs, and Remington Rand and the Industry They Created, 1865– 1956 by James W. Cortada (review). Technology and Culture, 35(2):422–423, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887937/pdf.

Kidwell:1996:LLI

[Kid96] Peggy Aldrich Kidwell. Learning, Language and Invention: Essays Presented to Francis Maddison ed. by W. D. Hackmann, A. J. Turner (review). Technology and Culture, 37(2):344–346, April 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887474/pdf.

Kiehl:1990:AWV

[Kie90a] J. T. Kiehl. Appropriating the Weather: Vilhelm Bjerknes and the Construction of a Modern Meteorology by Robert Marc Friedman (review). Technology and Culture, 31(4):888–889, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901672/pdf.

Kielbowicz:1990:MND

[Kie90b] Richard B. Kielbowicz. Maps with the News: The Development of American Journalistic Cartography by Mark Monmonier (review). Technology and Culture, 31(3):553–554, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888869/pdf.

Kielbowicz:1992:ITC

[Kie92] Richard B. Kielbowicz. The Impact of the Telegraph on Contract Law by Lester G. Lindley (review). Technology and Culture, 33(3):628– 630, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888276/pdf.

Kielbowicz:1993:VPO

[Kie93] Richard B. Kielbowicz. The Victorian Post Office: The Growth of a Bureaucracy by C. R. Perry (review). Technology and Culture, 34(4):940–941, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888055/pdf.

Kingery:1993:PMI

[Kin93] W. David Kingery. Painterly maiolica of the Italian Renaissance. Technology and Culture, 34(1):28–48, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888456/pdf.

Kinney:1999:WHW

[Kin99] Thomas A. Kinney. When Horses Walked on Water: Horse-Powered Ferries in Nineteenth-Century America (review). Technology and Culture, 40(4):892–894, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33445.

Kipnis:1997:OAE

[Kip97] Nahum Kipnis. Optics in the Age of Euler: Conceptions of the Nature of Light, 1700–1795 by Casper Hakfoort (review). Technology and Culture, 38(4):968–969, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887225/pdf.

Kirsch:1995:HEA

[Kir95] David A. Kirsch. History of the Electric Automobile: Battery-Only Powered Cars by Ernest H. Wakefield (review). Technology and Culture, 36(3):710–712, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887821/pdf.

Kryzhanovsky:1992:OLR

[KK92] Leonid N. Kryzhanovsky and Yuri G. Krivchenko. Ornamental lightning rods. Technology and Culture, 33(4):858–860, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888119/pdf.

Klein:1991:ADP

[Kle91] Maury Klein. Alfred I. du Pont: The Man and His Family by Joseph Frazier Wall (review). Technology and Culture, 32(4):1119–1120, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888665/pdf.

Klein:1995:ARF

[Kle95] Maury Klein. The American Railroad Freight Car: From the Wood-Car Era to the Coming of Steel by John H. White, Jr. (review). Technology and Culture, 36(1):191–192, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888021/pdf.

Kleinman:1999:EAI

[Kle99] Daniel Lee Kleinman. Enabling American Innovation: Engineering and the National Science Foundation (review). Technology and Culture, 40(4):917–919, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33448.

Kline:1990:FBE

[Kli90] Ronald R. Kline. Ferranti and the British Electrical Industry, 1864–1930 by J. F. Wilson (review). Technology and Culture, 31 (1):169–170, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888949/pdf.

Kline:1991:PTE

[Kli91] Ronald R. Kline. The Papers of Thomas A. Edison. Vol. 1: The Making of an Inventor: February 1847–June 1873 ed. by Reese V. Jenkins, et al. (review). Technology and Culture, 32(1):155–156, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888809/pdf.

Kline:1997:CSE

[Kli97a] Ronald Kline. The Creation of Scientific Effects: Heinrich Hertz and Electric Waves by Jed Z. Buchwald (review). Technology and Culture, 38(4):971–973, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887227/pdf.

Kline:1997:ISS

[Kli97b] Ronald R. Kline. Ideology and social surveys: Reinterpreting the effects of “Laborsaving” technology on American farm women. Technology and Culture, 38(2):355–385, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887325/pdf.

Kranzberg:1992:CRR

[KM92] Melvin Kranzberg and Leo Marx. Comment and response on the review of In Context. Technology and Culture, 33(2):406–407, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

888335/pdf. See [Mar91].

Knight:1995:ETI

[Kni95] Thomas J. Knight. Education, Technology and Industrial Performance in Europe, 1850–1939 ed. by Robert Fox and Anna Guagnini (review). Technology and Culture, 36(2):403–404, April 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887890/pdf.

Koch:1993:ISN

[Koc93] Ellen B. Koch. In the image of science? Negotiating the development of diagnostic ultrasound in the cultures of surgery and radiology. Technology and Culture, 34(4):858–893, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888042/pdf.

Koch:1998:MNI

[Koc98] Ellen Koch. The Machine in the Nursery: Incubator Technology and the Origins of Newborn Intensive Care by Jeffrey P. Baker (review). Technology and Culture, 39(1):177–179, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887205/pdf.

Kohlstedt:1996:PPC

[Koh96] Sally Gregory Kohlstedt. Patronage, Practice, and the Culture of American Science: Alexander Dallas Bache and the U.S. Coast Survey by Hugh R. Slotten (review). Technology and Culture, 37 (1):183–185, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887551/pdf.

Konig:1996:SBI

[Kön96] Wolfgang König. Science-based industry or industry-based science? Electrical engineering in Germany before World War I. Technology and Culture, 37(1):70–101, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887537/pdf.

Konvitz:1997:MCC

[Kon97] Josef W. Konvitz. Maps and Civilization: Cartography in Culture and Society by Norman J. W. Thrower (review). Technology and Culture, 38(4):979–980, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887234/pdf.

Konvitz:1998:LEM

[Kon98] Josef W. Konvitz. Law and the Emergence of Modern Dublin by W. N. Osborough (review). Technology and Culture, 39(2):321–322, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887143/pdf.

Konvitz:1999:MP

[Kon99] Josef W. Konvitz. Maps and politics. Technology and Culture, 40(3):650–652, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33381.

Koppes:1991:WCL

[Kop91] Clayton R. Koppes. War and Cinema: The Logistics of Perception by Paul Virilio (review). Technology and Culture, 32(2 (part 1)):447– 448, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888776/pdf.

Kline:1996:UAT

[KP96] Ronald Kline and Trevor Pinch. Users as agents of technological change: The social construction of the automobile in the rural United States. Technology and Culture, 37(4):763–795, October 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887403/pdf.

Krammer:1990:DPH

[Kra90] Arnold Krammer. Divide and Prosper: The Heirs of I. G. Farben under Allied Authority 1945–1951 by Raymond G. Stokes (review). Technology and Culture, 31(3):537–539, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888860/pdf.

Krammer:1992:OPW

[Kra92] Arnold Krammer. Operation PLUTO: a wartime partnership for petroleum. Technology and Culture, 33(3):441–466, July 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888240/pdf.

Kraemer:1993:BSS

[Kra93] Sylvia K. Kraemer. Blueprint for Space: Science Fiction to Science Fact ed. by Frederick I. Ordway, Randy Liebermann (review). Technology and Culture, 34(2):461–462, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888448/pdf.

Kraft:1994:MHW

[Kra94a] James P. Kraft. Musicians in Hollywood: Work and technological change in entertainment industries, 1926–1940. Technology and Culture, 35(2):289–314, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887919/pdf.

Kragh:1994:KCI

[Kra94b] Helge Kragh. The krarup cable: Invention and early development. Technology and Culture, 35(1):129–157, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887957/pdf.

Kranakis:1994:EME

[Kra94c] Eda Kranakis. Engineering and the Mind’s Eye by Eugene S. Ferguson (review). Technology and Culture, 35(2):402–404, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887925/pdf.

Kraft:1995:SRC

[Kra95] James Kraft. Selling Radio: The Commercialization of American Broadcasting, 1920–1934. By Susan Smulyan (review). Technology and Culture, 36(3):715–717, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887824/pdf.

Kranzberg:1996:KG

[Kra96] Melvin Kranzberg. T. Keith Glennan (1905–1995). Technology and Culture, 37(3):659–662, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887466/pdf.

Kragh:1998:AIT

[Kra98] Helge Kragh. Arrhenius: From Ionic Theory to the Greenhouse Effect by Elisabeth Crawford (review). Technology and Culture, 39(2):327–328, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887146/pdf.

Krige:1992:LHL

[Kri92] John Krige. Lawrence and His Laboratory: a History of the Lawrence Berkeley Laboratory by J. L. Heilbron, Robert W. Seidel (review). Technology and Culture, 33(1):170–172, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888359/pdf.

Kriger:1997:MAS

[Kri97] Colleen E. Kriger. Metalworking in Africa South of the Sahara: an Annotated Bibliography ed. by Ibironke O. Lawal (review). Technology and Culture, 38(1):241–242, January 1997. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887379/pdf.

Kriger:1998:AMC

[Kri98] Colleen Kriger. African Material Culture ed. by Mary Jo Arnoldi, et al. (review). Technology and Culture, 39(1):142–144, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887185/pdf.

Kriger:1999:ITE

[Kri99] Colleen E. Kriger. Iron technology in East Africa: Symbolism, science, and archaeology. Technology and Culture, 40(2):401–402, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33355.

Konvitz:1990:TC

[KRT90] Josef W. Konvitz, Mark H. Rose, and Joel A. Tarr. Technology and the city. Technology and Culture, 31(2):284–294, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888881/pdf.

Kryzhanovsky:1990:LRC

[Kry90] Leonid N. Kryzhanovsky. The lightning rod in 18th-Century St. Petersburg: a note on the occasion of the bicentennial of the death of Benjamin Franklin. Technology and Culture, 31(4):813–817, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901651/pdf.

Kryzhanovsky:1992:BFS

[Kry92a] Leonid N. Kryzhanovsky. Benjamin Franklin’s Science by I. Bernard Cohen (review). Technology and Culture, 33(2):358–359, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888305/pdf.

Kryzhanovsky:1992:BSJ

[Kry92b] Leonid N. Kryzhanovsky. Boris Semyonovich Jacobi by Anatoly V. Yarotsky (review). Technology and Culture, 33(1):162, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888353/pdf.

Kryzhanovsky:1992:HTE

[Kry92c] Leonid N. Kryzhanovsky. History of Technology ed. by Graham Hollister-Short, Frank James (review). Technology and Culture, 33 (4):851–852, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888117/pdf.

Kubicek:1990:DSD

[Kub90] Robert V. Kubicek. The design of shallow-draft steamers for the British empire, 1868–1906. Technology and Culture, 31(3):427–450, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888830/pdf.

Kucher:1992:SVA

[Kuc92] Michael Kucher. Shipbuilders of the Venetian Arsenal: Workers and Workplace in the Preindustrial City by Robert C. Davis (review). Technology and Culture, 33(3):607–608, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888264/pdf.

Kucher:1994:TUD

[Kuc94] Michael Kucher. Turin, 1564–1680: Urban Design, Military Culture, and the Creation of the Absolutist Capital by Martha D. Poliak, and: Military Architecture, Cartography and the Representation of the Early Modern European City: a Checklist of Treatises on Fortification in the Newberry Library by Martha D. Poliak (review). Technology and Culture, 35(1):185–187, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887966/pdf.

Kuehl:1999:EAW

[Kue99] Daniel T. Kuehl. Ethics and airpower in World War II: The British bombing of German cities. Technology and Culture, 40(2):448–450, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33338.

Kunkle:1995:TSW

[Kun95] Gregory C. Kunkle. Technology in the seamless Web: “Success” and “Failure” in the history of the electron microscope. Technology and Culture, 36(1):80–103, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888005/pdf.

Kynell:1999:MST

[Kyn99] Teresa C. Kynell. From Millwrights to Shipwrights to the Twenty-First Century: Explorations in a History of Technical Communication in the United States (review). Technology and Culture, 40 (4):879–881, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33446.

Lowood:1990:CBH

[LAC^{+}90] Henry Lowood, Judith A. Adams, Stephen H. Cutcliffe, Jane Morley, Christine M. Roysdon, and Ian Winship. Current bibliography in the history of technology (1988). Technology and Culture, 31(3):561– 723, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888874/pdf.

Lowood:1991:CBH

[LAC^{+}91] Henry Lowood, Judith A. Adams, Stephen H. Cutcliffe, Jane Morley, Christine M. Roysdon, and Ian Winship. Current bibliography in the history of technology (1989). Technology and Culture, 32(3):659– 776, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888733/pdf.

Lowood:1992:CBH

[LAC^{+}92] Henry Lowood, Judith A. Adams, Stephen H. Cutcliffe, Jane Morley, Christine M. Roysdon, and Ian Winship. Current bibliography in the history of technology (1990). Technology and Culture, 33(Supplement):1–138, ???? 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888033/pdf.

Lowood:1993:CBH

[LAC^{+}93] Henry Lowood, Judith A. Adams, Stephen H. Cutcliffe, Katalin Harkányi, and Ian Winship. Current bibliography in the history of technology (1991). Technology and Culture, 34(Supplement):1– 135, ???? 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887999/pdf.

Lowood:1994:CBH

[LAC^{+}94] Henry Lowood, Judith A. Adams, Stephen H. Cutcliffe, Katalin Harkányi, and Ian Winship. Current bibliography in the history of technology (1992). Technology and Culture, 35(Supplement):1– 150, ???? 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887702/pdf.

LaFollette:1991:EML

[LaF91] Marcel C. LaFollette. The Exploratorium: The Museum as Laboratory by Hilde Hein (review). Technology and Culture, 32(3):653– 655, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888731/pdf.

Laird:1991:NIS

[Lai91] Pamela Walker Laird. New and Improved: The Story of Mass Marketing in America by Richard S. Tedlow (review). Technology and Culture, 32(3):622–624, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888713/pdf.

Laird:1992:LAG

[Lai92] Pamela Walker Laird. Lord of Attention: Gerald Stanley Lee and the Crowd Metaphor in Industrializing America by Gregory W. Bush (review). Technology and Culture, 33(3):600–601, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888259/pdf.

Laird:1993:BNP

[Lai93] Pamela Walker Laird. Buy Now, Pay Later: Advertising, Credit, and Consumer Durables in the 1920s by Martha L. Olney (review). Technology and Culture, 34(2):450–452, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888441/pdf.

Laird:1996:YRT

[Lai96a] Pamela Walker Laird. 100 Years on the Road: The Traveling Salesman in American Culture by Timothy B. Spears (review). Technology and Culture, 37(2):374–375, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887490/pdf.

Laird:1996:AOS

[Lai96b] Pamela Walker Laird. Auto Opium: a Social History of Automobile Design by David Gartman (review). Technology and Culture, 37(2):377–379, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887492/pdf.

Laird:1996:CSW

[Lai96c] Pamela Walker Laird. “The car without a single weakness”: Early automobile advertising. Technology and Culture, 37(4):796–812, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887404/pdf.

Laird:1997:SNA

[Lai97] Pamela Walker Laird. Spreading the News: The American Postal System from Franklin to Morse by Richard R. John (review). Technology and Culture, 38(2):512–514, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887342/pdf.

Laird:1998:PH

[Lai98] Pamela Walker Laird. The public’s historians. Technology and Culture, 39(3):474–482, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33215.

LaLande:1990:HTP

[LaL90] Jeff LaLande. Hard Times in Paradise: Coos Bay, Oregon, 1850– 1986 by William G. Robbins (review). Technology and Culture, 31 (4):875–877, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901664/pdf.

Lamoreaux:1993:BOM

[Lam93] Naomi R. Lamoreaux. Business Organization and the Myth of the Market Economy by William Lazonick (review). Technology and Culture, 34(4):951–953, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888061/pdf.

Langins:1990:AIA

[Lan90a] Janis Langins. Architectes et Ingénieurs au Siècle des Lumières by Antoine Picon (review). Technology and Culture, 31(4):870–871, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901661/pdf.

Lankford:1990:JLM

[Lan90b] John Lankford. James Lick’s Monument: The Saga of Captain Richard Floyd and the Building of the Lick Observatory by Helen Wright, and: A Man Who Loved the Stars: The Autobiography of John A. Brashear by John A. Brashear (review). Technology and Culture, 31(4):890–891, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901673/pdf.

Landes:1991:ETC

[Lan91a] David S. Landes. Empires of Time: Calendars, Clocks, and Cultures by Anthony F. Aveni (review). Technology and Culture, 32(3):596– 599, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888698/pdf.

Langins:1991:STP

[Lan91b] Janis Langins. Sciences et techniques en perspective. Vol. 15: La mutation de l’enseignement scientifique en France (1750–1810) et le rôle des écoles centrales: L’Exemple de Nantes by Pierre Lamandé (review). Technology and Culture, 32(1):176–177, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888822/pdf.

Landes:1993:PTG

[Lan93] David S. Landes. The Pulse of Time: Galileo Galilei, the Determination of Longitude, and the Pendulum Clock by Silvio A. Bedini (review). Technology and Culture, 34(3):670–672, July 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888137/pdf.

Landels:1994:RAW

[Lan94a] J. G. Landels. Roman Aqueducts and Water Supply by A. Trevor Hodge (review). Technology and Culture, 35(2):404–405, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887926.

Langins:1994:PPA

[Lan94b] Janis Langins. Public and Private Art Education in France, 1648– 1793 by Reed Benhamou (review). Technology and Culture, 35 (4):901–902, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887746/pdf.

Lankton:1994:MAH

[Lan94c] Larry Lankton. “Made in America” at the Henry Ford Museum. Technology and Culture, 35(2):389–395, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887923/pdf.

Langins:1995:LIM

[Lan95a] Janis Langins. L’Invention de l’ingénieur moderne: L’École des Ponts et Chaussées, 1747–1851 by Antoine Picon (review). Technology and Culture, 36(1):181–183, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888016/pdf.

Lankton:1995:TIA

[Lan95b] Larry Lankton. The Texture of Industry: an Archaeological View of the Industrialization of North America by Robert B. Gordon and Patrick M. Malone (review). Technology and Culture, 36(2):406– 408, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887892/pdf.

Landis:1998:MSO

[Lan98a] Leo Landis. To Make a Spotless Orange: Biological Control in California (review). Technology and Culture, 39(4):801–802, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33256.

Langins:1998:LAA

[Lan98b] Janis Langins. L’École d’application de l’artillerie et du génie de Metz, 1802–1870: Enseignement et recherches. Actes de la journée, d’étude du 2 novembre 1995 ed. by Bruno Belhoste and Antoine Picon (review). Technology and Culture, 39(1):148–149, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887188/pdf.

Landis:1999:MHS

[Lan99] Leo Landis. Mixed Harvest: The Second Great Transformation in the Rural North, 1870–1930 (review). Technology and Culture, 40 (4):889–890, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33447.

Larson:1991:SSR

[Lar91] James Larson. Science in Sweden: The Royal Swedish Academy of Sciences 1739–1989 ed. by Tore Frängsmyr (review). Technology and Culture, 32(1):172–173, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888819/pdf.

Larson:1998:NRE

[Lar98] John Lauritz Larson. The National Road ed. by Karl Raitz, and: A Guide to the National Road ed. by Karl Raitz (review). Technology and Culture, 39(2):307–309, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887136/pdf.

Laudan:1995:NAF

[Lau95] Rachel Laudan. Natural alliance or forced marriage? Changing relations between the histories of science and technology. Technology and Culture, 36(2S):S17–S30, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887707/pdf.

Laux:1998:LAD

[Lau98a] James M. Laux. L’Automobile autrefois: Une difficile adoption by Pierre-Lucien Pouzet (review). Technology and Culture, 39 (1):163–164, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887197/pdf.

Laux:1998:MBL

[Lau98b] James M. Laux. Marius Berliet by Louis Muron (review). Technology and Culture, 39(2):345–346, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887155/pdf.

Launius:1999:HAI

[Lau99] Roger D. Launius. A history of aerodynamics and its impact on flying machines. Technology and Culture, 40(3):688–690, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33374.

Lavenir:1993:RCX

[Lav93] Catherine Bertho Lavenir. Les révolutions de la communication XIX^{e}–XX^{e} siècle by Pascal Griset, and: Une histoire de la communication moderne: Espace public et vie privée by Patrice Flichy, and: La communication par la bande: Introduction aux sciences de l’information et de la communication by Daniel Bougnoux (review). Technology and Culture, 34(2):457–458, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888445/pdf.

Law:1991:TNH

[Law91] John Law. Theory and narrative in the history of technology: Response. Technology and Culture, 32(2 (part 1)):377–384, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888743/pdf.

Law:1992:OEC

[Law92] John Law. The Olympus 320 engine: a case study in design, development, and organizational control. Technology and Culture, 33(3):409–440, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888239/pdf.

Layton:1991:EHE

[Lay91] Edwin T. Layton, Jr. The Engineers: a History of the Engineering Profession in Britain, 1750–1914 by R. A. Buchanan (review). Technology and Culture, 32(1):143–144, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888802/pdf.

Layton:1992:SAS

[Lay92a] Edwin T. Layton, Jr. Science, American Style by Nathan Reingold (review). Technology and Culture, 33(3):636–637, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888281/pdf.

Layton:1992:CRH

[Lay92b] Edwin T. Layton, Jr. Comment on the review of History of Heat Transfer. Technology and Culture, 33(4):861–862, October 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888120/pdf. See

[Kay91].

Layton:1993:EIT

[Lay93] Edwin T. Layton. Essays in Industry and Technology in the Eighteenth Century: England and France by John R. Harris (review). Technology and Culture, 34(3):685–686, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888163/pdf.

Layton:1996:NSM

[Lay96] Edwin T. Layton. A Nation of Steel: The Making of Modern America by Thomas J. Misa (review). Technology and Culture, 37(3):635– 636, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887453/pdf.

Layne:1997:ECI

[Lay97] Linda Layne. The Empty Cradle: Infertility in America from Colonial Times to the Present by Margaret Marsh, Wanda Ronner (review). Technology and Culture, 38(4):976–979, October 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887232/pdf.

Lazzaro:1999:TAG

[Laz99] Claudia Lazzaro. Territorial ambitions and the Gardens of Versailles. Technology and Culture, 40(3):660–662, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33401.

Lowood:1995:CBH

[LCH^{+}95] Henry Lowood, Stephen H. Cutcliffe, Katalin Harkányi, Roger Launius, and Ian Winship. Current bibliography in the history of technology (1993). Technology and Culture, 36(Supplement):1–146, ???? 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887714/pdf.

Lowood:1996:AI

[LCH^{+}96a] Henry Lowood, Stephen H. Cutcliffe, Katalin Harkányi, Roger Launius, and Ian Winship. Author index (1994). Technology and Culture, 37(Supplement):184–221, ???? 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887398/pdf.

Lowood:1996:CBH

[LCH^{+}96b] Henry Lowood, Stephen H. Cutcliffe, Katalin Harkányi, Roger Launius, and Ian Winship. Current bibliography in the history of technology (1994). Technology and Culture, 37(Supplement):1–183, ???? 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887397/pdf.

Lowood:1996:SI

[LCH^{+}96c] Henry Lowood, Stephen H. Cutcliffe, Katalin Harkányi, Roger Launius, and Ian Winship. Subject index (1994). Technology and Culture, 37(Supplement):222–284, ???? 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887399/pdf.

Lowood:1997:CBH

[LCH^{+}97] Henry Lowood, Stephen H. Cutcliffe, Katalin Harkányi, Patrick Harshbarger, Roger Launius, and Ian Winship. Current bibliography in the history of technology (1995). Technology and Culture, 38(Supplement):1–167, ???? 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887209/pdf.

Leary:1992:TPE

[Lea92] Thomas E. Leary. Technology and Power in the Early American Cotton Industry: James Montgomery, the Second Edition of His “Cotton Manufacture” (1840), and the Justitia Controversy about Relative Power Costs by David J. Jeremy (review). Technology and Culture, 33(1):185–187, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888368/pdf.

Leary:1998:FQI

[Lea98] Thomas E. Leary. The Factory Question and Industrial England, 1830–1860 by Robert Gray (review). Technology and Culture, 39 (1):157–159, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887193/pdf.

Leary:1999:BCM

[Lea99] Thomas E. Leary. The Boott Cotton Mills Museum and the American Textile History Museum. Technology and Culture, 40(2):363–368, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33368.

LeCain:1994:MMC

[LeC94] Timothy LeCain. Mill and Mine: The CF&I in the Twentieth Century by H. Lee Scamehorn (review). Technology and Culture, 35(2):433–435, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887944/pdf.

LeCain:1999:WTL

[LeC99] Timothy J. LeCain. Western Technological Landscapes (review). Technology and Culture, 40(4):891–892, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33449.

Lederer:1991:CWN

[Led91] Norman Lederer. A Celebration of Work by Norman Best (review). Technology and Culture, 32(2 (part 1)):432–434, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888767/pdf.

Lederer:1993:MAS

[Led93] Susan E. Lederer. Medicine in America: a Short History by James H. Cassedy (review). Technology and Culture, 34(4):929–930, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888048/pdf.

Lee:1996:BBH

[Lee96] Jean B. Lee. Back of the Big House: The Architecture of Plantation Slavery by John Michael Vlach (review). Technology and Culture, 37(4):840–841, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887413/pdf.

Lee:1997:FMS

[Lee97] J. Lawrence Lee. Flight Management Systems: The Evolution of Avionics and Navigation Technology by Samuel B. Fishbein (review). Technology and Culture, 38(1):265–267, January 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887393/pdf.

Lenger:1991:SMN

[Len91] Friedrich Lenger. The Skilled Metalworkers of Nuremberg: Craft and Class in the Industrial Revolution by Michael J. Neufeld (review). Technology and Culture, 32(4):1108–1110, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888658/pdf.

Leonard:1992:DWS

[Leo92] Eileen Leonard. Divided We Stand: Re-defining Politics, Technology and Social Choice by Michiel Schwarz and Michael Thompson (review). Technology and Culture, 33(2):389–391, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888326/pdf.

Leonard:1996:LAM

[Leo96] William Leonard. Local Attachments: The Making of an American Urban Neighborhood, 1850 to 1920 by Alexander von Hoffman (review). Technology and Culture, 37(2):367–368, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887487/pdf.

Lerman:1997:PDP

[Ler97] Nina E. Lerman. “Preparing for the duties and practical business of life”: Technological knowledge and social structure in mid-19th-Century Philadelphia. Technology and Culture, 38(1):31–59, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887364/pdf.

Lerman:1998:AAC

[Ler98] Nina E. Lerman. American Artisans: Crafting Social Identity, 1750– 1850 ed. by Howard B. Rock, et al. (review). Technology and Culture, 39(2):303–305, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887134/pdf.

Leshikar:1990:SBS

[Les90a] Margaret E. Leshikar. The Skin Boats of Saint Lawrence Island, Alaska by Stephen R. Braund (review). Technology and Culture, 31 (4):877–878, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901665/pdf.

Leslie:1990:ESN

[Les90b] Stuart W. Leslie. Entrepreneurial Science: New Links between Corporations, Universities, and Government by Robert F. Johnston, Christopher G. Edwards (review). Technology and Culture, 31(1):147–149, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888935/pdf.

Leslie:1993:LAJ

[Les93] Stuart W. Leslie. The Laser in America, 1950–1970 by Joan Lisa Bromberg (review). Technology and Culture, 34(3):709–711, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888164/pdf.

Leslie:1994:OSV

[Les94] Stuart W. Leslie. Osiris, 2d ser., vol. 7: Science after ’40 ed. by Arnold Thackray (review). Technology and Culture, 35(4):881–883, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887735/pdf.

Lessoff:1999:GGW

[Les99] Alan Lessoff. Galveston and the Great West (review). Technology and Culture, 40(4):895–897, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33450.

Leunig:1998:STB

[Leu98] Tim Leunig. Science, Technology, and the British Industrial “Decline,” 1870–1970 (review). Technology and Culture, 39(3):571–572, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33186.

Levin:1991:ANF

[Lev91a] Miriam R. Levin. Art Nouveau in Fin-de–Siècle France: Politics, Psychology, and Style by Debora L. Silverman (review). Technology and Culture, 32(3):601–602, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888700/pdf.

Levin:1991:TSH

[Lev91b] Miriam R. Levin. La Technologie, science humaine: Recherches d’histoire et d’ethnologie des techniques by André-Georges Haudricourt (review). Technology and Culture, 32(1):175–176, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888821/pdf.

Levine:1991:BTC

[Lev91c] George Levine. Beyond the Two Cultures: Essays on Science, Technology, and Literature ed. by Joseph W. Slade, Judith Y. Lee (review). Technology and Culture, 32(4):1146–1148, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888680/pdf.

Levin:1996:WFS

[Lev96] Miriam R. Levin. What the French have to say about the history of technology. Technology and Culture, 37(1):158–168, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887542/pdf.

Levinson:1997:EEP

[Lev97a] Mark Levinson. Engineers and Engineering: Papers of the Rolt Fellows ed. by R. Angus Buchanan (review). Technology and Culture, 38 (4):1016–1017, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887261/pdf.

Levy:1997:APH

[Lev97b] Eugene Levy. The aesthetics of power: High-voltage transmission systems and the American landscape. Technology and Culture, 38(3):575–607, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887268/pdf.

Lewis:1990:PAR

[Lew90] W. David Lewis. Pacific Air Race by Robert H. Scheppler (review). Technology and Culture, 31(4):894–895, October 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/901676/pdf.

Lewis:1993:PIA

[Lew93] W. David Lewis. The Politics of International Aviation by Eugene Sochor (review). Technology and Culture, 34(4):960–962, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888067/pdf.

Lewis:1994:OW

[Lew94] M. J. T. Lewis. The origins of the wheelbarrow. Technology and Culture, 35(3):453–475, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887834/pdf.

Lewenstein:1996:TYS

[Lew96] Bruce Lewenstein. Twenty Years of Science in the Public Interest: a History of the Congressional Science and Engineering Fellowship Program by Jeffrey K. Stine (review). Technology and Culture, 37 (1):211–213, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887567/pdf.

Lewis:1997:ISC

[Lew97] W. David Lewis. Iron and Steel: Class, Race, and Community in Birmingham, Alabama, 1875–1920 by Henry M. McKiven Jr. (review). Technology and Culture, 38(4):984–986, October 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887239/pdf.

Lewis:1999:FEM

[Lew99] Susan I. Lewis. The female economy: The millinery and dressmaking trades, 1860–1930. Technology and Culture, 40(1):175–176, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33282.

Lantink-Ferguson:1993:GSU

[LF93] Annie T. Lantink-Ferguson. Die Geschichte der Stunde: Uhren und moderne Zeitordnung by Gerhard Dohrn-van Rossum (review). Technology and Culture, 34(3):673–674, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888138/pdf.

Levinson:1999:C

[LHTA99] Mark Levinson, Matthias Heymann, Andrew D. Todd, and Ken Alder. Communications. Technology and Culture, 40(4):928–931, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33432.

Liebenau:1990:ASP

[Lie90] Jonathan Liebenau. Academic Scientists and the Pharmaceutical Industry: Cooperative Research in Twentieth-Century America by John P. Swann (review). Technology and Culture, 31(1):144–145, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888933/pdf.

Liebenau:1991:MLP

[Lie91] Jonathan Liebenau. My Life with the Printed Circuit by Paul Eisler (review). Technology and Culture, 32(4):1124–1126, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888668/pdf.

Liebhold:1992:WPR

[Lie92] Peter Liebhold. “The Working People of Richmond: Life and Labor in an Industrial City, 1865–1920” at the Valentine Museum, Richmond, Virginia. Technology and Culture, 33(3):564–570, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888245/pdf.

Liebhold:1993:ADE

[Lie93] Peter Liebhold. America and the Daguerreotype ed. by John Wood, and: The Daguerreotype: Nineteenth-Century Technology and Modern Science by M. Susan Barger, William B. White (review). Technology and Culture, 34(2):425–427, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888426/pdf.

Liebhold:1998:GEB

[Lie98] Peter Liebhold. George Eastman: a Biography (review). Technology and Culture, 39(4):784–785, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33231.

Lightfoot:1997:HTO

[Lig97a] Dale R. Lightfoot. History and Technology of Olive Oil in the Holy Land by Rafael Frankel, et al. (review). Technology and Culture, 38(1):242–244, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887380/pdf.

Lightfoot:1997:QLH

[Lig97b] Dale R. Lightfoot. Qanats in the Levant: Hydraulic technology at the periphery of early empires. Technology and Culture, 38(2):432– 451, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887328/pdf.

Light:1999:WCW

[Lig99] Jennifer S. Light. When computers were women. Technology and Culture, 40(3):455–483, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33396.

Linn:1991:PAC

[Lin91] Karen Linn. The Piano in America, 1890–1940 by Craig H. Roell (review). Technology and Culture, 32(3):624–626, July 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888714/pdf.

Lindner:1993:GLU

[Lin93a] Stephan Lindner. “Der glorreiche Lebenslauf unserer Fabrik”: Zur Geschichte von Dorf und Baumwollspinnerei Kuchen ed. by Christel Köhle–Hezinger, Walter Ziegler (review). Technology and Culture, 34(2):429–430, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888428/pdf.

Lindner:1993:ETI

[Lin93b] Stephan Lindner. Die Entwicklung der Technologie und ihre Stellung in der Wissenschaft während der gesellschaftlichen Umwälzung an der Wende vom 18. zum 19. Jahrhundert by Tadaaki Kimoto (review). Technology and Culture, 34(1):144–145, January 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888467/pdf.

Lindgren:1994:HCV

[Lin94a] Uta Lindgren. The History of Cartography. Vol. 2, bk. 1: Cartography in the Traditional Islamic and South Asian Societies ed. by J. B. Harley and David Woodward (review). Technology and Culture, 35(3):607–608, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887849/pdf.

Lindner:1994:CSP

[Lin94b] Stephan Lindner. Le coton et ses patrons: France, 1760–1840 by Serge Chassagne (review). Technology and Culture, 35(2):408–410, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887929.

Lindee:1997:RLM

[Lin97a] M. Susan Lindee. Refiguring Life: Metaphors of Twentieth-Century Biology by Evelyn Fox Keller (review). Technology and Culture, 38(2):497–499, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887335/pdf.

Lindstrom:1997:TIG

[Lin97b] Diane Lindstrom. Technological Innovation and the Great Depression by Rick Szostak (review). Technology and Culture, 38(3):776– 777, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887299/pdf.

Lindee:1998:SPP

[Lin98] Susan Lindee. Science, Politics, and the Pharmaceutical Industry: Controversy and Bias in Drug Regulation by John Abraham (review). Technology and Culture, 39(2):336–338, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887151/pdf.

Lipartito:1997:CIH

[Lip97] Kenneth Lipartito. Coordination and Information: Historical Perspectives on the Organization of Enterprise ed. by Naomi R. Lamoreaux, Daniel M. G. Raff (review). Technology and Culture, 38 (4):998–1000, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887250/pdf.

Lipartito:1998:IRI

[Lip98] Kenneth Lipartito. Industrial Research and Innovation in Business (review). Technology and Culture, 39(4):777–778, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33240.

Liu:1997:MIG

[Liu97] Tessie P. Liu. Manufacturing Inequality: Gender Division in the French and British Metalworking Industries, 1914–1939 by Laura Lee Downs (review). Technology and Culture, 38(3):784–786, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887304/
pdf.

Livingston:1992:FTP

[Liv92] James Livingston. Figured Tapestry: Production, Markets, and Power in Philadelphia Textiles, 1885–1941 by Philip Scranton (review). Technology and Culture, 33(1):147–149, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888345/pdf.

Levin:1991:ICH

[LM91] Miriam R. Levin and Arwen Mohun. The International Committee for the History of Technology Meeting, Paris, July 1990. Technology and Culture, 32(3):594–595, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888697/pdf.

Lerman:1997:SWS

[LMO97a] Nina E. Lerman, Arwen Palmer Mohun, and Ruth Oldenziel. The shoulders we stand on and the view from here: Historiography and directions for research. Technology and Culture, 38(1):9–30, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887362/pdf.

Lerman:1997:VTG

[LMO97b] Nina E. Lerman, Arwen Palmer Mohun, and Ruth Oldenziel. Versatile tools: Gender analysis and the history of technology. Technology and Culture, 38(1):1–8, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887360/pdf.

Logsdon:1991:ARM

[Log91] John M. Logsdon. Apollo: The Race to the Moon by Charles Murray, Catherine Cox (review). Technology and Culture, 32(3):634– 636, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888721/pdf.

Logsdon:1992:SSD

[Log92] John M. Logsdon. The Space Station Decision: Incremental Politics and Technological Choice by Howard E. McCurdy (review). Technology and Culture, 33(2):385–386, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888323/pdf.

Logsdon:1995:MSR

[Log95] John M. Logsdon. The Missile and Space Race by Alan J. Levine (review). Technology and Culture, 36(4):1058–1060, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887784/pdf.

Long:1991:IAI

[Lon91a] Pamela O. Long. Invention, authorship, “Intellectual property,” and the origin of patents: Notes toward a conceptual history. Technology and Culture, 32(4):846–884, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888645/pdf.

Long:1991:OKI

[Lon91b] Pamela O. Long. The openness of knowledge: an ideal and its context in 16th-Century writings on mining and metallurgy. Technology and Culture, 32(2 (part 1)):318–355, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888740/pdf.

Long:1993:FGC

[Lon93a] Pamela O. Long. Francesco di Giorgio: Checklist and History of Manuscripts and Drawings in Autographs and Copies from ca. 1470 to 1687 and Renewed Copies (1764–1839) by Gustina Scaglia (review). Technology and Culture, 34(3):674–675, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888139/pdf.

Long:1993:GMA

[Lon93b] Pamela O. Long. Griechische Mechanik und antike Gesellschaft: Studien zur staatlichen Förderung einer technischen Wissenschaft by Astrid Schürmann (review). Technology and Culture, 34(4):935–936, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888052/pdf.

Long:1995:RRH

[Lon95a] Pamela O. Long. Renaissance and Revolution: Humanists, Scholars, Craftsmen and Natural Philosophers in Early Modern Europe ed. by J. V. Field, Frank A. J. L. James (review). Technology and Culture, 36(3):723–725, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887828/pdf.

Long:1995:ADA

[Lon95b] Pamela O. Long. The Architectural Drawings of Antonio da Sangallo the Younger and His Circle, vol. 1: Fortifications, Machines, and Festival Architecture ed. by Christoph L. Frommel and Nicholas Adams (review). Technology and Culture, 36(2):397–399, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887887/pdf.

Long:1996:SSN

[Lon96a] Pamela O. Long. Science and the Secrets of Nature: Books of Secrets in Medieval and Early Modern Culture by William Eamon (review). Technology and Culture, 37(1):170–172, January 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887544/pdf.

Long:1996:SRH

[Lon96b] Pamela O. Long. The Scientific Revolution: a Historiographical Inquiry by H. Floris Cohen (review). Technology and Culture, 37(3):619–621, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887444/pdf.

Long:1997:LVS

[Lon97] Pamela O. Long. Leonardo da Vinci’s Sforza Monument Horse: The Art and the Engineering ed. by Diane Cole Ahl (review). Technology and Culture, 38(4):1020–1022, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887263/pdf.

Lowood:1991:HES

[Low91] Henry Lowood. The History of Engineering Science: an Annotated Bibliography by David F. Channell (review). Technology and Culture, 32(1):150–151, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888806/pdf.

Lowood:1992:EHT

[Low92] Henry Lowood. An Encyclopaedia of the History of Technology ed. by Ian McNeil (review). Technology and Culture, 33(1):183–184, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888366/pdf.

Lowen:1993:MSH

[Low93a] Rebecca S. Lowen. Millikan’s School: a History of the California Institute of Technology by Judith R. Goodstein (review). Technology and Culture, 34(1):188–189, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888492/pdf.

Lowood:1993:STP

[Low93b] Henry Lowood. Scientific and Technical Periodicals of the Seventeenth and Eighteenth Centuries: a Guide by David A. Kronick (review). Technology and Culture, 34(3):676–677, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888140/pdf.

Lowen:1996:STE

[Low96] Rebecca Lowen. Science, Technology, and the Environment: Multidisciplinary Perspectives ed. by James Rodger Fleming and Henry A. Gemery (review). Technology and Culture, 37(3):656–658, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887465/pdf.

Lowen:1997:RLC

[Low97] Rebecca S. Lowen. The Road to Love Canal: Managing Industrial Waste before EPA by Craig E. Colten and Peter N. Skinner (review). Technology and Culture, 38(1):269–270, January 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887395/pdf.

Lowood:1999:CBHa

[Low99a] Henry Lowood. Current bibliography in the history of technology (1996–1997). Technology and Culture, 40(3):1–4, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/168998.

Lowood:1999:CBHb

[Low99b] Henry Lowood. Current bibliography in the history of technology (1996–1997). Technology and Culture, 40(Supplement):1–4, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/38024.

Lubar:1991:CTC

[Lub91a] Steven Lubar. Control through Communication: The Rise of System in American Management by JoAnne Yates (review). Technology and Culture, 32(3):619–621, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888711/pdf.

Lubar:1991:TAP

[Lub91b] Steven Lubar. The transformation of Antebellum patent law. Technology and Culture, 32(4):932–959, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888648/pdf.

Lubar:1995:RP

[Lub95] Steven Lubar. Representation and power. Technology and Culture, 36(2S):S54–S82, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887709/pdf.

Luegenbiehl:1990:PTI

[Lue90] Heinz C. Luegenbiehl. Philosophy and Technology II: Information Technology and Computers in Theory and Practice ed. by Carl Mitcham, Alois Huning (review). Technology and Culture, 31 (1):196–198, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888964/pdf.

Lukitsh:1999:EVT

[Luk99] Joanne Lukitsh. The engine of visualization: Thinking through photography. Technology and Culture, 40(3):701–702, July 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33399.

Lundgreen:1990:TSD

[Lun90a] Peter Lundgreen. Technik und Sicherheit in der deutschen Industriegesellschaft 1850 bis 1930 by Wolfhard Weber, and: Arbeitssicherheit: Historische Beispiele — aktuelle Analysen by Wolfhard Weber (review). Technology and Culture, 31(2):352–353, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888915/pdf.

Lunt:1990:LGM

[Lun90b] Richard D. Lunt. The Lobster Gangs of Maine by James M. Acheson (review). Technology and Culture, 31(1):156–157, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888940/pdf.

Lundgren:1992:SHR

[Lun92] Anders Lundgren. Solomon’s House Revisited: The Organization and Institutionalization of Science ed. by Tore Frängsmyr (review). Technology and Culture, 33(3):640–642, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888284/pdf.

Lupold:1991:PAS

[Lup91] John Lupold. Proceedings of the 150th Anniversary Symposium on Technology and Society — Southern Technology: Past, Present, and Future ed. by Howard L. Hartman (review). Technology and Culture, 32(1):148–150, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888805/pdf.

Luthy:1998:FLM

[Lüt98] Christoph Lüthy. The Fabric of Life: Microscopy in the Seventeenth Century by Marian Fournier (review). Technology and Culture, 39(2):325–327, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887145/pdf.

Levinson:1995:CRR

[LV95] Mark Levinson and Walter G. Vincenti. Comment and response on “The Retractable Airplane Landing Gear”. Technology and Culture, 36(2):451–454, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887916/pdf. See [Vin94].

Lynn:1998:FBR

[Lyn98] Michael Lynn. The French Book: Religion, Absolutism, and Readership, 1585–1715 by Henri-Jean Martin (review). Technology and Culture, 39(1):137–139, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887182/pdf.

Mitcham:1997:NMT

[MA97] Carl Mitcham and Andoni Alonso. Nuevas meditaciones sobre la técnica ed. by Fernando Broncano (review). Technology and Culture, 38(3):808–809, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887317/pdf.

MacLeod:1991:PPI

[Mac91] Christine MacLeod. The paradoxes of patenting: Invention and its diffusion in 18th- and 19th-Century Britain, France, and North America. Technology and Culture, 32(4):885–910, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888646/pdf.

Mack:1992:SCE

[Mac92] Pamela E. Mack. The Sputniks Crisis and Early United States Space Policy: a Critique of the Historiography of Space by Rip Bulkeley (review). Technology and Culture, 33(4):839–841, October 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888110/pdf.

MacKenzie:1993:LEB

[Mac93a] Donald MacKenzie. From the luminiferous ether to the Boeing 757: a history of the laser gyroscope. Technology and Culture, 34(3):475– 515, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888122/pdf.

MacLeod:1993:SIT

[Mac93b] Christine MacLeod. Shaping Invention: Thomas Blanchard’s Machinery and Patent Management in Nineteenth-Century America by Carolyn C. Cooper (review). Technology and Culture, 34(1):164–165, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888479/pdf.

Macagno:1996:GSW

[Mac96] Enzo Macagno. Out of Galileo: The Science of Waters, 1628–1718 by Cesare S. Maffioli (review). Technology and Culture, 37(2):348– 349, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887476/pdf.

Magney:1994:TTC

[Mag94] John Magney. The Timetables of Technology: a Chronology of the Most Important People and Events in the History of Technology by Bryan Bunch and Alexander Hellemans (review). Technology and Culture, 35(4):889–891, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887740/pdf.

Magney:1995:DEL

[Mag95] John Magney. Designing Engineers by Louis L. Bucciarelli (review). Technology and Culture, 36(4):1030–1031, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887768/pdf.

Mahaney:1990:REI

[Mah90] Timothy R. Mahaney. Rock Eras: Interpretations of Music and Society 1954–1984 by Jim Curtis (review). Technology and Culture, 31(2):343–344, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888910/pdf.

Mahoney:1997:EHD

[Mah97] Michael S. Mahoney. The Early History of Data Networks by Gerard J. Holzman, Björn Pehrson (review). Technology and Culture, 38 (4):1004–1006, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887254/pdf.

Maier:1996:WSI

[Mai96] Helmut Maier. Werner von Siemens: Inventor and International Entrepreneur by Wilfried Feldenkirchen (review). Technology and Culture, 37(2):381–383, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887494/pdf.

Maines:1998:LME

[Mai98] Rachel Maines. Landmarks in Mechanical Engineering (review). Technology and Culture, 39(4):770–772, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33226.

Malone:1991:HDA

[Mal91] Patrick Malone. A Historical Dictionary of American Industrial Language ed. by William H. Mulligan, Jr. (review). Technology and Culture, 32(1):153–155, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888808/pdf.

Malone:1998:BLW

[Mal98] Patrick M. Malone. The Baldwin Locomotive Works, 1831–1915: a Study in American Industrial Practice (review). Technology and Culture, 39(4):774–777, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33232.

Mantovani:1999:DSS

[Man99] Roberto Mantovani. Due secoli di strumenti geomagnetici in Italia, 1740–1971 (two centuries of geomagnetic instruments in Italy, 1740– 1971). Technology and Culture, 40(2):404–406, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33329.

Marcus:1990:WFC

[Mar90] Irwin Marcus. And the Wolf Finally Came: The Decline of the American Steel Industry by John P. Hoerr (review). Technology and Culture, 31(1):141–142, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888931/pdf.

Marx:1991:CHH

[Mar91] Leo Marx. In Context: History and the History of Technology — Essays in Honor of Melvin Kranzberg ed. by Stephen H. Cutcliffe, Robert C. Post (review). Technology and Culture, 32(2 (part 1)):394– 396, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

888745/pdf. See comment and response [KM92].

Martin:1993:MIH

[Mar93a] Steven C. Martin. Medical Innovations in Historical Perspective ed. by John V. Pickstone (review). Technology and Culture, 34 (4):923–924, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888044/pdf.

Martin:1993:CSC

[Mar93b] Steven C. Martin. Chiropractic and the social context of medical technology, 1895–1925. Technology and Culture, 34(4):808–834, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888040/pdf.

Markel:1996:SPM

[Mar96] Howard Markel. Science and the Practice of Medicine in the Nineteenth Century by W. F. Bynum (review). Technology and Culture, 37(2):351–353, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887478/pdf.

Martin:1998:WTW

[Mar98] Ph. D. Schönfeld Martin. Was there a Western inventor of porcelain? Technology and Culture, 39(4):716–727, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33270.

Matthewson:1990:NAA

[Mat90] Tim Matthewson. Native American Architecture by Peter Nabokov, Robert Easton (review). Technology and Culture, 31(2):305–306, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888888/pdf.

Matthies:1991:MW

[Mat91] Andrea L. Matthies. The medieval wheelbarrow. Technology and Culture, 32(2 (part 1)):356–364, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888741/pdf.

Matthewson:1992:SSC

[Mat92a] Timothy M. Matthewson. “Smoke Signals: Cigarettes, Advertising, and The American Way of Life” at the Valentine Museum, Richmond, Virginia. Technology and Culture, 33(3):560–563, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888244/pdf.

Matthies:1992:MTA

[Mat92b] Andrea L. Matthies. Medieval treadwheels: Artists’ views of building construction. Technology and Culture, 33(3):510–547, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888242/pdf.

Matsumoto:1999:CMS

[Mat99a] Eiju Matsumoto. The Chiba Museum of Science and Technology. Technology and Culture, 40(1):102–106, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33319.

Matsumoto:1999:RJI

[Mat99b] Miwao Matsumoto. Reconsidering Japanese industrialization: Marine turbine transfer at Mitsubishi. Technology and Culture, 40 (1):74–97, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33320.

Matthews:1999:TFN

[Mat99c] Glenna Matthews. Technology as freedom: The New Deal and the electrical modernization of the American home. Technology and Culture, 40(1):179–180, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33314.

Mauskopf:1990:CCJ

[Mau90] Seymour H. Mauskopf. Chemistry and cannon: J.-L. Proust and gunpowder analysis. Technology and Culture, 31(3):398–426, July 1990.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888829/pdf.

Maule:1991:BTH

[Mau91] R. William Maule. Behind, the Tube: a History of Broadcasting Technology and Business by Andrew F. Inglis (review). Technology and Culture, 32(3):621–622, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888712/pdf.

Mauskopf:1993:SPC

[Mau93] Seymour H. Mauskopf. Science as Public Culture: Chemistry and Enlightenment in Britain, 1760–1820 by Jan Golinski (review). Technology and Culture, 34(3):679–681, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888142/pdf.

Mayr:1991:ETG

[May91] Otto Mayr. The Evolution of Technology by George Basalla (review). Technology and Culture, 32(1):130–131, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888794/pdf.

Mayr:1992:VMS

[May92] Otto Mayr. The Vital Machine: a Study of Technology and Organic Life by David F. Channell (review). Technology and Culture, 33 (4):800–801, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888086/pdf.

Mayr:1994:PMI

[May94] Otto Mayr. The Power of the Machine: The Impact of Technology from 1700 to the Present Day by R. A. Buchanan (review). Technology and Culture, 35(1):190–192, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887969/pdf.

Mayr:1998:EGF

[May98] Otto Mayr. The Enola Gay fiasco: History, politics, and the museum. Technology and Culture, 39(3):462–473, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33216.

Mazuzan:1990:EFG

[Maz90a] George T. Mazuzan. Energy and the Federal Government: Fossil Fuel Policies, 1900–1946 by John G. Clark (review). Technology and Culture, 31(3):543–544, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888863/pdf.

Mazuzan:1990:NHA

[Maz90b] George T. Mazuzan. Nucleus: The History of Atomic Energy of Canada Limited by Robert Bothwell (review). Technology and Culture, 31(2):349–350, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888913/pdf.

Mazuzan:1993:ICI

[Maz93] George Mazuzan. Invisible Connections: Instruments, Institutions, and Science ed. by Robert Bud, Susan Cozzens (review). Technology and Culture, 34(4):962–963, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888068/pdf.

Mazuzan:1994:AEC

[Maz94] George T. Mazuzan. The Atomic Energy Commission under Nixon: Adjusting to Troubled Times by Glenn T. Seaborg, Benjamin S. Loeb (review). Technology and Culture, 35(1):217–219, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887984/pdf.

Mori:1992:CRE

[MB92] Sadahiko Mori and Harold Belofsky. Comment and response on “Engineering Drawing — a Universal Language in Two Dialects”. Technology and Culture, 33(4):853–857, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888118/pdf. See [Bel91].

McBride:1992:SDT

[McB92] William M. McBride. Strategic determinism in technology selection: The electric battleship and U.S. naval–industrial relations. Technology and Culture, 33(2):248–277, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888292/pdf.

McBride:1997:UDS

[McB97] William M. McBride. The unstable dynamics of a strategic technology: Disarmament, unemployment, and the interwar battleship. Technology and Culture, 38(2):386–423, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887326/pdf.

McCutcheon:1993:AES

[McC93] Robert McCutcheon. An American Engineer in Stalin’s Russia: The Memoirs of Zara Witkin, 1932–1934 by Zara Witkin (review). Technology and Culture, 34(2):435–436, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888432/pdf.

McCraw:1995:HE

[McC95] Thomas K. McCraw. The hubris of the engineers. Technology and Culture, 36(4):1007–1014, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887761/pdf.

McEvoy:1995:WEE

[McE95] Arthur F. McEvoy. Working environments: an ecological approach to industrial health and safety. Technology and Culture, 36(2S):S145– S173, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887712/pdf.

McGee:1995:MME

[McG95] David McGee. Making up mind: The early sociology of invention. Technology and Culture, 36(4):773–801, October 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887753/pdf.

McGaw:1997:IOG

[McG97a] Judith A. McGaw. Inventors and other great women: Toward a feminist history of technological luminaries. Technology and Culture, 38(1):214–231, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887373/pdf.

McGovern:1997:HIP

[McG97b] Arthur F. McGovern, S.J. History and the Idea of Progress ed. by Arthur M. Melzer, et al. (review). Technology and Culture, 38(3):806–808, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887316/pdf.

McGee:1999:CDN

[McG99] David McGee. From craftsmanship to draftsmanship: Naval architecture and the three traditions of early modern design. Technology and Culture, 40(2):209–236, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33370.

McMahon:1994:MHV

[McM94a] Michal McMahon. From Memex to Hypertext: Vannevar Bush and the Mind’s Machine ed. by James M. Nyce, Paul Kahn (review). Technology and Culture, 35(2):441–442, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887948/pdf.

McMath:1994:SFE

[McM94b] Robert C. McMath, Jr. MIT: Shaping the Future ed. by Kenneth R. Manning (review). Technology and Culture, 35(1):209–210, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887979/pdf.

McNeill:1990:RDO

[McN90] William H. McNeill. Retreat from Doomsday: The Obsolescence of Major War by John Mueller (review). Technology and Culture, 31 (1):191–192, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888961/pdf.

McNeill:1993:STW

[McN93] William H. McNeill. Surviving Trench Warfare: Technology and the Canadian Corps, 1914–1918 by Bill Rawling (review). Technology and Culture, 34(4):957–958, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888065/pdf.

McShane:1998:HTR

[McS98] Clay McShane. Horse Trails to Regional Rails: The Story of Public Transit in Greater Cleveland, and: Cleveland’s Transit Vehicles: Equipment and Technology (review). Technology and Culture, 39(3):582–583, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33205.

McWilliam:1995:EHJ

[McW95] Robert C. McWilliam. The Engineer in History by John Rae, Rudi Volti (review). Technology and Culture, 36(4):1028–1029, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887767/pdf.

Means:1998:WRU

[Mea98] Robert C. Means. Why Regulate Utilities?: The New Institutional Economics and the Chicago Gas Industry, 1849–1924, and: Gas Pipelines and the Emergence of America’s Regulatory State: a History of Panhandle Eastern Corporation, 1928–1993 (review). Technology and Culture, 39(4):797–799, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33261.

Meikle:1990:ADA

[Mei90] Jeffrey L. Meikle. The American Design Adventure 1940–1975 by Arthur J. Pulos (review). Technology and Culture, 31(2):312–314, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888892/pdf.

Meikle:1993:PRA

[Mei93] Jeffrey L. Meikle. The Portable Radio in American Life by Michael Brian Schiffer (review). Technology and Culture, 34(1):171–173, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888483/pdf.

Meikle:1995:ATS

[Mei95] Jeffrey L. Meikle. American Technological Sublime by David E. Nye (review). Technology and Culture, 36(4):1021–1022, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887763/pdf.

Meikle:1996:BBB

[Mei96] Jeffrey L. Meikle. Of Bicycles, Bakelites, and Bulbs: Toward a Theory of Sociotechnical Change by Wiebe E. Bijker (review). Technology and Culture, 37(4):853–855, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887420/pdf.

Meiksins:1998:EEI

[Mei98a] Peter Meiksins. Engineers as Executives: an International Perspective by William Aspray (review). Technology and Culture, 39(2):360– 362, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887163/pdf.

Meilinger:1998:FTB

[Mei98b] Phillip S. Meilinger. Fokker: a Transatlantic Biography (review). Technology and Culture, 39(4):791–792, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33238.

Meikle:1999:MA

[Mei99a] Jeffrey L. Meikle. The motel in America. Technology and Culture, 40(2):431–432, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33341.

Meilinger:1999:WFL

[Mei99b] Phillip S. Meilinger. Wingless Flight: The Lifting Body Story (review). Technology and Culture, 40(4):921–922, October 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33452.

Melosi:1996:CDA

[Mel96] Martin V. Melosi. Chasing Dirt: The American Pursuit of Cleanliness by Suellen Hoy (review). Technology and Culture, 37(3):639– 641, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887455/pdf.

Meldrum:1999:RMW

[Mel99] Marcia Meldrum. Rationalizing medical work: Decision-support techniques and medical practices. Technology and Culture, 40 (3):719–720, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33378.

Mernitz:1990:GRC

[Mer90] Kenneth S. Mernitz. Governmental research and the corporate state: The Rittman refining process. Technology and Culture, 31(1):83–113, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888923/pdf.

Merritt:1995:BAB

[Mer95] Raymond H. Merritt. Building Air Bases in the Negev: The U.S. Army Corps of Engineers in Israel, 1979–1982 by Frank N. Schubert (review). Technology and Culture, 36(2):432–433, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887907/pdf.

Merritt:1996:SSW

[Mer96] Raymond H. Merritt. Structures in the Stream: Water, Science, and the Rise of the U. S. Army Corps of Engineers by Todd Shallat (review). Technology and Culture, 37(4):842–843, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887414/pdf.

Mernitz:1997:MRU

[Mer97] Kenneth S. Mernitz. Die Mechanisierung des Ruhrbergbaus, 1890– 1930 by Uwe Burghardt (review). Technology and Culture, 38 (2):519–521, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887346/pdf.

Mernitz:1998:CGC

[Mer98] Kenneth S. Mernitz. Competition and Growth: a Contemporary History of the Continental AG (review). Technology and Culture, 39 (4):788–789, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33239.

Meulen:1998:AAI

[Meu98] Jacob A. Meulen. The American Aerospace Industry: From Workshop to Global Enterprise (review). Technology and Culture, 39 (3):589–591, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33179.

Michie:1994:LAI

[Mic94] Ranald Michie. London in the Age of Industrialisation: Entrepreneurs, Labour Force and Living Conditions, 1700–1850 by L. D. Schwarz (review). Technology and Culture, 35(2):410–411, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887930.

Millbrooke:1991:VTO

[Mil91a] Anne Millbrooke. Vertical Transportation in Old Back Bay: a Museum Case Study — the Acquisition of a Small Residential Hydraulic Elevator by Robert M. Vogel (review). Technology and Culture, 32(3):611–612, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888707/pdf.

Miller:1991:SFB

[Mil91b] Lillian B. Miller. Samuel F. B. Morse by Paul J. Staiti (review). Technology and Culture, 32(3):602–604, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888701/pdf.

Millard:1993:BNE

[Mil93] Andre Millard. Before the Nickelodeon: Edwin S. Porter and the Edison Manufacturing Company by Charles Musser (review). Technology and Culture, 34(1):166–167, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888480/pdf.

Miles:1994:CAC

[Mil94] Ellen G. Miles. The Counter-Arts Conspiracy: Art and Industry in the Age of Blake by Morris Eaves (review). Technology and Culture, 35(3):636–637, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887867/pdf.

Millbrooke:1995:PIV

[Mil95] Anne Millbrooke. A Practical Introduction to Videohistory: The Smithsonian Institution and Alfred P. Sloan Foundation Experiment ed. by Terri A. Schorzman (review). Technology and Culture, 36(2):443–444, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887914/pdf.

Millard:1997:PTE

[Mil97] Andre Millard. The Papers of Thomas A. Edison. Vol. 3, Menlo Park: The Early Years ed. by Robert A. Rosenberg, et al. (review). Technology and Culture, 38(3):763–765, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887293/pdf.

Mindell:1995:CBF

[Min95a] David A. Mindell. “The clangor of that blacksmith’s fray”: Technology, war, and experience aboard the USS Monitor. Technology and Culture, 36(2):242–270, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887880/pdf.

Miner:1995:RCC

[Min95b] Craig Miner. Railroad Crossing: Californians and the Railroad, 1850–1910 by William Deverell (review). Technology and Culture, 36(3):707–708, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887819/pdf.

Miner:1995:PTM

[Min95c] Craig Miner. The Passenger Train in the Motor Age: California’s Rail and Bus Industries, 1910–1941 by Gregory Lee Thompson (review). Technology and Culture, 36(2):418–420, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887899/pdf.

Mindell:1996:CLT

[Min96] David A. Mindell. Changing Large Technical Systems ed. by Jane Summerton (review). Technology and Culture, 37(3):649–651, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887461/pdf.

Mindell:1997:QGB

[Min97] David A. Mindell. A Quest for Glory: a Biography of Rear Admiral John A. Dahlgren by Robert J. Shneller Jr. (review). Technology and Culture, 38(3):761–763, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887292/pdf.

Mindell:1999:OBW

[Min99] David A. Mindell. The one best way: Frederick Winslow Taylor and the enigma of efficiency. Technology and Culture, 40(1):156–158, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33293.

Misa:1990:MCT

[Mis90] Thomas J. Misa. From Monopoly to Competition: The Transformations of Alcoa, 1888–1986 by George David Smith (review). Technology and Culture, 31(1):137–139, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888929/pdf.

Misa:1993:WMS

[Mis93] Thomas Misa. When the Machine Stopped: a Cautionary Tale from Industrial America by Max Holland (review). Technology and Culture, 34(1):173–175, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888484/pdf.

Misa:1997:SHT

[Mis97] Thomas J. Misa. A Social History of Truth by Steven Shapin (review). Technology and Culture, 38(1):238–240, January 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887378/pdf.

Mitcham:1991:NED

[Mit91] Carl Mitcham. Narrative Experiments: The Discursive Authority of Science and Technology by Gayle L. Ormiston, Raphael Sassower (review). Technology and Culture, 32(2 (part 1)):459–460, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888782/pdf.

Mitcham:1992:IRI

[Mit92] Carl Mitcham. Instrumental Realism: The Interface between Philosophy of Science and Philosophy of Technology by Don Ihde (review). Technology and Culture, 33(3):652–653, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888290/pdf.

Mitchell:1995:DGC

[Mit95] Allan Mitchell. “A dangerous game”: The crisis of locomotive manufacturing in France before 1914. Technology and Culture, 36(1):29– 45, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888003/pdf.

Mitcham:1996:TWP

[Mit96] Carl Mitcham. Technology in the Western Political Tradition ed. by Arthur M. Melzer, et al. (review). Technology and Culture, 37 (1):175–178, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887547/pdf.

Mitcham:1997:DTR

[Mit97] Carl Mitcham. Democracy and Technology by Richard E. Sclove (review). Technology and Culture, 38(2):531–533, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887352/pdf.

Miyoshi:1994:AEI

[Miy94] Karen Miyoshi. The American Environment: Interpretations of Past Geographies ed. by Lary M. Dilsaver, Craig E. Colten (review). Technology and Culture, 35(2):430–431, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887942/pdf.

Mike:1993:RMI

[MKR93] Valerie Miké, Alfred N. Krauss, and Gail S. Ross. Reflections on a medical innovation: Transcutaneous oxygen monitoring in neonatal intensive care. Technology and Culture, 34(4):894–922, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888043/pdf.

Morrison-Low:1998:PSN

[ML98] A. D. Morrison-Low. The Practice of Science in the Nineteenth Century: Teaching and Research Apparatus in the Teyler Museum (review). Technology and Culture, 39(3):563–564, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33207.

Meade:1992:ABN

[MM92] Edmund P. Meade and Jane Morley. Architects and Builders in North Carolina: a History of the Practice of Building by Catherine W. Bishir, et al. (review). Technology and Culture, 33(3):623– 624, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888273/pdf.

Morgan:1993:TBT

[MM93] B. Morgan and Pauletta Morse. The Telephone Book: Technology, Schizophrenia, Electric Speech by Avital Ronell (review). Technology and Culture, 34(2):458–459, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888446/pdf.

Marx:1996:EM

[MMH96] Leo Marx, Joel Moses, and Thomas P. Hughes. Elting morison, 1909–1995. Technology and Culture, 37(4):864–879, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887426/pdf.

Moen:1994:APP

[Moe94] Eli Moen. The American Pulp and Paper Industry, 1900–1940: Mill Survival, Firm Structure, and Industry Relocation by Nancy K. Ohanian (review). Technology and Culture, 35(4):870–871, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887728/pdf.

Mohun:1990:CDM

[Moh90] Arwen Mohun. A Culture for Democracy: Mass Communication and the Cultivated Mind in Britain between the Wars by D. L. LeMahieu (review). Technology and Culture, 31(1):170–172, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888950/pdf.

Mohun:1994:WTA

[Moh94] Arwen Mohun. Women and Technology: an Annotated Bibliography by Cynthia Gay Bindocci (review). Technology and Culture, 35 (4):891–892, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887741/pdf.

Mohun:1995:LSO

[Moh95] Arwen Mohun. The Limits of Safety: Organizations, Accidents, and Nuclear Weapons by Scott D. Sagan (review). Technology and Culture, 36(2):433–435, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887908/pdf.

Mohun:1997:GTR

[Moh97a] Arwen Palmer Mohun. The Gender-Technology Relation: Contemporary Theory and Research ed. by Keith Grint and Rosalind Gill (review). Technology and Culture, 38(1):232–233, January 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887374/pdf.

Mohun:1997:LCT

[Moh97b] Arwen Palmer Mohun. Laundrymen construct their world: Gender and the transformation of a domestic task to an industrial process. Technology and Culture, 38(1):97–120, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887369/pdf.

Molella:1991:MMB

[Mol91] Arthur P. Molella. The museum that might have been: The Smithsonian’s National Museum of Engineering and Industry. Technology and Culture, 32(2 (part 1)):237–263, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888737/pdf.

Molella:1995:ECE

[Mol95a] Arthur P. Molella. Edward Clinton Ezell (1939–1993). Technology and Culture, 36(2):445–450, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887915/pdf.

Molella:1995:TW

[Mol95b] Arthur P. Molella. Tilting at windmills. Technology and Culture, 36 (4):1000–1006, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887760/pdf.

Moon:1998:TSS

[Moo98] Suzanne M. Moon. Takeoff or self-sufficiency? Ideologies of development in Indonesia, 1957–1961. Technology and Culture, 39(2):187– 212, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887127/pdf.

Morley:1990:MBH

[Mor90] Jane Morley. The Most Beautiful House in the World by Witold Rybczynski (review). Technology and Culture, 31(2):310–311, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888891/
pdf.

Morison:1991:AGC

[Mor91a] Elting E. Morison. American Genesis: a Century of Invention and Technological Enthusiasm 1870–1970 by Thomas P. Hughes (review). Technology and Culture, 32(1):127–130, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888793/pdf.

Morley:1991:SCS

[Mor91b] Jane Morley. Second Century of the Skyscraper ed. by Lynn S. Beedle, and: Hongkong Bank: The Building of Norman Foster’s Masterpiece by Stephanie Williams (review). Technology and Culture, 32(3):615–617, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888709/pdf.

Morris:1991:STR

[Mor91c] Peter J. T. Morris. Science, Technology, and Reparations: Exploitation and Plunder in Postwar Germany by John Gimbel (review). Technology and Culture, 32(4):1122–1124, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888667/pdf.

Morrow:1991:SSH

[Mor91d] John H. Morrow, Jr. Strike from the Sky: The History of Battlefield Air Attack 1911–1945 by Richard P. Hallion (review). Technology and Culture, 32(2 (part 1)):445–447, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888775/pdf.

Morrow:1993:CSB

[Mor93] John H. Morrow, Jr. To Command the Sky: The Battle for Air Superiority over Germany, 1942–1944 by Stephen L. McFarland, Wesley P. Newton (review). Technology and Culture, 34(4):958–960, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888066/pdf.

Morton:1997:OVI

[Mor97] Alan Morton. Osiris. Vol. 9, Instruments ed. by Albert van Helden, Thomas Hankins (review). Technology and Culture, 38(3):753–755, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887287/pdf.

Morton:1998:ARF

[Mor98] David Morton. Armour Research Foundation and the wire recorder: How academic entrepreneurs fail. Technology and Culture, 39(2):213– 244, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887128/pdf.

Morris:1999:CCI

[Mor99] Peter Morris. Chemistry and the chemical industry in the 19th Century: The Henrys of Manchester and other studies. Technology and Culture, 40(2):410–411, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33336.

Mouat:1999:BBL

[Mou99] Jeremy Mouat. Beyond the Boundaries: Life and Landscape at the Lake Superior Copper Mines, 1840–1875 (review). Technology and Culture, 40(4):887–889, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33453.

Moy:1994:ASP

[Moy94] Timothy Moy. Arms and the State: Patterns of Military Production and Trade by Keith Krause (review). Technology and Culture, 35 (1):210–212, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887980/pdf.

Moy:1996:AJH

[Moy96] Timothy Moy. Andrew Jackson Higgins and the Boats that Won World War II by Jerry E. Strahan (review). Technology and Culture, 37(4):858–859, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887423/pdf.

Morris-Suzuki:1992:SOJ

[MS92] Tessa Morris-Suzuki. Sericulture and the origins of Japanese industrialization. Technology and Culture, 33(1):101–121, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888339/pdf.

Muendel:1990:MME

[Mue90] John Muendel. The Mills of Medieval England by Richard Holt (review). Technology and Culture, 31(3):508–512, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888844/pdf.

Mueller:1991:BSR

[Mue91a] Milton Mueller. The Bell System, and Regional Business: The Telephone in the South, 1877–1920 by Kenneth Lipartito (review). Technology and Culture, 32(4):1114–1115, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888662/pdf.

Muendel:1991:IFC

[Mue91b] John Muendel. The internal functions of a 14th-Century Florentine flour factory. Technology and Culture, 32(3):498–520, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888686/pdf.

Muendel:1995:PWH

[Mue95] John Muendel. Power from Wind: a History of Windmill Technology by Richard L. Hills (review). Technology and Culture, 36(3):692– 694, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887811/pdf.

Muendel:1999:TRU

[Mue99] John Muendel. Technology and resource use in Medieval Europe: Cathedrals, mills, and mines. Technology and Culture, 40(3):659– 660, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33406.

Multhauf:1990:DSI

[Mul90a] Robert P. Multhauf. Das deutsche Salinenwesen 1550–1650: Invention — Innovation — Diffusion by Peter Piasecki (review). Technology and Culture, 31(3):515–517, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888847/pdf.

Multhauf:1990:JGB

[Mul90b] Robert P. Multhauf. John G. Burke (1917–1989). Technology and Culture, 31(4):921–924, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901688/pdf.

Multhauf:1994:ISA

[Mul94] Robert P. Multhauf. Iron and Steel in Ancient China by Donald B. Wagner (review). Technology and Culture, 35(3):605–606, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887848/pdf.

Multhauf:1995:SNN

[Mul95] Robert P. Multhauf. Science Has No National Borders: Harry C. Kelly and the Reconstruction of Science and Technology in Postwar Japan by Hideo Yoshikawa, Joanne Kauffman (review). Technology and Culture, 36(4):1055–1057, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887782/pdf.

Mulder:1996:MML

[Mul96a] Karel Mulder. Man-Made Lowlands: History of Water Management and Land Reclamation in the Netherlands ed. by G. P. van de Ven (review). Technology and Culture, 37(1):210–211, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887566/pdf.

Multhauf:1996:SCC

[Mul96b] Robert P. Multhauf. Science and Civilisation in China. Vol. 5: Chemistry and Chemical Technology. Pt. 6: Military Technology: Missiles and Sieges by Joseph Needham, Robin D. S. Yates (review). Technology and Culture, 37(3):621–622, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887445/pdf.

Multhauf:1996:JN

[Mul96c] Robert P. Multhauf. Joseph Needham (1900–1995). Technology and Culture, 37(4):880–891, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887427/pdf.

Myllyntaus:1991:TET

[Myl91] Timo Myllyntaus. The transfer of electrical technology to Finland, 1870–1930. Technology and Culture, 32(2 (part 1)):293–317, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888739/pdf.

Myllyntaus:1997:LGE

[Myl97] Timo Myllyntaus. Den lysende gas: Etableringen af det danske gassystem 1800–1890 by Ole Hyldtoft (review). Technology and Culture, 38(4):1002–1004, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887253/pdf.

Nahum:1997:TST

[Nah97] Andrew Nahum. Two-stroke or turbine? The Aeronautical Research Committee and British aero engine development in World War II. Technology and Culture, 38(2):312–354, April 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887324/pdf.

Nash:1990:GCE

[Nas90] Gerald D. Nash. Growth in a Changing Environment: a History of Standard Oil Company (New Jersey) 1950–1972 and Exxon Corporation 1972–1975 by Bennett H. Wall (review). Technology and Culture, 31(1):142–143, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888932/pdf.

Nash:1991:ASR

[Nas91] Gerald D. Nash. The American Synthetic Rubber Research Program by Peter J. T. Morris (review). Technology and Culture, 32 (4):1120–1122, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888666/pdf.

Neal:1998:CLD

[Nea98] Valerie Neal. The Challenger Launch Decision: Risky Technology, Culture, and Deviance at NASA (review). Technology and Culture, 39(4):816–818, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33262.

Nebeker:1990:PSR

[Neb90] Frederik Nebeker. Portraits in Silicon by Robert Slater (review). Technology and Culture, 31(1):198–199, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888965/pdf.

Needell:1996:PSI

[Nee96] Allan A. Needell. Politics and Space: Image Making by NASA by Mark E. Brynes (review). Technology and Culture, 37(2):392–394, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887500/pdf.

Neis:1998:CVP

[Nei98] Hajo Neis. CyberCities: Visual Perception in the Age of Electronic Communication (review). Technology and Culture, 39(3):598–600, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33180.

Nelson:1990:WII

[Nel90] Daniel Nelson. Worktime and Industrialization: an International History ed. by Gary Cross (review). Technology and Culture, 31 (2):319–321, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888896/pdf.

Nelson:1992:TTS

[Nel92] Daniel Nelson. Taylorism Transformed: Scientific Management Theory since 1945 by Stephen P. Waring (review). Technology and Culture, 33(4):812–813, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888093/pdf.

Neufeld:1990:WCF

[Neu90] Michael J. Neufeld. Weimar culture and futuristic technology: The rocketry and spaceflight fad in Germany, 1923–1933. Technology and Culture, 31(4):725–752, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901648/pdf.

Neufeld:1992:EAE

[Neu92] John L. Neufeld. Electricity in the American Economy: Agent of Technological Progress by Sam H. Schurr, et al. (review). Technology and Culture, 33(2):395–397, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888329/pdf.

Neufeld:1993:NFG

[Neu93] Michael J. Neufeld. A Nation of Fliers: German Aviation and the Popular Imagination by Peter Fritzsche (review). Technology and Culture, 34(2):440–441, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888435/pdf.

Neufeld:1994:WBA

[Neu94] Michael J. Neufeld. Wernher von Braun — Aufbruch in den Weltraum: Die Biographie by Ernst Stuhlinger, Frederick I. Ordway (review). Technology and Culture, 35(1):212–214, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887981/pdf.

Neufeld:1996:IWR

[Neu96a] Michael J. Neufeld. Die Illusion der Wunderwaffen: Die Rolle der Düsenflugzeuge und Flugabwehrraketen in der Rüstungspolitik des Dritten Reiches by Ralf Schabel (review). Technology and Culture, 37(3):642–644, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887457/pdf.

Neufeld:1996:VMT

[Neu96b] Michael J. Neufeld. V-Missiles of the Third Reich: The V-1 and V-2 by Dieter Hölsken (review). Technology and Culture, 37(1):199–201, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887560/pdf.

Neushul:1998:MCS

[Neu98] Peter Neushul. Marie C. Stopes and the popularization of birth control technology. Technology and Culture, 39(2):245–272, April 1998.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887129/pdf.

Newell:1991:CPM

[New91] Alan S. Newell. The Conrey Placer Mining Company: a Pioneer Gold-dredging Enterprise in Montana, 1897–1922 by Clark C. Spence (review). Technology and Culture, 32(3):609–610, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888705/pdf.

Newton:1994:SIA

[New94] Wesley Phillips Newton. Storm over Iraq: Air Power and the Gulf War by Richard P. Hallion (review). Technology and Culture, 35(3):635–636, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887866/pdf.

Newell:1997:APB

[New97] Edmund Newell. Atmospheric pollution and the British copper industry, 1690–1920. Technology and Culture, 38(3):655–689, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887271/pdf.

Nickles:1999:TDU

[Nic99a] David Nickles. Telegraph diplomats: The United States’ relations with France in 1848 and 1870. Technology and Culture, 40(1):1–25, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33321.

Nickles:1999:AHC

[Nic99b] Shelley Nickles. America at home: a celebration of Twentieth-Century housewares. Technology and Culture, 40(1):180–182, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33296.

Noble:1997:TPK

[Nob97] David W. Noble. Technology and the Politics of Knowledge ed. by Andrew Feenberg, Alastair Hannay (review). Technology and Culture, 38(3):809–811, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887318/pdf.

Norberg:1990:HTC

[Nor90] Arthur L. Norberg. High-technology calculation in the early 20th century: Punched card machinery in business and government. Technology and Culture, 31(4):753–779, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901649/pdf.

Newell:1994:DTC

[NP94] Dianne Newell and Kathleen Paulsen. Dynamic traditions: “Cannery Days” at Vancouver’s Museum of Anthropology. Technology and Culture, 35(4):864–867, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887726/pdf.

Nuwer:1990:ARW

[Nuw90] Michael Nuwer. American Rubber Workers and Organized Labor, 1900–1941 by Daniel Nelson (review). Technology and Culture, 31 (1):164–166, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888946/pdf.

Nye:1990:PPT

[Nye90] David E. Nye. Pioneers of Photography: Their Achievements in Science and Technology ed. by Eugene Ostroff (review). Technology and Culture, 31(2):340–342, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888908/pdf.

Nye:1992:TEM

[Nye92] David E. Nye. Thomas A. Edison and the Modernization of America by Martin V. Melosi (review). Technology and Culture, 33(2):361– 362, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888307/pdf.

Nye:1993:EFT

[Nye93a] David E. Nye. Electrifying Finland: The Transfer of a New Technology into a Late Industrialising Economy by Timo Myllyntaus (review). Technology and Culture, 34(2):438–439, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888434/pdf.

Nye:1993:PPP

[Nye93b] Mary Jo Nye. Patriotism, Profit, and the Promotion of Science in the German Enlightenment: The Economic and Scientific Societies, 1760–1815 by Henry E. Lowood (review). Technology and Culture, 34(1):142–144, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888466/pdf.

Nye:1995:PEI

[Nye95] David E. Nye. The Photographic Experience, 1839–1914: Images and Attitudes by Heinz K. Henisch and Bridget A. Henisch (review). Technology and Culture, 36(2):415–416, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887897/pdf.

Nye:1997:TPP

[Nye97] David E. Nye. Technology, Pessimism, and Postmodernism ed. by Yaron Ezrahi, et al. (review). Technology and Culture, 38(2):533– 535, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887353/pdf.

Nye:1998:BUD

[Nye98a] David E. Nye. Building the Ultimate Dam: John S. Eastwood and the Control of Water in the West by Donald C. Jackson (review). Technology and Culture, 39(2):313–315, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887139/pdf.

Nye:1998:MSR

[Nye98b] David E. Nye. A moment of synthesis: Recent textbooks in the history of technology. Technology and Culture, 39(2):292–298, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887131/pdf.

OBannon:1991:DPS

[O’B91] Patrick W. O’Bannon. The Development of the Pacific Salmon-Canning Industry: a Grown Man’s Game ed. by Dianne Newell (review). Technology and Culture, 32(2 (part 1)):416–417, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888757/pdf.

Ochs:1992:DOE

[Och92a] Kathleen Ochs. Despite the Odds: Essays on Canadian Women and Science ed. by Marianne G. Ainley (review). Technology and Culture, 33(1):152–154, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888348/pdf.

Ochs:1992:RAM

[Och92b] Kathleen H. Ochs. The rise of American mining engineers: a case study of the Colorado School of Mines. Technology and Culture, 33(2):278–301, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888293/pdf.

ODonnell:1994:TIN

[O’D94a] Brian O’Donnell. Technology and Industry: a Nordic Heritage ed. by Jan Hult, Bengt Nyström (review). Technology and Culture, 35 (1):189–190, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887968/pdf.

ODonnell:1994:TRM

[O’D94b] Brian O’Donnell. Technology and Religion in Medieval Sweden by Anna Götlind (review). Technology and Culture, 35(3):610–611, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887851/pdf.

ODonnell:1996:CHR

[O’D96a] Brian O’Donnell, S.J. Cities of the Heartland: The Rise and Fall of the Industrial Midwest by Jon C. Teaford (review). Technology and Culture, 37(2):368–370, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887488/pdf.

ODonnell:1996:MHF

[O’D96b] Brian O’Donnell, S.J. Memory and hope: Four local museums in the mill towns of the Industrial Northeast. Technology and Culture, 37(4):817–827, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887406/pdf.

ODonnell:1997:HFM

[O’D97] Brian O’Donnell, S.J. Henry Ford: Mass Production, Modernism and Design by Ray Batchelor (review). Technology and Culture, 38 (1):258–259, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887389/pdf.

Ogan:1990:ECI

[Oga90] Christine Ogan. The Electronic Commonwealth: The Impact of New Media Technologies on Democratic Politics by Jeffrey B. Abramson, et al. (review). Technology and Culture, 31(1):195–196, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888963/pdf.

OHar:1996:FFT

[O’H96] George M. O’Har. Farmers and Fishermen: Two Centuries of Work in Essex County, Massachusetts, 1630–1850 by Daniel Vickers (review). Technology and Culture, 37(2):353–354, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887479/pdf.

OHar:1998:AHM

[O’H98] George M. O’Har. Aldous Huxley and the Mysticism of Science (review). Technology and Culture, 39(4):825–827, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33237.

Oldenziel:1992:MDW

[Old92] Ruth Oldenziel. Machinery of Dominance: Women, Men, and Technical Know-How by Cynthia Cockburn (review). Technology and Culture, 33(1):151–152, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888347/pdf.

Oldenziel:1997:BTT

[Old97] Ruth Oldenziel. Boys and their toys: The Fisher Body Craftsman’s Guild, 1930–1968, and the making of a male technical domain. Technology and Culture, 38(1):60–96, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887367/pdf.

Oldenziel:1998:CST

[Old98] Ruth Oldenziel. Changing Sex: Transsexualism, Technology, and the Idea of Gender by Bernice L. Hausman (review). Technology and Culture, 39(1):179–181, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887206.

Olien:1990:PHS

[Oli90] Diana Davids Olien. Pattillo Higgins and the Search for Texas Oil by Robert W. McDaniel and Henry C. Dethloff (review). Technology and Culture, 31(4):883–884, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901669/pdf.

Olien:1998:BPH

[Oli98] Diana Davids Olien. Bless the Pure and Humble: Texas Lawyers and Oil Regulation, 1919–1936 by Nicholas George Malavis (review). Technology and Culture, 39(2):309–311, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887137/pdf.

Olsson:1997:TDI

[Ols97] Lars O. Olsson. Technological Development and Industrial Exhibitions, 1850–1914: Sweden in an International Perspective by Göran Ahlström (review). Technology and Culture, 38(1):251–253, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887385/pdf.

Olsson:1998:SHT

[Ols98] Lars O. Olsson. “To See How Things Were Done in a Big Way”: Swedish Naval architects in the United States, 1890–1915. Technology and Culture, 39(3):434–456, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33218.

Olwell:1996:DMC

[Olw96] Russell Olwell. “Detroit — Motor City,” at the Detroit Historical Museum. Technology and Culture, 37(4):813–816, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887405/pdf.

Olwell:1998:NRW

[Olw98] Russell Olwell. Nuclear Rites: a Weapons Laboratory at the End of the Cold War by Hugh Gusterson (review). Technology and Culture, 39(1):133–135, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887180/pdf.

OMalley:1990:TFS

[O’M90] Michael O’Malley. Time: The Familiar Stranger by J. T. Fraser (review). Technology and Culture, 31(2):358–359, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888925/pdf.

OMalley:1993:TMH

[O’M93] Michael O’Malley. The Time Museum Historical Catalogue of American Pocket Watches by Donald R. Hoke (review). Technology and Culture, 34(3):713–714, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888160/pdf.

Omissi:1992:FAW

[Omi92] David Omissi. The First Air War, 1914–1918 by Lee Kennett (review). Technology and Culture, 33(2):379–380, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888319/pdf.

Opie:1990:DNI

[Opi90] John Opie. Disenchanted Night: The Industrialization of Light in the Nineteenth Century by Wolfgang Schivelbusch (review). Technology and Culture, 31(2):335–336, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888905/pdf.

Orsi:1997:EMM

[Ors97] Jared Orsi. Energy and the Making of Modern California by James C. Williams (review). Technology and Culture, 38(3):797– 799, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887311/pdf.

Overland:1997:DMA

[Øve97] Christian W. Øverland. Designing Modernity: The Arts of Reform and Persuasion, 1885–1945 ed. by Wendy Kaplan (review). Technology and Culture, 38(2):475–477, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887331/pdf.

Owens:1991:DNS

[Owe91a] Larry Owens. In Defense of Naval Supremacy: Finance, Technology and British Naval Policy, 1889–1914 by Jon Tetsuro Sumida (review). Technology and Culture, 32(2 (part 1)):443–445, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888774/pdf.

Owens:1991:PFA

[Owe91b] Larry Owens. Patents, the “Frontiers” of American invention, and the Monopoly Committee of 1939: Anatomy of a discourse. Technology and Culture, 32(4):1076–1093, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888653/pdf.

Ozdural:1998:MSA

[Ozd98] Alpay Ozdural. A mathematical sonata for architecture: Omar Khayyam and the Friday Mosque of Isfahan. Technology and Culture, 39(4):699–715, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33269.

Pacey:1995:WHP

[Pac95] Arnold Pacey. White Heat: People and Technology by Carroll Pursell (review). Technology and Culture, 36(3):680–681, July 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887806/pdf.

Palmieri:1991:TXQ

[Pal91] Richard P. Palmieri. Tibetan xylography and the question of movable type. Technology and Culture, 32(1):82–90, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888787/pdf.

Palladino:1993:BCS

[Pal93] Paolo Palladino. Between craft and science: Plant breeding, Mendelian genetics, and British universities, 1900–1920. Technology and Culture, 34(2):300–323, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888411/pdf.

Panzer:1995:CIP

[Pan95] Mary Panzer. Contesting Images: Photography and the World’s Columbian Exposition by Julie K. Brown (review). Technology and Culture, 36(4):1045–1047, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887777/pdf.

Pandora:1998:KSR

[Pan98a] Katherine Pandora. Knowledge and Society: Research in Science and Technology Studies. Vol. 10, Material Culture (review). Technology and Culture, 39(3):541–543, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33224.

Pandora:1998:DSB

[Pan98b] Katherine Pandora. The Disunity of Science: Boundaries, Contexts, and Power (review). Technology and Culture, 39(3):545–546, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33187.

Pannabecker:1998:RMA

[Pan98c] John R. Pannabecker. Representing mechanical arts in Diderot’s Encyclopédie. Technology and Culture, 39(1):33–73, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887170/pdf.

Paquette:1997:SSF

[Paq97] Robert L. Paquette. Sugar and Slavery, Family and Race: The Letters and Diary of Pierre Dessalles, Planter In Martinique, 1808– 1856 ed. by Elborg Forster, Robert Forster (review). Technology and Culture, 38(4):982–984, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887237/pdf.

Parayil:1992:GRI

[Par92] Govindan Parayil. The green revolution in India: a case study of technological change. Technology and Culture, 33(4):737–756, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888079/pdf.

Parkinson:1994:MMT

[Par94] E. Malcolm Parkinson. Medieval Military Technology by Kelly DeVries (review). Technology and Culture, 35(1):183–185, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887965/pdf.

Parr:1997:WMW

[Par97] Joy Parr. What makes washday less blue? Gender, nation, and technology choice in Postwar Canada. Technology and Culture, 38 (1):153–186, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887371/pdf.

Paskoff:1996:EAT

[Pas96] Paul Paskoff. Early American Technology: Making and Doing Things from the Colonial Era to 1850 ed. by Judith A. McGaw (review). Technology and Culture, 37(2):356–358, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887481/pdf.

Patton:1995:ZEN

[Pat95] Phil Patton. Zipper: an Exploration in Novelty by Robert Friedel (review). Technology and Culture, 36(4):1022–1024, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887764/pdf.

Payne:1995:ETG

[Pay95] Peter L. Payne. Enterprise and Technology: The German and British Steel Industries, 1865–1895 by Ulrich Wengenroth (review). Technology and Culture, 36(2):405–406, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887891/pdf.

Perry:1990:PTC

[Per90] Charles R. Perry. The Process of Technological Change: New Technology and Social Choice in the Workplace by Jon Clark, et al. (review). Technology and Culture, 31(1):152–154, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888938/pdf.

Pershey:1991:AMD

[Per91] Edward Jay Pershey. “American moviemakers: The dawn of sound” at the museum of modern art. Technology and Culture, 32(1): 106–113, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888791/pdf.

Pershey:1993:WWR

[Per93] Edward Jay Pershey. We Will Rise in Our Might: Workingwomen’s Voices from Nineteenth-Century New England by Mary H. Blewett (review). Technology and Culture, 34(2):427–429, April 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888427/pdf.

Perdue:1999:TGF

[Per99a] Peter C. Perdue. Technology and gender: Fabrics of power in Late Imperial China. Technology and Culture, 40(1):173–175, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33277.

Pershey:1999:KAW

[Per99b] Edward Jay Pershey. Karaoke around the world: Global technology, local singing. Technology and Culture, 40(3):711–713, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33400.

Peterson:1990:LEH

[Pet90] Joyce Shaw Peterson. On the Line: Essays in the History of Auto Work ed. by Nelson Lichtenstein, Stephen Meyer (review). Technology and Culture, 31(2):317–319, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888895/pdf.

Peters:1992:SLC

[Pet92a] Tom F. Peters. The Sky’s the Limit: a Century of Chicago Skyscrapers ed. by Pauline A. Saliga (review). Technology and Culture, 33 (1):195–196, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888374/pdf.

Peterson:1992:RDJ

[Pet92b] Charles E. Peterson. The Royal Dockyards 1690–1850 by Jonathan G. Coad (review). Technology and Culture, 33(1):156–158, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888350/pdf.

Peters:1996:IAC

[Pet96] Tom F. Peters. Inca Architecture and Construction at Ollantaytambo by Jean-Pierre Protzen (review). Technology and Culture, 37 (1):169–170, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887543/pdf.

Peterson:1997:CPA

[Pet97] Jane Peterson. Ceramic Production in the American Southwest ed. by Barbara J. Mills, Patricia L. Crown (review). Technology and Culture, 38(3):747–748, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887283/pdf.

Peters:1998:BWF

[Pet98] Tom F. Peters. Brucken am Weg: Fruhe Brucken aus Eisen und Beton in Deutschland und Frankreich (review). Technology and Culture, 39(4):768–769, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33260.

Pfaffenberger:1990:HFH

[Pfa90] Bryan Pfaffenberger. The harsh facts of hydraulics: Technology and society in Sri Lanka’s colonization schemes. Technology and Culture, 31(3):361–397, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888828/pdf.

Pfaffenberger:1991:CHC

[Pfa91] Bryan Pfaffenberger. Computers in the Human Context: Information Technology, Productivity, and People ed. by Tom Forester (review). Technology and Culture, 32(4):1132–1135, October 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888673/pdf.

Pfaffenberger:1992:PPT

[Pfa92] Bryan Pfaffenberger. Priests and Programmers: Technologies of Power in the Engineered Landscape of Bali by J. Stephen Lansing (review). Technology and Culture, 33(3):613–614, July 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888267/pdf.

Pfatteicher:1997:IES

[Pfa97] Sarah K. A. Pfatteicher. The Introspective Engineer by Samuel C. Florman, and: The Existential Pleasures of Engineering by Samuel C. Florman (review). Technology and Culture, 38(4):1022–1025, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887264/pdf.

Pfaffenberger:1998:NTW

[Pfa98] Bryan Pfaffenberger. New Technology and the Workers’ Response: Microelectronics, Labour, and Society ed. by Amiya Kumar Bagchi and Ranabir Samaddar (review). Technology and Culture, 39(2):358– 360, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887162/pdf.

Phillips:1991:EMT

[Phi91] Almarin Phillips. The Electronic Media and the Transformation of Law by M. Ethan Katsh (review). Technology and Culture, 32 (1):140–141, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888800/pdf.

Phillips:1993:BWV

[Phi93] Mona Phillips. Before Writing, vol. 1: From Counting to Cuneiform by Denise Schmandt-Besserat (review). Technology and Culture, 34(3):690–691, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888147/pdf.

Phillips:1999:LHW

[Phi99] Gervase Phillips. Longbow and hackbutt: Weapons technology and technology transfer in early modern England. Technology and Culture, 40(3):576–593, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33418.

Pinch:1990:MLK

[Pin90] Trevor Pinch. Mind at Large: Knowing in the Technological Age by Paul Levinson (review). Technology and Culture, 31(2):357–358, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888922/pdf.

Pinch:1991:LTE

[Pin91] Trevor Pinch. Lifeworld and Technology ed. by Timothy Casey, Lester Embree (review). Technology and Culture, 32(4):1140–1141, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888676/pdf.

Pinch:1992:AHS

[Pin92] Trevor Pinch. From Artifact to Habitat: Studies in the Critical Engagement of Technology ed. by Gayle L. Ormiston (review). Technology and Culture, 33(1):207–209, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888381/pdf.

Pisano:1996:PWA

[Pis96] Dominick A. Pisano. A Passion for Wings: Aviation and the Western Imagination, 1908–1918 by Robert Wohl (review). Technology and Culture, 37(1):194–196, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887557/pdf.

Plax:1996:ASE

[Pla96] Julie-Anne Plax. Artful Science: Enlightenment Entertainment and the Eclipse of Visual Education by Barbara M. Stafford (review). Technology and Culture, 37(1):172–174, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887545/pdf.

Pumarada-ONeill:1999:LQL

[PO99] Luis Pumarada-O’Neill. Una luz que llego para quedarse: Comienzos del alumbrado electrico y su introduccion en Cuba. Technology and Culture, 40(2):414–415, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33328.

Polella:1990:JUN

[Pol90] Arthur P. Polella. John U. Nef (1899–1988). Technology and Culture, 31(4):916–920, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901687/pdf.

Pollock:1991:BAF

[Pol91] Norman H. Pollock. British Agriculture in the First World War by P. E. Dewey (review). Technology and Culture, 32(1):162–163, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888813/pdf.

Poos:1999:EEM

[Poo99] Lawrence R. Poos. The emergence of estate maps: Christ Church, Oxford, 1600 to 1840. Technology and Culture, 40(1):132–134, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33280.

Porter:1991:EEB

[Por91] Theodore M. Porter. Energy and Empire: a Biographical Study of Lord Kelvin by Crosbie Smith, M. Norton Wise (review). Technology and Culture, 32(1):145–147, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888804/pdf.

Porter:1993:ONS

[Por93] Charlotte M. Porter. The Origins of Natural Science in America: The Essays of George Brown Goode ed. by Sally G. Kohlstedt (review). Technology and Culture, 34(4):954–955, October 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888063/pdf.

Porter:1994:SCQ

[Por94] Charlotte M. Porter. Seeds of Change: a Quincentennial Commemoration ed. by Herman J. Viola, Carolyn Margolis (review). Technology and Culture, 35(1):177–179, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887962/pdf.

Porter:1997:ERI

[Por97] Glenn Porter. Exhibit review: The inaugural exhibition at the wolfsonian. Technology and Culture, 38(2):467–474, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887330/pdf.

Porter:1998:MRQ

[Por98] Theodore M. Porter. The Measure of Reality: Quantification and Western Society, 1250–1600 (review). Technology and Culture, 39(3):555–557, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33185.

Post:1995:PB

[Pos95] Robert C. Post. Passing the baton. Technology and Culture, 36 (4):735–738, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887751/pdf.

Post:1998:FGW

[Pos98a] Robert C. Post. Fourteenth and G, Washington, D.C., Summer of 41. Technology and Culture, 39(4):728–732, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33267.

Post:1998:HRC

[Pos98b] Robert C. Post. Hot rods and customs: The men and machines of California’s car culture, at the Oakland Museum of California. Technology and Culture, 39(1):116–121, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887173/pdf.

Post:1999:NMT

[Pos99] Robert C. Post. No mere technicalities: How things work and why it matters. Technology and Culture, 40(3):607–622, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33419.

Potamianos:1999:CIM

[Pot99] George Potamianos. Cinema and the invention of modern life. Technology and Culture, 40(3):707–709, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33383.

Pretzer:1997:TCS

[Pre97] William S. Pretzer. Technology and Creativity by Subrata Dasgupta (review). Technology and Culture, 38(4):966–968, October 1997. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887224/pdf.

Pretzer:1998:RPH

[Pre98] William S. Pretzer. Reviewing public history in light of the enola gay. Technology and Culture, 39(3):457–461, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33220.

Price:1995:WWC

[Pri95] B. B. Price. A World without Women: The Christian Clerical Culture of Western Science by David F. Noble (review). Technology and Culture, 36(2):394–396, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887885/pdf.

Priest:1998:SPB

[Pri98] Tyler Priest. For Social Peace in Brazil: Industrialists and the Remaking of the Working Class in São Paulo, 1920–1964 by Barbara Weinstein (review). Technology and Culture, 39(2):356–357, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887161/pdf.

Pryke:1993:STM

[Pry93] K. G. Pryke. Science, Technology and Medicine in Canada’s Past: Selections from “Scientia Canadensis” ed. by Richard A. Jarrell, James P. Hull (review). Technology and Culture, 34(4):931–933, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888050/pdf.

Pulos:1992:OSC

[Pul92] Arthur J. Pulos. Olivetti: a Study of the Corporate Management of Design by Sibylle Kicherer (review). Technology and Culture, 33(2):400–402, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888332/pdf.

Pursell:1990:MWT

[Pur90] Carroll Pursell. Making Weapons, Talking Peace: a Physicist’s Odyssey from Hiroshima to Geneva by Herbert F. York (review). Technology and Culture, 31(4):901–903, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901681/pdf.

Pursell:1993:LEO

[Pur93a] Carroll Pursell. “Am I a lady or an engineer”? The origins of the Women’s Engineering Society in Britain, 1918–1940. Technology and Culture, 34(1):78–97, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888458/pdf.

Pursell:1993:RFA

[Pur93b] Carroll Pursell. The rise and fall of the appropriate technology movement in the United States, 1965–1985. Technology and Culture, 34(3):629–637, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888127/pdf.

Pursell:1997:BDH

[Pur97] Carroll Pursell. Biographical Dictionary of the History of Technology ed. by Lance Day, Ian McNeil (review). Technology and Culture, 38 (4):1014–1015, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887260/pdf.

Pursell:1999:BEH

[Pur99] Carroll W. Pursell. Beyond engineering: How society shapes technology. Technology and Culture, 40(2):395–397, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33351.

Pyenson:1990:CSA

[Pye90] Lewis Pyenson. The Commonwealth of Science: ANZAAS and the Scientific Enterprise in Australasia 1888–1988 ed. by Roy MacLeod (review). Technology and Culture, 31(2):355–356, April 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888919/pdf.

Pyenson:1992:ISN

[Pye92] Lewis Pyenson. International Science and National Scientific Identity: Australia between Britain and America ed. by Roderick W. Home, Sally G. Kohlstedt (review). Technology and Culture, 33(3): 640, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888283/pdf.

Quinn:1993:IEA

[Qui93] M.-L. Quinn. Industry and environment in the Appalachian copper basin, 1890–1930. Technology and Culture, 34(3):575–612, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888125/pdf.

Quivik:1994:BUO

[Qui94] Fredric L. Quivik. Bisbee: Urban Outpost on the Frontier ed. by Carlos A. Schwantes (review). Technology and Culture, 35(2):435– 436, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887945.

Quivik:1995:EHR

[Qui95] Fredric L. Quivik. Environmental History Review. Vol. 18, no. 1: Special Issue on Technology, Pollution, and the Environment ed. by Joel A. Tarr, Jeffrey K. Stine, and: Journal of Urban History. Vol. 20, no. 3: Special Issue Devoted to the Environment and the City ed. by Christine M. Rosen, Joel A. Tarr (review). Technology and Culture, 36(4):1038–1041, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887773/pdf.

Quirke:1998:WCN

[Qui98] Viviane M. Quirke. Wealth of Contrasts: Nyegaard & Co. — a Norwegian Pharmaceutical Company, 1874–1985 by Rolv Petter Amdam and Knut Sogner (review). Technology and Culture, 39(2):350– 352, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887158/pdf.

Raber:1996:MAM

[Rab96] Michael S. Raber. Making Arms in the Machine Age: Philadelphia’s Frankford Arsenal, 1816–1870 by James J. Farley (review). Technology and Culture, 37(1):185–186, January 1996. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887552/pdf.

Raber:1998:CCN

[Rab98] Michael Scott Raber. Culture Change and the New Technology: an Archaeology of the Early American Industrial Era (review). Technology and Culture, 39(3):577–579, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33203.

Randolph:1997:UES

[Ran97] Robert Randolph. Unglued Empire: The Soviet Experience with Communications Technologies by Gladys D. Ganley (review). Technology and Culture, 38(4):1006–1008, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887255/pdf.

Rasmussen:1990:TWF

[Ras90] Wayne D. Rasmussen. Toward a Well-Fed World by Don Paarlberg (review). Technology and Culture, 31(1):181–183, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888956/pdf.

Rasmussen:1992:BTB

[Ras92a] Wayne Rasmussen. Bull Threshers and Bindlestiffs: Harvesting and Threshing on the North American Plains by Thomas D. Isern (review). Technology and Culture, 33(1):187–188, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888369/pdf.

Rassweiler:1992:SSS

[Ras92b] Anne D. Rassweiler. Science and the Soviet Social Order ed. by Loren R. Graham (review). Technology and Culture, 33(3):642– 644, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888285/pdf.

Rasmussen:1999:WMW

[Ras99] Nicolas Rasmussen. What moves when technologies migrate? “Software” and hardware in the transfer of biological electron microscopy to postwar Australia. Technology and Culture, 40(1):47–73, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33322.

Rau:1998:CTS

[Rau98] Erik P. Rau. Complex Technical Systems (review). Technology and Culture, 39(3):538–540, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33195.

Reeds:1998:PKH

[Ree98] Karen Reeds. Picturing Knowledge: Historical and Philosophical Problems Concerning the Use of Art in Science, and: Fetal Positions: Individualism, Science, Visuality (review). Technology and Culture, 39(4):760–762, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33228.

Reeds:1999:JGM

[Ree99] Karen Reeds. Johann Gutenberg: The man and his invention. Technology and Culture, 40(2):403–404, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33342.

Reingold:1991:PFB

[Rei91] Nathan Reingold. The Pasteurization of France by Bruno Latour, and: Discovering: Inventing and Solving Problems at the Frontiers of Scientific Knowledge by Robert Scott Root-Bernstein (review). Technology and Culture, 32(1):177–179, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888823/pdf.

Reich:1993:FFA

[Rei93a] Leonard S. Reich. Farther and Faster: Aviation’s Adventuring Years, 1909–1939 by Terry Gwynn-Jones (review). Technology and Culture, 34(2):441–442, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888436/pdf.

Reingold:1993:DAP

[Rei93b] Nathan Reingold. The Development of American Pharmacology: John J. Abel and the Shaping of a Discipline by John Parascandola (review). Technology and Culture, 34(4):930–931, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888049/pdf.

Reich:1995:SSL

[Rei95a] Leonard S. Reich. From the spirit of st. Louis to the SST: Charles lindbergh, technology, and environment. Technology and Culture, 36(2):351–393, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887884/pdf.

Reingold:1995:ISV

[Rei95b] Nathan Reingold. Information and Secrecy: Vannevar Bush, Ultra, and the Other Memex by Colin Burke (review). Technology and Culture, 36(4):1051–1053, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887780/pdf.

Reid:1996:TPL

[Rei96] Donald Reid. Transforming Paris: The Life and Labors of Baron Haussmann by David P. Jordan, and: The Nation and Its City: Politics, “Corruption,” and Progress in Washington, D.C., 1861– 1902 by Alan Lessoff (review). Technology and Culture, 37(2):370– 373, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887489/pdf.

Reich:1999:SDP

[Rei99a] Leonard S. Reich. Ski-dogs, Pol-cats, and the mechanization of winter: The development of recreational snowmobiling in North America. Technology and Culture, 40(3):484–516, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33420.

Reid:1999:LHP

[Rei99b] Donald Reid. L’eau, l’hygiene publique et les infrastructures. Technology and Culture, 40(1):185–186, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33306.

Reuss:1991:CNJ

[Reu91] Martin Reuss. The Control of Nature by John McPhee (review). Technology and Culture, 32(2 (part 1)):405–407, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888751/pdf.

Reuss:1993:GWS

[Reu93a] Martin Reuss. Die Geschichte des Wasserbaus in der Schweiz by Niklaus Schnitter (review). Technology and Culture, 34(4):936–938, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888053/pdf.

Reutter:1993:BHP

[Reu93b] Mark Reutter. The Battle for Homestead, 1880–1892: Politics, Culture, and Steel by Paul Krause (review). Technology and Culture, 34(4):947–949, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888059/pdf.

Reuss:1995:WNE

[Reu95] Martin Reuss. The Wealth of Nature: Environmental History and the Ecological Imagination by Donald Worster (review). Technology and Culture, 36(3):685–687, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887808/pdf.

Reuss:1999:ASP

[Reu99a] Martin Reuss. The art of scientific precision: River research in the United States Army Corps of Engineers to 1945. Technology and Culture, 40(2):292–323, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33371.

Reuss:1999:BNC

[Reu99b] Martin Reuss. Building the Nineteenth Century. Technology and Culture, 40(1):145–146, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33304.

Reynolds:1990:MIS

[Rey90a] Terry S. Reynolds. Medieval Ironwork in Sweden by Lennart Karlsson (review). Technology and Culture, 31(3):514–515, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888846/pdf.

Reynolds:1990:TTS

[Rey90b] Terry S. Reynolds. Technology in Transition: The “Soo” Ship Canal, 1889–1985 by Robert W. Passfield (review). Technology and Culture, 31(4):880–881, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901667/pdf.

Reynolds:1991:PFE

[Rey91] Terry S. Reynolds. A Promising Field: Engineering at Alabama, 1837–1987 by Robert J. Norrell (review). Technology and Culture, 32(3):607–608, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888704/pdf.

Reynolds:1994:LUH

[Rey94] Terry S. Reynolds. Lehigh University: a History of Education in Engineering, Business, and the Human Condition by W. Ross Yates (review). Technology and Culture, 35(1):207–209, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887978/pdf.

Reynolds:1998:CCC

[Rey98] Terry R. Reynolds. Ceramic Commodities and Common Containers: Production and Distribution of White Mountain Red Ware in the Grasshopper Region, Arizona (review). Technology and Culture, 39(3):552–553, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33206.

Reynard:1999:UMM

[Rey99] Pierre-Claude Reynard. Unreliable mills: Maintenance practices in early modern papermaking. Technology and Culture, 40(2):237–262, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33372.

Rhees:1994:VTA

[Rhe94] David Rhees. “The Virtuoso Tribe of Arts and Sciences”: Studies in the Eighteenth-Century Work and Membership of the London Society of Arts ed. by D. G. C. Allan, John L. Abbott (review). Technology and Culture, 35(2):411–413, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887931/pdf.

Rhees:1998:APC

[Rhe98] David Rhees. American Plastic: a Cultural History by Jeffrey L. Meikle (review). Technology and Culture, 39(1):169–171, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887201/pdf.

Rhodes:1993:RPT

[Rho93] Benjamin D. Rhodes. Road to Power: The Trans-Siberian Railroad and the Colonization of Asian Russia, 1850–1917 by Steven G. Marks (review). Technology and Culture, 34(2):434, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888431/pdf.

Riney-Kehrberg:1997:WTR

[RK97] Pamela Riney-Kehrberg. Women, technology, and rural life: Some recent literature. Technology and Culture, 38(4):942–953, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887216/pdf.

Robb:1990:PAO

[Rob90a] Frances Robb. A Photo Album of Ohio’s Canal Era, 1825–1913 by Jack Gieck (review). Technology and Culture, 31(4):879, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901666/pdf.

Robinson:1990:IIR

[Rob90b] Eric Robinson. Inventing the Industrial Revolution: The English Patent System, 1660–1800 by Christine MacLeod (review). Technology and Culture, 31(3):519–520, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888849/pdf.

Robb:1992:CNC

[Rob92a] Frances C. Robb. Canals for a Nation: The Canal Era in the United States, 1790–1860 by Ronald E. Shaw (review). Technology and Culture, 33(2):372–373, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888314/pdf.

Robinson:1992:TGW

[Rob92b] Michael C. Robinson. Thirst for Growth: Water Agencies as Hidden Government in California by Robert Gottlieb, Margaret FitzSimmons (review). Technology and Culture, 33(4):829–830, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888103/pdf.

Robb:1993:BMR

[Rob93] Frances Robb. Beautiful Machine: Rivers and the Republican Plan, 1755–1825 by John Seelye (review). Technology and Culture, 34 (2):424–425, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888425/pdf.

Robinson:1994:TDS

[Rob94a] Eric Robinson. Technological Development and Science in the Industrial Age: New Perspectives on the Science-Technology Relationship ed. by Peter Kroes and Martijn Bakker (review). Technology and Culture, 35(3):616–618, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887855/pdf.

Robinson:1994:RDW

[Rob94b] Michael C. Robinson. To Reclaim a Divided, West: Water; Law, and Public Policy, 1848–1902 by Donald J. Pisani (review). Technology and Culture, 35(4):871–872, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887729/pdf.

Roberts:1995:EMD

[Rob95] Lissa Roberts. The Enlightenment of Matter: The Definition of Chemistry from Agricola to Lavoisier by Marco Beretta (review). Technology and Culture, 36(2):401–403, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887889/pdf.

Robb:1996:SVR

[Rob96a] Frances Robb. Southwestern Virginia’s Railroad: Modernization and the Sectional Crisis by Kenneth W. Noe (review). Technology and Culture, 37(2):360–361, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887483/pdf.

Robinson:1996:LST

[Rob96b] John Robinson. London’s smaller technology collections: a sampler. Technology and Culture, 37(1):151–157, January 1996. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887541/pdf.

Rocke:1994:RMO

[Roc94] Alan J. Rocke. The Rainbow Makers: The Origins of the Synthetic Dyestuffs Industry in Western Europe by Anthony S. Travis (review). Technology and Culture, 35(2):405–407, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887927.

Rogers:1995:MBP

[Rog95] Richard Rogers. Managing British public opinion of the Channel Tunnel. Technology and Culture, 36(3):636–640, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887794/pdf.

Roland:1990:AMH

[Rol90] Alex Roland. Of Arms and Men: a History of War, Weapons, and Aggression by Robert L. O’Connell (review). Technology and Culture, 31(4):900–901, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901680/pdf.

Roland:1992:STW

[Rol92] Alex Roland. Secrecy, technology, and war: Greek fire and the defense of Byzantium, 678–1204. Technology and Culture, 33(4): 655–679, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888076/pdf.

Roland:1993:TWH

[Rol93a] Alex Roland. Technology and war: The historiographical revolution of the 1980s. Technology and Culture, 34(1):117–134, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888461/pdf.

Rolando:1993:ITA

[Rol93b] Victor R. Rolando. Industry and Technology in Antebellum Tennessee: The Archaeology of Bluff Furnace by R. Bruce Council, et al. (review). Technology and Culture, 34(4):945–947, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888058/pdf.

Rolando:1994:LCD

[Rol94] Victor R. Rolando. Loading charcoal at danby station, Vermont. Technology and Culture, 35(2):383–388, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887922/pdf.

Roland:1995:FML

[Rol95a] Alex Roland. Feeding Mars: Logistics in Western Warfare from the Middle Ages to the Present ed. by John A. Lynn (review). Technology and Culture, 36(1):172–173, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888010/pdf.

Roland:1995:STW

[Rol95b] Alex Roland. Science, technology, and war. Technology and Culture, 36(2S):S83–S100, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887710/pdf.

Roland:1997:WHK

[Rol97] Alex Roland. What hath Kranzberg wrought? Or, does the history of technology matter? Technology and Culture, 38(3):697– 713, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887273/pdf.

Roland:1998:PWN

[Rol98a] Alex Roland. Postmodern War: The New Politics of Conflict (review). Technology and Culture, 39(4):811–813, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33243.

Roland:1998:VM

[Rol98b] Alex Roland. Voices in the museum. Technology and Culture, 39(3):483–488, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33221.

Rome:1997:SPS

[Rom97a] Adam Rome. A Sense of Place, a Sense of Time by John Brinckerhoff Jackson (review). Technology and Culture, 38(3):801–802, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887313/pdf.

Rome:1997:SUS

[Rom97b] Adam Rome. The Search for the Ultimate Sink: Urban Pollution in Historical Perspective by Joel A. Tarr (review). Technology and Culture, 38(3):781–783, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887302/pdf.

Ronnberg:1999:SBP

[Ron99] Erik A. R. Ronnberg. Ships’ bilge pumps: a history of their development, 1500–1900. Technology and Culture, 40(1):139–140, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33301.

Rorabaugh:1991:ATM

[Ror91] W. J. Rorabaugh. Apprenticeship: From Theory to Method and Back Again ed. by Michael W. Coy (review). Technology and Culture, 32(2 (part 1)):427–429, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888764/pdf.

Rose:1992:SRE

[Ros92] Mark H. Rose. Suburbia Re-examined ed. by Barbara M. Kelly (review). Technology and Culture, 33(1):190–192, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888371/pdf.

Rosen:1993:CUC

[Ros93] Christine M. Rosen. Constructing Urban Culture: American Cities and City Planning, 1800–1920 by Stanley K. Schultz (review). Technology and Culture, 34(2):448–450, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888440/pdf.

Rosenberg:1994:APA

[Ros94a] Robert Rosenberg. Alexanderson: Pioneer in American Electrical Engineering by James E. Brittain (review). Technology and Culture, 35(4):875–876, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887731/pdf.

Rosenberg:1994:ISP

[Ros94b] Robert Rosenberg. Innovation as a Social Process: Elihu Thomson and the Rise of General Electric, 1870–1900 by W. Bernard Carlson (review). Technology and Culture, 35(2):418–420, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887935/pdf.

Rosenband:1998:MPI

[Ros98] Leonard N. Rosenband. Machines à papier: innovation et transformations de l’industrie papetière en France, 1798–1860 (review). Technology and Culture, 39(3):566–567, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33177.

Rothenberg:1992:ESL

[Rot92] Marc Rothenberg. Eye on the Sky: Lick Observatory’s First Century by Donald E. Osterbrock, et al., and: Astronomical Centers of the World by Kevin Krisciunas (review). Technology and Culture, 33(1):169, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888358/pdf.

Rothenberg:1993:BHS

[Rot93] Marc Rothenberg. Beyond History of Science: Essays in Honor of Robert E. Schofield ed. by Elizabeth Garber (review). Technology and Culture, 34(1):192–193, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888495/pdf.

Rothenberg:1994:PPR

[Rot94a] Marc Rothenberg. Pauper and Prince: Ritchey, Hale, and Big American Telescopes by Donald E. Osterbrock (review). Technology and Culture, 35(4):876–877, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887732/pdf.

Rothman:1994:SLT

[Rot94b] Hal.K. Rothman. The sky’s the limit? Technology and the American West. Technology and Culture, 35(1):168–173, January 1994.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887959/pdf.

Rothschild:1995:GTM

[Rot95] Joan Rothschild. Gender and Technology in the Making by Cynthia Cockburn, Susan Ormrod, and: Bringing Technology Home: Gender and Technology in a Changing Europe ed. by Cynthia Cockburn, Ruža Fürst–Dilić (review). Technology and Culture, 36(3):681– 685, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887807/pdf.

Rothenberg:1997:PCM

[Rot97] David Rothenberg. The Prometheus Complex: Man’s Obsession with Superior Technology by Hans van de Braak (review). Technology and Culture, 38(1):235–237, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887376/pdf.

Roth:1999:MHE

[Rot99a] Matthew W. Roth. Mulholland highway and the engineering culture of Los angeles in the 1920s. Technology and Culture, 40(3):545–575, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33421.

Rothenberg:1999:SCD

[Rot99b] Marc Rothenberg. Scientific communities in the developing world. Technology and Culture, 40(1):171–172, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33281.

Rotter:1999:PUD

[Rot99c] Andrew Jon Rotter. Prompt and Utter Destruction: Truman and the Use of Atomic Bombs Against Japan (review). Technology and Culture, 40(4):912–913, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33454.

Roush:1996:LWF

[Rou96] Wade Roush. 1939: The Lost World of the Fair by David Gelernter (review). Technology and Culture, 37(4):832–834, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887409/pdf.

Rozen:1990:NTP

[Roz90] Frieda Shoenberg Rozen. New Technology and the Process of Labour Regulation by Eric Batstone, et al. (review). Technology and Culture, 31(1):151–152, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888937/pdf.

Rozwadowski:1996:OPN

[Roz96] Helen M. Rozwadowski. “Ocean Planet” at the National Museum of Natural History, Smithsonian Institution. Technology and Culture, 37(2):330–339, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887471/pdf.

Rozwadowski:1998:IMW

[Roz98] Helen M. Rozwadowski. Iron Men, Wooden Women: Gender and Seafaring in the Atlantic World, 1700–1920 ed. by Margaret S. Creighton and Lisa Norling (review). Technology and Culture, 39 (2):305–307, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887135/pdf.

Rudig:1992:WMM

[Rüd92] Wolfgang Rüdig. Wissenschaft für Macht und Markt: Kernforschung und Mikroelektronik in der Bundesrepublik Deutschland by Michael Eckert, Maria Osietzki (review). Technology and Culture, 33(1): 179–181, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888364/pdf.

Ruegamer:1990:BDL

[Rue90] Lana Ruegamer. Building Domestic Liberty: Charlotte Perkins Gilman’s Architectural Feminism by Polly Wynn Allen (review). Technology and Culture, 31(2):308–309, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888890/pdf.

Rugh:1996:MCB

[Rug96] Susan Sessions Rugh. Making the Corn Belt: a Geographical History of Middle-Western Agriculture by John C. Hudson (review). Technology and Culture, 37(2):362–363, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887484/pdf.

Rumm:1992:KWH

[Rum92] John C. Rumm. Keeping Watch: a History of American Time by Michael O’Malley (review). Technology and Culture, 33(2):366–367, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888310/pdf.

Rumm:1997:WFC

[Rum97] John C. Rumm. World of Fairs: The Century-of-Progress Expositions by Robert W. Rydell (review). Technology and Culture, 38(3):773–775, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887298/pdf.

Russell:1999:SCD

[Rus99] Edmund Russell. The strange career of DDT: Experts, Federal capacity, and environmentalism in World War II. Technology and Culture, 40(4):770–796, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33455.

Rycroft:1992:PEW

[Ryc92] Robert W. Rycroft. The Political Economy of World Energy: a Twentieth-Century Perspective by John G. Clark (review). Technology and Culture, 33(4):806–808, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888090/pdf.

Rycroft:1993:LTI

[Ryc93] Robert W. Rycroft. Losing Time: The Industrial Policy Debate by Otis L. Graham, Jr. (review). Technology and Culture, 34(2):467– 469, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888452/pdf.

Ryden:1998:STC

[Ryd98] Goran Ryden. Skill and technical change in the Swedish iron industry, 1750–1860. Technology and Culture, 39(3):383–407, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33222.

Salay:1992:ACW

[Sal92a] David Salay. Another Civil War: Labor, Capital, and the State in the Anthracite Regions of Pennsylvania, 1840–68 by Grace Palladino (review). Technology and Culture, 33(2):375–376, April 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888316/pdf.

Salsbury:1992:ANG

[Sal92b] Stephen Salsbury. American Narrow Gauge Railroads by George W. Hilton (review). Technology and Culture, 33(4):817–820, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888096/pdf.

Salsbury:1995:UMH

[Sal95] Stephen Salsbury. The Unforgiving Minute: How Australia Learned to Tell the Time by Graeme Davison (review). Technology and Culture, 36(3):712–714, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887822/pdf.

Sampson:1999:PEP

[Sam99] Gary D. Sampson. Picturing Empire: Photography and the Visualization of the British Empire and: Camera Indica: The Social Life of Indian Photographs (review). Technology and Culture, 40 (4):902–904, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33456.

Sandberg:1990:WCM

[San90] Lars G. Sandberg. What’s a Coal Miner to Do? The Mechanization of Coal Mining by Keith Dix (review). Technology and Culture, 31 (1):163–164, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888945/pdf.

Sanabria:1992:LAD

[San92] Sergio L. Sanabria. The Logic of Architecture: Design, Computation, and Cognition by William J. Mitchell (review). Technology and Culture, 33(3):625–627, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888274/pdf.

Sanabria:1994:ATS

[San94] Sergio L. Sanabria. Architectural Technology up to the Scientific Revolution: The Art and Structure of Large-Scale Buildings ed. by Robert Mark (review). Technology and Culture, 35(4):896–899, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887744/pdf.

Sanabria:1997:STC

[San97] Sergio L. Sanabria. Studies in Tectonic Culture: The Poetics of Construction in Nineteenth and Twentieth Century Architecture by Kenneth Frampton (review). Technology and Culture, 38(4):992–995, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887245/pdf.

Schmandt-Besserat:1995:HPW

[SB95] Denise Schmandt-Besserat. The History and Power of Writing by Henri-Jean Martin (review). Technology and Culture, 36(4): 1047–1049, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887778/pdf.

Scarpino:1999:OMR

[Sca99] Philip V. Scarpino. The organic machine: The remaking of the Columbia River. Technology and Culture, 40(2):419–420, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33363.

Schubert:1991:WEL

[Sch91] Frank N. Schubert. Where Eagles Land: Planning and Development of U.S. Army Airfields, 1910–1941 by Jerold E. Brown (review). Technology and Culture, 32(3):631–632, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888719/pdf.

Schatzberg:1992:BSE

[Sch92a] Eric Schatzberg. Barons of the Sky: From Early Flight to Strategic Warfare — the Story of the American Aerospace Industry by Wayne Biddle (review). Technology and Culture, 33(4):838–839, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

888109/pdf. See comment [BS94].

Schatzberg:1992:TSS

[Sch92b] Eric Schatzberg. Technologies and Society: The Shaping of People and Things by Ron Westrum (review). Technology and Culture, 33(2):391–393, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888327/pdf.

Scharff:1993:PDE

[Sch93a] Virginia Scharff. Possible Dreams: Enthusiasm for Technology in America ed. by John L. Wright (review). Technology and Culture, 34(2):417–419, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888421/pdf.

Schiffer:1993:CIP

[Sch93b] Michael Brian Schiffer. Cultural imperatives and product development: The case of the shirt-pocket radio. Technology and Culture, 34(1):98–113, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888459/pdf.

Schatzberg:1994:ITC

[Sch94a] Eric Schatzberg. Ideology and technical choice: The decline of the wooden airplane in the United States, 1920–1945. Technology and Culture, 35(1):34–69, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887954/pdf.

Schroeder:1994:EUT

[Sch94b] Fred E. H. Schroeder. The Evolution of Useful Things by Henry Petroski (review). Technology and Culture, 35(1):192–194, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887970/pdf.

Schubert:1994:NBG

[Sch94c] Frank N. Schubert. The Neck of the Bottle: George W. Goethals and the Reorganization of the U.S. Army Supply System, 1917–1918 by Phyllis A. Zimmerman (review). Technology and Culture, 35 (1):205–206, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887977/pdf.

Schwoch:1994:TMM

[Sch94d] James Schwoch. Telecommunications, Mass Media, and Democracy: The Battle for the Control of U.S. Broadcasting, 1928–1935 by Robert W. McChesney (review). Technology and Culture, 35 (4):873–874, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887730/pdf.

Schwoch:1995:NWT

[Sch95] James Schwoch. News over the Wires: The Telegraph and the Flow of Public Information in America, 1844–1897 by Menahem Blondheim (review). Technology and Culture, 36(3):714–715, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887823/pdf.

Schatzberg:1999:IDH

[Sch99a] Eric Schatzberg. Invention by design: How engineers get from thought to thing. Technology and Culture, 40(1):152–154, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33305.

Schramm:1999:SDM

[Sch99b] Jeff Schramm. From Steam to Diesel: Managerial Customs and Organizational Capabilities in the Twentieth-Century American Locomotive Industry (review). Technology and Culture, 40(4):898–900, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33457.

Scranton:1990:TTT

[Scr90] Philip Scranton. Textiles in Transition: Technology, Wages, and Industry Relocation in the U.S. Textile Industry, 1880–1930 by Nancy Frances Kane (review). Technology and Culture, 31(1):157–158, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888941/pdf.

Scranton:1991:SSD

[Scr91a] Philip Scranton. Scale and Scope: The Dynamics of Industrial Capitalism by Alfred D. Chandler, Jr. (review). Technology and Culture, 32(4):1102–1104, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888655/pdf.

Scranton:1991:WCA

[Scr91b] Philip Scranton. Working-Class Americanism: The Politics of Labor in a Textile City, 1914–1960 by Gary Gerstle (review). Technology and Culture, 32(2 (part 1)):431–432, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888766/pdf.

Scranton:1991:TNH

[Scr91c] Philip Scranton. Theory and narrative in the history of technology: Comment. Technology and Culture, 32(2 (part 1)):385–393, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888744/pdf.

Scranton:1994:MDP

[Scr94] Philip Scranton. Manufacturing diversity: Production systems, markets, and an American consumer society, 1870–1930. Technology and Culture, 35(3):476–505, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887835/pdf.

Scranton:1995:DIH

[Scr95] Philip Scranton. Determinism and indeterminacy in the history of technology. Technology and Culture, 36(2S):S31–S53, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887708/pdf.

Scranton:1996:BIM

[Scr96] Philip Scranton. Bond of Iron: Master and Slave at Buffalo Forge by Charles B. Dew (review). Technology and Culture, 37(3):633– 635, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887452/pdf.

Scranton:1998:SCR

[Scr98a] Philip Scranton. Socializing Capital: The Rise of the Large Industrial Corporation in America (review). Technology and Culture, 39 (4):772–774, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33254.

Scranton:1998:TCM

[Scr98b] Philip Scranton. Technological Change: Methods and Themes in the History of Technology ed. by Robert Fox (review). Technology and Culture, 39(1):124–126, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887175/pdf.

Scranton:2000:MTC

[Scr00] Philip Scranton. Missing the target? A comment on Edward Constant’s “Reliable Knowledge and Unreliable Stuff”. Technology and Culture, 41(4):752–764, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33638. See [Con99].

Smith-Divita:1998:WMU

[SD98] Stephanie Smith-Divita. Writing in a Milieu of Utility: The Move to Technical Communication in American Engineering Programs, 1850–1950 (review). Technology and Culture, 39(4):769–770, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33249.

Seely:1990:DAS

[See90] Bruce E. Seely. The Decline of American Steel: How Management, Labor, and Government Went Wrong by Paul A. Tiffany (review). Technology and Culture, 31(1):139–140, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888930/pdf.

Seely:1993:RES

[See93] Bruce Seely. Research, engineering, and science in American engineering colleges: 1900–1960. Technology and Culture, 34(2):344– 386, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888413/pdf.

Seely:1995:YSS

[See95a] Bruce Seely. 200 Years of Soot and Sweat: The History and Archeology of Vermont’s Iron, Charcoal, and Lime Industries by Victor R. Rolando (review). Technology and Culture, 36(2):408–410, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887893/pdf.

Seely:1995:CAT

[See95b] Bruce E. Seely. Critical Assembly: a Technical History of Los Alamos during the Oppenheimer Years, 1943–1945 by Lillian Hoddeson, et al. (review). Technology and Culture, 36(4):1053–1055, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887781/pdf.

Seely:1995:SHT

[See95c] Bruce E. Seely. SHOT, the history of technology, and engineering education. Technology and Culture, 36(4):739–772, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887752/pdf.

Seely:1998:AIR

[See98a] Bruce E. Seely. American Iron, 1607–1900 by Robert Gordon (review). Technology and Culture, 39(2):301–303, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887133/pdf.

Seely:1998:ERU

[See98b] Bruce E. Seely. Education in a Research University ed. by Kenneth J. Arrow, et al. (review). Technology and Culture, 39(1):151–153, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887190/pdf.

Seely:1999:CCW

[See99a] Bruce Edsall Seely. Creating the Cold War university: The transformation of Stanford. Technology and Culture, 40(1):199–201, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33295.

Seely:1999:RFA

[See99b] Bruce Edsall Seely. The rational factory: Architecture, technology, and work in America’s Age of Mass Production. Technology and Culture, 40(3):677–679, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33380.

Segal:1990:MMP

[Seg90] Howard P. Segal. “Made in Maine”: a professor’s perspective. Technology and Culture, 31(3):463–468, July 1990. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888833/pdf.

Segal:1994:TPP

[Seg94] Howard P. Segal. TVA’s Public Planning: The Vision, the Reality by Walter L. Creese (review). Technology and Culture, 35(2):428– 430, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887941.

Seidel:1993:OSD

[Sei93] Robert W. Seidel. The Origins of SDI, 1944–1983 by Donald R. Baucom (review). Technology and Culture, 34(4):973–975, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888074/pdf.

Seidel:1996:ECA

[Sei96a] Robert W. Seidel. Elements of Controversy: The Atomic Energy Commission and Radiation Safety in Nuclear Weapons Testing, 1947–1974 by Barton C. Hacker (review). Technology and Culture, 37(2):389–390, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887498/pdf.

Seidel:1996:CUS

[Sei96b] Robert W. Seidel. The Computer in the United States: From Laboratory to Market, 1930 to 1960 by James W. Cortada (review). Technology and Culture, 37(1):205–207, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887563/pdf.

Seidel:1998:CHI

[Sei98] Robert W. Seidel. Computer: a History of the Information Machine (review). Technology and Culture, 39(4):820–821, October 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33233.

Seligmann:1990:FCC

[Sel90] Linda J. Seligmann. Food, Conquest, and Colonization in Sixteenth-Century Spanish America by John C. Super (review). Technology and Culture, 31(1):172–174, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888951/pdf.

Selk:1991:PTF

[Sel91] Eugene E. Selk. Philosophy of Technology by Frederick Ferré (review). Technology and Culture, 32(2 (part 1)):457–459, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888781/pdf.

Sella:1992:BLC

[Sel92] Domenico Sella. Di bianco lin Candida prole: La manifattura della carta in Età, Moderna e il caso toscano by Renzo Sabbatini (review). Technology and Culture, 33(2):353–354, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888302/pdf.

Selman:1997:EAS

[Sel97] Jan Robert Selman. East Asian Science: Tradition and Beyond ed. by Hashimoto Keizô, et al. (review). Technology and Culture, 38(2):499–501, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887336/pdf.

Selman:1999:LWT

[Sel99] J. R. Selman. Lightning wires: The telegraph and China’s technological modernization, 1860–1890. Technology and Culture, 40(3):671– 673, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33375.

Servos:1992:PFS

[Ser92] John W. Servos. Pure Food: Securing the Federal Food and Drugs Act of 1906 by James Harvey Young (review). Technology and Culture, 33(1):166–167, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888356/pdf.

Servos:1994:CPM

[Ser94] John W. Servos. Changing partners: The Mellon Institute, private industry, and the Federal patron. Technology and Culture, 35(2):221– 257, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887917/pdf.

Servos:1996:EBA

[Ser96] John W. Servos. Engineers, businessmen, and the academy: The beginnings of sponsored research at the University of Michigan. Technology and Culture, 37(4):721–762, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887402/pdf.

Sethia:1999:CCT

[Set99] Tara Sethia. Colonialism, chemical technology, and industry in Southern India, 1880–1937. Technology and Culture, 40(3):673–675, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33410.

Shallat:1990:NWH

[Sha90a] Todd Shallat. The National Waterway: a History of the Chesapeake and Delaware Canal, 1769–1985 by Ralph D. Gray (review). Technology and Culture, 31(4):882–883, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/901668/pdf.

Shallat:1990:BWS

[Sha90b] Todd Shallat. Building waterways, 1802–1861: Science and the United States Army in early public works. Technology and Culture, 31(1):18–50, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888918/pdf.

Sharrer:1991:BBH

[Sha91] G. Terry Sharrer. The Business of Breeding: Hybrid Corn in Illinois, 1890–1940 by Deborah Fitzgerald (review). Technology and Culture, 32(4):1117–1119, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888664/pdf.

Sharrer:1993:RCW

[Sha93] G. Terry Sharrer. Robert Cole’s World: Agriculture and Society in Early Maryland by Lois G. Carr, et al. (review). Technology and Culture, 34(3):694–696, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888150/pdf.

Shayt:1994:SVH

[Sha94] David Shayt. The Staircase, vol. 1: History and Theories; vol. 2: Studies of Hazards, Falls, and Safer Design by John Templer (review). Technology and Culture, 35(4):892–894, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887742/pdf.

Shapiro:1995:MLS

[Sha95a] Carolyn G. Shapiro. Medical Lives and Scientific Medicine at Michigan, 1891–1969 ed. by Joel D. Howell (review). Technology and Culture, 36(3):729–731, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887831/pdf.

Shapiro:1995:LHA

[Sha95b] Henry D. Shapiro. The Lawn: a History of an American Obsession by Virginia Scott Jenkins, and: Front Yard America: The Evolution and Meanings of a Vernacular Domestic Landscape by Fred E. H. Schroeder (review). Technology and Culture, 36(4):1042–1044, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887775/pdf.

Sharrer:1995:APA

[Sha95c] G. Terry Sharrer. An Anxious Pursuit: Agricultural Innovation and Modernity in the Lower South, 1730–1815 by Joyce E. Chaplin (review). Technology and Culture, 36(1):177–179, January 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888013/pdf.

Shallat:1996:PFC

[Sha96a] Todd Shallat. Planning the French Canals: Bureaucracy, Politics, and Enterprise under the Restoration by Reed G. Geiger (review). Technology and Culture, 37(1):181–183, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887550/pdf.

Shayt:1996:GGR

[Sha96b] David H. Shayt. The Great Guano Rush: Entrepreneurs and American Overseas Expansion by Jimmy M. Skaggs (review). Technology and Culture, 37(2):363–365, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887485/pdf.

Sherwood:1993:AAH

[She93] Morgan Sherwood. Alaskan Aviation History. Vol. 1: 1897–1928. Vol. 2: 1929–1930 by Robert W. Stevens (review). Technology and Culture, 34(1):167–169, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888481/pdf.

Shedel:1994:TOM

[She94a] James Shedel. Time and Order in Metropolitan Vienna: a Seizure of Schedules by Robert Rotenberg (review). Technology and Culture, 35(1):187–188, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887967/pdf.

Sherow:1994:GTC

[She94b] James E. Sherow. The Great Thirst: Californians and Water, 1770s– 1990s by Norris Hundley (review). Technology and Culture, 35 (1):174–176, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887960/pdf.

Shields:1997:RNT

[Shi97] Mark A. Shields. Resistance to New Technology: Nuclear Power, Information Technology, and Biotechnology ed. by Martin Bauer (review). Technology and Culture, 38(2):535–538, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887354/pdf.

Silverman:1993:SPD

[Sil93] Robert J. Silverman. The stereoscope and photographic depiction in the 19th century. Technology and Culture, 34(4):729–756, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888037/pdf.

Silverman:1999:WCA

[Sil99] M. Joshua Silverman. The writing on the cloud: American culture confronts the atomic bomb. Technology and Culture, 40(1):192–194, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33310.

Simms:1990:HSW

[Sim90] D. L. Simms. The Health and Safety of Workers: Case Studies in the Politics of Professional Responsibility ed. by Ronald Bayer (review). Technology and Culture, 31(1):149–150, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888936/pdf.

Simms:1991:GAB

[Sim91] D. L. Simms. Galen on Archimedes: Burning mirror or burning pitch? Technology and Culture, 32(1):91–96, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888788/pdf.

Simms:1994:KFR

[Sim94] Dennis Simms. Keepers of the Flame: The Role of Fire in American Culture, 1775–1925 by Margaret H. Hazen, Robert M. Hazen (review). Technology and Culture, 35(1):194–195, January 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887971/pdf.

Simms:1995:PTE

[Sim95a] D. L. Simms. Pan’s Travail: Environmental Problems of the Ancient Greeks and Romans by J. Donald Hughes (review). Technology and Culture, 36(2):396–397, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887886/pdf.

Simms:1995:SNE

[Sim95b] Dennis L. Simms. Science and Nature: Essays in the History of the Environmental Sciences ed. by Michael Shortland, and: The Norton History of the Environmental Sciences by Peter J. Bowler (review). Technology and Culture, 36(1):196–197, January 1995. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888024/pdf.

Singer:1993:SPE

[Sin93] Bayla Singer. Science and the Past ed. by Sheridan Bowman (review). Technology and Culture, 34(2):413–415, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888419/pdf.

Sinclair:1995:RMB

[Sin95] Bruce Sinclair. The road to Madison and back: Notes from a traveler. Technology and Culture, 36(2S):S3–S16, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887706/pdf.

Singer:1996:EWD

[Sin96] Bayla Singer. The Electronic Word: Democracy, Technology, and the Arts by Richard A. Lanham (review). Technology and Culture, 37(1):174–175, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887546/pdf.

Singer:1997:VRT

[Sin97] Bayla Singer. Virtual Realities and Their Discontents ed. by Robert Markley (review). Technology and Culture, 38(4):959–960, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887220/pdf.

Singer:1999:HCS

[Sin99a] Bayla Singer. The historiography of contemporary science and technology. Technology and Culture, 40(2):394–395, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33359.

Singer:1999:TNU

[Sin99b] Bayla Singer. Trapped in the net: The unintended consequences of computerization. Technology and Culture, 40(1):201–202, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33308.

Singh:1999:USC

[Sin99c] J. P. Singh. Universal Service: Competition, Interconnection, and Monopoly in the Making of the American Telephone System, and: Telecommunications Competition: The Last Ten Miles. Technology and Culture, 40(2):450–452, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33349.

Skramstad:1991:HMU

[Skr91] Harold K. Skramstad, Jr. History Museums in the United States: a Critical Assessment ed. by Warren Leon, Roy Rosenzweig (review). Technology and Culture, 32(3):651–653, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888730/pdf.

Skramstad:1993:CEM

[Skr93] Harold K. Skramstad, Jr. Cultural Excursions: Marketing Appetites and Cultural Tastes in Modern America by Neil Harris (review). Technology and Culture, 34(2):410–411, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888417/pdf.

Slade:1991:NUE

[Sla91] Joseph W. Slade. Notes on the Underground: an Essay on Technology, Society, and the Imagination by Rosalind Williams (review). Technology and Culture, 32(4):1148–1150, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888681/pdf.

Slade:1992:PFT

[Sla92] Joseph W. Slade. The Place of Fiction in the Time of Science: a Disciplinary History of American Writing by John Limon (review). Technology and Culture, 33(1):204–206, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888379/pdf.

Slade:1995:FIM

[Sla95] Joseph W. Slade. Future Imperfect: The Mixed Blessings of Technology in America by Howard P. Segal (review). Technology and Culture, 36(4):1026–1028, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887766/pdf.

Slade:1997:PST

[Sla97] Joseph W. Slade. Postmodern Sublime: Technology and American Writing from Mailer to Cyberpunk by Joseph Tabbi (review). Technology and Culture, 38(4):957–959, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887219/pdf.

Slade:1998:GMT

[Sla98] Joseph W. Slade. Getting the Message Through: a Branch History of the U.S. Army Signal Corps (review). Technology and Culture, 39(3):591–593, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33201.

Slade:1999:JJT

[Sla99] Joseph W. Slade. James Joyce’s techno-poetics. Technology and Culture, 40(2):444–446, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33361.

Slotten:1995:REF

[Slo95] Hugh Richard Slotten. Radio engineers, the Federal Radio Commission, and the social shaping of broadcast technology: Creating “Radio Paradise”. Technology and Culture, 36(4):950–986, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887758/pdf.

Slotten:1996:RSF

[Slo96] Hugh Richard Slotten. “Rainbow in the sky”: FM radio, technical superiority, and regulatory decision-making. Technology and Culture, 37(4):686–720, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887401/pdf.

Slotten:1998:SAC

[Slo98] Hugh Richard Slotten. Selling the Air: a Critique of the Policy of Commercial Broadcasting in the United States by Thomas Streeter (review). Technology and Culture, 39(1):173–175, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887203/pdf.

Small:1995:LDC

[Sma95] James S. Small. Landmarks in Digital Computing: a Smithsonian Pictorial History by Peggy A. Kidwell, Paul E. Ceruzzi (review). Technology and Culture, 36(3):717–719, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887825/pdf.

Smith:1990:TEG

[Smi90a] Denis Smith. Transitions in Engineering: Guillaume Henri Dufour and the Early 19th Century Cable Suspension Bridges by Tom F. Peters (review). Technology and Culture, 31(3):530–531, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888856/pdf.

Smith:1990:FBS

[Smi90b] George David Smith. The Fall of the Bell System: a Study in Prices and Politics by Peter Temin, Louis Galambos (review). Technology and Culture, 31(1):133–137, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888928/pdf.

Smith:1990:STA

[Smi90c] John Kenly Smith, Jr. The scientific tradition in American industrial research. Technology and Culture, 31(1):121–131, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888926/pdf.

Smith:1991:CCF

[Smi91a] Denis Smith. Claudius Crozet: French Engineer in America, 1790– 1864 by Robert F. Hunter, Edwin L. Dooley, Jr. (review). Technology and Culture, 32(2 (part 1)):411–412, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888754/pdf.

Smith:1991:ITL

[Smi91b] Merritt Roe Smith. Industry, technology, and the “Labor question” in 19th-Century America: Seeking synthesis. Technology and Culture, 32(3):555–570, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888688/pdf.

Smith:1991:ACB

[Smi91c] Robert W. Smith. Astronomer by Chance by Bernard Lovell (review). Technology and Culture, 32(3):636–637, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888722/pdf.

Smith:1991:CTA

[Smi91d] Cecil O. Smith, Jr. Chronique de trente années d’équipement nucléaire à Électricité de France by Georges Lamiral (review). Technology and Culture, 32(2 (part 1)):454–457, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888780/pdf.

Smith:1992:HTX

[Smi92a] Cecil Smith. Histoire des techniques aux XIX^{e} et XX^{e} siècles by Alain Beltran, Pascal Griset (review). Technology and Culture, 33(3):604–605, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888262/pdf.

Smith:1992:DCR

[Smi92b] Karen L. Smith. Damming the Colorado: The Rise of the Lower Colorado River Authority, 1933–1939 by John A. Adams, Jr. (review). Technology and Culture, 33(3):620–621, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888271/pdf.

Smith:1993:CCP

[Smi93a] Bruce L. R. Smith. Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI by Gregg Herken (review). Technology and Culture, 34(2):462–464, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888449/pdf.

Smith:1993:NII

[Smi93b] Merritt Roe Smith. Nature Incorporated: Industrialization and the Waters of New England by Theodore Steinberg (review). Technology and Culture, 34(1):155–157, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888474/pdf.

Smith:1994:TWS

[Smi94] Karen L. Smith. Turning on Water with a Shovel: The Career of Elwood Mead by James R. Kluger (review). Technology and Culture, 35(1):176–177, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887961/pdf.

Smith:1995:BHD

[Smi95a] Karen L. Smith. Building Hoover Dam: an Oral History of the Great Depression by Andrew J. Dunar and Dennis McBride (review). Technology and Culture, 36(2):422–423, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887901/pdf.

Smith:1995:NBA

[Smi95b] Cecil O. Smith, Jr. Le noir et le bleu: 40 ans d’histoire de Gaz de France by Alain Beltran and Jean-Pierre Williot (review). Technology and Culture, 36(1):186–187, January 1995. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888018/pdf.

Smith:1997:SCC

[Smi97a] Carl Smith. Smoldering City: Chicagoans and the Great Fire, 1871–1874 by Karen Sawislak (review). Technology and Culture, 38(2):514–515, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887343/pdf.

Smith:1997:TWS

[Smi97b] Michael O. Smith. Trolley Wars: Streetcar Workers on the Line by Scott Molloy (review). Technology and Culture, 38(1):255–256, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887387/pdf.

Smith:1998:MDR

[Smi98] Pamela H. Smith. The Microscope in the Dutch Republic: The Shaping of Discovery (review). Technology and Culture, 39(4):762–764, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33255.

Smith:1999:MSC

[Smi99] Terry Smith. Machine in the Studio: Constructing the Postwar American Artist (review). Technology and Culture, 40(4):906–908, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33458.

Staudenmaier:1996:MMK

[SMP^{+}96] John M. Staudenmaier, S.J., Robert P. Multhauf, Carroll Pursell, R. Angus Buchanan, August W. Giebelhaus, and Robert C. Post. In memoriam: Melvin Kranzberg (1917–1995). Technology and Culture, 37(3):403–427, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887429/pdf.

Smulyan:1990:EDR

[Smu90] Susan Smulyan. The Early Days of Radio Broadcasting by George H. Douglas (review). Technology and Culture, 31(2):342–343, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888909/pdf.

Smulyan:1993:EAM

[Smu93] Susan Smulyan. Empire of the Air: The Men Who Made Radio by Tom Lewis (review). Technology and Culture, 34(3):701–702, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888154/pdf.

Smulyan:1994:DST

[Smu94] Susan Smulyan. Discovering science and technology through American history. Technology and Culture, 35(4):846–856, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887724/pdf.

Sokal:1993:MKH

[Sok93] Michael M. Sokal. Manufacturing Knowledge: a History of the Hawthorne Experiments by Richard Gillespie (review). Technology and Culture, 34(3):703–704, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888155/pdf.

Solovey:1998:MSP

[Sol98] Mark Solovey. Misunderstanding Science? The Public Reconstruction of Science and Technology ed. by Alan Irwin and Brian Wynne (review). Technology and Culture, 39(2):338–340, April 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887152/pdf.

Spears:1995:MLW

[Spe95] Timothy B. Spears. Magic Lands: Western Cityscapes and American Culture after 1940 by John M. Findlay (review). Technology and Culture, 36(1):194–195, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888023/pdf.

Spicer:1995:HWE

[Spi95] Dag Spicer. The Hubble Wars by Eric Chaisson (review). Technology and Culture, 36(2):435–436, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887909/pdf.

Shapiro-Shapin:1997:ACH

[SS97] Carolyn G. Shapiro-Shapin. American Cardiology: The History of a Specialty and Its College by W. Bruce Fye (review). Technology and Culture, 38(4):974–976, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887230/pdf.

Shapiro-Shapin:1999:HCP

[SS99] Carolyn G. Shapiro-Shapin. Health care policy in contemporary America. Technology and Culture, 40(2):452–454, April 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33345.

Stine:1998:IHT

[ST98] Jeffrey K. Stine and Joel A. Tarr. At the intersection of histories: Technology and the environment. Technology and Culture, 39 (4):601–640, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33271.

Starbuck:1990:BBH

[Sta90a] David R. Starbuck. Bricks and Brickmaking: a Handbook for Historical Archaeology by Karl Gurcke (review). Technology and Culture, 31(2):307–308, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888889/pdf.

Staudenmaier:1990:CPA

[Sta90b] John M. Staudenmaier, S.J. Conspicuous Production: Automobiles and Elites in Detroit, 1899–1933 by Donald Finlay Davis (review). Technology and Culture, 31(2):316–317, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888894/pdf.

Staudenmaier:1990:ASM

[Sta90c] John M. Staudenmaier, S.J. In the Age of the Smart Machine: The Future of Work and Power by Shoshana Zuboff (review). Technology and Culture, 31(1):192–194, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888962/pdf.

Stapleton:1992:EAS

[Sta92a] Darwin H. Stapleton. Early American Scientific and Technical Literature: an Annotated Bibliography of Books, Pamphlets, and Broadsides by Margaret W. Batschelet (review). Technology and Culture, 33(1):184–185, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888367/pdf.

StaudenmaierSJ:1992:CSC

[Sta92b] John M. Staudenmaier S.J. Charles Sheeler and the Cult of the Machine by Karen Lucic (review). Technology and Culture, 33 (4):799–800, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888085/pdf.

Stanley:1994:BTG

[Sta94] Amy Dru Stanley. Beyond the Typewriter: Gender, Class, and the Origins of Modern American Office Work, 1900–1930 by Sharon H. Strom (review). Technology and Culture, 35(2):420–422, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887936/pdf.

Stapleton:1996:CTV

[Sta96a] Darwin H. Stapleton. Channel Tunnel Visions, 1850–1945: Dreams and Nightmares by Keith Wilson (review). Technology and Culture, 37(4):855–856, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887421/pdf.

Staudenmaier:1996:TB

[Sta96b] John M. Staudenmaier, S.J. Taking the baton. Technology and Culture, 37(1):1–4, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887534/pdf.

Stapleton:1997:PWT

[Sta97a] Darwin H. Stapleton. Papers of William Thornton. Vol. 1, 1781– 1802 ed. by C. M. Harris (review). Technology and Culture, 38 (3):760–761, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887291/pdf.

Stapleton:1997:TKA

[Sta97b] Darwin H. Stapleton. Technical Knowledge in American Culture: Science, Technology and Medicine since the Early 1800s ed. by Hamilton Cravens, et al. (review). Technology and Culture, 38 (4):965–966, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887223/pdf.

Staudenmaier:1998:ENP

[Sta98] John M. Staudenmaier. Editor’s note: The prospect before us. Technology and Culture, 39(3):379–382, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33212.

Stanley:1999:C

[Sta99] Autumn Stanley. Communication. Technology and Culture, 40 (1):203–205, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33323.

Stephens:1989:MRT

[Ste89] Carlene Stephens. “The Most Reliable Time”: William Bond, the New England railroads, and time awareness in 19th-Century America. Technology and Culture, 30(1):1–24, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889109/pdf.

Steiner:1990:CCN

[Ste90] Frances H. Steiner. Creating Chicago’s North Shore: a Suburban History by Michael H. Ebner (review). Technology and Culture, 31(2):304–305, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888887/pdf.

Steinberg:1991:ERN

[Ste91a] Theodore Steinberg. Ecological Revolutions: Nature, Gender, and Science in New England by Carolyn Merchant (review). Technology and Culture, 32(2 (part 1)):398–399, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888747/pdf.

Stephens:1991:E

[Ste91b] Carlene E. Stephens. To the Editor. Technology and Culture, 32 (1):185–186, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888826/pdf.

Steele:1992:FYH

[Ste92] Brett D. Steele. The First 75 Years: a History of the Engineering Foundation by Lance E. Metz, Ivan M. Viest (review). Technology and Culture, 33(4):848–849, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888115/pdf.

Steinberg:1993:WFF

[Ste93] Theodore Steinberg. “That World’s Fair feeling”: Control of water in 20th-Century America. Technology and Culture, 34(2):401–409, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888416/pdf.

Steele:1994:MPB

[Ste94a] Brett D. Steele. Muskets and pendulums: Benjamin robins, leonhard Euler, and the ballistics revolution. Technology and Culture, 35(2):348–382, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887921/pdf.

Steinberg:1994:TAP

[Ste94b] Theodore Steinberg. The Trans-Alaska Pipeline Controversy: Technology, Conservation, and the Frontier by Peter A. Coates (review). Technology and Culture, 35(2):437–438, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887946/pdf.

Stewart:1994:KGI

[Ste94c] Larry Stewart. “The King George III Collection” at the Science Museum. Technology and Culture, 35(4):857–863, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887725/pdf.

Steele:1995:HTE

[Ste95] Brett Steele. History of Technology ed. by Graham Hollister-Short, Frank A. J. L. James (review). Technology and Culture, 36(3):732– 734, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887833/pdf.

Stewart:1996:SIS

[Ste96] Larry Stewart. Science and Instruments in Seventeenth-Century Italy by Silvio A. Bedini (review). Technology and Culture, 37 (2):346–347, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887475/pdf.

Steen:1998:CCP

[Ste98] Kathryn Steen. Crusading for Chemistry. The Professional Career of Charles Holmes Herty by Germaine M. Reed (review). Technology and Culture, 39(2):328–330, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887147/pdf.

Sterling:1999:TIH

[Ste99] Christopher H. Sterling. Television: an international history of the formative years. Technology and Culture, 40(2):443–444, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33332.

Stirling:1991:AGA

[Sti91] Dale Stirling. Aviation’s Golden Age: Portraits from the 1920s and 1930s ed. by William M. Leary (review). Technology and Culture, 32(1):167–169, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888816/pdf.

Stine:1998:PEC

[Sti98] Jeffrey K. Stine. The Politics of Expertise in Congress: The Rise and Fall of the Office of Technology Assessment (review). Technology and Culture, 39(4):815–816, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33230.

Stokes:1991:TWG

[Sto91] Raymond G. Stokes. Technology and the West German wirtschaftswunder. Technology and Culture, 32(1):1–22, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888783/pdf.

Stone:1997:SDP

[Sto97] Tammy Stone. The Social Dynamics of Pottery Style in the Early Puebloan Southwest by Michelle Hegmon (review). Technology and Culture, 38(3):748–749, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887284/pdf.

Stokes:1999:CPR

[Sto99] Raymond G. Stokes. Capitalism, politics, and railroads in Prussia, 1830–1870. Technology and Culture, 40(3):669–671, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33382.

Stranges:1990:SFT

[Str90a] Anthony N. Stranges. Synthetic Fuel Technology Development in the United States: a Retrospective Assessment by Michael Crow, et al. (review). Technology and Culture, 31(2):350–352, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888914/pdf.

Stratton:1990:ITS

[Str90b] Michael Stratton. Integrating technological and social history: “Engines of Change” at the National Museum of American History. Technology and Culture, 31(2):271–277, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888879/pdf.

Stranges:1991:CMB

[Str91] Anthony N. Stranges. Canada’s Mines Branch and its synthetic fuel program for energy independence. Technology and Culture, 32(3):521–554, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888687/pdf.

Stueland:1994:OSE

[Stu94] Samuel Stueland. The Otis steam excavator. Technology and Culture, 35(3):571–574, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887838/pdf.

Sullivan:1998:BNY

[Sul98a] Joseph P. Sullivan. Building New York’s Sewers: Developing Mechanisms of Urban Management (review). Technology and Culture, 39(3):579–580, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33189.

Sullivan:1998:SUH

[Sul98b] Woodruff Turner Sullivan. To See the Unseen: a History of Planetary Radar Astronomy (review). Technology and Culture, 39(3):593–594, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/33182.

Sumida:1991:GNB

[Sum91] Jon Tetsuro Sumida. The Great Naval Battle of Ottawa by David Zimmerman (review). Technology and Culture, 32(1):158–160, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888811/pdf.

Summers:1993:DIH

[Sum93] Mark Wahlgren Summers. Destroyer of the Iron Horse: General Joseph E. Johnston and Confederate Rail Transport, 1861–1865 by Jeffrey N. Lash (review). Technology and Culture, 34(1):149–150, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888470/pdf.

Summers:1996:RHL

[Sum96] Gregory Summers. Rivers of History: Life on the Coosa, Tallapoosa, Cahaba, and Alabama by Harvey H. Jackson III (review). Technology and Culture, 37(3):626–627, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887448/pdf.

Sumner:1997:AMV

[Sum97] Gregory D. Sumner. After the Machine: Visual Arts and the Erasing of Cultural Boundaries by Miles Orvell (review). Technology and Culture, 38(4):960–962, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887221/pdf.

Sundin:1997:ALE

[Sun97] Bo Sundin. Anden i lampan: Etnologiska perspektiv på ljus och mörker (The Genie of the Lamp: Ethnological Perspectives on Light and Darkness) by Jan Garnert (review). Technology and Culture, 38(4):970–971, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887226/pdf.

Sundt:1999:EMC

[Sun99] Richard Alfred Sundt. The engineering of medieval cathedrals. Technology and Culture, 40(3):657–658, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33422.

Suttmeier:1990:CBB

[Sut90] Richard R. Suttmeier. China Builds the Bomb by John Wilson Lewis and Xue Litai (review). Technology and Culture, 31(4):903–905, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
901682/pdf.

Sviedrys:1993:HET

[Svi93] Romualdas Sviedrys. The History of Electrical Technology: an Annotated Bibliography by Bernard S. Finn (review). Technology and Culture, 34(3):698–699, July 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888152/pdf.

Swenson:1992:WNM

[Swe92] Loyd S. Swenson, Jr. Where No Man Has Gone Before: a History of Apollo Lunar Exploration Missions by William David Compton (review). Technology and Culture, 33(2):383–385, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888322/pdf.

Swenson:1993:SNP

[Swe93] Loyd S. Swenson. A Spacefaring Nation: Perspectives on American Space History and Policy ed. by Martin J. Collins, Sylvia D. Fries (review). Technology and Culture, 34(4):971–973, October 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888073/pdf.

Szostak:1996:EFW

[Szo96] Rick Szostak. The End of the Future: The Waning of the High-Tech World by Jean Gimpel (review). Technology and Culture, 37 (4):831–832, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887408/pdf.

Talbott:1994:BBI

[Tal94] John E. Talbott. Big Business and Industrial Conflict in Nineteenth-Century France: a Social History of the Parisian Gas Company by Lenard R. Berlanstein (review). Technology and Culture, 35(3):615– 616, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887854/pdf.

Tango:1992:SNH

[Tan92] Gerardo G. Tango. Science and the Navy: The History of the Office of Naval Research by Harvey M. Sapolsky (review). Technology and Culture, 33(1):172–173, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888360/pdf.

Tann:1993:BLC

[Tan93] Jennifer Tann. Before the Luddities: Custom, Community and Machinery in the English Woollen Industry, 1776–1809 by Adrian Randall (review). Technology and Culture, 34(1):177–179, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888486/pdf.

Tapper:1999:DBT

[Tap99] Melbourne Tapper. Drawing blood: Technology and disease identity in Twentieth-Century America. Technology and Culture, 40(3):716– 718, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33412.

Tarver:1995:TTR

[Tar95] W. T. S. Tarver. The traction trebuchet: a reconstruction of an early medieval siege engine. Technology and Culture, 36(1):136–167, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888007/pdf.

Tashjian:1997:TWP

[Tas97] Dickran Tashjian. Trading Words: Poetry, Typography, and Illustrated Books in the Modern Literary Economy by Claire Hoertz Badaracco (review). Technology and Culture, 38(4):955–957, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887218/pdf.

Teich:1999:GAS

[Tei99] Albert H. Teich. This Gifted Age: Science and technology at the Millennium. Technology and Culture, 40(2):392–394, April 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33339.

Tenner:1992:IAN

[Ten92] Edward Tenner. “Information Age” at the National Museum of American History. Technology and Culture, 33(4):780–787, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888082/pdf.

Terry:1994:HSA

[Ter94] Ann Terry. Hagia Sophia: From the Age of Justinian to the Present ed. by Robert Mark and Ahmet Ş. Çakmak (review). Technology and Culture, 35(4):894–896, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887743/pdf.

Theerman:1990:HCV

[The90] Paul Theerman. The History of Cartography. Vol. 1: Cartography in Prehistoric, Ancient, and Medieval Europe and the Mediterranean ed. by J. B. Harley, David Woodward (review). Technology and Culture, 31(3):551–552, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888868/pdf.

Thomas:1990:NCT

[Tho90] Donald E. Thomas, Jr. Nazi “Coordination” of technology: The case of the Bavarian Polytechnical Society. Technology and Culture, 31(2):251–264, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888877/pdf.

Thomas:1991:NPO

[Tho91a] Donald Thomas. New Profession, Old Order: Engineers and German Society, 1815–1914 by Kees Gispen (review). Technology and Culture, 32(3):604–605, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888702/pdf.

Thomson:1991:CIT

[Tho91b] Ross Thomson. Crossover inventors and technological linkages: American shoemaking and the broader economy, 1848–1901. Technology and Culture, 32(4):1018–1046, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888651/pdf.

Thomas:1993:GEP

[Tho93] Donald Thomas. The German Experience of Professionalization: Modern Learned Professions and Their Organizations from the Early Nineteenth Century to the Hitler Era by Charles E. McClelland (review). Technology and Culture, 34(1):146–147, January 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888468/pdf.

Thompson:1994:WIS

[Tho94] Emily Thompson. Wireless Imagination: Sound, Radio, and the Avante-Garde ed. by Douglas Kahn, Gregory Whitehead (review). Technology and Culture, 35(2):425–426, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887939/pdf.

Thompson:1997:ARH

[Tho97a] Emily Thompson. America on Record: a History of Recorded Sound by Andre Millard (review). Technology and Culture, 38(1):256–258, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887388/pdf.

Thompson:1997:SSR

[Tho97b] Emily Thompson. Steinway and Sons by Richard K. Lieberman (review). Technology and Culture, 38(2):523–525, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887348/pdf.

Thompson:1998:SSM

[Tho98] Emily Thompson. Stage to Studio: Musicians and the Sound Revolution, 1890–1950 by James P. Kraft (review). Technology and Culture, 39(1):167–169, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887200/pdf.

Thrall:1991:HHC

[Thr91] Charles A. Thrall. Homework: Historical and Contemporary Perspectives on Paid Labor at Home ed. by Eileen Boris, Cynthia R. Daniels (review). Technology and Culture, 32(2 (part 1)):434–435, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888768/pdf.

Tilly:1990:OCM

[Til90] Louise A. Tilly. Once a Cigar Maker: Men, Women, and Work Culture in American Cigar Factories, 1900–1919 by Patricia A. Cooper (review). Technology and Culture, 31(1):159–161, January 1990. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888943/pdf.

Tiles:1994:NWN

[Til94] Mary Tiles. New Worlds, New Technologies, New Issues ed. by Stephen H. Cutcliffe, et al. (review). Technology and Culture, 35 (1):179–180, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887963/pdf.

Tilly:1997:EWP

[Til97] Louise A. Tilly. European Women and Preindustrial Craft ed. by Daryl M. Hafter (review). Technology and Culture, 38(1):249–251, January 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887384/pdf.

Tann:1996:TTD

[TJ96] Jennifer Tann and R. Glyn Jones. Technology and transformation: The diffusion of the roller mill in the British flour milling industry, 1870–1907. Technology and Culture, 37(1):36–69, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887536/pdf.

Tobey:1992:MSO

[Tob92] Ronald C. Tobey. Making Science Our Own: Public Images of Science 1910–1955 by Marcel C. LaFollette (review). Technology and Culture, 33(1):202–204, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888378/pdf.

Todd:1995:PTC

[Tod95] Edmund Todd. The Politics of Technological Change in Prussia: Out of the Shadow of Antiquity, 1809–1848 by Eric Dorn Brose (review). Technology and Culture, 36(4):1033–1034, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887770/pdf.

Todd:1996:IBP

[Tod96] Edmund N. Todd. Industrialisierung — Begriffe and Prozesse: Festschrift Akos Paulinyi zum 65. Geburtstag ed. by Volker Benad-Wagenhoff (review). Technology and Culture, 37(2):349–351, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/887477/
pdf.

Todd:1997:ICS

[Tod97] Edmund N. Todd. Industrial Constructions: The Sources of German Industrial Power by Gary Herrigel (review). Technology and Culture, 38(3):777–779, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887300/pdf.

Todd:1999:MAG

[Tod99] Edmund N. Todd. Mobilitat fur alle: Geschichte des offentlichen Personennahverkehrs in der Stadt zwischen technischem Fortschritt und sozialer Pflicht (review). Technology and Culture, 40(4):897–898, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33460.

Travis:1990:PMA

[Tra90] Anthony S. Travis. Perkins mauve: Ancestor of the organic chemical industry. Technology and Culture, 31(1):51–82, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888921/pdf.

Travis:1991:EPC

[Tra91] Anthony S. Travis. Engineering and politics: The Channel Tunnel in the 1880s. Technology and Culture, 32(3):461–497, July 1991.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888685/pdf.

Travis:1994:MMM

[Tra94] Anthony S. Travis. From Manchester to Massachusetts via mulhouse: The transatlantic voyage of aniline black. Technology and Culture, 35(1):70–99, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887955/pdf.

Travis:1995:OOP

[Tra95] Anthony S. Travis. Opting for Oil: The Political Economy of Technological Change in the West German Chemical Industry, 1945–1961 by Raymond G. Stokes (review). Technology and Culture, 36(4): 1057–1058, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887783/pdf.

Travis:1996:SME

[Tra96] Anthony S. Travis. Sidney Milton Edelstein (1912–1994). Technology and Culture, 37(1):216–219, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887569/pdf.

Tracy:1997:TDS

[Tra97a] James D. Tracy. Trade and Discovery: The Scientific Study of Artefacts from, Post-Medieval Europe and Beyond ed. by Duncan R. Hook, David R. M. Gaimster (review). Technology and Culture, 38(3):750–751, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887285/pdf.

Travis:1997:LEC

[Tra97b] Anthony S. Travis. Lavoisier in European Context: Negotiating a New Language for Chemistry ed. by Bernadette Bensaude-Vincent and Ferdinando Abbri (review). Technology and Culture, 38(2):508– 510, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887340/pdf.

Travis:1998:AGH

[Tra98a] Anthony S. Travis. “Ambitious and glory hunting … impractical and fantastic”: Heinrich caro at BASF. Technology and Culture, 39(1):105–115, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887172/pdf.

Travis:1998:EFC

[Tra98b] Anthony S. Travis. Edward Frankland: Chemistry, Controversy, and Conspiracy in Victorian England by Colin A. Russell (review). Technology and Culture, 39(1):155–157, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887192/pdf.

Tribe:1990:FBT

[Tri90a] Keith Tribe. From Farming to Biotechnology: a Theory of Agro-Industrial Development by David Goodman, et al. (review). Technology and Culture, 31(1):180–181, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888955/pdf.

Trimble:1990:CAD

[Tri90b] William F. Trimble. Cierva Autogiros: The Development of Rotary-Wing Flight by Peter W. Brooks (review). Technology and Culture, 31(4):893–894, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901675/pdf.

Trilling:1991:LWS

[Tri91] Leon Trilling. Light, Wind, and Structure: The Mystery of the Master Builders by Robert Mark (review). Technology and Culture, 32(2 (part 1)):408–409, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888752/pdf.

Trimble:1993:PAB

[Tri93] William F. Trimble. The Politics of Aircraft: Building an American Military Industry by Jacob A. Vander Meulen, and: Airshipmen, Businessmen, and Politics, 1890–1940 by Henry Cord Meyer (review). Technology and Culture, 34(2):443–445, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888437/pdf.

Trimble:1994:HTE

[Tri94] William F. Trimble. History of Technology ed. by Graham Hollister-Short and Frank A. J. L. James (review). Technology and Culture, 35(3):654–655, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887877/pdf.

Turner:1995:TMC

[Tur95] A. J. Turner. The Time Museum Catalogue of Chronometers by Anthony G. Randall (review). Technology and Culture, 36(1):202–204, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888028/pdf.

Tweedale:1992:SMV

[Twe92] Geoffrey Tweedale. Steel Metropolis: a view of Sheffield industry at Kelham Island Industrial Museum. Technology and Culture, 33(2):328–335, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888295/pdf.

Tympas:1999:IC

[Tym99] Aristotle Tympas. The invention of communication. Technology and Culture, 40(3):694–695, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33398.

Sun:1990:SCC

[tZS90] E tu Zen Sun. Science and Civilisation in China, vol. 5: Chemistry and Chemical Technology, pt. 9: Textile Technology: Spinning and Reeling by Dieter Kuhn (review). Technology and Culture, 31 (4):865–867, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901658/pdf.

Uber:1994:BFT

[Ube94] Terrence L. Uber. Breaking Frame: Technology and the Visual Arts in the Nineteenth Century by Julie Wosk (review). Technology and Culture, 35(3):638–639, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887868/pdf.

Uber:1999:ADT

[Ube99] Terrence L. Uber. American Design in the Twentieth Century: Personality and Performance (review). Technology and Culture, 40 (4):904–906, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33461.

Usselman:1990:YMM

[Uss90] Steven Usselman. 60 Years with Men and Machines by Fred H. Colvin (review). Technology and Culture, 31(2):322–324, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888898/pdf.

Usselman:1991:PPR

[Uss91] Steven W. Usselman. Patents purloined: Railroads, inventors, and the diffusion of innovation in 19th-Century America. Technology and Culture, 32(4):1047–1075, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888652/pdf.

Usselman:1993:DWP

[Uss93] Steven W. Usselman. Daniel Willard and Progressive Management on the Baltimore & Ohio Railroad by David M. Vrooman (review). Technology and Culture, 34(2):432–434, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888430/pdf.

Usselman:1997:CWC

[Uss97] Steven W. Usselman. The Closed World: Computers and the Politics of Discourse in Cold War America by Paul N. Edwards (review). Technology and Culture, 38(4):1008–1010, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887256/pdf.

VanGelder:1996:MKH

[Van96] Lawrence Van Gelder. Melvin Kranzberg, 78, historian of technology. Technology and Culture, 37(3):401–402, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887428/pdf.

VanderMeulen:1997:CAU

[Van97] Jacob Vander Meulen. Commuter Airlines of the United States by R. E. G. Davies and I. E. Quastler (review). Technology and Culture, 38(2):521–529, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887350/pdf.

Vastokas:1998:EPT

[Vas98] Joan M. Vastokas. The Emergence of Pottery: Technology and Innovation in Ancient Societies ed. by William K. Barnett and John W. Hoopes (review). Technology and Culture, 39(2):315–317, April 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887140/pdf.

Verbong:1997:GWM

[Ver97] Geert Verbong. Die Geschichte der Windenergienutzung, 1890–1990 by Matthias Heymann (review). Technology and Culture, 38(3):795– 797, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887310/pdf.

vonHippel:1996:MLD

[vHT96] Eric von Hippel and Margie Tyre. The mechanics of learning by doing: Problem discovery during process machine use. Technology and Culture, 37(2):312–329, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887470/pdf.

Victor:1992:FEW

[Vic92] Stephen K. Victor. “Factory” at the Eli Whitney Museum, Hamden, Connecticut. Technology and Culture, 33(2):336–341, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888296/pdf.

Viehe:1990:ICU

[Vie90] Fred W. Viehe. Instant Cities: Urbanization and the Rise of San Francisco and Denver by Gunther Barth (review). Technology and Culture, 31(2):302–304, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888886/pdf.

Vincenti:1993:NEA

[Vin93] Walter G. Vincenti. NASA Engineers and the Age of Apollo by Sylvia D. Fries (review). Technology and Culture, 34(4):969–971, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888072/pdf.

Vincenti:1994:RAL

[Vin94] Walter G. Vincenti. The retractable airplane landing gear and the Northrop “Anomaly”: Variation-selection and the shaping of technology. Technology and Culture, 35(1):1–33, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887953/pdf. See com-

ment [Con94a, LV95].

Vincenti:1997:ETM

[Vin97] Walter G. Vincenti. Engineering theory in the making: Aerodynamic calculation “Breaks the Sound Barrier”. Technology and Culture, 38 (4):819–851, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887212/pdf.

vanKeuren:1992:FGA

[vK92a] David K. van Keuren. The First Golden Age of Rocketry: Congreve and Hale Rockets of the Nineteenth Century by Frank H. Winter, and: Rockets into Space by Frank H. Winter (review). Technology and Culture, 33(4):836–838, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888108/pdf.

vankeuren:1992:SPM

[vk92b] David K. van keuren. Science, progressivism, and military preparedness: The case of the Naval research laboratory, 1915–1923. Technology and Culture, 33(4):710–736, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888078/pdf.

vanLente:1990:WVW

[vL90] Dick van Lente. Werkplaatsen van Wetenschap en Techniek: Industriële en Academische Laboratoria in Nederland, 1860–1940 ed. by R. P. W. Visser, C. Hakfoort (review). Technology and Culture, 31(1):145–147, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888934/pdf.

vonLindern:1999:FHB

[vL99] Celia von Lindern. Fritz Haber, 1868–1934: Eine Biographie (review). Technology and Culture, 40(4):886–887, October 1999. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33451.

Voeltz:1990:BTE

[Voe90] Richard A. Voeltz. British Technology and European Industrialization: The Norwegian Textile Industry in the Mid Nineteenth Century by Kristine Bruland (review). Technology and Culture, 31(3):524– 525, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888852/pdf.

Volti:1991:DTS

[Vol91a] Rudi Volti. Disappearing through the Skylight: Culture and Technology in the Twentieth Century by O. B. Hardison, Jr. (review). Technology and Culture, 32(4):1151–1152, October 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888683/pdf.

Volti:1991:TWA

[Vol91b] Rudi Volti. Technical Workers in an Advanced Society: The Work, Careers and Politics of French Engineers by Stephen Crawford (review). Technology and Culture, 32(1):141–142, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888801/pdf.

Volti:1991:THF

[Vol91c] Rudi Volti. Toyota: a History of the First 50 Years by Toyota Motor Corporation (review). Technology and Culture, 32(2 (part 1)):423– 424, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888761/pdf.

Volti:1993:DAA

[Vol93] Rudi Volti. Driving Ambitions: an Analysis of the American Hot Rod Enthusiasm by H. F. Moorhouse (review). Technology and Culture, 34(2):419–420, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888422/pdf.

Volti:1994:ACS

[Vol94a] Rudi Volti. America Calling: a Social History of the Telephone to 1940 by Claude S. Fischer (review). Technology and Culture, 35(2):424–425, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887938/pdf.

Volti:1994:WWH

[Vol94b] Rudi Volti. Ways of the World: a History of the World’s Roads and of the Vehicles That Used Them by Maxwell G. Lay (review). Technology and Culture, 35(3):608–609, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/887850/pdf.

Volti:1995:PAM

[Vol95] Rudi Volti. The Petersen Automotive Museum. Technology and Culture, 36(3):646–650, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887796/pdf.

Volti:1996:CA

[Vol96] Rudi Volti. A century of automobility. Technology and Culture, 37 (4):663–685, October 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887400/pdf.

Volti:1998:FTD

[Vol98a] Rudi Volti. Fordism Transformed: The Development of Production Methods in the Automobile Industry (review). Technology and Culture, 39(3):550–552, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33204.

Volti:1998:PUH

[Vol98b] Rudi Volti. Passage to Union: How the Railroads Transformed American Life, 1829–1929 (review). Technology and Culture, 39 (3):581–582, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33190.

VonLindern:1999:JLC

[Von99] Celia Von Lindern. Justus von Liebig: The chemical gatekeeper. Technology and Culture, 40(2):408–409, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33331.

vanRiemsdijk:1990:MLF

[vR90] John van Riemsdijk. La machine locomotive en France des origines au milieu du XIX^{e} siècle by Jacques Payen (review). Technology and Culture, 31(3):534–536, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888858/pdf.

Wachelder:1994:VBC

[Wac94] Joseph Wachelder. Van beroep “Chemiker”: De opkomst van de industriële chemicus en het polytechnische onderwijs in Duitsland (1790–1850) by Ernst Homburg (review). Technology and Culture, 35(3):611–613, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887852/pdf.

Wachtel:1999:HWC

[Wac99] Edward Wachtel. How writing came about. Technology and Culture, 40(1):128–129, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33309.

Walker:1991:RNN

[Wal91] J. Samuel Walker. Rickover and the Nuclear Navy: The Discipline of Technology by Francis Duncan (review). Technology and Culture, 32(2 (part 1)):453–454, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888779/pdf.

Walker:1992:ARB

[Wal92] Mark Walker. Atomic Rivals by Bertrand Goldschmidt (review). Technology and Culture, 33(1):177–179, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888363/pdf.

Walker:1994:MAN

[Wal94] J. Samuel Walker. Mobilizing against Nuclear Energy: a Comparison of Germany and the United States by Christian Joppke (review). Technology and Culture, 35(4):885–886, October 1994. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887737/pdf.

Walton:1995:WTV

[Wal95] Steven Walton. Words of technological virtue: “The battle of brunanburh” and Anglo-Saxon sword manufacture. Technology and Culture, 36(4):987–999, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887759/pdf.

Walker:1999:HNA

[Wal99a] J. Samuel Walker. Heisenberg and the Nazi Atomic Bomb Project: a Study in German Culture (review). Technology and Culture, 40 (4):910–911, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33462.

Walton:1999:AIA

[Wal99b] Steven A. Walton. Artistry and ingenuity in artificial stone: Indiana’s concrete bridges, 1900–1942. Technology and Culture, 40 (3):715–716, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33384.

Wang:1996:SSD

[Wan96] Zuoyue Wang. Scientists and the State: Domestic Structures and the International Context ed. by Etel Solingen (review). Technology and Culture, 37(1):213–215, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887568/pdf.

Warner:1991:VMS

[War91] Deborah Jean Warner. “Vulgar and Mechanick”: The Scientific Instrument Trade in Ireland, 1650–1921 by J. E. Burnett, A. D. Morrison-Low (review). Technology and Culture, 32(3):650–651, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888729/pdf.

Warner:1992:QSE

[War92a] Deborah Warner. The Quantifying Spirit in the Eighteenth Century ed. by Tore Frängsmyr, J. L. Heilbron, and Robin E. Rider (review). Technology and Culture, 33(2):354–356, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888303/pdf.

Warner:1992:PPS

[War92b] Deborah Jean Warner. From Pleasure and Profit to Science and Security: Etienne Lenoir and the Transformation of Precision Instrument-Making in France 1760–1830 by A. J. Turner, and:

Le Citoyen Lenoir: Scientific Instrument Making in Revolutionary France by J. A. Bennett (review). Technology and Culture, 33 (1):158–160, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888351/pdf.

Warner:1993:SDS

[War93] Deborah Warner. Storia delle scienze. Vol. 1: Gli strumenti ed. by Gerard L’E. Turner (review). Technology and Culture, 34(1): 184–185, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888490/pdf.

Warner:1997:TTL

[War97] Deborah Jean Warner. The Tradition of Technology: Landmarks of Western Technology in the Collections of the Library of Congress by Leonard C. Bruno (review). Technology and Culture, 38(3):755– 756, July 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887288/pdf.

Warner:1998:SOL

[War98a] Deborah Warner. At the Sign of the Oriental Lamp: The Musschenbroek Workshop in Leiden. 1660–1750 by Peter de Clercq (review). Technology and Culture, 39(1):141–142, January 1998. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887184/pdf.

Warner:1998:AIT

[War98b] Deborah Jean Warner. Astronomical Instruments and Their Users: Tycho Brahe to William Lassell (review). Technology and Culture, 39(3):557–558, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33183.

Warner:1999:LCP

[War99] Deborah Jean Warner. The Leiden cabinet of physics: a descriptive catalogue. Technology and Culture, 40(1):136–137, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33278.

Washburn:1991:HHH

[Was91] Wilcomb E. Washburn. Hearth and Home: a History of Material Culture by Norman J. G. Pounds (review). Technology and Culture, 32(1):132–133, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888795/pdf.

Wasserman:1992:BEB

[Was92] Mark Wasserman. British Enterprise in Brazil: The St. John d’el Rey Mining Company and the Morro Velho Gold Mine, 1830–1960 by Marshall C. Eakin (review). Technology and Culture, 33(4):803–804, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888088/pdf.

Washburn:1993:SWW

[Was93] Wilcomb E. Washburn. The Skulking Way of War: Technology and Tactics among the New England Indians by Patrick M. Malone (review). Technology and Culture, 34(1):147–149, January 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888469/pdf.

Wayman:1990:AET

[Way90] Michael L. Wayman. Archeomaterials ed. by Tamara Stech (review). Technology and Culture, 31(3):546–548, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888865/pdf.

Webb:1990:MAF

[Web90a] R. K. Webb. Manchester and the Age of the Factory: The Business Structure of Cottonopolis in the Industrial Revolution by Roger Lloyd-Jones, M. J. Lewis (review). Technology and Culture, 31 (3):520–522, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888850/pdf.

Webber:1990:VTA

[Web90b] Kimberley Webber. Victorian Things by Asa Briggs (review). Technology and Culture, 31(3):556–557, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888871/pdf.

Weber:1990:TIR

[Web90c] Wolfhard Weber. Technik und Industrielle Revolution ed. by Theo Pirker, et al. (review). Technology and Culture, 31(3):527–528, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888854/pdf.

Weber:1996:BHR

[Web96] Wolfhard Weber. Blauer Himmel über der Ruhr: Geschichte der Umwelt im Ruhrgebiet, 1840–1990 von Franz-Josef Brüggemeier, Thomas Rommelspacher (review). Technology and Culture, 37(1): 208–209, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887565/pdf.

Weiner:1994:ERJ

[Wei94] Douglas R. Weiner. The Electrification of Russia, 1880–1926 by Jonathan Coopersmith (review). Technology and Culture, 35(2):413– 415, April 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887932/pdf.

Wengenroth:1992:ETB

[Wen92] Ulrich Wengenroth. Erfindergeist und Technikkritik: Der Beitrag Amerikas zur Modernisierung und die Technikdebatte seit 1900 by Andreas Schüler (review). Technology and Culture, 33(1):209–210, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888382/pdf.

Wengenroth:1998:DTD

[Wen98] Ulrich Wengenroth. Does Technology Drive History?: The Dilemma of Technological Determinism (review). Technology and Culture, 39 (4):755–758, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33258.

Wermiel:1992:HTE

[Wer92] Sara Wermiel. History of Technology ed. by Graham Hollister-Short, Frank James (review). Technology and Culture, 33(4):849–851, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888116/pdf.

Wermiel:1997:HBC

[Wer97] Sara Wermiel. Historical Building Construction: Design, Materials, and Technology by Donald Friedman (review). Technology and Culture, 38(2):516–517, April 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887344/pdf.

Westbrook:1990:BUM

[Wes90] Jack H. Westbrook. The Beginnings of the Use of Metals and Alloys ed. by Robert Maddin (review). Technology and Culture, 31(3):544– 546, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888864/pdf.

Westfall:1991:SNL

[Wes91] Catherine Westfall. Sandia National Laboratories: The Postwar Decade by Necah S. Furman (review). Technology and Culture, 32(3):639–641, July 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888724/pdf.

Westbrook:1993:SIV

[Wes93a] Virginia Westbrook. “The Spirit of Invention: da Vinci to Vermont” at the Chaffee Art Center, Rutland. Technology and Culture, 34 (1):114–116, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888460/pdf.

Westfall:1993:BSG

[Wes93b] Catherine Westfall. Big Science: The Growth of Large-Scale Research ed. by Peter Galison, Bruce Hevly (review). Technology and Culture, 34(4):966–967, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888070/pdf.

Westwood:1995:GTA

[Wes95] J. N. Westwood. Guns for the Tsar: American Technology and the Small Arms Industry in Nineteenth-Century Russia by Joseph Bradley (review). Technology and Culture, 36(4):1034–1036, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887771/pdf.

Wesolowski:1999:ETB

[Wes99] Edward Wesolowski. Early television: a bibliographic guide to 1940. Technology and Culture, 40(2):441–443, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33357.

Westfall:1996:TSB

[WH96] Catherine Westfall and Lillian Hoddeson. Thinking small in big science: The founding of Fermilab, 1960–1972. Technology and Culture, 37(3):457–492, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887431/pdf.

White:1990:RSP

[Whi90] John H. White, Jr. Riding in style: Palace cars for the cattle trade. Technology and Culture, 31(2):265–270, April 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888878/pdf.

White:1992:RAA

[Whi92a] Roger B. White. Roadside America: The Automobile in Design and Culture ed. by Jan Jennings (review). Technology and Culture, 33 (1):198–200, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888376/pdf.

White:1992:NCR

[Whi92b] John H. White, Jr. The North Carolina Railroad, 1849–1871, and the Modernization of North Carolina by Allen W. Trelease (review). Technology and Culture, 33(2):378, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888318/pdf.

Whitney:1992:MAU

[Whi92c] Elspeth Whitney. Medieval Artisans: an Urban Class in Late Medieval England by Heather Swanson (review). Technology and Culture, 33(2):351–352, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888301/pdf.

White:1993:GBR

[Whi93] Roger B. White. Galloping Bungalows: The Rise and Demise of the American House Trailer by David A. Thornburg (review). Technology and Culture, 34(1):169–171, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888482/pdf.

Whiteside:1998:SFT

[Whi98] James Whiteside. Safety First: Technology, Labor, and Business in the Building of American Work Safety, 1870–1939 (review). Technology and Culture, 39(3):586–588, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33176.

Wigen:1996:HCV

[Wig96] Kären Wigen. The History of Cartography, vol. 2, bk. 2, Cartography in the Traditional East and Southeast Asian Societies ed. by J. B. Harley, David Woodward (review). Technology and Culture, 37(2):340–341, April 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887472/pdf.

Williams:1990:BLF

[Wil90] Michael R. Williams. “Beyond the Limits: Flight Enters the Computer Age” at the National Air and Space Museum. Technology and Culture, 31(4):838–845, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/901654/pdf.

Williams:1991:WET

[Wil91] Rosalind Williams. “When the Eiffel Tower was new” at the MIT Museum. Technology and Culture, 32(1):102–105, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888790/pdf.

Wilde:1992:SGM

[Wil92a] Mark Wilde. Satisfaction Guaranteed: The Making of the American Mass Market by Susan Strasser (review). Technology and Culture, 33(2):364–365, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888309/pdf.

Wilson:1992:JCN

[Wil92b] Richard Guy Wilson. J. C. Nichols and the Shaping of Kansas City: Innovation in Planned Residential Communities by William S. Worley (review). Technology and Culture, 33(1):189–190, January 1992.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888370/pdf.

Williams:1993:SWC

[Wil93a] Rosalind Williams. Strange Weather: Culture, Science, and Technology in the Age of Limits by Andrew Ross (review). Technology and Culture, 34(2):459–461, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888447/pdf.

Wilson:1993:CSS

[Wil93b] Lindsay Wilson. Colonialism and Science: Saint Domingue in the Old Regime by James E. McClellan (review). Technology and Culture, 34(4):933–935, October 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888051/pdf.

Wilkins:1995:LTC

[Wil95a] Mira Wilkins. Learning and Technological Change ed. by Ross Thomson (review). Technology and Culture, 36(1):173–175, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888011/pdf.

Williams:1995:EWH

[Wil95b] James C. Williams. Energy in World History by Vaclav Smil (review). Technology and Culture, 36(3):690–692, July 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/887810/pdf.

Williams:1995:HTE

[Wil95c] James C. Williams. History from Things: Essays on Material Culture ed. by Steven Lubar and W. David Kingery (review). Technology and Culture, 36(1):175–177, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888012/pdf.

Wilkie:1999:HAS

[Wil99a] Laurie A. Wilkie. Historical archaeology and the study of American culture. Technology and Culture, 40(3):654–655, July 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33386.

Williams:1999:CPS

[Wil99b] James C. Williams. Consuming power: a social history of American energies. Technology and Culture, 40(2):411–413, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33334.

Winters:1990:ECK

[Win90] Donald L. Winters. The Emergence of the Cotton Kingdom in the Old Southwest, 1770–1860 by John Hebron Moore (review). Technology and Culture, 31(1):177–178, January 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888953/pdf.

Winer:1998:FEW

[Win98] Daniel Winer. Fire Engines with Water Hoses and the Method of Fighting Fires Now Used in Amsterdam by Jan van der Heyden, Sr. and Jan van der Heyden, Jr. (review). Technology and Culture, 39 (1):139–141, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887183/pdf.

Wistoft:1998:PTP

[WN98] Birgitte Wistoft and Henry Nielsen. Painting technological progress: P. S. Kroyer’s The Industrialists. Technology and Culture, 39(3):408– 433, July 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33217.

Woodward:1992:EAE

[Woo92] David R. Woodward. England and the Aeroplane: an Essay on a Militant and Technological Nation by David Edgerton (review). Technology and Culture, 33(4):834–836, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888107/pdf.

Worthington:1990:PPM

[Wor90] William Worthington, Jr. The privy and the pump: The matthewman & Johnson excavating device. Technology and Culture, 31 (3):451–455, July 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888831/pdf.

Worthington:1991:CAE

[Wor91] William Worthington. The Colossus of 1812: an American Engineering Superlative by Lee H. Nelson (review). Technology and Culture, 32(2 (part 1)):412–414, April 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888755/pdf.

Worthington:1993:LTS

[Wor93a] William Worthington. Lighting the Town: a Study of Management in the North West Gas Industry, 1805–1880 by John F. Wilson (review). Technology and Culture, 34(4):938–940, October 1993. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888054/pdf.

Worthington:1993:TSS

[Wor93b] William Worthington. Tidewater by Steamboat: a Saga of the Chesapeake by David C. Holly (review). Technology and Culture, 34 (1):152–153, January 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888472/pdf.

Worthington:1994:FWI

[Wor94] William Worthington. Ferris Wheels: an Illustrated History by Norman D. Anderson (review). Technology and Culture, 35(1):201–203, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
887975/pdf.

Worthington:1995:CWF

[Wor95] William Worthington. A Century in the Works: Freese and Nichols Consulting Engineers, 1894–1994 by Simon W. Freese, Deborah Lightfoot Sizemore (review). Technology and Culture, 36(4): 1031–1033, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887769/pdf.

Wormbs:1998:MSS

[Wor98] Nina Wormbs. Making Sense of Space: The History of Norwegian Space Activities (review). Technology and Culture, 39(4):818–820, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
33235.

Wosk:1991:PCA

[Wos91] Julie Wosk. Postmodern Currents: Art and Artists in the Age of Electronic Media by Margot Lovejoy (review). Technology and Culture, 32(1):138–140, January 1991. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888799/pdf.

Wosk:1992:TIC

[Wos92] Julie Wosk. The Telling Image: The Changing Balance between Pictures and Words in a Technological Age by Duncan Davies, et al. (review). Technology and Culture, 33(4):845–846, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888113/pdf.

Wosk:1994:REV

[Wos94] Julie Wosk. The Reconfigured Eye: Visual Truth in the Postphotographic Era by William J. Mitchell (review). Technology and Culture, 35(3):640–642, July 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887869/pdf.

Wright:1992:GPA

[Wri92] Helena E. Wright. George Pullman and the Allen paper car wheel. Technology and Culture, 33(4):757–768, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888080/pdf.

Wright:1994:OPA

[Wri94] Helena E. Wright. The Origins of Photojournalism in America by Michael L. Carlebach (review). Technology and Culture, 35 (1):197–199, January 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887973/pdf.

Wright:1996:NMS

[Wri96] Thomas Wright. The National Museum of Science and Industry: an overview. Technology and Culture, 37(1):147–150, January 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/887540/pdf.

Yates:1992:AIP

[Yat92] JoAnne Yates. The Age of Information: The Past Development and Future Significance of Computing and Communications by Stephen Saxby (review). Technology and Culture, 33(3):627–628, July 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888275/pdf.

Yates:1996:EBB

[Yat96] JoAnne Yates. Exploring the Black Box: Technology, Economics, and History by Nathan Rosenberg (review). Technology and Culture, 37(3):617–619, July 1996. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887443/pdf.

Yeager:1990:LLF

[Yea90] Mary A. Yeager. Livestock Legacy: The Fort Worth Stockyards, 1887–1987 by J’Nell L. Pate (review). Technology and Culture, 31 (4):884–886, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/901670/pdf.

Hsu:1998:SCC

[yH98] Cho yun Hsu. Science and Civilisation in China. Vol. 6, Biology and Biological Technology. Pt. 3, Agro-Industries and Forestry by Christian Daniels and Nicholas K. Menzies (review). Technology and Culture, 39(1):162–163, January 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887196/pdf.

Young:1993:HPR

[You93] Terence Young. Hard Places: Reading the Landscape of America’s Historic Mining Districts by Richard V. Francaviglia (review). Technology and Culture, 34(2):422–424, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888424/pdf.

Youngflesh:1995:OTA

[You95] V. Kay Youngflesh. The Ordering of Time: From the Ancient Computus to the Modern Computer by Arno Borst (review). Technology and Culture, 36(2):437–438, April 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887910/pdf.

Yu:1999:NST

[Yu99] Jui-Chen Yu. The National Science and Technology Museum of Taiwan. Technology and Culture, 40(1):107–113, January 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33302.

Zachary:1998:HDD

[Zac98] G. Pascal Zachary. Hellions of the Deep: The Development of American Torpedoes in World War II (review). Technology and Culture, 39(4):809–811, October 1998. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/33242.

Zachmann:1999:PZS

[Zac99] Karin Zachmann. Vom Pianotyp zur Zeilensetzmaschine: Setzmaschinenentwicklung und Geschlechterverhaltnis 1840–1900. Technology and Culture, 40(2):434–436, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33354.

Zaknic:1992:SMB

[Zak92] Ivan Zaknic. Skyscraper: The Making of a Building by Karl Sabbagh (review). Technology and Culture, 33(1):196–198, January 1992. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888375/pdf.

Zeitlin:1995:FMP

[Zei95] Jonathan Zeitlin. Flexibility and mass production at war: Aircraft manufacture in Britain, the United States, and Germany, 1939– 1945. Technology and Culture, 36(1):46–79, January 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888004/pdf.

Zeleza:1997:PPA

[Zel97] Paul Tiyambe Zeleza. People of the Plow: an Agricultural History of Ethiopia, 1800–1990 by James C. McCann (review). Technology and Culture, 38(4):980–982, October 1997. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/887235/pdf.

Zembala:1992:PMH

[Zem92] Dennis Zembala. “Power in motion” at the Henry Ford Museum. Technology and Culture, 33(2):342–346, April 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888297/pdf.

Zembala:1993:IYR

[Zem93] Dennis Zembala. Ingenious Yankees: The Rise of the American System of Manufactures in the Private Sector by Donald R. Hoke (review). Technology and Culture, 34(2):416–417, April 1993. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888420/pdf.

Ziegler:1992:MDP

[Zie92] Charles A. Ziegler. The Moby Dick Project: Reconnaissance Balloons over Russia by Curtis Peebles (review). Technology and Culture, 33 (4):833–834, October 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/888106/pdf.

Ziegler:1994:WDC

[Zie94] Charles A. Ziegler. Weapons development in context: The case of the World War I balloon bomber. Technology and Culture, 35 (4):750–767, October 1994. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/887721/pdf.